A Complete Bibliography of Publications in Computational Statistics & Data Analysis (2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.02

Title word cross-reference

(d, r) [192]. 1 [972]. 2 [116]. 2^{n−k} [41]. 2 × 2 [268, 642, 2170, 1235, 1286, 1911, 888]. 2 × c [1044]. 3 [1311, 1574, 425, 1482, 1748]. ^{(}R) [1925, 1854, 1852]. _{50} [138]. _{90} [280]. A [1453, 700, 963]. α [1202, 2133, 905]. ARH(1) [1381]. B [656]. β [1526, 849]. C [1693, 883, 1949]. C_{p} [1769, 901]. D [32, 1954, 168, 700, 636, 1913, 963]. d > 2 [2151]. E [700]. ℓ_{1} [793]. ϵ [1121]. F [94, 1486, 246, 2038, 1802]. G(λ) [1154]. Γ [115, 1031]. GARCH(1, 1) [1140]. I [2185, 2119, 814, 993, 1476]. ∞ [1426]. K [98, 1117, 1047, 1860, 247, 1371, 1460, 974, 771, 785, 1226, 2148, 2231, 1907, 802, 144, 1355]. L [1342, 1905, 1641, 415, 1078, 1663]. L^{2} [647]. L_{1} [149, 880, 276, 1097, 84, 213, 701]. L_{2} [1760, 2314]. L_{p} [2237]. L_{q} [1494]. M [791, 712, 2256, 1696]. W [2281]. m × k [1057]. O(N) [2193]. P [2146, 852, 100, 1193, 1702, 1498, 1464, 2247, 1953, 1050, 1574, 1690, 626, 102, 1295, 2311, 2293]. P (X < Y) [2257]. ϕ [409, 1089]. πps [1282]. Pr(X < Y) [1770]. Pr{Y > X} [1145]. Q [149]. Q_{n} [2137]. R [452, 1389, 907]. R^{2} [390, 75, 1658, 1640]. r × c [1223]. S [1141]. S^{2} [1045]. s^{n} [957]. t [411, 1222, 1899, 1762, 55, 2295, 2006, 2, 1411, 1869]. τ [1182, 152]. T L [1342].

-additive [785]. -centroids [1047]. -circulant [1226, 2148]. -components [1907]. -D [1311]. -distribution [1526, 1411]. -distributions [1869, 1141].

-divergence [409, 1089]. -estimates [791, 712, 1389]. -estimation [1905]. -expectation [849]. -factor [907]. -means [1371, 883, 771, 144, 1693, 1860]. -measures [75]. -minimax [115, 963]. -mode [149]. -moments [1342, 1641, 415, 1078, 1663]. -nearest-neighbor [247]. -nearest-neighbour [1460]. -nomial [1574]. -norm [880]. -normal [1899]. -optimal [1949, 1453, 32, 1954, 168, 700, 1913]. -out [1690]. -partitioning [1057]. -ranks [802]. -Sample [1355, 1696, 2231]. -spline [1702]. -Splines [1953, 852, 1193, 2247]. -squared [452]. -stable [1202, 2133, 905]. -stationary [1154]. -tests [1762, 2038]. -type [2237, 1640]. -value [100]. -values [1574, 626, 1498, 102, 1295, 2293, 2146, 1050].

1 [1638]. 100K [1925, 2345]. 1990s [1106].

2000 [1111, 291]. 2003 [679, 787]. 2007 [1851, 1850, 1855, 1854, 1852]. 21st [4]. 2nd [790, 1516, 1580].

31 [30]. 33 [277]. 34 [120].

46 [767]. 46/1 [767]. 47 [693, 1303, 939]. 47/1 [693]. 49 [826, 827]. 49/4 [827].

50 [1304]. 51 [2347]. 52 [2348, 2345]. 53 [2344, 2067, 2346].

7 [1361]. 74 [2345].

9/11 [2243]. 978 [1361]. 978-0-387-30369-7 [1361].

A-LMSV [1727]. abrupt [245]. absence [1618]. absolute [687]. Absorption [1322]. abuse [1930]. accelerated [2152, 2051, 2168, 1337]. Accelerating [1121]. Acceleration [1281]. accident [1100]. according [1560]. Accounting [2112]. Accuracy [1898, 1109, 679, 1522, 351, 291, 787, 1851, 1995, 1852]. Accurate [1175, 2146, 1073, 188, 1708, 1306]. ACD [1167]. achieve [521]. acid [263]. Acknowledgement [806]. across [230]. ACT [1321, 551]. action [433]. activity [1739, 1305]. actual [1989]. actually [7]. acupuncture [1991]. acute [2003]. AdaBoost [1189]. adapting [352]. Adaptive [2058, 2167, 1765, 1784, 1836, 1572, 1821, 64, 2155, 1452, 1331, 1193, 860, 1494, 1614, 95, 1424, 221, 2247, 746, 1916]. Addendum [767]. addition [1097]. Additional [1157, 1698]. Additive [2242, 2057, 1150, 808, 1219, 141, 1932, 1359, 1919, 852, 205, 1884, 785, 1757, 1953, 2147, 1940, 2211]. additivity [564]. adequacy [2076]. ADF [2334]. adherence [1791]. Adjusted [1640, 1257, 1908, 984, 1295, 2171]. adjusting [1623]. adjustment [1624, 1280, 920, 703, 212, 1429, 29, 282, 1860]. Adjustments [75]. admixture [2091]. Advances [1401]. aerosol [708]. affected [857]. affects [1447]. affine [743]. affinity [2087]. Affymetrix [2345, 1925]. Africa [2232]. AFT [886]. after [1997, 2003, 2243]. against [778, 1205, 2131, 416, 1151]. age [1105, 1832]. age-dependent [1105]. agent [370]. Agglomerative [559]. aggregate [1739]. Aggregating [645]. Aggregation [1829, 732, 2145, 1591, 15, 57]. agreement [1976, 1483, 1117, 1924, 2024]. AIC [1703]. aid [1526]. AIDS [1975]. air [1330, 320]. Akaike [945, 884, 1665]. alcohol [921]. algebra [306]. algebraic [1316]. algorithm [508, 613, 159, 2110, 2201, 1281, 341, 1419, 465, 289, 2253, 106, 90, 1208, 815, 1520, 529, 717, 949, 91, 150, 2128, 796, 656, 261, 1177, 1624, 1279, 1118, 342, 2012, 1701, 1521, 1049, 152, 2018, 1121, 938, 624, 697, 658, 856, 2148, 188, 615, 566, 166, 830, 921, 2053, 1188, 235, 343, 115, 374, 2304, 1428, 1835, 1950, 286, 1042, 1122, 956, 213, 373, 1142, 282, 1911, 1223, 24, 277, 1860, 832, 746, 313, 256, 1225, 1216, 1512, 2016, 1996, 265, 819, 459, 934, 1660, 1805, 1044]. algorithmic [1915]. Algorithms [1519, 1192, 1528, 1884, 1564, 1573, 132, 141, 614, 1574, 1035, 226, 892, 611, 1241, 640, 1728, 1954, 439, 2294, 1518, 1412, 493, 1607, 308, 1226, 1043, 40, 1221, 674, 1432, 1600, 660, 1152, 740, 741, 617, 1870, 1869, 894, 1125, 1248, 1440, 850, 612, 192, 749, 1068, 1198]. alias [224]. aligned [387]. alignment [1816, 1970]. allele [1240]. allelic [2047]. allocation [2200, 1231]. allowing [2288]. Alternating [81, 308, 1043]. Alternative [1253, 886, 699, 2008, 778, 1263, 91, 656, 2131, 2205, 1151, 1769, 1337]. alternatives [1205, 978]. AMCMC [1424]. amino [263]. among [1662, 1454, 2180]. Anal [2344, 120, 826, 2067, 939, 2347, 2348, 2346, 277, 2345]. analyses [163, 1037, 1431, 503, 1475, 2236, 1329, 1661, 462, 2082, 1325]. Analysing [126, 1187, 1327, 1586]. Analysis [1963, 1745, 375, 693, 1303, 1728, 949, 516, 1117, 937, 1304, 1115, 460, 827, 888, 2339, 2002, 108, 629, 211, 219, 1519, 930, 25, 1102, 353, 1330, 16, 2110, 1753, 1051, 2203, 1722, 677, 6, 647, 946, 2287, 1171, 1232, 1810, 1345, 98, 1623, 2117, 2262, 765, 90, 149, 880, 3, 1013, 1612, 1988, 733, 133, 2255, 1887, 766, 1754, 73, 354, 1029, 972, 655, 310, 1024, 94, 728, 543, 2170, 200, 2349, 278, 1404, 30, 1130, 2259, 2319, 1739, 797, 1116, 1266, 1991, 1034, 1022, 1620, 2204, 1027, 1361, 1055, 1705, 1472, 2166, 1292, 502, 911, 1417]. analysis [83, 1154, 2341, 1118, 2343, 456, 2196, 1410, 182, 1583, 1003, 1826, 1994, 570, 1565, 301, 938, 846, 821, 515, 224, 1047, 1669, 798, 276, 899, 1771, 2301, 1703, 2306, 1559, 240, 1156, 1097, 1379, 2055, 1568, 186, 17, 687, 147, 1752, 337, 2244, 2073, 487, 269, 1016, 1784, 1444, 2137, 2101, 236, 1134, 1759, 928, 1240, 1806, 1986, 619, 780, 1147, 2059, 2205, 2076, 102, 1748, 1138, 31, 1421, 1870, 42, 1133, 768, 2047, 1644, 1615, 2328, 701, 2279, 983, 8, 1007, 769, 1315, 1815, 1777, 157, 933, 924, 2097, 758, 803, 1548, 2209, 445, 1391, 1550, 275, 1283]. analysis [1593, 489, 430, 2100, 1321, 2272, 144, 2212, 462, 1020, 313, 1763, 2199, 2207, 1385, 2030, 128, 1356, 1434, 2083, 915, 1357, 1970, 858, 1339, 2210, 934, 1201, 1786, 1148, 792, 2213, 2206, 1396]. analysts [2063]. Analytic [1912]. analyzers [326, 1411, 1805]. Analyzing [573, 2085, 1429, 986]. and/or [2336, 1557]. Andersen [1889]. angle [1536]. angular [791]. animal [189]. animals [455]. annealing [159, 500, 1897]. annotated [2297]. anomaly [487]. ANOVA [33, 603, 1446, 1910, 1678]. applicability [1169]. Application [2340, 2201, 2058, 1270, 1123, 1709, 1329, 661, 1444, 1427, 1817, 1122, 1305, 1572, 2100, 1466, 2002, 129, 1531, 2110, 2159, 1281, 365, 1307, 2338, 1872, 1832, 1988, 335, 916, 283, 94, 2032, 538, 1246, 1991, 2028, 1733, 796, 845, 99, 1482, 911, 2295, 1063, 1764, 2037, 555, 1463, 836, 1997, 70, 517, 2077, 542, 2309, 856, 1840, 1990, 920, 921, 1089, 468, 1095, 2003, 619, 1662, 1103, 15, 887, 2021, 1582, 1108, 1708, 673, 1259, 2096, 324, 1686, 2269, 1098, 1939]. Applications [1965, 1516, 83, 162, 166, 1380, 1609, 786, 610, 1930, 932, 1693, 1528, 167, 997, 465, 841, 820, 405, 85, 304, 432, 311, 1571, 1737, 1267, 1923, 976, 884, 1577, 2322, 727, 1160, 632, 799, 1975, 134, 52, 691, 2318, 1301, 2138, 1794, 281, 758, 1775, 1384, 2156, 294, 1050, 1552]. Applied [1745, 1276, 2117, 1266, 251, 707, 1699, 76, 1748, 286, 1687]. applying [831, 2064]. appraising [2087]. approach [653, 238, 189, 211, 268, 1173, 2162, 112, 271, 1396, 6, 1901, 1785, 1488, 647, 1720, 1553, 190, 2062, 1758, 2118, 556, 1013, 133, 73, 1743, 283, 248, 795, 1143, 2321, 446, 1736, 1579, 590, 2028, 1021, 1984, 1504, 1823, 524, 1000, 2123, 1431, 1619, 564, 865, 544, 1377, 1893, 1521, 2085, 669, 831, 1261, 1446, 676, 2277, 992, 1871, 1202, 1673, 1052, 922, 2320, 1699, 101, 337, 1702, 2073, 135, 1915, 1774, 1945, 1240, 1265, 2035, 1099, 1429, 480, 1427, 2274, 617, 1324, 620, 956, 755, 93, 1815, 1694, 530, 1461, 804, 1742, 255, 2082, 469, 809]. approach [1715, 1796, 1325, 967, 968]. approaches [1330, 43, 2298, 1253, 1751, 280, 1711, 886, 401, 1309, 1081, 1158, 445, 2083, 419]. appropriate [1777, 521]. Approximate [941, 1718, 840, 1758, 1308, 2182, 2037, 978, 254, 1962, 1581, 1054, 1528, 1578, 1310, 590, 492, 2260, 2088, 1708]. approximately [963]. Approximating [776, 1262, 1802, 2211]. Approximation [279, 1812, 889, 1764, 555, 666, 702, 2338, 590, 55, 1898, 2158, 2248, 1594, 2, 246, 2007, 1990, 1188, 495, 263, 1059, 1542, 1306, 1616]. approximations [1222, 85, 193, 1491, 1073, 1912, 1413, 905, 1670, 1287, 1716, 1686]. arbitrage [1177]. arbitrage-free [1177]. Arbitrarily [1298, 500]. arbitrary [465, 2121, 1511]. ARCH [1139, 1187]. ARCH-M [1187]. archaeology [20, 524]. Archaeometry [691]. Archimedean [1778, 1906, 1489]. area [1204, 2189, 2196, 2254, 902]. areas [2190]. ARFIMA [1081]. ARIMA [1747]. arise [587]. arising [301, 1629]. ARL [1045]. ARMA [1005, 1674, 1480, 1829]. Arnold [1124]. arranging [24, 277]. Array [2345, 1925]. arrays [1915]. Arslan [1214]. artificial [1519]. ARX [1257]. Ascent [1063]. ascertained [2097]. aspect [2287, 479, 546]. aspects [385, 611, 372, 312, 1282, 552]. assay [1892]. assess [2041, 948, 147, 2063]. Assessing [873, 1924, 587, 1842, 1662, 853, 1976, 1483, 909, 2331, 116]. Assessment [666, 1332, 1109, 1143, 1537, 727, 560, 585, 1601]. asset [2113, 618, 177]. associated [1714, 360]. association [508, 608, 907, 2194, 1754, 208, 1353, 449, 252, 2092, 1864, 1561, 1142, 2094, 462, 1041]. associations [1488, 186, 1240]. assumed [138]. Assumption [2076]. assumptions [59, 569, 2168]. astronomical [839]. Asymmetric [1102, 2122, 61, 15, 378, 919, 1166, 1641, 815, 1002, 241, 1730, 1861, 1262]. asymmetrically [2097]. asymmetry [784, 827]. Asymptotic [167, 497, 1481, 67, 343, 757, 848, 1985, 332, 508, 642, 227, 1408, 2237, 19, 811, 1060, 1254, 754]. Asymptotical [268, 1362]. asymptotics [295]. atmosphere [1104]. atmospheric [1376]. attribute [2236]. attributes [1561]. attributing [189]. auctions [1063]. augmentation [1510]. augmented [1280]. Australian [1276]. Author [105, 140, 173, 210, 244, 298, 349, 394, 436, 482, 540, 594, 650, 709, 788, 1010, 1675, 2348, 31]. auto [1280]. auto-logistic [1280]. autocorrelated [137, 1293, 700]. autocorrelation [2126, 2189, 176, 952]. autocovariances [1737]. autologistic [1935]. Automatic [622, 1059, 1772, 1076, 920, 638, 703, 96, 769]. autoradiographic [2341]. autoregressions [1407, 117, 1688]. autoregressive [863, 614, 1928, 1271, 1655, 372, 1165, 999, 1782, 1931, 1147, 1455, 2147, 93, 417, 1338, 2197, 192, 1322]. Auxiliary [1267, 1856, 180, 1426, 1961, 536]. available [1635]. average [372, 1762, 1100, 1147, 1593, 36]. averaging [1720, 835, 2076]. Avoiding [1797]. axes [514, 772]. axial [232]. Azen [438].

B [1138, 1794]. B-spline [1794]. B-splines [1138]. background [456]. BACON [64, 1049]. bagged [2000]. Bagging [2042, 199, 1545, 777, 197]. balance [1453, 454]. balanced [1247, 849, 579, 1677]. Balancing [2078]. bands [1483, 2303, 1308, 61]. Bandwidth [123, 992, 127, 1722, 136, 499, 1695, 599, 1451, 1772, 968]. bank [1522]. Bartholomew [52]. base [266, 146]. base-line [266]. based [189, 1064, 461, 2162, 1959, 378, 2251, 1023, 1396, 1633, 1051, 2233, 990, 1744, 1488, 852, 947, 1345, 2179, 2027, 2310, 1532, 1676, 1846, 813, 1958, 133, 1626, 283, 1352, 1679, 311, 1040, 1797, 1090, 910, 1789, 1374, 1556, 2298, 1073, 370, 1513, 1613, 884, 2095, 1175, 2294, 539, 564, 191, 481, 2290, 1118, 2196, 1974, 669, 898, 1261, 2033, 726, 1065, 559, 785, 515, 1082, 1669, 1475, 992, 607, 534, 1492, 1378, 1935, 1673, 952, 2280, 409, 1089, 905, 180, 1847, 1387, 347, 86, 35, 386, 980, 1304, 568, 1761, 1095, 1793, 1862, 1101, 1467]. based [1070, 1839, 1099, 237, 2024, 617, 1381, 743, 1455, 396, 1341, 549, 767, 2088, 1496, 1554, 1296, 530, 1625, 1391, 1093, 2228, 1651, 1046, 1957, 1572, 2100, 1321, 1041, 1539, 1287, 686, 931, 450, 962, 2240, 270, 1792, 2258, 2153, 2034, 1524, 1796, 934, 1355, 2060, 749, 1198, 2293]. baseline [1628, 362, 346, 2230]. baseline-category [1628]. baseline-risk [346]. bases [795]. basis [2023, 737, 1501]. Baumgartner [1387, 386]. Bayes [238, 2296, 1232, 1001, 1229, 1493, 2191, 2308, 2071, 1652, 2190]. Bayesian [827, 1327, 219, 2108, 2072, 353, 1105, 1199, 1638, 760, 1770, 711, 1901, 852, 223, 1899, 2310, 2062, 2112, 698, 813, 1196, 1578, 2270, 339, 1824, 65, 1965, 1310, 204, 2170, 46, 1090, 447, 150, 1991, 2028, 2298, 1642, 1613, 2232, 706, 2315, 524, 356, 2289, 1705, 1619, 1624, 1776, 2182, 191, 481, 1069, 1353, 1893, 1741, 2341, 1193, 110, 1410, 669, 1997, 2288, 1671, 1656, 1489, 2056, 1931, 1873, 2009, 230, 317, 134, 1202, 2191, 2055, 1052, 174, 1413, 2320, 1865, 1155, 1300, 2017, 1224, 1904, 1474, 2101, 236, 1134, 76, 1101, 784, 1688, 780, 835, 1779]. Bayesian [1147, 1429, 2138, 2271, 2274, 1421, 636, 71, 743, 2247, 488, 2047, 2099, 1962, 1907, 1315, 1158, 281, 2268, 1228, 1708, 924, 2302, 1790, 1338, 649, 2228, 1593, 1742, 1306, 2161, 1710, 469, 1339, 1909, 2090, 1325, 1389, 968, 1786, 1505, 1098, 1973]. before [2173, 522]. behavioral [1847]. behaviour [1638, 432, 208, 171]. Behrens [1398]. benchmark [84]. benchmarking [2234]. bent [305]. Berlin [1361]. Bernoulli [1751, 1639]. bespoke [1845]. Bessel [303]. Best [216, 7, 1518, 750, 1101, 1259]. Beta [566, 44, 1838, 2084, 624, 1079, 2347]. beta-binomial [2084]. Beta-distributed [566]. Better [2208, 307]. between [682, 2194, 972, 1803, 1679, 795, 407, 1117, 1969, 1983, 1680, 519, 1561, 1309, 445, 1934]. BHEP [2026]. bi [479, 546]. bi-aspect [479, 546]. Bianco [453]. Bias [1257, 1335, 77, 106, 1310, 1261, 1935, 1052, 885, 1992, 1867, 1157, 956, 512, 520]. Bias-adjusted [1257]. bias-corrected [1261]. bias-reduction [1935]. Biased [527, 1200]. biases [757]. BIB [2238]. bibliography [2297]. biclustering [667]. bidiagonalization [543]. bimodality [2015]. binary [129, 1747, 1982, 181, 1264, 1633, 990, 1403, 2310, 2338, 145, 1943, 717, 1246, 1351, 1335, 1863, 1417, 23, 2037, 2093, 1813, 60, 1983, 134, 467, 1291, 1006, 186, 2053, 885, 1995, 2022, 780, 1561, 2047, 49, 673, 869, 1913, 1314, 959, 2199, 1710, 705, 2090, 523, 792]. binary-time-series [134]. binned [1042]. Binomial [1981, 2002, 990, 501, 1235, 2084, 1889, 2343, 1060, 528, 1082, 993, 1402, 754, 1822, 561, 511, 1894, 1450, 643, 2029]. bins [2256]. bioassays [1349]. biology [2070]. Biplot [1321]. biplots [2208]. Birnbaum [1304, 2066, 1885, 1714, 1443, 1669, 1354, 2330, 414, 980, 855, 646, 1356]. birth [538]. Birthday [438]. bites [1544]. Bivariate [1437, 1392, 1832, 19, 505, 475, 2295, 2006, 2018, 1347, 351, 719, 994, 1124, 1474, 1662, 1103, 584, 550, 1645, 1909, 1141, 1685, 749]. bivariate- [2006]. blind [1430, 1874]. blinded [705]. Block [1751, 1730, 238, 1718, 266, 791, 168, 796, 1077, 126]. block-angular [791]. block-circulant [796]. blockade [769]. blocked [64]. Blocking [1087]. blocks [579, 957, 1651, 24, 277, 2030]. BLUP [441]. Board [472, 483, 490, 526, 273, 288, 299, 315, 316, 322, 350, 367, 384, 395, 403, 413, 422, 437, 463, 507, 554, 567, 581, 595, 609, 621, 635, 651, 665, 678, 692, 710, 729, 748, 770, 789, 807, 828, 847, 862, 875, 890, 903, 917, 936, 964, 989, 541, 1011, 1038, 1112, 1161, 1180, 1209, 1249, 1274, 1312, 1364, 1399, 1422, 1514, 1562, 1610, 1647, 1683, 1724, 1755, 1799, 1818, 1848, 1878, 1895, 1926, 1947, 1978, 2004, 2068, 2105, 2143, 2176, 2221, 2245, 2265, 2283]. Boca [1766]. bond [1173, 882]. Book [2216, 1361, 2220, 2217, 1766, 2219, 2218, 2120]. Boolean [492]. Boosting [823, 1953, 2147, 1470, 2042, 199, 198, 2057, 196, 756, 826, 1350, 1760, 197, 1542, 1660]. Bootstrap [1993, 129, 1868, 1271, 1021, 1903, 176, 689, 1974, 627, 241, 1490, 926, 1665, 1962, 1653, 1650, 2109, 7, 1718, 877, 1720, 1, 1307, 106, 1086, 1252, 1310, 2052, 735, 1762, 1294, 736, 1912, 539, 1000, 1938, 2123, 83, 37, 836, 898, 1261, 2267, 1446, 571, 1673, 740, 86, 1761, 1265, 127, 1839, 1464, 1902, 1187, 1081, 1968, 119, 2276, 696, 2153, 504]. Bootstrap-based [1974, 898]. Bootstrapping [1349, 735, 2312, 2135, 931]. bootstraps [2334]. both [808, 659]. boundaries [1521]. boundary [908, 703]. Bounded [494, 1192, 734, 61]. bounded-influence [1192]. Bounding [910, 1758]. bounds [259]. Bowker [1316]. Box [1288, 952]. boxplot [1908, 606]. brain [401]. branching [1613]. breakdown [1383]. breaking [1568]. breaks [1890]. breast [365, 283, 1711, 1709, 1662]. Bregman [1057]. Breslow [1534]. Breusch [1000]. bridge [972]. Brownian [2119]. budget [339]. Building [1535, 1279, 995]. bumping [86, 836]. Bundling [777]. Burr [2152, 525, 1615, 1795]. business [1587, 741]. busy [1638]. BV [1437].

Calculation [1384, 889, 1291, 868, 256]. Calculations [52, 2185]. Calibrated [251, 905, 2159]. Calibration [2127, 1729, 982, 769]. Call [596]. Canadian [739]. cancer [1092, 365, 2310, 2062, 283, 2298, 2195, 1129, 1711, 1709, 1662, 768, 2089, 1996, 2090]. Candecomp [802]. Candecomp/Parafac [802]. candidate [1480]. Canonical [1321, 16, 972, 797, 803, 915]. capture [1402, 2172, 2286, 641, 1505]. CAR [2192]. carcinogenicity [189]. cardiovascular [1990]. caries [1246]. Carlo [226, 97, 695, 53, 13, 1722, 2338, 813, 1260, 948, 2124, 695, 1649, 1814, 351, 2135, 1448, 2331, 893, 1341, 1358, 708]. Carlo-based [1341]. Carroll [972]. CART [538, 2201]. case [1963, 38, 1846, 2183, 1298, 283, 1256, 1438, 721, 739, 1671, 1284, 899, 2092, 336, 58, 754, 1661, 1386, 331, 1981, 1228, 2163]. case-cohort [1284]. case-control [1228]. case-series [899]. cases [157, 901, 2228]. CAT [1327]. catch [907, 434]. categorical [1753, 1667, 1404, 91, 447, 1031, 1279, 83, 1118, 47, 251, 831, 153, 427, 1557, 1179, 1834, 985, 818]. categories [1425]. category [1628]. Causal [1131, 2112, 1972, 1132, 983]. causality [1255, 1584]. cause [2175, 365]. cause-specific [2175, 365]. caused [2092]. cautionary [100]. CCC [1976]. CDF [2303]. cDNA [2074]. Celebration [438]. cell [2058, 716, 1622]. cell-cycle [1622]. Censored [1637, 576, 2242, 1502, 871, 2251, 593, 136, 1810, 1245, 1208, 1523, 1288, 191, 481, 151, 872, 977, 1065, 2288, 937, 2104, 1347, 1837, 1443, 812, 2169, 235, 980, 1304, 2299, 1122, 2014, 1795, 1445, 103, 855, 1692, 2171, 1792, 2258, 63, 1909, 1507, 680]. censoring [2152, 2051, 1717, 1959, 458, 1447, 1811, 1767, 1780, 725, 2249, 2328, 2014, 683, 1968]. center [99, 346]. central [2195, 2007]. Centre [1630]. centroid [1062]. centroids [1047]. century [4]. cereal [435]. certain [1453]. Chain [550, 112, 948, 2331, 2100, 13, 504]. chains [1718, 751, 1183, 1468]. change [1146, 155, 1989, 356, 161, 319, 66, 1229, 1604, 240, 668, 1759, 1468, 1207, 286, 2021, 1104, 1821, 1185, 967, 2213]. change-point [161, 66, 1185]. change-points [1146, 1989, 1468, 286]. changepoints [1478, 2228]. Changes [1376, 845, 245, 443, 1686]. Chapman [1766]. characteristic [1789, 2290, 726, 93, 2139]. characteristics [958]. characterization [1681, 1243]. characterized [107]. Characterizing [1663, 2297]. chart [2027, 1293, 188, 2167, 2164, 2060]. chartjunk [1854]. charts [1037, 1833, 1749, 2311, 1971]. chatter [1749]. CHD [578]. Chebyshev [495]. check [130]. checking [2270, 89]. chemical [1453]. Chen [1386]. Chi [1299, 1488, 1882, 702, 246, 2007]. Chi-square [1882, 702, 246, 2007]. Chi-squared [1299, 1488]. Chisquare [1826, 411]. Choice [629, 1138, 813, 974, 627, 1426, 1599, 2026, 2256]. choices [717]. Choosing [1635, 7, 341, 1053, 342, 750, 1777]. chromosomal [2094]. chromosomes [96]. circles [1921]. circulant [796, 1226, 2148]. circular [1456, 25, 1116, 1865, 601, 1631, 1772]. circular-linear [25]. city [1077]. city-block [1077]. Class [304, 434, 1551, 2332, 1914, 340, 91, 118, 992, 2291, 691, 1629, 2035, 237, 996, 2215, 338, 1763, 2199, 858, 1996, 1170, 967, 2211]. classes [2008, 587, 1125]. classical [506, 1232, 142, 1907, 1858]. Classification [1479, 1092, 1620, 1347, 1871, 337, 1706, 1418, 476, 891, 841, 2041, 247, 1241, 947, 1553, 2253, 2310, 2062, 2324, 881, 1655, 1372, 1750, 1965, 1195, 305, 538, 974, 118, 1094, 1129, 2267, 60, 707, 860, 1056, 632, 800, 519, 180, 1847, 568, 2049, 605, 1444, 1101, 1114, 96, 2297, 1698, 1042, 512, 2089, 8, 1554, 1314, 338, 2083, 1552, 1660]. classified [1390]. classifiers [645, 392, 773, 777, 1845, 587, 762]. Classifying [1380, 2090]. climate [1104]. clinical [760, 167, 1980, 1991, 99, 565, 872, 517, 2174, 1107, 1444, 2001, 1318, 986]. clipped [65]. close [1510]. CLUES [1539]. Cluster [133, 1055, 1537, 607, 1982, 853, 1051, 3, 1298, 1220, 30, 1556, 478, 1863, 1469, 1047, 1429, 31, 1677]. Cluster-based [607]. cluster-randomized [1429]. Cluster-wise [1537]. clustered [129, 1264, 2319, 1994, 1300, 1798, 2279, 1385, 294]. Clustering [854, 1484, 2336, 1473, 1862, 619, 2212, 961, 1064, 1538, 1753, 1555, 1570, 406, 1655, 883, 1626, 1884, 1024, 2058, 1487, 1336, 1751, 771, 1025, 302, 2294, 539, 865, 330, 839, 1063, 182, 559, 2173, 391, 800, 1492, 1924, 2054, 2280, 1095, 694, 410, 867, 1561, 829, 15, 2089, 1046, 715, 1860, 1414, 1539, 612, 934, 1476, 2211, 2293]. clusters [2181, 878, 998, 500, 1351, 292, 23, 2318, 2202]. Clusterwise [718, 1029, 2214]. clutter [296]. co [1997, 1568]. co-breaking [1568]. co-evolution [1997]. coarse [1016]. coauthor [1675, 2348]. Cobra [1568]. code [146, 116]. codes [1866]. codifference [2337]. coefficient [1982, 411, 965, 54, 120, 137, 1030, 999, 1482, 2123, 1346, 1933, 848, 887, 1877, 1627]. coefficients [1196, 1139, 1974, 722, 1150, 1467, 900, 745, 397]. cognition [2230]. cognitive [356]. cohort [1284]. coincide [802]. cointegrated [733, 1586]. cointegration [1868, 2110, 310]. Coleman [30]. collapsed [833]. colleagues [2345]. collected [1340]. Collective [2024]. Collective-agreement-based [2024]. collide [2070]. color [1159]. colors [2223]. colour [821]. column [2039]. columns [2336, 659]. combination [1574, 2006, 1002, 1508, 857]. combinations [1349]. Combinatorial [1077]. Combined [655, 1561, 859, 1971]. Combining [538, 1707, 70, 2082, 1738, 626]. Comment [1213, 1214]. Comments [30, 31]. common [1955, 1586, 976, 2023, 1975, 1490, 2308, 1588, 95, 1457, 1441, 42, 1877]. comovements [1736]. Comp [767]. compact [794, 1573, 2016]. Comparative [128, 268, 476, 1502, 400, 439, 1290, 1969, 675, 1236, 216, 1509, 103, 986]. compare [1886, 1995]. compared [643]. Comparing [642, 2160, 1972, 885, 1441, 893, 1946, 1149, 1930, 514, 1633, 1220, 977, 399]. Comparison [1297, 943, 1987, 387, 1922, 1241, 1166, 1976, 228, 1220, 948, 473, 574, 1751, 1573, 565, 1889, 1218, 1998, 534, 253, 1158, 1934, 2311, 63, 523, 749, 219, 181, 113, 698, 1641, 2052, 248, 795, 1374, 1034, 1273, 2237, 1348, 66, 669, 444, 707, 1873, 230, 626, 579, 519, 670, 1388, 1806, 1839, 1977, 1309, 49, 744, 2163, 87, 894, 346, 2171, 2030, 2075]. Comparisons [997, 816, 1892, 1967, 1776, 391, 834, 553, 1904, 694, 584, 1450, 143, 513, 962, 36]. compass [229]. compatibility [260]. competing [570, 2299, 1200]. complete [1566, 1430, 690]. Complex [1793, 1340, 223, 97, 224, 449, 1866, 1132, 2047, 1687]. Complex-valued [1793]. Complexity [2086, 2314, 1334]. compliance [1291, 1831, 573]. complicated [672]. Component [375, 516, 930, 919, 98, 149, 880, 1612, 973, 1034, 1923, 1273, 276, 1379, 343, 2013, 582, 2205, 1870, 1953, 1644, 2278, 983, 1315, 445, 1391, 275, 2212, 792]. component-wise [1953]. components [906, 1299, 611, 878, 1679, 1256, 175, 2288, 1049, 1074, 329, 1588, 153, 48, 661, 928, 2239, 772, 331, 1907, 325, 809, 1357, 1940]. componentwise [674]. composite [778, 2159, 2031]. Comput [2344, 120, 826, 2067, 939, 2347, 2348, 2346, 277, 2345]. Computation [791, 1500, 1045, 14, 2070, 1883, 1571, 1523, 1452, 706, 2178, 2225, 2236, 713, 1132, 263, 72, 1821, 1440, 1287, 1692, 1590, 469]. Computational [1726, 730, 1251, 2107, 385, 611, 693, 1303, 372, 1980, 1266, 1284, 1282, 1224, 740, 1304, 1277, 827, 3, 1278, 312, 30, 2010, 573, 2244, 699, 31]. Computationally [492, 37, 64, 1184]. computations [790, 1967, 92]. compute [188]. computed [381]. Computer [1822, 146, 1866, 1847, 116, 255, 485]. computer-intensive [255]. computerized [1709]. computers [9]. Computing [411, 1576, 227, 44, 51, 712, 466, 1949, 737, 1646, 2151, 371, 1766, 132, 259, 1574, 1023, 666, 1211, 1954, 2124, 474, 1518, 697, 668, 830, 1882, 102, 373, 2193, 1869, 967]. concave [1484, 1420]. concentration [2332, 544, 2282]. Concept [2138, 1682, 2076]. concepts [1014]. concerning [2002, 593, 46]. concerns [1171]. CONCOR [938]. concordance [1030]. condition [802]. Conditional [2324, 1019, 1721, 1246, 1002, 912, 863, 260, 123, 1269, 1054, 1928, 919, 1750, 39, 557, 1165, 1233, 717, 2129, 2023, 565, 66, 2093, 1004, 1709, 67, 1779, 201, 1265, 1187, 914, 1694, 924, 1338, 2096, 103, 522, 888, 2197, 809, 1956]. Conditional-mean [1721]. conditionally [1432]. conditionals [1073, 1685]. conditions [130, 573, 1085]. Confidence [990, 965, 773, 2303, 1416, 231, 1975, 1680, 58, 1773, 1498, 549, 767, 1618, 2229, 2346, 1877, 2269, 2109, 840, 62, 648, 1080, 1303, 1872, 2324, 1433, 734, 578, 1308, 398, 1673, 740, 76, 254, 61, 561, 511, 1894, 1450, 143, 1223, 1985, 1653, 2171, 1085, 2029]. Configural [462]. configurations [307]. confined [1288]. Confounding [1242, 224]. conjoint [1691]. conjugate [1333]. conquer [831]. Consequences [59]. Consistent [1713, 301, 604, 1338]. Constant [2152, 2126, 520]. Constant-partially [2152]. Constrained [1531, 1412, 892, 712, 1123, 1493, 1097, 701, 1911, 803, 1775, 1225, 2206]. constraining [625]. constraints [318, 1584, 697, 1097, 409, 1089, 213, 275, 1319]. Constructing [431, 2148, 1915, 605, 115, 1125]. Construction [509, 467, 1432, 582, 1141, 62, 1223, 1985, 805, 419]. consumer [208]. contact [1510]. containing [1644, 2278]. contaminated [117]. contamination [475]. Contents [1012, 1039, 1113, 1162, 1181, 1210, 1250, 1275, 1313, 1365, 1400, 1423, 1515, 1563, 1611, 1648, 1684, 1725, 1756, 1800, 1819, 1849, 1879, 1896, 1927, 1948, 1979, 2005, 2069, 2106, 2144, 2177, 2222, 2246, 2266, 2284]. context [611, 2214, 1138, 747]. contexts [1732]. contingency [608, 514, 465, 2179, 2170, 2019, 2023, 99, 2336, 389, 1353, 689, 1044, 1911, 462, 2210]. Continuous [79, 958, 411, 220, 1744, 545, 1667, 1682, 1733, 466, 861, 737, 898, 1713, 559, 2056, 444, 518, 953, 1662, 1309, 1582, 1809, 1318, 1946, 1715, 1185, 1678, 2213]. continuous-time [1582]. Continuum [1785, 1530, 587]. contrasted [643]. contribution [2262]. contributions [769]. Control [2060, 2159, 2086, 2027, 163, 1513, 1037, 1293, 2092, 188, 2167, 1749, 1228, 2311, 1971, 184]. controlled [27, 1661]. controller [769]. controlling [633, 1617]. convenient [1215]. Convergence [1460, 1043, 863, 948, 1121, 2331, 549, 767, 2021]. convex [2162]. convexity [811]. convolution [2264]. convolutions [188]. coordinate [433]. coordinates [878, 430]. COPD [673]. Copula [1761, 1069, 1221, 1489, 2132]. copula-GARCH [2132]. copulas [2115, 1778, 1352, 2124, 1732, 1906, 505, 1218, 1103, 1794, 1141]. corners [586]. corrected [884, 1261, 2]. correction [1867, 77, 1917, 1524]. correlated [159, 1872, 1921, 2048, 498, 1417, 441, 1410, 1247, 1840, 1286, 566, 2136, 236, 1439, 780, 957, 537, 49, 1788, 1946, 888, 1434]. Correlation [1357, 1982, 508, 2188, 411, 62, 227, 2226, 107, 797, 736, 1253, 1119, 1000, 837, 566, 848, 582, 1839, 1867, 1301, 1700, 803, 1877, 1216, 915]. correlations [1623, 1503, 2022, 450]. Correspondence [1240, 2205, 2203, 2204, 1559, 2209, 2207, 2206]. Corrigendum [2344, 2067, 827]. coskewness [2113]. cost [2037, 589]. cost-efficient [2037, 589]. costs [1840, 88]. Count [1833, 1407, 765, 1504, 1178, 1671, 2192, 2044, 1960, 2263, 2227, 2314, 1334]. counterparts [1597]. countries [1106, 1608]. counts [2002, 2189]. coupled [2030]. course [2072, 2079, 1706, 1622, 934]. Covariance [1431, 508, 763, 2233, 1623, 478, 1323, 2309, 2079, 1793, 1506, 1659, 29, 1716, 1875]. Covariate [2169, 1273, 100, 861, 1004, 2320, 1324, 1309, 984]. covariate-adjusted [984]. covariates [2125, 1914, 1404, 2166, 2093, 174, 914, 1184, 641, 985]. covariogram [303]. cover [287, 314, 321, 393, 402, 412, 421, 434]. Coverage [1929, 1362]. Cox [529, 1288, 2160, 952, 377, 63]. Coxian [2320]. Cramér [2344, 2275, 534]. CRC [1766]. credit [1263, 860, 620]. Cressie [979]. criminal [1207]. crises [1110]. crisp [354]. Criteria [1136, 1061, 2108, 882, 565, 1517, 2088, 1930, 749]. criterion [945, 820, 884, 1826, 798, 86, 1665, 2099, 901]. Critical [1967, 951, 1490]. crop [649]. cross [1635, 2226, 1690, 1307, 1000, 1938, 1383, 2267, 510, 329, 602, 2050, 2088, 589]. cross-correlation [2226]. cross-over [1635, 589]. cross-sectional [1307]. cross-validation [1690, 1383, 2267]. crosses [1285]. crossroads [4]. crude [2175]. cubic [365, 2016]. cumulants [171]. cumulated [1534]. cumulative [2175, 1326, 51, 1478, 505]. cure [362, 334, 1632]. cured [337]. currency [1170]. current [1828]. curricula [728]. curse [2261]. curvature [808, 162]. curve [1993, 1901, 640, 556, 138, 249, 1063, 739, 28, 2054, 835, 29, 2325, 902]. curve-clustering [1063]. Curves [446, 1392, 1372, 1374, 548, 1071, 977, 534, 1709, 1379, 1806, 1653, 1946]. customer [860]. CUSUM [1045, 2167, 1836, 1821, 1971, 2322, 188]. CUSUM-type [1821]. cutoff [1305]. cutpoints [1488]. cycle [2058, 741, 1622]. cycles [1587]. cyclical [1586, 100].

D [30, 1311, 425, 1482, 1748, 116]. D. [30]. DACE [1278]. daily [1459, 2194, 1074, 1103, 2141]. DALASS [1283]. damped [121]. Data [1745, 1171, 693, 1303, 2344, 120, 206, 1527, 908, 1024, 30, 208, 911, 826, 169, 2067, 1366, 631, 939, 1304, 2347, 1101, 2348, 203, 767, 2346, 277, 2345, 2002, 129, 1456, 1499, 906, 1939, 1747, 1479, 942, 1442, 211, 1566, 25, 1340, 45, 1538, 760, 970, 2188, 1264, 2251, 944, 1753, 2201, 1495, 1359, 7, 1276, 1633, 167, 220, 891, 990, 425, 545, 6, 202, 2115, 841, 1555, 946, 1363, 593, 136, 1810, 2292, 991, 2338, 405, 1245, 1532, 1976, 2262, 764, 1667, 1921, 765, 1544, 1208, 205, 146, 2270, 2183, 1371, 1368, 603, 1086, 766, 1754, 354]. data [883, 498, 1626, 1298, 1682, 1884, 304, 1462, 655, 283, 248, 1943, 5, 728, 200, 529, 910, 2349, 2057, 278, 1404, 285, 428, 1989, 538, 949, 91, 1556, 2319, 535, 1246, 473, 97, 447, 1351, 1452, 1991, 2028, 2298, 1021, 1230, 516, 1034, 117, 1504, 1030, 1923, 1237, 1361, 1117, 1689, 1025, 161, 261, 1055, 1000, 1431, 1863, 1833, 2260, 191, 1031, 262, 2166, 481, 2130, 330, 690, 1377, 1577, 1889, 897, 1323, 737, 83, 1331, 1178, 1009, 2196, 727, 872, 162, 657, 2085, 831, 977, 1129, 1003, 2288, 1994, 1565, 1671, 1998, 2173, 70, 1604]. data [1473, 1781, 937, 676, 2104, 1160, 2056, 1347, 1469, 449, 707, 873, 1837, 2192, 1443, 252, 632, 988, 1449, 1871, 2309, 2301, 1975, 1711, 2009, 812, 134, 1924, 1936, 2054, 2306, 1378, 1075, 1291, 429, 659, 14, 326, 1006, 573, 660, 2044, 1828, 347, 487, 1026, 1630, 1132, 235, 1016, 1784, 661, 1798, 154, 126, 1945, 1134, 886, 76, 1277, 460, 2299, 784, 827, 1575, 1706, 2318, 2239, 2081, 1986, 780, 454, 1631, 1561, 2205, 1867, 1451, 1301, 2063, 1835, 1960, 1950, 2263, 2271, 2001, 2274, 982, 2227, 2071, 1042, 1122, 1652, 1341, 1644, 2278, 2328, 2279, 1809, 1108, 8]. data [1962, 1795, 537, 1777, 901, 1418, 1708, 673, 1259, 424, 869, 1788, 2172, 1651, 1046, 715, 1859, 1622, 489, 832, 667, 2314, 427, 1321, 103, 1557, 746, 2215, 1414, 144, 1020, 552, 1203, 2199, 178, 1692, 1844, 2082, 2171, 686, 850, 1710, 1385, 2197, 2339, 2030, 426, 2257, 1334, 448, 363, 1184, 2202, 705, 63, 294, 2083, 1510, 1909, 1632, 1639, 934, 1552, 1876, 1881, 680, 1686, 2269, 1505, 792, 1678, 1200, 2075, 2293]. data-adaptive [1331]. data-augmentation [1510]. data-based [686]. Data-driven [908]. database [1130, 644]. datasets [2186, 292]. dating [741, 443]. day [1735]. Deal [1441]. dealer [1172]. dealing [704]. death [336]. deblurring [1874]. debt [1110]. decision [1023, 836, 779, 2049, 616]. decisions [1272]. decomposable [2019, 2043]. Decomposition [923, 796, 186, 1529, 1687, 1476, 792]. decompositions [311]. deconvolution [499, 908]. Decoupling [163]. decreasing [2329, 1801]. deepest [72]. defective [2311]. defects [1495]. defined [124]. definite [2007]. Definition [1131]. definitions [118]. degeneracy [213]. degree [408, 1765]. degrees [588, 1940]. Delaunay [1191]. delayed [1107]. Deletion [1091, 2150, 1097]. demand [845, 921, 1259]. dementia [2230]. demerit [1758]. demonstrated [667]. denoising [930, 405, 1098]. dense [1711]. densities [1064, 1438, 1380, 1817, 10, 2268, 1338]. Density [2231, 455, 352, 1696, 123, 2115, 2046, 1899, 136, 2292, 1690, 907, 1846, 133, 499, 1825, 1234, 839, 2264, 1261, 2056, 1153, 1186, 905, 599, 1437, 81, 197, 221, 1772, 1046, 1946, 1909, 968, 902]. density-based [1046]. dental [2028]. departures [628, 34]. dependence [1719, 1392, 271, 1307, 1743, 1938, 1472, 1031, 2050, 2207]. dependencies [248]. dependent [1105, 2187, 2214, 1462, 2170, 2127, 1351, 676, 2104, 2320, 928, 1688, 1451, 77, 1203, 962]. depends [649]. Depth [1378, 1191, 259, 1888, 1637, 1384]. Depth-based [1378]. derivative [1341]. Derivatives [563, 1260]. Derived [1049, 1720]. Deriving [1737]. descent [1173, 2095]. Describing [99]. Description [995, 1608]. descriptors [659]. Design [1035, 410, 496, 1340, 433, 973, 1144, 2121, 1543, 821, 1435, 1981, 2080, 471, 282, 2238, 805]. designing [488]. designs [1061, 1963, 2342, 159, 1635, 387, 1453, 32, 1954, 1830, 795, 309, 1964, 1513, 168, 724, 1087, 1949, 232, 2037, 2093, 1691, 1226, 41, 1247, 2148, 467, 267, 1432, 700, 575, 126, 1439, 602, 2039, 399, 1511, 636, 1088, 957, 589, 1913, 963, 24, 277, 1159, 1041, 1971]. detect [1841, 2181, 1055, 1983, 2190, 809]. Detecting [1495, 577, 1146, 233, 1468, 1207, 1309, 1406, 2260, 670, 1759, 1686]. Detection [1116, 245, 1749, 1847, 2262, 1816, 1298, 500, 1258, 1595, 1533, 1478, 1285, 478, 1845, 842, 1709, 429, 2064, 487, 1095, 1662, 286, 488, 2268, 857, 1821, 128, 1476]. determinant [478]. determinants [495]. Determination [562, 1767, 134, 725, 1228, 1710]. determine [1521, 1817]. Determining [175, 1995, 2318, 772, 585, 1305]. deterministic [1974, 468]. Developing [1860, 2076]. Development [1037, 111]. developmental [1143]. Developments [112, 323]. Deviance [2099]. deviation [840, 2164, 755]. deviations [687]. device [672]. diabetes [2309]. diagnosis [1996, 2090]. Diagnostic [28, 2121, 1994, 337, 885, 1995, 1945, 1992, 1618, 1946, 1385, 1930]. Diagnostics [1370, 2207, 1356, 2327, 764, 1838, 340, 1096, 1178, 1443, 327, 361, 2045, 699, 2305, 563, 2240, 294, 1466]. diagonal [2023]. diagonalization [1381]. diallel [602]. dialysis [1466]. DIC [2099]. dichotomous [1320]. difference [506, 1576, 682, 12, 1433, 1174, 1837, 992, 67, 1814, 1944, 2067, 1680, 1236, 1629, 254, 1894, 42, 1450, 643, 1985, 346, 888]. difference-based [992]. differences [477, 1077, 670, 2040]. different [1297, 1939, 1751, 578, 1346, 723, 1100, 2003, 374, 886, 1662, 49, 850, 1930]. differentiable [2117]. Differential [867, 613, 1733, 2078, 1117, 1055, 2260, 1522, 2079, 17, 2076, 1081]. differentially [1406, 822, 1001, 1197, 2202]. differentiation [859]. differing [2030]. diffusion [731, 486, 1642, 1308, 1482, 1817]. digraph [907, 1341]. digraphs [434]. Dimension [932, 1558, 305, 1692, 1048, 607, 2165, 2301, 2151, 568, 1506, 116, 1550, 1796, 1956, 2149]. Dimension-reduced [1692]. dimension/small [1550]. dimensional [906, 1963, 1479, 1555, 304, 1649, 1998, 1781, 2281, 326, 644, 1207, 72, 1046, 1557, 1844, 2083, 1552, 1940, 2293]. Dimensionality [653, 1546, 1679, 79, 958, 1314, 1076]. dimensions [1363, 1643]. Direct [1820, 920]. directed [1205]. direction [1436]. directional [455]. directions [232]. Dirichlet [1923, 1417, 833, 14, 1783, 1616]. disaggregation [400]. disclosure [2313]. Discontinuities [916]. discontinuous [1244, 1260]. discordant [255]. discount [1008]. Discovering [1834, 2090]. Discovery [1201, 545, 633, 434, 1427, 1041]. discrepant [2139]. Discrete [339, 616, 1499, 2175, 260, 411, 1744, 688, 344, 845, 737, 1898, 1416, 775, 2009, 518, 1629, 620, 1929, 1678]. discrete/continuous [1744]. Discriminant [647, 181, 1051, 90, 1887, 1292, 502, 1156, 1752, 1115, 1138, 1870, 157, 1550, 1283, 2083]. Discriminating [407]. Discrimination [1585, 661, 944, 882, 446, 717, 1709, 518, 2020]. Discussion [1939, 1360]. disease [941, 345, 600, 391, 988, 2092, 2191, 1995, 694, 1992, 714, 2097, 2190, 1510]. disease-related [2097]. disease-specific [1992]. Disentangling [1239]. Disequilibrium [2102, 857, 2096]. disjoint [2212]. disparity [531, 212, 1440]. Dispersion [178, 2259, 408, 41, 547, 399]. dispersional [1329]. dispersions [2227]. display [451, 672]. displays [46, 428, 1573, 109, 426]. dissimilarities [1033, 1863]. dissimilarity [1670, 1859]. Distance [539, 491, 1733, 2312, 329, 347, 1140, 2140, 686]. Distance-based [539, 347]. distributed [1442, 405, 1107, 566, 572, 264, 2199, 88]. Distribution [690, 27, 601, 2081, 2313, 2152, 2051, 22, 634, 2008, 1668, 1526, 1311, 1342, 1770, 112, 2251, 1359, 411, 328, 1299, 1922, 167, 2233, 1222, 702, 1966, 1804, 1758, 533, 44, 879, 85, 776, 1360, 2235, 1641, 51, 26, 1258, 1217, 466, 344, 501, 987, 1119, 2237, 19, 2336, 404, 191, 889, 170, 481, 2084, 457, 1394, 2031, 1843, 1193, 2012, 1663, 1768, 781, 723, 1215, 1656, 1714, 2018, 1343, 1900, 1669, 2077, 1354, 2, 246, 1004, 2225, 2007, 188, 1075, 360, 1425, 166, 830, 1880, 2055, 2231, 1411, 1699, 215, 719, 994, 1124, 597, 939, 414, 980, 1304, 1992]. distribution [362, 201, 1428, 2234, 742, 1960, 1103, 1477, 1652, 913, 868, 525, 1615, 1625, 417, 1801, 1461, 2142, 103, 2276, 855, 1802, 2257, 448, 646, 1792, 2034, 1685]. distribution-free [2237, 170]. Distribution-preserving [2313]. distributional [2104]. distributions [2002, 1333, 260, 2066, 1054, 388, 1392, 411, 840, 124, 688, 2329, 630, 1484, 1955, 2275, 2344, 1208, 1167, 1883, 1462, 1723, 2125, 407, 976, 357, 290, 987, 1362, 1908, 1123, 2295, 2006, 1893, 1416, 1969, 2033, 1254, 2010, 679, 1221, 775, 1780, 837, 1202, 2134, 566, 241, 1413, 2320, 905, 336, 953, 597, 939, 153, 1882, 2022, 1629, 2045, 981, 2013, 2299, 1746, 1631, 2059, 237, 584, 286, 1421, 2224, 562, 1302, 2229, 2346, 1415, 1869, 1645, 1125, 1262, 1742, 1485, 708, 36, 1852, 1141].

DIVCLUS [1570]. DIVCLUS-T [1570]. divergence [896, 884, 409, 1089]. divergences [1057, 979]. diversity [56]. divide [831]. divide-and-conquer [831]. divisive [1570]. DLL [461]. DNA [1089, 633, 2101, 434, 1551, 2100]. DO [551]. DO-ACT [551]. document [302, 800]. Does [1823]. domain [2165, 253, 928]. domains [734]. dominant [2207]. Dong [30]. dose [1245, 125, 155, 138, 249, 542, 2306, 496, 768]. dose-response [542]. doses [249, 720, 1475]. Double [2042, 1868, 1, 1252, 1767, 305, 37, 740, 2334]. Double-Bagging [2042]. doubly [946, 1222, 2321, 1792]. down [1617]. downdating [311]. DPLS [660]. drawing [10]. drift [898, 1713]. drilling [1749]. driven [1728, 908, 631, 2142]. Drop [1981, 1431]. drop-outs [1431]. Drop-the-losers [1981]. dropout [2338, 147]. Drug [585, 1791]. Dual [1027, 2341, 2215]. duration [1701, 924]. during [1997]. dyadic [1006]. dye [1197]. dye-swapped [1197]. Dynamic [80, 1173, 1581, 1592, 1743, 736, 1253, 2289, 1791, 1439, 622]. dynamical [1278, 460]. dynamics [1172, 1263, 1189, 532, 1122].

each [1521]. EAR [916]. Early

[1110, 198, 565, 1709]. Easily [2022]. easy [1854]. ECM [1225]. ecological [2201, 2191, 2244]. Econometric [925, 1171, 2117, 1256, 1266, 967, 2116]. Econometrics [1163, 730, 1251, 2107, 440]. economic [167, 43, 1739, 1036]. Economical [2342]. economics [2269]. economies [1608]. Ecosystem [1790]. Ed [1766, 138, 280]. edge [842, 1280, 714, 1917]. edges [1595]. Editorial [273, 288, 299, 315, 316, 322, 350, 367, 384, 395, 403, 413, 422, 437, 463, 472, 483, 490, 507, 526, 554, 567, 581, 595, 609, 621, 635, 651, 665, 678, 692, 710, 729, 748, 770, 789, 807, 828, 847, 862, 875, 890, 903, 917, 936, 964, 989, 368, 323, 484, 652, 541, 1011, 1038, 1112, 1161, 1180, 1209, 1249, 1274, 1312, 1364, 1399, 1422, 1514, 1562, 1610, 1647, 1683, 1724, 1755, 1799, 1818, 1848, 1878, 1895, 1926, 1947, 1978, 2004, 2068, 2105, 2143, 2176, 2221, 2245, 2265, 2283]. EEG [460]. Effect [1982, 878, 198, 428, 2061, 1500, 1126, 2023, 861, 1463, 583, 517, 1100, 1658, 2252, 1727, 1757, 1466]. effective [1245, 249, 720, 1501]. effectiveness [1740, 1475]. Effects [1605, 2230, 1297, 634, 1841, 1987, 1367, 476, 1330, 2188, 2338, 2112, 12, 817, 1244, 2124, 590, 2121, 168, 1126, 59, 2095, 99, 1000, 1619, 456, 1701, 2158, 1768, 774, 503, 1983, 542, 1242, 41, 1280, 28, 553, 1840, 1936, 592, 186, 1131, 1774, 1662, 714, 2305, 1122, 2279, 156, 1007, 57, 397, 1445, 178, 1248, 1512, 809, 1184, 1091, 324]. Efficiency [846, 634, 2061, 227, 640, 695, 738, 496]. Efficient [2111, 1518, 821, 775, 2132, 1286, 215, 84, 868, 1708, 2302, 1775, 1445, 1590, 1996, 1685, 1876, 64, 465, 18, 474, 492, 37, 2037, 1951, 1866, 115, 1560, 1731, 2139, 589, 1184]. efforts [432]. EGARCHX [177]. EGARCHX-processes [177]. eigenanalysis [799]. eigenproblem [1598]. eigenvalue [555]. eigenvalues [1716]. Eight [442]. elasticity [2126]. element [801]. element-wise [801]. elemental [1536]. elements [1644, 2278]. elimination [1156]. ellipsoidal [1541, 1977]. elliptical [1965, 1893, 2134, 1332, 2059, 2305, 1716]. elliptically [572]. EM-type [1869]. embedded [562]. Embeddings [1018]. emerging [2181]. emphasis [1117, 1379]. Empirical [238, 1681, 1232, 1360, 2127, 1004, 1259, 2190, 2240, 1109, 1887, 1789, 1393, 1001, 2290, 726, 2191, 2308, 2071, 1677, 887, 2082, 2034, 1507]. Enabling [2063]. endophenotype [2097]. endpoint [1449]. endpoints [2040, 1318, 705]. energy [1744]. enfuvirtide [1997]. enhanced [1351]. ensemble [198, 2064, 2297]. ensembles [1479, 2024, 2042]. enterprise [1201]. entries [1404]. entropy [881, 14, 1779]. entry [522]. Enumerating [2043]. envelope [1822]. environment [1582]. environmental [2201, 319, 1609, 2243]. epidemic [1569]. epidemiologic [2306]. epidemiological [473]. epidemiology [13]. epileptic [401]. epistemology [2070]. equal [58]. equality [763, 1326, 1323, 1493, 528, 1503, 153, 1882, 1506, 1788, 1093, 1716]. equation [491, 318, 2259, 444, 873, 1873, 757, 442, 1875]. equations [613, 1633, 2117, 1733, 1254, 993, 135, 1945, 1867]. equilibrium [2086]. equity [1276, 1741, 1103]. Equivalence [725, 1803, 1461]. equivalency [33, 1083]. Erratum [693, 1303, 120, 826, 939, 1304, 2347, 2346, 277]. error [470, 2324, 1195, 588, 2028, 1642, 1204, 1000, 92, 1394, 1002, 2267, 301, 604, 1604, 993, 1779, 955, 1602, 1603, 417, 623, 1934, 270, 1524, 1601]. Errors [944, 1596, 159, 704, 1359, 1369, 1965, 137, 1912, 2078, 441, 1473, 1247, 1936, 569, 700, 572, 236, 1439, 637, 1704, 2000, 1243, 1605, 1597, 2327]. Errors-in-variables [1596, 1369, 637, 1704, 1597]. Erythematosus [1939]. Escaping [2223]. Estimate [2033, 544, 1843, 1229, 697, 329, 343, 374, 955, 69, 1742]. Estimate-based [2033]. estimated [1993, 2159, 2164, 1961, 1920, 768]. estimates [1502, 2159, 167, 791, 813, 1220, 712, 910, 1500, 1393, 1705, 737, 899, 569, 631, 2017, 713, 981, 81, 1158, 2139, 56, 256, 363, 1389]. Estimating [352, 2175, 1359, 2296, 734, 1217, 716, 2065, 861, 2031, 2267, 121, 2018, 517, 358, 571, 1402, 362, 68, 1511, 331, 2141, 1185, 985, 906, 508, 491, 1633, 1349, 1086, 239, 370, 151, 2084, 1218, 988, 993, 1224, 53, 135, 1773, 633, 154, 1945, 1779, 201, 1867, 745, 2139, 1338, 1775, 1306]. Estimation [2051, 22, 1367, 2251, 628, 1750, 1462, 1258, 1516, 1204, 1237, 1119, 1078, 1994, 781, 1604, 834, 2053, 1344, 2308, 242, 235, 1992, 333, 825, 984, 683, 1625, 2096, 1645, 2238, 103, 522, 819, 1297, 238, 1145, 1456, 527, 1668, 613, 2154, 1719, 1164, 1696, 1199, 1770, 1264, 458, 123, 531, 2111, 1447, 1192, 971, 2159, 470, 2187, 62, 2115, 1257, 877, 1899, 1, 136, 2155, 2292, 1690, 1811, 266, 598, 693, 1349, 682, 1048, 1905, 1846, 1028, 296, 133, 1824, 499, 908, 815, 1310, 1830, 916, 74, 1778, 372, 1825, 1256, 1040, 2125, 1942, 89]. estimation [1090, 155, 1989, 137, 548, 303, 150, 2128, 1666, 1613, 578, 1234, 1177, 191, 1584, 1903, 138, 280, 249, 1353, 1123, 493, 66, 245, 839, 1118, 2264, 721, 1178, 456, 1002, 37, 1261, 1713, 2288, 1481, 774, 301, 604, 904, 1489, 2056, 444, 1082, 1873, 607, 1004, 1242, 775, 2011, 317, 2132, 2134, 2133, 1709, 1831, 1814, 2174, 1186, 1005, 1674, 966, 1152, 215, 575, 599, 58, 376, 1100, 597, 939, 1131, 1132, 414, 980, 1304, 1437, 1730, 2045, 1079, 2347, 2304, 1448, 1426, 835, 440, 1099, 1451, 264, 586, 1891, 237, 480, 197, 1950, 1427, 2274, 293, 982, 1194]. estimation [221, 743, 2326, 887, 1602, 1341, 1700, 1157, 956, 1603, 525, 1794, 1036, 213, 2279, 782, 1836, 1007, 1081, 1907, 93, 1694, 49, 1827, 1140, 1731, 2302, 1302, 1415, 1941, 2140, 1772, 1877, 1295, 1957, 1572, 2254, 2314, 2091, 1897, 708, 855, 1225, 1216, 1512, 1892, 2161, 2285, 1616, 610, 2257, 1334, 192, 646, 1792, 2258, 1050, 1524, 1909, 2325, 685, 968, 1337, 364, 1068, 88, 680, 2269, 2029, 1505, 1200]. estimations [723, 1656, 76, 1070]. estimator [132, 1712, 1534, 2159, 1576, 997, 345, 214, 2074, 1667, 222, 453, 1733, 896, 478, 170, 2312, 671, 1734, 409, 1089, 1188, 1540, 2137, 547, 1441, 1961, 637, 455, 1358, 1485]. Estimators [1052, 220, 997, 1232, 1369, 600, 1958, 82, 1263, 262, 720, 475, 738, 816, 992, 1075, 351, 1798, 759, 212, 216, 536, 1652, 2000, 549, 767, 2021, 373, 744, 1262, 959, 346, 686, 1384, 523, 1390]. European [1106]. evaluate [1995]. Evaluating [2046, 747, 858, 232, 399, 2091]. Evaluation [942, 1958, 1442, 2146, 1872, 600, 1573, 262, 1607, 1900, 846, 1153, 1814, 1490, 1761, 714, 1227, 1318, 1176]. eve [4]. even [832]. event [2183, 1298, 1989, 1889, 1784, 2035, 1876]. event-related [1876]. events [1832, 1989, 2236, 2317, 522, 1148]. eventual [1682]. evidence [942, 1442, 285, 177]. evolution [1997, 1522, 867]. evolutionary [1035]. evolving [1133, 424]. EWMA [2164]. Exact [506, 260, 1717, 1574, 32, 23, 739, 2225, 1005, 966, 1044, 143, 1499, 642, 159, 1488, 1690, 106, 404, 528, 1998, 697, 1286, 754, 386, 1894, 643, 1227]. example [25, 1276, 339, 2091]. Examples [634, 758]. Excel [1111, 679, 291, 787, 1851, 1850, 1854, 1852, 1212, 291, 1853, 1855]. Excess [2092]. exchangeable [252, 1639]. executables [1241]. Executive [2243]. Existence [899, 1843]. exogenous [1688]. exotic [1009]. expansion [848]. expectation [849]. expectations [109]. Expected [2249, 1173]. expectile [2307]. expensive [1797]. experiences [425]. experiment [573]. experimental [159, 1285, 1144, 1949, 1439]. Experiments [612, 108, 1406, 2072, 822, 228, 1943, 509, 1504, 457, 1691, 821, 224, 1197, 2073, 2087, 2080, 282, 1159, 178, 1317]. explains [7]. explanatory [1258]. Exploiting [2165, 550]. Exploration [2104]. Exploratory [1538, 1754, 1022, 262, 2073, 633, 430]. Exploring [1183, 1226, 2339]. exponent [1344]. Exponential [2247, 2051, 1770, 1717, 1299, 682, 1641, 2032, 407, 357, 481, 2012, 121, 1215, 1656, 1780, 2225, 1186, 1152, 994, 725, 2224, 2168, 2257, 1685]. exponentially [1442]. exponentials [868]. exponentiated [2051, 1699, 683]. exponentiated-Weibull [683]. exponents [27]. exposure [768]. exposures [2195]. expressed [1406, 822, 2058, 1001, 1197, 2202]. expression [211, 2074, 2078, 1055, 2260, 2084, 2085, 1129, 2104, 2077, 1492, 1924, 2079, 1706, 2318, 2081, 2076, 434, 1137, 934, 1552, 2075]. expressions [1567]. Extended [1623, 1264, 2159, 1914, 151, 2295, 1463, 358, 2333, 758, 1805]. Extending [1025, 182, 1434]. extensible [424]. Extension [668, 1411, 1122, 895, 815, 80, 1628, 1615]. Extensions [1870, 1190, 1107, 454, 1678]. extensive [707]. external [655]. extra [765]. Extraction [1580, 43, 935, 922, 48, 925, 918, 1589]. extrapolation [125, 637]. Extremal [1176, 1666]. Extreme [489, 1811, 94, 1395]. Extreme-value [489, 1811]. extremely [2146]. extremes [1311, 2277, 638]. eyes [9].

F [1361]. face [1203]. Factor [1374, 1957, 1753, 907, 2319, 1493, 1826, 515, 326, 1411, 1752, 2088, 1777, 2210, 1805]. Factor-based [1374]. Factorial [144, 108, 228, 795, 821, 41, 957, 178]. Factorization [1803, 1528]. factors [2296, 163, 535, 615, 618, 1777]. facts [1168]. Faddeev [796]. failure [871, 2329, 191, 1474, 2279, 1801, 2258, 1337]. failure-censored [191]. failures [2249]. false [1993, 990, 2084, 2092, 633, 1427, 1617, 1041]. false-positive [990]. families [2224]. family [1841, 125, 1462, 398, 720, 953, 1503, 1631, 683, 1639]. Fast [2187, 794, 1689, 290, 1177, 502, 1383, 1118, 2250, 1600, 263, 93, 2094, 696, 1868, 997, 1252, 91, 2150, 1063, 169, 940, 2334, 1902, 2016, 1881]. fast-food [169]. Faster [1674, 261, 832]. fat [396]. Faults [1111]. Favorability [1846]. Fay [1912]. Feature [1825, 1835, 1241, 1156, 1787, 2268, 1860]. features [1999, 1586, 672, 292]. feedbacks [738]. Feedforward [48]. fences [1289]. Ferraty [1361]. fertility [1377]. fertilizer [649]. few [1635]. FFT [905]. FFT-based [905]. fiducial [2034]. FIEGARCH [1727]. field [1578, 1583, 1496]. fields [2187, 2226, 1721, 684, 65, 2323, 1689, 1227]. Fieller [249]. fifth [290]. fifth-order [290]. file [287, 314, 321, 366, 383, 393, 402, 412, 421]. filter [1591, 1498, 1957]. Filtering [1580, 2118, 1728, 1595, 918, 1589, 1381]. filters [927, 369]. FIM [1124]. Finance [1726]. Financial [1163, 1171, 1868, 677, 1168, 1266, 619, 1262, 1176]. find [1561]. Finding [618, 983, 544, 1331, 496, 1240]. fine [2194, 2095]. fine-mapping [2095]. Finite [1333, 2113, 335, 1938, 324, 2109, 1409, 2303, 1073, 1405, 1412, 1063, 342, 1254, 675, 1413, 1224, 115, 1839, 1891, 536, 1421, 1677, 331, 585, 931, 1050, 1805]. finite-sample [675]. First [1817, 1832, 262, 2258]. first-failure-censored [2258]. First-passage-time [1817]. Fisher [1959, 1156, 1398, 1044, 1287]. fit [218, 1340, 1299, 1922, 879, 2275, 2344, 1676, 1368, 909, 1352, 26, 109, 564, 2290, 2033, 993, 1075, 750, 1186, 1699, 1296, 417, 2026, 2276, 1715, 2034]. fits [520]. Fitting [614, 1921, 94, 1405, 161, 505, 1075, 329, 336, 1702, 334, 160, 1901, 223, 640, 1721, 2335, 11, 1607, 86, 1630, 1540, 84, 900, 894, 746]. fix [623]. Fixed [1446, 1658, 476, 608, 465, 2118, 2121, 1126, 292, 2158, 1983, 592, 960]. Fixed-effect [1658]. fixed-effects [2121, 1983]. Fixes [1111]. flexibility [1102]. Flexible [1403, 2125, 1619, 981, 1318, 1137, 1901, 2183, 2260, 1631, 1138]. flight [1491]. flow [544, 1615]. fMRI [1876]. follows [287, 314, 321, 366, 383, 393, 402, 412, 421]. food [169]. footprint [2244]. force [163]. Forecast [1273, 1064]. Forecasting [1747, 1612, 1294, 1074, 379, 1023, 1735, 1175, 1377, 2254]. forecasts [1738, 1272, 747]. Forest [1790, 1681, 2000]. form [497, 320, 950, 923, 861, 1153]. formative [1454]. formative-reflective [1454]. forms [2046, 2007, 371]. formula [731, 1582]. formulae [699]. formulations [2121]. Forward [1396, 2129, 1535]. four [1840]. four-part [1840]. Fourier [303, 1734, 1657]. fourth [2107, 166]. fourth-order [166]. fraction [1832]. Fractional [1890, 2119, 1237, 1594, 41, 940, 1324, 995, 1081, 178]. fractionally [372, 825]. Fracture [1243]. frailty [1832, 284, 1410, 256, 522, 819, 1337]. framework [2118, 206, 1586, 1316, 1565, 1960, 1108, 424, 2090]. franchise [169]. fraud [1845]. Fréchet [572]. free [2237, 1177, 170, 1320, 317, 1138, 96]. free-knot [1320, 317]. freedom [588, 408]. Frequency [1152, 1753, 1171, 1154, 253, 928, 369, 1615, 462, 1639]. frequent [2150]. Fridley [2345]. Friedman [387]. Fully [2190, 2320]. function [238, 971, 214, 776, 1905, 51, 356, 889, 151, 811, 1229, 1416, 898, 1713, 726, 1714, 1153, 215, 1702, 212, 1426, 511, 1817, 2337, 69, 745, 1501, 445, 2015, 2034, 685, 2029]. Functional [2120, 1373, 999, 1366, 1939, 1747, 932, 2108, 1367, 1173, 946, 1369, 1672, 1370, 1371, 1368, 603, 1086, 320, 446, 2349, 2057, 1361, 1375, 2123, 861, 1377, 897, 1871, 2054, 1378, 1380, 1784, 2239, 2271, 1973]. functions [969, 1326, 365, 810, 1846, 1086, 125, 1174, 1789, 1308, 1292, 1458, 1807, 2290, 756, 826, 1395, 2231, 468, 599, 1706, 201, 914, 373, 1645, 1622]. fund [1872]. fundamental [301]. funds [619]. further [133]. fusion [1566, 2048, 203, 1582]. fuzzification [1360]. Fuzzy [881, 1029, 1032, 1036, 363, 1693, 1035, 1019, 406, 1013, 1028, 73, 354, 883, 1024, 516, 1034, 1022, 1037, 1025, 1031, 1018, 1026, 1016, 15, 1020, 2211]. fuzzy/crisp [354].

G [1638, 1132, 1426]. G-computation [1132]. GACV [843]. galaxies [305]. games [935]. Gamma [819, 1333, 630, 2031, 1215, 2308, 1124, 361, 1625, 2247]. gaps [1813, 1148]. GARCH [1102, 969, 1199, 731, 1459, 289, 1737, 1634, 2122, 1175, 811, 2132, 926]. gasoline [1259]. Gauss [815, 510, 1287]. Gaussian [1199, 341, 2187, 639, 1553, 38, 1721, 65, 2323, 1409, 1856, 1937, 1689, 1732, 1410, 1768, 781, 2248, 1594, 1866, 468, 1842, 1746, 940, 1960, 1059, 1227, 1731, 1093, 2103, 1137]. Gaussians [1412]. GeD [1778]. GEE [993, 1096, 779, 1798]. gene [211, 2253, 2310, 2062, 2074, 1055, 2084, 2085, 1129, 2104, 2077, 1492, 1924, 2079, 1706, 2318, 2081, 2076, 434, 1551, 961, 934, 1552, 2075]. GeneChip [1925, 2345]. General [454, 2154, 476, 969, 895, 849, 2280, 754, 606, 1827, 1802, 1085, 805]. Generalised [1459, 762, 2339, 654, 797]. generalization [98, 501, 142, 1477, 551]. Generalized [852, 2155, 2275, 2344, 205, 1346, 1768, 723, 1656, 2038, 397, 591, 1014, 685, 22, 2342, 1841, 353, 1342, 1770, 2066, 1966, 533, 2235, 2036, 972, 26, 576, 1217, 1405, 407, 1308, 987, 1812, 1331, 1663, 1215, 1160, 779, 1669, 408, 2281, 230, 993, 2250, 1075, 1880, 2055, 1699, 186, 1702, 1857, 135, 1945, 1793, 361, 1448, 1787, 699, 1960, 2263, 913, 2088, 900, 1302, 2229, 2346, 803, 1093, 1319, 783, 2276, 2257, 294, 2241, 1091]. generate [1579, 624]. generated [1043]. Generating [1804, 2248, 1127, 465, 290, 1215, 1221, 566]. Generation [615, 1822, 940]. generators [1853]. generic [864]. genes [1406, 822, 2058, 2078, 1001, 1197, 2089, 2090]. Genetic [1519, 141, 1487, 177, 613, 614, 2255, 1520, 2128, 2095, 2093, 1521, 1226, 1152, 1240, 1835, 617, 1870, 859, 956, 13, 192, 680]. genetics [2070]. genome [1999, 2094, 1041]. genome-wide [1999, 1041]. genomic [2146, 1999]. genomics [2070, 1531]. genotype [1881]. genotypes [1925, 2345]. genotyping [2093]. geoadditive [1065]. geographic [433]. geographical [2003]. Geometric [958, 6, 630, 1607, 2011, 1652, 1769, 786]. geometrical [814]. geometry [2226, 390]. geostatistical [1743]. getting [341]. GGobi [424]. GI [1638]. GI/G/1 [1638]. Gibbs [2338, 150, 1823, 833, 111]. Gill [1889]. Gini [1554]. given [1804, 2022, 2142]. GLMMs [2191]. Global [306, 1526, 2185, 2155, 1136, 370, 2294, 1063, 1104, 1897, 1434, 1996]. glog [2077]. glog-normal [2077]. GMM [1738]. goal [981]. goals [425]. gold [1618]. good [309, 1854, 1127]. Goodman [1031]. Goodness [1340, 26, 2290, 1186, 2275, 2344, 1676, 1368, 1352, 564, 2033, 993, 1075, 750, 1699, 1296, 417, 2026, 2276, 1715, 2034]. Goodness-of-fit [1340, 26, 2290, 2275, 2344, 1676, 564, 2033, 993, 1075, 750, 1699, 1296, 417, 2026, 1715, 2034]. gp120 [1997]. gp41 [1997]. gradient [1173, 196, 804]. GRAFT [1566]. grafting [203]. grain [1379]. grain-size [1379]. Gram [311]. Granger [1255]. graph [1579, 1142, 2043]. graph-intersection [1142]. graphical [942, 202, 447, 109, 232, 669, 212, 617, 2096, 2103]. graphics [2223]. Graphing [1182]. graphs [2049, 399, 1854, 2096]. Graybill [1441]. grids [794]. Gröbner [795]. Group [2255, 228, 1984, 156, 1321, 1916]. grouped [1299, 2056, 1950]. grouping [870]. groups [1220, 1374, 230, 1662, 2227, 962, 1510, 2213]. growing [190]. growth [556, 28, 700, 29]. Grubbs [2045]. GS [953]. GS-distributions [953]. GSA [1524]. GSA-based [1524]. GTU [1014]. guarantee [682]. guaranteed [2008]. guess [352]. guide [1006]. guided [224].

Haar [963]. Haar-wavelet [963]. Hadamard [309, 509]. Haemolytic [2228]. Half [1261, 22, 1157]. Half-life [1261, 1157]. half-logistic [22]. Hall [1766]. Hall/CRC [1766]. Handbook [1766]. handling [1171, 1279]. haplotype [2095, 1881]. hard [1854]. Hardin [30]. Hardy [2086]. harmonic [1592]. Hastings [219, 150, 674, 532]. hazard [2108, 1534, 628, 365, 1229, 1416, 1065, 1714, 1451, 1694]. hazards [1297, 2175, 1832, 1150, 1828, 362, 1941]. health [1330, 1244, 2198, 1469, 1930, 2269]. heavy [1205, 1634, 2136, 2045, 2337, 2071]. heavy-tailed [1205, 2045, 2337, 2071]. heavy-tails [2136]. hedging [1740]. Hellinger [1733, 2140]. Henderson [1590]. Herriot [1912]. heterogeneity [929, 284, 1110, 2092, 978, 327, 1093, 346]. Heterogeneous [845, 1336, 99, 582]. heteroscedastic [328, 1002, 1601]. heteroscedasticity [1080, 2112, 557, 2129, 1386, 1083, 2240, 36]. heteroskedastic [1928, 1166, 1735, 735, 1862]. heteroskedasticity [2188, 1269, 497, 1165, 736, 950, 1187, 1465]. heteroskedasticity-robust [736]. heuristic [370, 620, 1525, 1897, 1068]. Heuristics [1516, 610]. Hidden [751, 1583, 1748, 1168, 339, 1654, 795, 1183, 2017, 2101, 1496, 1820]. Hierarchical [1105, 1711, 1774, 869, 629, 1581, 1570, 1220, 1436, 1143, 1705, 2196, 1997, 559, 80, 2306, 981, 1114, 824, 1059, 1496, 1296, 1651, 1414, 2286, 523]. Hierarchical-likelihood [1774]. High [1555, 1781, 72, 1552, 906, 1963, 1479, 274, 1171, 304, 1643, 1383, 544, 1998, 2281, 326, 644, 1207, 1550, 1305, 1046, 2190, 1844, 2083, 1940, 2293]. high-breakdown [1383]. High-dimensional [1555, 1781, 72, 1552, 906, 1963, 1479, 304, 1998, 2281, 326, 644, 1207, 1046, 2083]. high-risk [2190]. high-throughput [1305]. higher [830, 171, 295]. higher-order [830, 295]. highest [341, 1261]. Hill [1853]. HIV [1791, 1936, 1122]. HJM [2118]. HMM [934]. HMM-based [934]. Hodrick [1591]. hold [2267]. hold-out [2267]. Hollander [19]. Homogeneity [2316, 687, 476, 2255, 477, 1060, 1944, 2067, 409, 17, 2227, 331, 325]. Homogeneous [1938, 1688, 2272, 625]. honest [190]. Honour [438]. horizontally [1452]. hospital [324]. host [1847]. household [988, 1106]. housing [1270]. httpd [487]. Huber [791]. Human [1925, 2345, 1428]. humoral [2202]. Hurdle [335]. Hurst [517, 1344]. Hybrid [703, 927, 751, 865, 937, 620, 1445, 256, 1198]. hybrids [410]. Hyper [47]. Hyper-EM [47]. hyperbolic [913]. hyperboloids [1764]. hypercube [2039]. hypergeometric [85]. hyperlink [302]. hyperparameters [1073]. hypotheses [754, 633, 1628]. Hypothesis [1606, 85, 1174, 1022, 897, 2131, 1490, 1700, 2238, 1203, 902].

I-censored [1347]. I-Scal [1033]. IBF [869]. ICA [1793]. ICC [1976]. ideas [1014]. Identifiability [1914]. identifiable [1208]. Identification [1744, 841, 1245, 535, 41, 2202, 141, 113, 1643, 872, 240, 1300, 575, 858]. Identification-robust [1744]. Identify [1651, 5, 1556]. Identifying [822, 1197, 954, 1289, 255, 1241, 2058, 1001, 699, 606, 2094, 2228]. identity [2095]. identity-by-descent [2095]. IDL [810]. IFS [1834]. ignorable [873, 1936, 573]. ignoring [2230]. II [1304, 2152, 1717, 2027, 1811, 2119, 937, 980, 2249, 683, 1792]. III [1615]. illustration [169, 377]. illustrations [385]. image [1495, 1797, 1116, 1583, 1874]. image-based [1797]. imagery [296]. images [1327, 2074, 11, 2341, 842, 429, 824]. imaging [1482]. imbalance [1982]. IML [1096]. Impact [2309, 969, 2198]. Impartial [1371]. imperfect [573]. Implementation [295, 160, 2066, 84, 2094, 1550]. Implementing [453, 674, 967, 1253]. implications [1170]. implicit [604]. Implied [2114]. Importance [2145, 102, 71, 1681, 2111, 991, 1336, 1903, 1951, 2138]. imposing [318]. imprecise [202, 1021, 1272, 363]. improper [1098]. improve [695, 1522]. Improved [808, 296, 1126, 1354, 1703, 1135, 2157, 1079, 2347, 369, 667, 646, 1085, 945, 1299, 688, 2158, 816, 833, 713, 694, 561, 1920, 511, 536, 1259, 803, 1223]. improvement [1044]. improvements [733]. Improving [863, 1712, 1447, 199, 1252, 1954, 2186, 1523, 1951, 2064, 1324, 1550]. Imputation [1409, 1633, 1958, 1460, 1310, 529, 1577, 768, 1306, 850, 504]. impute [1962]. imputing [1682]. incidence [1326, 899, 1773, 1510]. incidences [2175]. incident [1220]. included [1496]. Including [1546, 705, 555, 1960]. income [2118, 1258, 1106]. incomplete [45, 1017, 910, 91, 690, 1323, 831, 579, 784, 827, 1777, 1223, 888, 1636]. incompletely [447, 1621]. inconsistency [625]. incorporate [1411, 1698]. incorporating [1791]. incorrectly [138]. Increasing [1919, 1715]. independence [1246, 1230, 2023, 689, 23, 2196, 1657, 1089, 584, 2337, 1911, 2096, 2153]. Independent [375, 1752, 1391, 626, 1450, 983]. independently [647]. Index [21, 50, 78, 1554, 105, 140, 173, 210, 244, 298, 349, 394, 436, 482, 540, 594, 650, 709, 788, 1010, 1666, 1160, 1670, 644, 1920, 829, 1509, 1790, 1941, 1485, 2161, 1627, 874]. Indexing [1803, 1529]. indicator [1968, 1970]. indicators [400]. Indirect [2134, 2133, 916, 2250, 920, 762]. individual [12, 1914, 1077, 157, 2091, 270]. individuals [2093, 337]. inducer [2064]. industrial [25]. inequalities [793, 2116]. inequality [1493]. infant [336]. infarction [2003]. infected [485]. infection [1466]. infectious [1510]. Inference [593, 1464, 381, 2224, 1499, 941, 2108, 1367, 1581, 1717, 378, 711, 1744, 18, 630, 1578, 1208, 497, 1885, 2321, 1767, 1737, 1268, 1733, 1642, 1126, 976, 1235, 2123, 1065, 1714, 1354, 1202, 1378, 2191, 360, 592, 174, 2320, 1972, 2157, 1224, 1132, 1467, 1688, 742, 2138, 1708, 673, 1461, 1622, 783, 1179, 183, 1636, 1876, 1881]. Inferences [2002, 450, 1999, 46, 319, 2063]. inferential [688]. inferiority [2040]. infinite [459]. Inflated [250, 353, 335, 184, 2241]. inflation [2162, 2044]. Influence [1838, 1292, 503, 1443, 147, 361, 2305, 2272, 294, 1466, 1192, 2259, 681, 1431, 1458, 1807, 873, 240, 569, 979, 572, 1332, 2045, 1842, 955, 286, 157, 445, 1601]. Influential [901, 1722, 954, 699, 1920]. influenza [1686]. Information [1213, 851, 1214, 2138, 2108, 1959, 945, 1019, 330, 690, 1517, 559, 785, 2165, 401, 536, 1698, 1665, 2099, 2100, 2153]. information-theoretic [2100]. Informational [206]. Informative [2088]. inhomogeneous [711]. initial [352, 342, 307]. initials [251]. injury [70]. innovation [141]. innovative [577]. Input [1549, 354, 363]. instability [823]. instantaneous [738]. instrument [1053]. instrumental [1612]. instruments [1053]. insurance [1708]. intact [2230]. integer [1061]. integer-valued [1061]. Integral [2237]. integrals [1764]. Integrated [470, 2062, 372, 1174, 1873, 1837, 1734, 825, 1440]. integration [2296, 1890, 1120, 1237]. integrity [306]. intelligence [2244]. intensity [2104]. intensity-dependent [2104]. intensive [255]. Interacting [1649, 1497, 711]. interaction [841, 1239, 1309, 2193, 426]. interactions [387, 1983, 1317]. interactive [424]. intercept [1403, 1232, 2028]. interclass [1119]. interdependence [1736]. interest [1682, 1733, 2128, 1126]. interesting [1754]. interface [1424]. interior [84]. intermittency [167]. intermittent [949]. Internal [1999, 655]. international [177]. Internet [1009, 856]. interpolation [1650, 1765]. interpretable [2020]. Interpretation [1268, 664, 814, 212, 958, 1769].

Interpreting [164]. intersection [1142]. Interval [138, 1082, 2174, 1792, 1050, 88, 1145, 840, 965, 62, 1811, 648, 1080, 1303, 1872, 1230, 1034, 578, 1033, 280, 249, 720, 977, 1065, 2288, 849, 1814, 58, 1630, 980, 1304, 1773, 1079, 2347, 1498, 511, 1223, 1985, 2096, 1877, 1692, 2257, 646, 419, 1909, 2269, 2029]. interval-censored [977, 1065, 2288, 1692]. interval-valued [1230]. intervals [1650, 2109, 990, 2324, 1271, 1433, 2323, 82, 773, 398, 1416, 1975, 2167, 1673, 1680, 740, 1238, 254, 61, 561, 1894, 1450, 549, 767, 143, 1618, 2229, 2346, 1653, 1929, 2171]. intervention [622]. intra [1459, 1735]. intra-daily [1459]. intra-day [1735]. intraclass [1503]. intracluster [1982]. introduced [625]. Introduction [918]. intrusion [1095, 488]. invariant [216]. Inverse [756, 826, 2074, 548, 477, 1410, 781, 2248, 67, 360, 1960, 1551, 1659, 1415, 1985, 1093, 1198]. inverse-variance [2074]. inversion [864]. investigated [1317]. Investigating [956]. investigation [185]. involving [1671, 336, 1480, 1159, 1557]. irregularly [1520]. ISBN [1361]. islets [1428]. isolation [2087, 1821]. Isomap [1530]. Isotonic [644, 1571, 1952]. Israel [20]. Issue [1163, 1726, 1251, 1564, 1516, 438, 1580, 790, 730, 2107, 2203, 918]. issues [3, 432, 30, 573, 31]. Italy [1587]. item [1320]. items [871]. Iterated [991, 106, 922, 180, 859, 792]. Iterative [2019, 160, 1607, 308, 2234, 2228, 1216]. iteratively [491, 1440]. IV [738, 1254].

J [30, 1766]. Jackknife [1666]. Jackpot [257]. Jacobian [2193]. Japanese [1890]. job [1444]. joinpoint [2298]. joinpoints [1185]. Joint [365, 1436, 2035, 1717, 1744, 1221, 1900, 1449, 2249, 2210]. July [30]. Jump [2013, 364, 286]. jumps [317, 2136].

Kalman [1381, 1957]. Kendall [1447, 1182, 152]. Keno [257]. Kernel [1541, 681, 1314, 1295, 130, 123, 2310, 1846, 499, 1825, 1335, 1234, 2264, 1004, 1156, 599, 1839, 81, 1451, 264, 1464, 2089, 1548, 1046, 968]. kernel-based [1839]. kidney [1466]. killing [1892]. kind [191]. Kinderman [623]. KNN [1046]. KNN-kernel [1046]. Knot [1350, 1919, 1320, 317]. knots [494]. knowledge [1017]. known [1359, 566, 1602]. Kohonen [182]. Kolmogorov [51, 558]. Kontoghiorghes [1766]. Kotz [1119]. Kotz-type [1119]. Koul [1502]. kriging [1689, 468]. Kruskal [1031, 802]. Krylov [1707]. Kullback [884]. kurtosis [1712].

L [30, 659]. L-PLSR [659]. laboratory [2001]. LAD [137, 975, 680]. lag [951]. lagged [738, 77]. Lagrange [1253, 2129, 1765]. lambda [1342, 1217, 987, 1663, 1302, 2229, 2346]. Lanczos [543]. Langenberg [1107]. language [1567]. Laplace [1213, 1214, 1922, 1220, 1898, 1701, 851, 1262]. Laplace-EM [1701]. Large [5, 1716, 545, 365, 1824, 794, 2186, 1689, 1025, 292, 1518, 280, 1053, 657, 47, 831, 2033, 935, 580, 1425, 660, 1206, 1277, 454, 1561, 1506, 755, 93, 1418, 469, 1881]. large-sample [280]. large-scale [1518]. largest [113, 648, 1080, 1303, 1872]. LASS [933]. Lasso [1782, 1547, 1775, 1916, 1940, 1283]. Latent [1803, 340, 691, 338, 1264, 1578, 339, 1914, 91, 1624, 570, 515, 988, 615, 659, 661, 434, 2035, 1099, 1677, 1461, 2172, 1529, 1319, 2215, 313, 1454, 2199, 858, 2211]. later [1880]. Latin [2039]. lattice [2193, 931]. law [1262, 1325]. Lawrance [994]. laws [1699]. layer [473, 1203]. layout [564]. LD [1651]. leading [1754]. Learning [1581, 1546, 202, 1527, 2009, 18, 1544, 200, 447, 207, 1560, 1859]. learnings [1530]. Least [876, 1028, 1596, 132, 1502, 1576, 1721, 73, 543, 2312, 1843, 671, 308, 1043, 1598, 697, 658, 801, 1536, 568, 1779, 1599, 746, 664, 1440, 850, 625]. least-squares [1721, 73, 543, 658, 801, 664, 850, 625]. least-trimmed [671, 697]. leave [1690]. leave- [1690]. leaves [247]. left [593, 1122, 2328]. left-censored [1122]. length [721, 188, 324, 1200]. length-biased [1200]. Lepage [1387]. leptokurtosis [396]. lesion [2002]. lethality [189]. letter [1573]. Level [1599, 942, 223, 1943, 1711, 2148, 52, 1915, 2279, 649, 1008, 805]. level-1 [223]. levels [426]. Levene [35]. Leverage [2136, 2252, 1727]. Leverrier [796]. Lévy [1728, 1491, 2142]. Lévy-driven [1728, 2142]. Lewis [994]. libraries [2146]. library [385]. life [2152, 2051, 2008, 1743, 1462, 1261, 1652, 1157, 2168]. life-testing [2168]. Lifetime [1669, 1966, 2329, 1343, 14, 886, 537, 1801, 2171]. lift [1864]. Likelihood [1770, 378, 1504, 1449, 1771, 753, 1717, 1264, 341, 345, 223, 1667, 222, 1654, 498, 1641, 372, 2025, 910, 590, 1393, 356, 492, 92, 2260, 544, 811, 1346, 1123, 493, 66, 1118, 2264, 23, 2158, 774, 1160, 444, 2011, 2191, 1005, 1674, 966, 601, 1131, 1540, 1774, 2304, 1441, 1628, 212, 263, 586, 480, 1861, 525, 782, 1059, 93, 1302, 673, 1820, 783, 1179, 1897, 256, 1225, 1287, 1512, 1692, 2082, 2015, 1616, 1802, 2240, 612, 183, 1524, 523, 1355, 2060, 1076, 332]. Likelihood-based [378, 1118]. Likelihood-ratio [753]. likelihood-ratio-test [1628]. likelihoods [1856, 97]. likeliood [713]. limit [2120]. limitation [2313]. limitations [552]. limited [1584]. limiting [1937]. limits [1362]. Line [1469, 266, 1188]. Linear [1592, 354, 1893, 584, 1794, 870, 1061, 1841, 2154, 25, 1719, 45, 1502, 614, 654, 1574, 1722, 895, 1720, 1369, 306, 598, 693, 732, 1958, 1028, 2235, 752, 73, 1029, 793, 1965, 909, 1220, 248, 1040, 89, 1233, 1116, 473, 1230, 474, 1695, 2121, 1204, 1937, 1405, 1328, 1949, 175, 1619, 1812, 1292, 1394, 2006, 1486, 110, 1002, 1535, 912, 2010, 1768, 1604, 1160, 1983, 697, 658, 834, 230, 849, 1097, 592, 1329, 2157, 186, 1702, 1857, 572, 1630, 1783, 154, 1658, 1332, 1448, 1628, 61, 2317, 637, 699, 511, 1589, 216, 1950, 1870, 622, 1757, 1602]. linear [213, 701, 79, 958, 1582, 1549, 1694, 900, 901, 1941, 1259, 1548, 1550, 963, 1319, 397, 1445, 783, 2272, 1605, 1934, 469, 128, 2240, 363, 294, 1091, 2083, 2060, 1068, 2029, 985]. linearity [1155, 57]. linearization [1225]. link [1702, 1509, 1505]. linkage [2255, 2097, 857, 2096]. linked [2102]. links [1621, 1561]. liquidity [1500]. litter [1639]. LMS [1188]. LMSV [1727]. load [1907]. load-sharing [1907]. Local [1719, 2259, 1160, 154, 955, 1081, 445, 1601, 818, 2185, 1035, 136, 1407, 1623, 1958, 1695, 1431, 1472, 1292, 1458, 1395, 873, 978, 703, 1332, 1070, 1207, 1427, 1108, 1694, 933, 1008, 1539, 1660]. Localized [1887]. Locally [1780, 2247, 1585, 221]. Locating [1497]. location [2200, 475, 479, 546, 351, 518, 35, 547, 1817, 72, 1389]. location-scale [35]. locations [961]. loci [1497, 2098, 2102, 2082]. Loevinger [853]. Log [1810, 1795, 2193, 2109, 1484, 2323, 248, 94, 1420, 976, 151, 912, 1508, 1443, 361, 495, 2333, 2268, 1302, 2015, 1802, 1356, 88]. log- [94]. log-Birnbaum [1443, 1356]. Log-Burr [1795]. log-concave [1484, 1420]. log-determinants [495]. log-extended [2333]. log-gamma [361]. log-Gaussian [2323]. Log-Jacobian [2193]. log-likelihood [2015, 1802]. log-linear [248, 912]. Log-modified [1810]. log-normal [976]. log-normally [88]. log-rank [151, 1508]. log-spectral [2268]. logarithmic [26]. Logical [1211]. logistic [22, 906, 1939, 1340, 2300, 214, 1846, 453, 1267, 717, 1204, 1120, 706, 2166, 138, 249, 1129, 70, 542, 1280, 979, 866, 1277, 560, 1787, 1986, 399, 416, 1296, 1314, 1775, 1715, 1552, 1471]. logistic-normal [1120]. Logit [219, 2175, 1373, 278, 377, 1628, 1505]. LogitBoost [1704]. loglinear [538, 2286, 2184]. lognormal [2251, 879, 1308, 1969, 2055, 1897, 2269]. Logrank [977]. Logrank-type [977]. Lomax [719]. Long [1791, 1164, 1638, 271, 1459, 1740, 1762, 262, 911, 1329, 1588, 2135, 1759, 2304, 1842, 2252, 333, 1727, 705]. long- [705]. long-memory [1459, 911, 1329, 1842, 2252, 1727]. long-range [271]. long-run [1762, 262]. long-tailed [1638]. Long-term [1791, 333]. Longitudinal [970, 2301, 1747, 1327, 1633, 2338, 556, 1626, 1965, 2025, 949, 1991, 1504, 1431, 457, 1475, 1449, 2309, 1131, 1132, 126, 1332, 2035, 1867, 1913]. look [7]. Looking [197]. losers [1981]. loss [815, 1787, 602, 1742]. lottery [845]. Low [125, 1615, 2311, 2293]. Low-dose [125]. low-flow [1615]. Lower [259, 1770, 2257]. LR [406, 1028, 1716]. LTS [696]. lumps [197]. Luo [1386]. Lupus [1939]. Lyapunov [27].

M [30, 1426, 1187]. M/G/ [1426]. Machine [1527, 1706, 2089, 1815, 2088]. machine-based [2088]. machines [1543, 1494]. macroeconometric [737]. macroeconomic [1266]. macros [578]. magnetic [824]. main [186]. main-effects [186]. Maintenance [2150]. major [1044]. majorization [308, 2149]. make [1854]. Making [319, 9]. malaria [2232]. malicious [1241]. Mallows [1769]. malware [2064]. management [1742]. Mandelbrot [1123]. manifold [1530]. Mann [1331]. MANOVA [1812]. many [772, 1480]. map [2130, 182, 1463, 714, 914]. MAPLE [193]. Mapping [1925, 2232, 2345, 941, 345, 2098, 655, 1285, 2095, 2191, 2099, 2190]. maps [600]. Marginal [1856, 465, 820, 566, 1059, 2142, 1742, 888, 1216, 523, 749]. marginally [1432]. marginals [2124]. margins [608, 913]. mark [1239]. mark/point [1239]. marked [102, 1239, 1133, 1058]. marked-point [1239]. marker [2102, 1992, 2089, 1318]. market [2118, 1270, 1741, 1463, 1103, 747]. marketing [644]. markets [1744, 2114]. Markov [112, 1264, 1718, 2187, 1168, 1345, 1721, 1654, 948, 1737, 1736, 751, 1183, 2023, 1689, 1583, 1990, 2017, 2101, 1688, 2331, 1468, 1808, 1748, 2274, 1341, 1227, 1820, 2100, 13, 504]. Markov-switching [1737]. Markov/semi [751]. Markov/semi-Markov [751]. Markovian [166, 830, 1836]. Marshall [2018]. martingale [1174]. masked [1652]. masquerade [1816]. mass [1532, 1904]. matched [583]. Matching [375, 1958, 1190, 1577, 1428]. maternity [1832]. mathematical [645, 1567]. matrices [1213, 763, 892, 1679, 1130, 303, 309, 509, 2121, 796, 1323, 938, 837, 2079, 851, 1214, 1793, 495, 1506, 1391, 1875]. Matrix [1803, 790, 1528, 2233, 107, 308, 301, 80, 224, 659, 186, 1716, 469, 1390, 2211]. matrix-variate [80]. matter [2194]. MAVE [1844]. Maximally [1488, 404]. Maximin [1913]. maximization [1820]. Maximizing [1741, 97, 152]. Maximum [1264, 1667, 92, 1123, 2264, 836, 774, 444, 2011, 2304, 1779, 586, 1287, 1512, 345, 222, 1641, 372, 910, 544, 493, 66, 2158, 1075, 1005, 14, 1674, 966, 713, 1131, 1441, 263, 1861, 525, 93, 1302, 1897, 256, 1225, 1216, 1692, 2285, 1616, 612, 1524, 332, 782]. MCMC [1054, 892, 2296, 1268, 1624, 1776, 775, 1951, 1202, 376, 2013, 1070, 1424, 286, 1731, 1957]. MCMC-based [1070, 1957]. MDS [1861]. Mean [1578, 982, 778, 465, 946, 1955, 648, 1080, 1303, 1872, 1721, 46, 1204, 976, 1308, 1002, 816, 1975, 2038, 188, 2320, 1670, 2017, 1441, 1659, 1836, 143, 857, 2171, 1085, 1956]. Mean-field [1578]. means [614, 1693, 33, 1371, 883, 574, 1021, 771, 1967, 1156, 241, 1680, 58, 670, 1904, 1236, 2013, 2227, 1093, 1860, 144, 1083, 183, 88]. measurabilities [1018]. measure [2041, 54, 120, 1195, 1031, 449, 572, 1100, 1301, 79]. measured [1359, 1925, 2345]. measurement [2028, 92, 162, 301, 604, 1604, 1473, 1247, 1936, 955, 1603, 1605, 1934, 1601]. measurements [647, 266, 320, 1679, 1797, 2037, 837, 1291]. Measures [218, 1963, 2332, 1681, 853, 838, 452, 681, 785, 28, 2160, 1379, 519, 75, 979, 1945, 1986, 2138, 1640, 1652, 156, 1142, 1091, 1786]. Measuring [145, 515, 157, 2180]. mechanism [177, 1860]. median [1668, 876, 927]. medians [1388, 1149, 962]. mediation [1242]. medical [1840, 1444, 88]. membership [829]. memory [1164, 1459, 1740, 1584, 911, 1329, 1588, 2135, 2304, 1842, 2252, 1727]. merging [1095]. message [721]. Meta [1068, 2306, 42, 1007]. meta-analysis [2306, 42, 1007]. Meta-heuristic [1068]. method [514, 62, 1, 688, 998, 2062, 1570, 1430, 1062, 146, 1368, 1750, 239, 543, 447, 1949, 1094, 232, 1001, 528, 2173, 1221, 358, 2301, 2250, 1935, 2308, 1100, 597, 939, 53, 713, 1630, 1882, 605, 127, 2234, 1560, 1861, 2071, 606, 10, 2193, 281, 2139, 803, 623, 715, 1306, 1539, 1844, 686, 1616, 1184, 805, 1510, 1337, 1881, 680]. methodologies [948, 2091]. methodology [1693, 1023, 761, 1666, 814, 1331, 515, 1866, 782, 1321]. Methods [1726, 1546, 400, 2178, 1382, 130, 1987, 1054, 387, 2203, 666, 199, 223, 2098, 752, 1641, 1260, 761, 1024, 948, 1556, 1034, 1613, 356, 2182, 1383, 83, 721, 727, 1218, 2010, 723, 1284, 1594, 1449, 1924, 626, 1329, 1657, 1084, 1388, 1783, 2003, 84, 1115, 1561, 925, 2076, 1589, 1609, 2297, 1815, 49, 1525, 1775, 2228, 1957, 786, 1892, 1385, 13, 418, 63, 2210, 1198]. methylation [2085]. metric [947, 3, 1655, 30, 1077, 307, 572, 31]. Metropolis [219, 150, 1765, 532]. MEV [846]. MFACT [514]. MH [1823]. micro [2145]. micro-aggregation [2145]. microarray [1406, 211, 2072, 841, 2310, 2074, 822, 1055, 2260, 2084, 457, 1003, 2173, 1160, 821, 707, 632, 1197, 1924, 2073, 633, 568, 2071, 1815, 715, 1159, 1622, 667, 2082, 2202, 2083]. microarray-based [568]. microarrays [456, 434, 1551]. Microsoft [1111, 679, 291, 787, 1851, 1853, 1850, 1854, 1852]. microstructure [1734]. migration [1263]. minima [818]. minimax [1040, 115, 963]. Minimizing [1243, 1779]. Minimum [721, 1961, 1140, 1440, 491, 531, 1245, 1136, 1733, 478, 100, 952, 409, 1089, 410, 212, 1435, 2140, 686]. Mining [206, 1527, 860, 425, 1544, 205, 208, 169, 76, 1101, 1561, 1301]. Minkowski [6]. mis [625]. mis-constraining [625]. misclassification [1276, 990, 1481, 1429, 513, 1228, 270]. misclassified [780, 2274]. Mises [2344, 2275, 534, 336]. miss [1431]. miss-specification [1431]. Missing [1577, 267, 760, 944, 991, 2262, 1211, 1404, 2259, 949, 2166, 330, 911, 844, 873, 1975, 1936, 1831, 573, 236, 1779, 982, 768, 1644, 2278, 1962, 1968, 1306, 850, 448, 1184]. missingness [1972]. Missouri [1790]. misspecification [634]. misspecifications [1841]. misspecified [129, 2124, 1394, 1475, 1259, 119]. misspecifying [59]. Mixed [1285, 1065, 1841, 1987, 1367, 1269, 1667, 222, 2273, 588, 2025, 590, 474, 1204, 1126, 1328, 59, 1619, 330, 1094, 1394, 1893, 1898, 2343, 1768, 774, 2277, 503, 2192, 1983, 849, 1936, 2148, 592, 2079, 2157, 518, 691, 1915, 1784, 1774, 1658, 1448, 2305, 1122, 1757, 1665, 673, 589, 591, 1445, 783, 1248, 2015, 1184, 294, 2241, 1091, 805, 1678]. mixed-effects [1367, 1126, 503, 1983, 1936, 1774, 2305, 1122, 1184]. mixed-level [2148, 1915, 805]. mixing [1406, 1723, 833, 2331, 868]. MIXMOD [1051]. Mixture [1928, 328, 1401, 335, 1404, 330, 1415, 2156, 1632, 1333, 211, 1327, 1199, 1753, 341, 2086, 891, 323, 1419, 1553, 1208, 205, 766, 1409, 1195, 1267, 1856, 2319, 1268, 1634, 1751, 2122, 357, 1055, 1412, 1438, 1417, 1768, 1931, 800, 856, 2169, 1413, 1411, 337, 2073, 343, 333, 81, 824, 1427, 2224, 331, 1418, 325, 585, 1305, 2314, 1414, 1225, 1512, 2285, 1334, 448, 1050, 324, 1805, 332]. Mixtures [1244, 347, 1213, 704, 411, 470, 1899, 1403, 1407, 1484, 339, 1420, 1233, 1408, 1073, 344, 1405, 175, 1063, 342, 1416, 1078, 726, 2288, 1402, 326, 631, 329, 336, 851, 1214, 1540, 2013, 1688, 1114, 1421, 868]. ML [2111, 256]. ML-EM [256]. MLE [1935]. MLEs [1869]. modal [1372]. mode [220, 997, 149, 1308, 1730, 1664, 1128, 2215]. Model [130, 1051, 318, 1626, 1556, 724, 1348, 1543, 11, 1945, 560, 1748, 1088, 1297, 218, 1747, 527, 211, 219, 353, 1696, 1173, 1199, 970, 1497, 2188, 1502, 731, 1264, 945, 2300, 2086, 891, 486, 666, 1419, 1257, 820, 1720, 1553, 2310, 1811, 32, 12, 698, 813, 817, 1832, 205, 1028, 2270, 766, 883, 2048, 882, 1436, 1278, 284, 1373, 576, 2025, 1942, 2126, 1258, 1139, 340, 312, 2127, 538, 2028, 1500, 1285, 208, 2121, 1204, 1912, 2232, 1328, 2315, 492, 2023, 1175, 2289, 2294, 1812, 151, 330, 814, 1094, 1438, 1394, 1889, 1279, 1417, 441]. model [2312, 1741, 1150, 1535, 1410, 669, 2010, 1065, 1768, 1671, 301, 604, 1604, 904, 2277, 444, 1873, 1983, 542, 988, 1771, 230, 1280, 28, 134, 1840, 1936, 2306, 240, 668, 1107, 1935, 2160, 592, 2079, 2280, 1413, 1411, 518, 2053, 2308, 468, 242, 1828, 2073, 1857, 376, 644, 725, 86, 1630, 1474, 1761, 2101, 2003, 886, 2045, 1079, 2347, 1862, 1628, 784, 827, 835, 1631, 699, 1920, 1480, 2263, 2271, 2274, 636, 1861, 1902, 1122, 743, 1652, 15, 995, 331, 116, 1665, 69, 2279, 2333, 1158, 281, 1496, 1104, 744, 924, 1968, 673, 77, 2172, 894, 1305, 1820, 1557]. model [1605, 2215, 1414, 1934, 2199, 2168, 1710, 2197, 485, 13, 128, 1627, 1050, 184, 1524, 324, 902, 1337, 1805, 1686, 1918, 2211, 985, 1471]. Model-based [1051, 1626, 1556, 2294, 1065, 2280, 1862]. Model-robust [724, 636]. Model-sensitive [1088, 724]. Modeling [1745, 1363, 355, 684, 1320, 2001, 1596, 485, 1050, 629, 760, 328, 1966, 318, 556, 5, 1143, 2186, 1270, 356, 1624, 262, 1279, 1178, 583, 856, 2191, 1866, 2280, 14, 1006, 1089, 1155, 757, 2081, 1114, 782, 8, 663, 1454, 2156, 2230, 1875, 1876]. Modelling [1163, 946, 1890, 1407, 2198, 1743, 248, 1964, 1634, 1516, 2189, 326, 1457, 1727, 1105, 271, 365, 289, 1196, 204, 2120, 344, 1619, 912, 1588, 1100, 177, 401, 824, 2035, 1960, 714, 617, 1455, 2326, 1036, 2096, 1243, 430, 610, 1137]. Models [1401, 250, 863, 629, 1061, 129, 1939, 508, 942, 808, 1841, 1327, 2242, 1519, 2154, 1102, 1719, 2072, 1987, 1367, 1340, 969, 1219, 1581, 614, 1606, 378, 2111, 2113, 1932, 1928, 271, 1299, 628, 341, 1919, 1722, 323, 1744, 1035, 2200, 1459, 202, 895, 919, 223, 1168, 1403, 1810, 1345, 2338, 598, 693, 1166, 1735, 732, 929, 1846, 222, 1578, 406, 2235, 2036, 1988, 1728, 752, 335, 1654, 1167, 1965, 909, 916, 1409, 1195, 1220, 74, 372, 557, 94, 588, 1256, 89, 1767, 735, 1914, 439, 1737, 1989, 1267, 1856, 2259, 91, 382, 590, 447]. models [150, 1268, 1634, 1733, 1230, 370, 474, 736, 950, 1253, 923, 1642, 1120, 1751, 578, 1937, 1823, 1027, 1126, 2065, 59, 999, 2195, 175, 161, 1055, 1705, 1482, 452, 2123, 1619, 1518, 1569, 389, 2182, 1348, 1791, 911, 1353, 811, 1346, 1123, 1486, 844, 1893, 737, 1898, 2343, 738, 1193, 110, 2196, 1517, 1974, 2037, 722, 1701, 47, 251, 1169, 831, 898, 1713, 1997, 1583, 1813, 1493, 1074, 570, 1446, 1671, 774, 70, 121, 833, 676, 2104, 1160, 80, 1931, 1469, 2178, 444, 503, 873, 2192, 1443, 2330, 800, 1449, 2309, 834, 230, 978, 849, 993, 2132, 2133, 467, 2169, 580]. models [2160, 1005, 1646, 174, 1972, 966, 2157, 2291, 186, 380, 75, 52, 700, 377, 575, 691, 1702, 979, 572, 2017, 347, 2136, 53, 713, 1784, 343, 1774, 1798, 126, 236, 1238, 1730, 374, 1658, 361, 1332, 981, 1842, 1448, 926, 362, 334, 1265, 2252, 637, 333, 1099, 1480, 893, 955, 1808, 480, 1640, 1427, 399, 1662, 617, 1324, 2305, 622, 1757, 1953, 913, 1341, 2047, 1603, 1187, 1081, 1795, 1059, 93, 2193, 29, 49, 900, 1418, 1140, 1731, 2302, 1296, 530, 1941, 1259, 417, 589, 2140, 2141, 869, 585, 2096, 963, 1008, 1319, 397, 591, 2100, 1445]. models [960, 783, 119, 338, 2103, 256, 522, 1946, 664, 1248, 1225, 1512, 2161, 2285, 2015, 469, 2240, 192, 363, 1184, 294, 1466, 1356, 1715, 2327, 2241, 1091, 1434, 858, 2286, 1170, 819, 1185, 1632, 442, 1390, 1389, 685, 1176, 1507, 1786, 2184, 1068, 1940, 680, 874, 641, 1697, 1601, 2213, 2116, 1322, 332]. modification [1534, 318, 583]. modifications [1693, 1938, 1316]. Modified [414, 2334, 855, 2083, 214, 1966, 1810, 74, 311, 796, 249, 1279, 1387, 886, 1802, 819]. moment [130, 414, 855, 1796]. moment-based [1796]. moments [1342, 193, 1641, 1883, 415, 1078, 1663, 555, 2250, 371]. monetary [745]. monitor [1311]. monitoring [451, 1833, 1293, 517, 2164, 434, 622, 2060]. monothetic [1570]. monotone [944, 1412, 844, 756, 826]. monotonic [1488, 1832, 746]. Monte [1722, 226, 2338, 813, 1260, 948, 2124, 97, 695, 1649, 1814, 351, 2135, 53, 1448, 2331, 893, 1341, 1358, 708, 13]. Moran [2185, 1476]. Morris [1232]. mortality [1105, 2194, 1377, 2003, 768, 537]. mosaicplots [431]. Most [778, 1754]. motion [301]. motions [2119]. motivated [1003]. motor [163, 70]. movement [455]. moving [372, 1147, 1593]. MRI [2002]. MSAE [269]. MSE [1961]. MTAR [827, 784]. MTD [328]. Multi [2287, 2117, 205, 473, 1021, 99, 546, 2064, 1664, 2326, 620, 2279, 1321, 346, 1763, 1203, 1844, 1996]. Multi-aspect [2287]. multi-center [99, 346]. multi-class [1763, 1996]. multi-dimensional [1844]. multi-inducer [2064]. multi-layer [473, 1203]. multi-level [2279]. multi-mixture [205]. multi-objective [620]. multi-partitioning [1664]. multi-sample [1021, 546]. Multi-step [2117]. multi-table [1321]. multi-taper [2326]. multiblocks [551]. multicenter [284]. multichannel [2321, 928]. Multiclass [2253, 1129]. multicollinear [906, 661]. Multidimensional [1745, 728, 1033, 12, 1146, 2214, 794, 1032, 798, 307, 914, 1858]. multilabel [605]. Multilayer [392, 552]. Multilevel [1270, 17, 629, 223, 1232, 569, 1006, 2044, 560, 480, 960, 442]. multilinear [1584]. multimode [1557]. multinomial [2175, 214, 896, 398, 1123, 2033, 231, 2011, 580, 1425, 53, 713, 1882, 954, 1421, 1786]. multiperiod [1513]. Multiple [1753, 670, 768, 986, 1497, 411, 1633, 1901, 1363, 1196, 822, 1298, 655, 543, 1130, 1556, 2078, 884, 478, 161, 185, 319, 1776, 1486, 1764, 1129, 515, 230, 2225, 553, 579, 2308, 1300, 1388, 1904, 2317, 286, 2099, 2040, 143, 513, 901, 87, 1295, 450, 128, 36, 1050, 1185, 1930, 2210, 504, 1918, 2213, 818]. multiple-comparison [1388]. multiple-treatment [822]. multiplexed [1892]. multiplicative [1954, 1432, 1125]. multiplier [1253, 2129]. multiply [1962]. Multiresolution [375, 1711]. multiresponse [1549, 530]. Multiscale [1759, 1134, 2268]. multisensor [1582]. multistage [1668]. multitype [1613]. Multivariable [995, 722, 809]. Multivariate [930, 1269, 2273, 2032, 1255, 1732, 671, 837, 1588, 1114, 221, 913, 8, 1008, 653, 527, 942, 2340, 1538, 969, 1886, 944, 2113, 113, 341, 545, 2115, 647, 451, 1667, 85, 2275, 2344, 1196, 881, 2270, 684, 1778, 1165, 1825, 1571, 1139, 1404, 535, 1246, 1642, 1912, 2122, 884, 1055, 290, 987, 1812, 1412, 1293, 2312, 816, 2288, 1604, 779, 860, 1056, 1771, 2038, 1411, 1657, 966, 1749, 1857, 572, 597, 939, 547, 2022, 177, 2035, 2271, 622, 2302, 1869, 1046, 255, 2199, 1710, 2180, 1796, 1956, 1632, 968, 1786]. mutual [690, 559, 785, 619, 2153]. MV [467]. MV- [467]. myocardial [2003].

Nadarajah [1213]. Nasdaq [1172]. national [1608]. nations [2244]. nature [844]. NDSCAL [814]. near [260, 1915, 1875]. near-orthogonal [1915]. Nearest [850, 247, 2179, 2062, 1460, 974, 1339]. nearest-neighbor [1339]. Negative [2343, 2002, 1803, 501, 811, 1889, 2007, 2022, 1822, 1421]. negative-multinomial [1421]. negativist [381]. neighbor [2342, 247, 2179, 2062, 974, 1339]. neighborhood [1280, 2197]. neighbors [2043]. neighbour [1460]. neighbours [850]. neonatal [324]. nested [1094, 441, 2012, 658, 377, 2279]. network [1519, 773, 1294, 544, 164, 2322, 1463, 676, 1095, 63, 1044]. networks [677, 18, 1804, 1672, 339, 1487, 1695, 182, 2009, 1621, 48, 76, 1101, 1779, 1675, 2348, 71, 1501]. Neural [18, 1695, 1519, 1294, 164, 676, 48, 76, 1187, 63]. neural-network [164]. neuro [2244]. neuro-computational [2244]. neuromuscular [769]. neurophysiological [727]. Newton [97, 189, 815, 1287]. Newton-based [189]. NHPP [1784, 1808]. nine [1290]. NLMIXED [1006]. no [3, 30, 564, 31]. noise [1921, 1958, 1237, 1594, 1734, 1746, 940]. Noisy [60, 266, 842]. nomial [1574]. nominal [1786]. nominators [64]. Non [1803, 1420, 1233, 1613, 1559, 2231, 2228, 506, 969, 704, 345, 1488, 38, 600, 817, 1832, 1965, 529, 89, 1856, 2232, 987, 1732, 564, 2196, 2093, 70, 658, 1936, 2007, 1291, 1831, 253, 1155, 377, 1398, 928, 1688, 772, 2234, 1059, 1158, 2040, 900, 1731, 87, 1539]. non-additivity [564]. non-Bayesian [1158]. non-central [2007]. non-classical [506]. non-compliance [1291, 1831]. non-Gaussian [38, 1856, 1732, 1059, 1731]. non-genetic [2093]. non-hierarchical [2196]. non-homogeneous [1688]. non-ignorable [1936]. non-inferiority [2040]. Non-iterative [2228]. Non-linear [1233, 1965, 658, 900]. non-linearity [1155]. non-monotonic [1488, 1832]. Non-negative [1803, 2007]. non-nested [377]. non-normal [704, 987, 87]. Non-parametric [1420, 1613, 2231, 969, 345, 600, 817, 89, 70, 1398, 2234, 1539]. non-response [529]. non-stationary [2232, 253, 928]. Non-symmetric [1559]. non-trivial [772]. noncentral [411, 1222, 2046, 44]. noncompliance [1814, 1944, 2067, 2174]. Nonconservative [1499]. nongaussianity [1807]. nonhomogeneous [1339, 1325]. nonidentifiable [113]. Nonignorable [2338, 2259, 1972, 1962, 1911, 1184]. Nonlinear [1163, 1172, 1512, 1987, 1407, 1672, 2117, 2118, 222, 2036, 355, 1136, 2052, 2259, 590, 1294, 1642, 1126, 59, 2289, 2335, 1843, 40, 774, 444, 503, 873, 2330, 978, 837, 1936, 2306, 1774, 126, 2035, 2305, 1122, 2147, 1314, 1572, 1445, 1248, 1844, 1184, 2327, 1170, 685, 1322]. nonlinearities [1634]. nonlinearity [285]. nonmetric [90]. nonnegative [1528]. nonnested [1166]. nonnormal [840, 1288, 290, 597, 939, 1645]. nonnormality [848]. nonnull [227]. Nonparametric [971, 1, 2292, 1690, 1433, 39, 2349, 303, 839, 1994, 1709, 579, 1388, 2317, 859, 887, 1694, 1382, 1909, 2325, 504, 1200, 1987, 1886, 62, 199, 385, 289, 2027, 54, 120, 1958, 82, 446, 2232, 2315, 2289, 1177, 583, 898, 1671, 992, 627, 479, 546, 380, 215, 1874, 1662, 1794, 1338, 1622, 1225, 1692, 1148, 1940, 332, 1361]. nonresponse [1911]. Nonstandard [1212, 1517]. nonstationarity [271, 784, 827]. nonstationary [639, 1005, 922]. nonsymmetric [2209]. norm [880]. normal [704, 944, 1269, 1359, 470, 2046, 1899, 702, 1955, 648, 1080, 1303, 1872, 46, 574, 1634, 1120, 976, 2122, 1175, 1055, 987, 1438, 726, 2010, 2288, 2077, 2, 2038, 2007, 1153, 188, 2053, 52, 343, 371, 2013, 521, 1461, 585, 87, 1869, 1083, 36, 2327, 2247]. normal-GARCH [1175]. normality [274, 1676, 1205, 1094, 1397, 2281, 755, 521, 753]. normalization [2104, 1815]. normally [1107, 88]. normals [1403]. note [464, 155, 55, 100, 2012, 386, 2252, 512, 1216, 1916, 2015, 2257, 2029]. Notes [1831, 1814, 525]. notice [2348]. novel [1094, 2104, 1796, 934]. NPMLE [868]. nuclear [2326]. nuisance [1397]. null [778, 1430, 588, 2028, 457, 2010, 2225, 633]. number [998, 1679, 175, 2084, 2288, 121, 935, 1425, 166, 830, 1853, 329, 638, 2249, 772, 331, 1617, 1777, 325, 585, 2256]. numbers [1351, 1127]. Numerical [306, 1120, 1302, 2146, 834, 1153, 171, 263, 804, 749]. Numerically [310, 381].

O. [1214]. Objective [2315, 620]. observation [1144, 455, 1148]. observations [971, 312, 911, 80, 1097, 1107, 267, 954, 236, 699, 1920, 602, 957, 255, 1636]. observed [1642, 2093, 1621]. Obtaining [2010]. occasions [1511]. occurring [1598]. ocean [1104]. oceanology [1380]. odds [716, 2065, 2174, 1986, 986]. Off [1845]. Off-the-peg [1845]. oil [2114]. Olkin [2018]. OLS [1254]. OLSE [77]. omitted [956]. One [1686, 1963, 25, 476, 1447, 1433, 2078, 34, 1967, 1812, 52, 135, 254, 602, 69, 156, 2149]. one-parameter [69]. one-sided [1967, 135, 156]. One-step [1686]. one-way [25, 476, 1812, 52]. Online [856, 2137, 1063]. only [1447]. onto [16]. operating [2325]. operational [1022, 1742]. operator [1212, 15]. operators [1211]. opsonophagocytic [1892]. optima [1472]. Optimal [159, 648, 1080, 1872, 732, 1584, 2093, 1691, 2131, 580, 2307, 957, 29, 1041, 1453, 32, 1303, 1954, 309, 168, 1272, 1949, 1331, 1053, 1780, 615, 700, 494, 1444, 636, 1088, 282, 1913, 963, 2339]. optimisation [1517, 867, 1525]. Optimising [771]. Optimization [1516, 1857, 1206, 1971, 1526, 370, 2294, 308, 799, 460, 620, 1216, 610]. optimize [373]. Optimizing [2080, 1125]. Optimum [1231, 1439, 973, 974, 467]. option [731, 747]. options [1260, 1764, 1729, 747, 1170]. Order [360, 2340, 2162, 193, 1278, 2120, 716, 951, 290, 1353, 2295, 2006, 232, 1900, 166, 830, 1282, 742, 295]. order-restricted [716, 1353]. Ordered [278, 628, 910, 1031, 2011, 579, 670, 1388, 1904, 1151, 744, 1763, 2207]. ordering [428, 983]. orders [1147, 282]. ordinal [99, 689, 1475, 1771, 1559, 1798, 1986, 1461, 1179, 1710]. ordinalized [1682]. ordinary [613]. organizing [182, 1463, 914]. orientation [1243]. origin [876]. Ornstein [2142]. Orthogonal [2039, 2292, 1559, 1915, 870]. Orthogonal-column [2039]. other [997, 1577, 1855]. otolaryngologic [1788]. out-of-sample [239, 1255, 1858]. outbreaks [988]. outcome [2270]. outcomes [1982, 1403, 1196, 1244, 2198, 1682, 2273, 2259, 2093, 1831, 1006, 2053, 376, 2035, 523]. Outlier [1643, 478, 113, 64, 38, 577, 2262, 117, 2130, 429, 468, 128]. outliers [141, 451, 2225, 1300, 416, 606, 1289, 1644, 2278, 448]. outlying [233]. outperform [1823]. output [2338, 354, 1029, 363]. outputs [1776]. outs [1431]. over- [452]. over-complete [1430]. over-dispersion [408]. Overall [2179]. overdispersed [1960]. overdispersion [2162, 2189, 713]. overlap [145, 58]. overlapping [54, 120, 887]. overview [1687]. Ozark [1790]. ozone [2282].

P [1361]. package [1569, 2288, 1568, 1986]. packages [1211, 771, 1290]. Pagan [1000]. Pair [1458, 1807, 2102, 1793]. Pair-perturbation [1458, 1807]. paired [248, 669, 252]. pairs [735, 2097, 857]. Pairwise [2025, 783, 1179, 962, 711, 2179, 1967, 480]. pancreatic [1428]. panel [2188, 1264, 1307, 929, 1000, 1938, 441, 2131, 1784, 1341, 2050, 1259]. papers [596]. Paradigm [206, 488]. PARAFAC [894, 802]. Parallel [264, 1227, 430, 1766, 698, 433, 570, 1907, 2193, 459]. Parallelizing [1189]. Parameter [2274, 513, 1319, 2251, 990, 124, 533, 879, 74, 1767, 1478, 262, 1741, 2031, 1843, 781, 359, 1354, 569, 627, 242, 414, 980, 1304, 115, 374, 956, 525, 1615, 69, 1081, 1827, 1801, 2026, 1897, 855, 1616, 646, 1068]. Parameterisation [1731]. parameters [22, 508, 613, 1932, 682, 107, 1217, 1762, 1204, 1705, 1362, 1123, 1043, 2018, 231, 579, 2164, 235, 1867, 1138, 683, 513, 1618, 1775, 1593, 1572, 686, 1465, 192, 363, 1434]. Parametric [1330, 2098, 1989, 1671, 35, 2163, 1145, 969, 1606, 387, 345, 1672, 600, 817, 1420, 1723, 89, 1613, 2315, 539, 280, 1218, 1446, 1998, 70, 2231, 922, 1972, 1398, 1474, 1761, 1070, 2234, 1427, 530, 1539, 1068, 874]. Parametrization [1988]. Pareto [1213, 1214, 388, 533, 1167, 1090, 191, 1075, 851, 1485, 2276, 1792]. Pareto-type [388, 1485]. Parsimonious [1219]. part [1840, 960, 874]. Partial [543, 1719, 889, 1486, 510, 658, 568, 1995, 829, 1529, 664, 2240]. Partially [1145, 2152, 598, 693, 1208, 834, 1005, 1941]. particle [867]. particulate [2194]. Partition [2293, 204, 240, 668]. partitional [867]. partitioned [1452, 938]. partitioning [1577, 1057, 2145, 1664, 1128]. partitions [1479]. partly [1722]. passage [1817]. past [319]. path [2127, 1279, 663, 1454]. path-dependent [2127]. pathological [1092]. paths [1525]. patients [1092, 365, 1466]. pattern [545, 1037, 2150, 1996]. patterns [854, 841, 907, 109, 224, 1880, 921, 661, 1133, 2094, 1622, 1072, 1917, 850, 1834, 931, 2180]. payment [2061]. payoffs [1260]. PC [1747, 1707]. PC-ARIMA [1747]. PCA [2262, 2321, 1458, 1383, 2130, 657, 1973]. peaked [2205]. peakedness [396]. peaks [239]. Pearson [939, 2235, 597]. pedigree [97]. peg [1845]. Peña [1135]. Penalized [498, 2319, 1375, 722, 1092, 2155, 1768, 2191, 2160, 1225, 2015]. penalizing [808]. penalties [1988]. penalty [1193, 1494]. penetrance [2047]. perceived [2243]. percentage [601]. percentile [1673, 127]. perceptron [473]. perceptrons [552, 1203]. Perfect [1058]. Performance [1232, 1607, 2003, 1677, 1712, 2110, 1310, 198, 163, 2186, 473, 262, 475, 675, 587, 380, 1661, 1783, 1187, 667, 63, 1201]. Performing [897]. Period [1472, 1638]. periodic [1928, 1074, 1688, 1829]. periodogram [2109, 947]. periodogram-based [947]. permanent [1832]. Permutation [812, 465, 2287, 1003, 1398, 95, 2318, 1142, 915]. perspective [1196, 2315, 532]. pertains [1682]. perturbation [2117, 1458, 1807, 954, 2043]. Perturbations [1097, 701, 786]. perturbed [1164]. PH [886]. pharmacodynamic [2285]. pharmacokinetic [2285]. pharmacokinetic/pharmacodynamic [2285]. Phase [171, 2027, 2320, 1194]. phase-type [2320]. phenotype [858]. Phi [896]. Phi-divergence [896]. Phillips [739]. photoidentification [1505]. phylogenetic [7, 1997]. Piecewise [556, 614, 746]. plane [1576]. Planning [1317]. plans [1311, 1959, 973, 481]. plasmodes [2091]. plate [1942]. players [935]. plot [387, 433, 1781, 1769, 521, 550, 1076]. plots [909, 1945, 560, 1920]. PLS [654, 1785, 1373, 656, 1279, 657, 1707, 660, 718, 662, 663, 1314, 1454]. PLS1 [1108]. PLSR [655, 659]. Plug [1451]. Plug-in [1451]. plus [1917, 295]. pneumococcal [1892]. Point [1191, 1811, 1705, 980, 1304, 854, 711, 2200, 155, 356, 161, 1967, 292, 1376, 2182, 66, 1229, 1604, 1469, 1711, 1904, 84, 1079, 2347, 1207, 102, 111, 1239, 1133, 2021, 1072, 1917, 2180, 2325, 1185, 1058]. point-level [1711]. points [1146, 1989, 240, 601, 1759, 1468, 286, 2180]. Poisson [353, 946, 1407, 405, 2316, 765, 335, 74, 576, 2321, 1478, 1823, 452, 1889, 250, 2012, 800, 75, 2017, 1236, 68, 1960, 513, 184, 2241, 1339, 1325]. Poisson-distributed [405]. Poisson-shifted [1960]. Polarity [96]. Polarity-free [96]. policy [745]. pollution [1311, 1330, 320]. Polya [1082]. polyhazard [174]. POLYMARS [355]. Polynomial [1941, 1719, 32, 1676, 290, 700, 575, 703, 84, 1324, 1757, 93, 1384, 1507]. polynomials [1559, 1765, 995]. polytomous [444]. polytope [6]. Pooled [1910]. POP [1381]. POP-based [1381]. popular [1060]. Population [2285, 2296, 1433, 2303, 2298, 2095, 2092, 1402, 1224, 1240, 1961, 532, 536, 2139, 1572, 1716, 1881]. population-based [2298, 1572]. populations [1717, 944, 1872, 2316, 1090, 574, 2249, 1677, 585, 87, 1357]. portable [1466]. portfolio [1764, 1176]. portfolios [620]. portmanteau [675, 1135]. Positioning [306]. positive [1993, 608, 990, 2115, 2092, 2022]. positively [566]. positives [1617]. possibilistic [202, 1027]. Possibility [1015]. possible [634, 1579, 992]. possibly [784, 827]. Posterior [570, 666, 813, 1073, 1082, 2017, 1730]. posteriori [836, 2285]. potency [1349]. potential [2264]. Poussin [1040]. Power [1102, 2185, 2036, 1130, 2204, 391, 993, 694, 2002, 642, 476, 274, 1360, 1641, 882, 125, 284, 1288, 357, 565, 1594, 978, 1490, 979, 2252, 1839, 584, 1757, 2097, 2339, 1715, 1325]. power-divergences [979]. power-transformation [1288]. power-transformed [2252]. powerful [778]. powers [1737, 1812]. pp [1361]. PPLS [660]. PPP [1307]. PQL [1220]. Practical [1017, 499, 1188, 758, 1715]. practice [2349, 443, 1361]. pragmatic [1324, 786]. precedence [1212, 2027]. precision [2158]. preconditioning [1707]. predictability [882, 1741]. Predicting [1621, 998]. Prediction [468, 1238, 653, 1638, 639, 1271, 684, 65, 2323, 82, 2057, 474, 1912, 578, 1791, 615, 926, 1551, 1625, 1790, 1259, 2282, 1929, 419, 1636, 2075]. predictions [808]. Predictive [1783, 1325, 2108, 1054, 698, 39, 2186]. predictor [2165, 269, 701, 1605]. predictors [2108, 871, 1720, 1204, 180, 2138, 1677, 1715]. Preface [300, 195, 423]. preference [1066, 2215]. preferences [766, 2214]. prefix [1952]. Preliminary [896, 1324]. Prescott [1591]. presence [2125, 168, 2189, 1237, 1734, 1052, 174, 386, 885, 1995, 1439, 982, 2227]. preserving [2313]. prevalence [485]. prevention [556]. price [845, 649]. pricing [731, 2113, 618, 747, 1170]. primary [1449]. Principal [516, 1379, 1644, 2278, 792, 906, 930, 514, 611, 98, 878, 149, 880, 1612, 1558, 1679, 1034, 1273, 276, 1972, 48, 928, 772, 2205, 1315, 445, 1653, 2212, 1357, 1014]. prior [1333, 1193, 240, 2101, 2224, 2247]. priori [3, 30, 31]. priors [1904, 1746, 1098]. Probabilistic [1803, 1017, 2262, 864, 773, 1025, 2049]. probabilities [666, 813, 716, 896, 2095, 1362, 2031, 2011, 52, 1779, 263, 643, 888, 1505]. Probability [186, 1019, 668, 1153, 188, 71, 10, 1929]. probit [2310, 1771, 1492, 2053, 377, 53, 480, 1786]. problem [1276, 791, 1195, 1789, 656, 555, 1057, 1837, 2077, 327, 801, 479, 546, 468, 337, 1398, 84, 410, 1441, 1550, 1858]. Problems [1111, 1516, 491, 2200, 991, 1654, 908, 311, 43, 524, 1375, 1518, 308, 785, 1994, 799, 1490, 1600, 1821, 1597, 610, 625, 1660]. PROC [1006]. procedure [1650, 1738, 763, 160, 533, 2214, 1040, 97, 558, 2335, 1438, 2104, 920, 377, 95, 1874, 216, 1338, 1225, 270, 902]. procedures [1297, 181, 810, 1532, 2315, 492, 1535, 812, 67, 579, 291, 787, 1851, 754, 670, 1773, 633, 1862, 1779, 782, 1836, 1617, 87, 2190, 1821, 128, 986]. Process [1478, 2200, 946, 639, 1797, 2186, 1037, 1308, 2182, 245, 1417, 456, 833, 2164, 1749, 147, 1783, 718, 662, 2087, 184, 1636, 1137, 364]. processes [711, 630, 1370, 1146, 2321, 1270, 1266, 1613, 1937, 1782, 1293, 1154, 624, 1990, 1329, 1866, 1588, 177, 2304, 1147, 1207, 102, 825, 111, 1239, 1817, 1829, 2142, 1593, 1339, 1325, 1058]. processing [1532, 1746]. processors [162]. Procrustes [655, 797, 2339]. produce [1854]. product [466, 240, 668]. productivity [2117]. products [719, 994]. proficiency [2059]. profile [356, 530, 1137, 1076]. profile-based [530]. profiles [1965, 1492, 2282, 2060]. profiling [2155, 2080]. prognosis [2032]. prognostic [535, 337]. program [1824, 1096]. programming [645, 761, 1231, 1040, 1949, 2011, 459]. programs [995]. progressive [2152, 1959, 458, 2249, 2258]. progressively [871, 2251, 937, 235, 2299]. Project [1790]. Projected [804]. Projection [1696, 1553, 880, 1368, 543, 795, 1807, 615, 101]. proofs [287, 314, 321, 393, 402, 412, 421]. propagation [160, 1797]. propensity [1624, 1475, 408]. Properties [1415, 2233, 1720, 296, 795, 1233, 1393, 1490, 700, 343, 1798, 2252]. proportion [1831, 633, 254, 1617]. proportional [1297, 160, 2019, 2065, 1828, 1986, 362, 1941]. proportionality [628]. proportions [506, 465, 398, 1060, 528, 1082, 2291, 254, 511, 1894, 1450, 1788, 2029]. proposal [286]. proposals [340]. proposed [55]. prostate [1092]. protein [1748]. proteomic [1532]. proteomics [1835]. proximity [891, 907, 519]. prune [2042]. pruning [198, 836, 2024]. PS [287, 314, 321, 366, 383, 393, 402, 412, 421]. PS-file [287, 314, 321, 366, 383, 393, 402, 412, 421]. Pseudo [452, 345, 222, 1449, 180, 1127, 1552]. pseudo-likelihood [345, 1449]. pseudo-logistic [1552]. pseudo-maximum [222]. pseudo-predictors [180]. pseudo-random [1127]. psychiatric [1923]. public [2063]. Publisher [464]. purchase [845]. Pursuit [375, 1553, 880, 1807, 101].

QTL [1285, 2099]. quadratic [2046, 1764, 301, 604, 1598, 2007, 1153, 371, 1622, 1287, 2149]. quadrature [510]. qualitative [529, 858]. quality [853, 279, 1293, 2063, 2087, 649, 2171]. quality-adjusted [2171]. quantifying [1984, 786]. Quantile [1392, 399, 1194, 2261, 265, 877, 593, 1523, 1903, 170, 1002, 1078, 1614, 237, 536, 2014, 914, 1815, 521, 843]. quantile-based [237]. quantiles [2112, 1637, 239, 242, 201, 1511, 396, 2229, 2346, 1918]. quantitative [1497, 1753, 2098, 2082, 1970]. quantities [1961]. quantized [971]. quarterly [400]. quartile [1576, 965]. Quasi [695, 1448, 1917, 498, 2023, 2191]. quasi-independence [2023]. quasi-likelihood [498, 2191]. Quasi-Monte [695, 1448]. Quasi-plus [1917]. Quest [457]. questionnaires [1306]. queueing [1638]. quick [1438]. Quintile [1475]. quotes [1172].

R [2345, 2066, 1405, 1569, 2288, 1986, 1424, 995, 1318]. radar [1372]. radial [1501]. radiation [1376]. rainfall [1108]. Ramage [623]. Rand [998]. Random [1016, 1358, 1297, 634, 1479, 1841, 1681, 2188, 2187, 1403, 1232, 2226, 1804, 907, 2338, 12, 1360, 1721, 1368, 1888, 1460, 684, 65, 2323, 1491, 229, 1487, 1139, 137, 1984, 466, 1022, 168, 1689, 1000, 1619, 1584, 481, 1577, 844, 1701, 555, 1583, 1768, 1446, 1215, 2248, 542, 28, 1840, 1202, 1018, 566, 1853, 52, 599, 572, 994, 866, 1026, 153, 1629, 582, 1757, 1700, 2000, 2279, 2014, 1007, 1496, 1227, 1072, 397, 1512, 1929, 1127, 809, 1466, 324, 1639]. random-effects [1841, 542]. randomisation [171]. randomised [1220, 1990]. Randomized [558, 1982, 420, 2166, 1291, 1831, 1814, 1944, 2067, 2174, 573, 126, 1429, 1318, 985, 1471]. randomly [1507]. randomness [675]. Range [1630, 271, 1735, 1759, 1083]. ranges [539]. Rank [34, 252, 1385, 1868, 387, 62, 227, 1066, 1623, 1533, 1144, 2121, 261, 185, 1732, 564, 404, 151, 1508, 584, 832, 1216]. rank-based [564]. Rank-one [34]. rank-two [34]. ranked [22, 1668, 2340, 1109, 1084, 216]. ranking [1109, 347, 1560]. rankings [669]. ranks [802]. Rao [1232]. Raphson [97]. rare [1561]. Rasch [2025]. rate [863, 2329, 2324, 1478, 1733, 1416, 2267, 988, 2092, 633, 379, 1427, 2021, 2040, 1694, 513, 1801, 1041, 2311]. rates [1993, 1890, 1220, 2128, 1377, 2343, 358, 993, 1773, 68, 549, 767, 1390]. rating [1263, 1522, 2215]. ratio [2074, 1654, 588, 1235, 2065, 2260, 1346, 23, 1286, 1831, 2174, 866, 601, 1628, 68, 1894, 1946, 1802, 183, 753, 902, 1355, 2060, 88]. ratios [2046, 716, 719, 994, 371, 986]. Raton [1766]. Rayleigh [723, 1699]. re [2207]. re-ordered [2207]. reaction [745]. reactor [2326]. Read [979]. reading [431]. real [1797]. real-time [1797]. realized [1677, 2141]. reappraisal [2211]. reasoning [283, 1015]. recalculation [705]. recapture [1402, 2172, 2286, 641]. Reciprocal [1071]. reconsideration [2119]. reconstruction [7, 1444, 1072]. record [1770, 2150, 1158, 1625, 2257, 1339]. records [432]. rectangular [1884]. Recurrence [2295, 2006]. recurrences [257]. recurrent [1989, 1889, 1784, 1466, 1148]. Recursive [440, 311, 2121, 1577, 47, 251, 831, 1156, 830, 809]. reduced [261, 1732, 832, 1692]. reduces [634]. Reducing [408, 1144]. Reduction [1546, 653, 932, 106, 1048, 1558, 1310, 305, 1335, 607, 2165, 2301, 1935, 568, 1314, 1557, 1796, 1956]. redundancy [1785, 758, 1548]. referees [806]. reference [16, 548]. reflective [1279, 1454]. Regarding [2345]. regenerative [1718]. regenerative-block [1718]. Regime [1167, 1686]. Regime-switching [1167]. region [1261]. Regional [909, 1587]. regions [39, 734, 544, 231, 2151, 2003, 1207, 381, 1085]. registration [2054]. Regression [388, 765, 1813, 224, 542, 659, 1026, 906, 132, 527, 2242, 932, 2175, 1719, 1340, 2109, 45, 1502, 876, 704, 654, 1932, 971, 2159, 628, 2300, 1722, 199, 1785, 852, 1720, 1592, 214, 1369, 1810, 1672, 1811, 791, 32, 764, 1370, 1483, 145, 1905, 1676, 817, 1196, 1958, 1028, 453, 752, 73, 354, 1029, 1965, 1136, 909, 82, 712, 543, 1373, 1174, 1838, 576, 89, 1571, 735, 312, 538, 717, 382, 548, 1579, 473, 823, 1452, 2298, 1500, 1285, 975, 1695, 736, 950, 1253, 706, 1823, 1923, 1027, 884, 1288, 1096, 1375, 261, 2335, 1335, 185, 1273]. regression [1732, 452, 92, 1518, 1833, 2166, 390, 151, 911, 1889, 1486, 441, 2312, 671, 1397, 1517, 1974, 756, 826, 1129, 1713, 1065, 785, 1049, 1707, 40, 70, 904, 676, 1160, 860, 1443, 2330, 1983, 992, 697, 1771, 2306, 1097, 627, 1600, 1536, 2291, 380, 2308, 75, 1300, 494, 1614, 979, 269, 86, 1380, 1630, 1798, 1783, 154, 236, 703, 759, 1238, 361, 1079, 2347, 1277, 560, 1961, 1806, 1986, 61, 718, 662, 399, 1662, 2263, 2271, 416, 995, 1757, 116, 1551, 984, 2278, 956, 1603, 1659, 69, 701, 520, 2014, 1108, 8, 1795, 2333, 1549, 1815, 49, 901, 1952]. regression [1296, 530, 1941, 77, 2172, 1775, 1008, 832, 959, 1470, 1572, 870, 103, 1605, 119, 313, 664, 2261, 2207, 1384, 1465, 128, 418, 363, 63, 1356, 1715, 2327, 2241, 324, 1796, 1956, 265, 1507, 1552, 1068, 874, 1198, 1918, 1697, 985, 1471, 818]. regression-based [1198]. regressions [1523, 439, 1405, 1002]. regularisation [1437]. Regularization [200, 518, 881, 1095]. Regularized [1548, 2209, 1918, 1598, 1551, 1775]. reintroduction [485]. reinvestigation [1891]. Reiss [638]. rejection [558, 1765, 1822, 381]. rejections [2084]. related [2203, 2336, 1609, 2097, 1876]. relating [2195]. relation [1602]. relational [1444]. relations [2295, 2006]. relationship [445, 2180]. relationships [1675, 2348, 1036, 1454, 1507]. Relative [136, 907, 537, 218, 227, 1349, 600, 477, 738, 899, 67, 1506]. Relaxed [1547]. release [358]. Relevance [785]. relevant [618, 1561]. Reliability [1767, 1252, 973, 976, 2065, 1290, 1808, 1652]. reliable [190]. remedies [634]. removal [1919]. renewal [889]. Repeatability [2345]. Repeated [927, 2267, 1986, 1963, 647, 266, 1118, 1771, 28, 837, 1291, 1945, 156, 869]. replacement [1575, 263]. replicated [186, 2071]. replication [60, 1159]. replications [1976, 564]. Reply [1214]. representation [2214, 1022]. representations [1017]. Representing [1525]. Reproducibility [1925, 2345]. reproductive [2259]. required [1983]. Resampling [89, 1084, 418, 1903, 1383, 1329]. research [760, 2146, 1980, 1246, 2156, 1930]. Residual [1496, 1810, 1370, 909, 952, 212, 1296]. Residual-based [1496]. residuals [2235, 1743, 2121, 1646, 1454]. resistance [1791]. Resistant [38, 86, 547, 759, 589]. resonance [824]. respect [1612, 661]. Responder [872]. Response [31, 782, 959, 1956, 388, 1028, 125, 761, 1943, 529, 155, 168, 1087, 2166, 138, 249, 232, 1898, 2037, 542, 2306, 467, 2169, 177, 701, 769, 1525, 744, 1653, 24, 277, 1445, 178, 2202, 985, 1471]. responses [1288, 845, 1791, 1417, 1320, 1118, 23, 1056, 1771, 1107, 49, 1461, 1913]. restaurant [169]. Restricted [752, 1603, 1606, 716, 574, 1353, 493, 2158, 1432, 1869]. restriction [2291, 374]. restrictions [2272]. result [1746]. results [1408, 771, 164, 2135, 2059, 1839]. RETRACTED [1675]. Retraction [2348]. retrieval [878]. return [882, 177]. returns [1173, 1233, 1739, 505, 926, 2252, 1103, 2141]. Reversible [2013, 317, 286]. Review [1361, 1766, 2216, 340, 2220, 2217, 2297, 116, 2219, 2218, 2075]. revisit [1441]. revisited [1344, 772, 445]. Reweighted [1234, 491, 1440]. RGBland [2223]. Ricker [358]. ridge [261, 955, 832, 1470]. Riemann [1358]. right [593, 1245, 2328, 1507]. right-censored [1245]. Risk [2003, 776, 600, 1750, 1143, 2124, 2232, 578, 1235, 1175, 727, 816, 570, 1286, 1814, 1944, 2067, 618, 1429, 42, 1985, 585, 1262, 2190, 1742, 346, 888, 1176]. risks [1100, 2299, 1200]. RJMCMC [265]. RLadyBug [1569]. RMM [782]. road [1100]. Robust [1456, 1406, 458, 1396, 2300, 220, 1722, 838, 2262, 764, 149, 1062, 1544, 1048, 1527, 1830, 46, 1205, 975, 1272, 2130, 1377, 1193, 1247, 842, 952, 1536, 1540, 1867, 293, 1902, 1007, 1597, 1382, 1334, 108, 1061, 763, 45, 1497, 1534, 997, 1744, 18, 2074, 557, 1595, 1336, 2028, 736, 724, 2129, 2335, 502, 475, 2085, 1535, 799, 2250, 241, 1600, 2076, 1891, 1950, 636, 1700, 2050, 963, 1485, 696, 418]. Robustified [1760]. robustly [94]. Robustness [357, 1394, 1734, 700, 602, 416, 1698, 1390, 1393, 162, 351, 757, 212, 1324, 2238, 786, 563]. ROC [1993, 1709, 1763, 1946, 2325, 902]. Rocke [30]. Rodriguez [1135]. role [1740, 1110, 2280]. root [1307, 420, 951, 1938, 966, 69, 313, 1067]. roots [2131, 2050]. rose [229]. Rosenblatt [1352]. rotatability [232]. Rotated [16]. rotation [1826, 664]. roughness [1193]. rounding [2061]. roundtrip [486]. rows [2336, 659]. Rukhin [750]. rule [853, 1864, 180]. rules [113, 38, 2027, 1754, 1954, 279, 772, 1561, 1698, 745]. run [1762, 262, 188, 282]. runs [112, 2027]. runs-type [2027]. RV [1933]. Ryzin [1502].

S [1213, 44, 295]. S-plus [295]. S-system [44]. Saddlepoint [1222, 1990]. SAEM [2053, 1122]. safety [1100]. same [2013]. Sample [951, 1983, 1291, 561, 2040, 1355, 1499, 1963, 2072, 2109, 1696, 1326, 2113, 411, 2287, 106, 1433, 733, 1654, 1641, 239, 1767, 1789, 1255, 1021, 558, 1938, 280, 1521, 1463, 1254, 1994, 675, 2277, 1837, 1197, 2225, 1935, 571, 592, 479, 546, 2231, 754, 1661, 725, 1882, 1995, 848, 1839, 536, 1511, 1450, 1506, 1228, 471, 1550, 1858, 1716, 183, 1792, 705, 1389, 2293]. sample-based [1935]. sample-size [1767]. sampled [278]. sampler [1823, 833, 1730, 869]. samplers [1649]. samples [871, 1886, 214, 1433, 1090, 2095, 1084, 980, 1304, 2039, 1891, 293]. Sampling [1906, 1221, 1575, 22, 1668, 1311, 1340, 2340, 2111, 226, 991, 1109, 698, 1883, 1797, 2303, 1267, 1856, 150, 477, 481, 695, 2031, 1521, 846, 1951, 67, 2167, 1402, 1591, 1282, 1536, 1765, 754, 1435, 102, 216, 1829, 1511, 1194, 71, 1227, 1985, 2228, 1917, 1203, 2258]. sampling-based [2228]. SAS [578, 1096, 1006, 699, 995, 1318]. SAS/IML [1096]. Saunders [1304, 2066, 1885, 1714, 1443, 1669, 1354, 2330, 414, 980, 855, 646, 1356]. Scal [1033]. scalability [1763]. scalar [1126]. scale [1233, 1518, 1397, 35, 1891, 1185]. scaled [1491]. scales [1336, 319]. scaling [1531, 12, 2019, 1033, 1032, 1077, 798, 307, 1858, 2339]. scan [1327, 2183, 1520, 102, 2094, 2184]. scatter [1391]. SCD [2111]. scenario [846]. scenarios [886]. scheme [683, 2043, 1454]. schemes [1281, 1045, 1282, 2030]. Schindler [386]. schizophrenia [2156]. Schmidt [311]. Schmidt-based [311]. science [4]. sciences [1609]. Score [250, 2241, 214, 1624, 1190, 1475, 408, 978, 2044, 1986]. scores [1306]. scoring [1130, 1287]. SCoTLASS [804]. scree [1076]. screening [1999, 228, 509, 2173, 1305, 270]. Search [1396, 1130, 100, 615, 616, 1572]. seasonal [420, 951, 825]. seasonally [733]. SEATS [920]. Second [730, 191, 232]. section [104, 114, 122, 131, 139, 148, 158, 165, 172, 179, 187, 287, 314, 321, 194, 209, 217, 225, 234, 243, 258, 272, 297, 348, 366, 383, 393, 402, 412, 421, 1850]. sectional [1307, 1000, 1938, 2050]. sector [1741]. Secure [1452]. Seeking [432, 18]. Seemingly [312, 439]. segmentation [1436, 1463, 1871, 2101]. segmented [1500]. Segmenting [824, 100]. segregation [907, 2179]. SELDI [1835]. SELDI-TOF [1835]. select [2093, 1501, 521]. selected [1488, 404]. Selecting [117, 2089, 2223, 686]. Selection [1519, 1969, 1940, 970, 1497, 123, 1932, 945, 1722, 820, 611, 1241, 136, 2253, 2310, 2062, 2048, 499, 882, 2319, 1695, 2315, 1782, 1348, 1543, 1069, 1438, 1517, 1535, 785, 2288, 676, 359, 1056, 1350, 992, 542, 1703, 1156, 2169, 2160, 2280, 1536, 1413, 1972, 101, 599, 376, 86, 638, 1658, 1115, 1787, 127, 1480, 1451, 1835, 1902, 1551, 1665, 512, 520, 1549, 1769, 2088, 281, 513, 157, 901, 1554, 1772, 649, 1283, 1605, 1844, 2161, 13, 418, 968, 1076, 1918]. selective [369]. selectively [278]. Self [1661, 914, 182, 1463, 376, 933, 1860]. self-adjustment [1860]. Self-controlled [1661]. Self-organizing [914, 182, 1463]. self-selection [376]. self-similarity [933]. Semantic [1803, 1529]. Semi [1672, 1723, 874, 1330, 1168, 1183, 1990, 1474, 1427, 1809, 459, 1068]. semi-continuous [1809]. semi-infinite [459]. semi-Markov [1168, 751, 1183, 1990]. Semi-parametric [1672, 1723, 874, 1330, 1474, 1427, 1068]. semicontinuous [1991]. SEMIFAR [271]. Semiparametric [1164, 2115, 1483, 43, 2298, 2166, 583, 1865, 2328, 2279, 1148, 641, 1406, 1696, 271, 1722, 1419, 820, 2098, 1218, 1449, 2309, 1936, 886, 334, 1589, 1509, 744, 256, 522, 902]. semiparametrically [591]. Semisupervised [1859]. sensible [457]. sensitive [724, 1088]. sensitivities [1930]. Sensitivity [1810, 798, 1097, 269, 701, 1431, 844, 1052, 147, 2101, 1070, 530, 2163, 1434]. Separability [960]. Separation [829, 1430, 416]. Sequence [1816, 1183, 1043, 1444, 2101, 2228, 1970]. sequences [11, 2322, 166, 1501, 2100]. Sequential [1729, 708, 1345, 2255, 1984, 1649, 2037, 636, 1088, 1289, 270, 749]. sequentially [69]. serial [736, 1253, 1839]. serially [441, 1247]. Seriation [2206, 524]. Series [193, 1361, 1868, 943, 1064, 1650, 1164, 141, 328, 1023, 167, 365, 639, 1168, 947, 577, 1407, 2292, 732, 1271, 929, 2036, 1655, 1612, 1586, 39, 355, 1830, 2052, 2058, 5, 557, 26, 43, 420, 1595, 1533, 1294, 1695, 923, 884, 999, 1833, 1178, 233, 1813, 1074, 675, 899, 134, 169, 253, 1585, 1186, 1646, 966, 468, 1661, 142, 154, 2137, 1134, 1457, 1759, 1862, 619, 1265, 1839, 1114, 1480, 2337, 2147, 1104, 417, 57, 1179, 1834, 1970, 364]. Serology [2342]. servers [2063]. services [1444]. Set [2345, 1191, 22, 1668, 2340, 577, 1109, 107, 1925, 1556, 117, 1053, 727, 1536, 1084, 1561, 216, 1894, 79, 1108]. sets [545, 98, 1144, 1117, 1689, 1025, 1053, 233, 657, 660, 1016, 1277, 454, 1418, 1414]. Setting [308, 1327, 478]. several [1955, 2316, 528, 95, 2227, 1318]. sexual [1930]. Shannon [881]. shape [124, 2287, 533, 897]. shaped [1298, 500, 1520]. shapes [1116, 241].

Shapiro [1397, 2281]. shared [284, 256, 522, 819]. sharing [1907]. Sharp [2149]. sheep [1327]. Shewhart [1971]. Shewhart-CUSUM [1971]. shift [1533]. shifted [1960]. shocks [1322]. short [1638, 1330, 233, 1759, 705]. short- [1638]. short-term [1330, 705]. shortest [2323]. shortly [2243]. show [455]. shrink [2079]. Shrinkage [2154, 682, 2048, 266, 1783, 703, 1549, 959, 1098]. shrinking [1539]. sib [2097, 857]. sided [1967, 135, 156, 270]. SIDS [335]. Sieve [1762, 1294, 2334]. sieves [2264]. sigmoid [1098]. Sign [799, 1351, 2050]. Signal [935, 1580, 1589, 266, 43, 922, 177, 925, 918, 178]. signal-response [178]. signaling [2027]. signals [672, 121, 1152, 928, 1746]. signed [1066, 1802]. Significance [545, 1825, 1776, 1933, 95, 2318, 42, 1501, 1622]. significant [1999, 672, 79, 958, 2090]. SIMEX [1481]. similar [2010]. similarity [2041, 1679, 933, 1996, 2211]. Simple [1268, 396, 2020, 2090, 527, 1785, 688, 1754, 1139, 91, 1254, 2173, 697, 1882, 759, 61, 1560, 1870, 606, 2043, 2091, 1216]. simplex [1949]. simplicial [1384]. simply [670, 1388, 1904]. Simulated [2140, 159, 222, 500, 935, 2022, 1897, 747]. Simulating [1425, 987, 597, 939, 1142]. Simulation [2110, 1345, 1513, 1099, 743, 2142, 1633, 1744, 1720, 400, 382, 1613, 164, 1649, 1969, 27, 359, 1100, 568, 2081, 637, 584, 1677, 1685, 1058, 2206]. Simulation-based [1345, 1513, 1099, 743, 1744, 1613]. simulations [476, 1206, 2331, 893, 2091]. Simultaneous [1932, 2062, 2054, 1673, 2030, 2210, 113, 398, 1254, 938, 800, 2040, 143, 1918]. simultaneously [2141]. Singapore [1270]. single [1823, 1160, 644, 2205, 1920, 1941, 2161, 1627, 874]. single-index [1160, 644, 1920, 1941, 1627, 874]. single-peaked [2205]. singular [2233, 1529, 1875, 792]. SiNos [1467]. siRNA [1501]. site [1823, 1790]. situation [1857]. Size [1757, 2287, 1767, 312, 951, 1463, 1197, 1291, 1379, 1490, 725, 1224, 1882, 1995, 2040, 1228, 1550, 708, 2030, 705, 1390, 2293]. size-classified [1390]. SiZer [1467, 1806]. Sizes [898, 1060, 1521, 1994, 1983, 571, 1402, 1503, 561, 1506]. skew [1899, 1965, 309, 509, 2295, 2006, 1893, 2327]. skew- [2295, 2006]. skew-elliptical [1965, 1893]. skew-normal [2327]. skewed [1908, 2130, 60]. skewness [220]. skin [2195]. skipped [450]. sky [839]. Sliced [548, 1659, 1551]. slope [808, 232, 2308]. slope-rotatability [232]. slow [455]. SM [385]. Small [733, 592, 1661, 1450, 183, 1499, 2146, 1886, 2287, 214, 106, 1641, 1220, 1351, 1204, 2189, 2196, 1994, 1197, 2073, 2039, 293, 1550, 24, 277, 2254]. small-area [2196]. Small-sample [733, 1661, 1450, 1499, 106]. smallest [1716]. Smirnov [51, 558]. SML [376]. Smooth [1395, 2192, 1067, 1697]. Smooth-CAR [2192]. Smoothed [1905, 127, 1296, 1968, 2153]. Smoothing [1369, 1863, 2277, 359, 1932, 385, 2155, 1728, 794, 1942, 2095, 319, 1346, 722, 1489, 627, 2234, 1464, 2307, 2026, 2254, 563, 1590, 2016, 843, 1940]. smoothness [275]. SMV [467]. SMV-optimum [467]. Snaer [1824]. Social [1675, 2348, 1487]. socioeconomic [2198]. soft [1023]. Software [1564, 1051, 1096, 771, 727, 1290, 1565, 1828, 1206, 1986, 699, 1808, 2116]. soil [1311]. solution [2117, 793, 804, 2207]. solutions [1375, 66, 1063]. solvers [471]. Solving [2116, 439, 717, 308, 1660]. Some [531, 1745, 85, 3, 30, 1190, 1107, 2291, 2135, 1798, 1839, 31, 536, 87, 871, 1922, 688, 2329, 1211, 1767, 1571, 340, 976, 1812, 555, 1713, 675, 240, 1490, 142, 1317, 2171, 448]. sorting [279]. source [1430]. sources [655, 1157]. sovereign [1110, 1263]. Space [320, 382, 2181, 945, 1430, 1048, 684, 150, 923, 1183, 1482, 1348, 1608, 1133, 1731, 2302, 1686]. Space-varying [382, 1482]. spaces [1606, 1707, 115]. Spanish [1103]. Sparse [1844, 1824, 1129, 2033, 580, 469]. Spatial [1311, 2200, 2194, 2178, 2326, 2282, 2199, 2184, 1105, 2332, 854, 2188, 2201, 907, 1244, 2198, 2183, 2214, 1298, 500, 1520, 1270, 1823, 2189, 2095, 2195, 2182, 865, 2196, 846, 1469, 2192, 1280, 495, 1277, 2318, 102, 111, 93, 2193, 1496, 1790, 1651, 1917, 783, 931, 2197, 2202, 1339]. spatial-temporal [111]. spatially [711, 2187, 817, 1196, 1244, 498, 1193, 236, 1203, 2199]. spatio [432]. spatio-temporal [432]. Spatiotemporal [2195]. SPD [1177]. Spearman [62, 227]. Special [1163, 1726, 790, 730, 1251, 2203, 1564, 1516, 1850, 438, 1580, 2107, 1379, 918, 331]. species [539, 327]. specific [2175, 365, 343, 1992]. specification [1723, 1144, 1431, 1193, 898, 1671, 380, 1496]. specifications [382, 240]. specificities [1930]. specificity [1052]. specified [1139, 528, 566, 582, 1317]. speckled [296]. Spectral [1942, 1707, 796, 292, 1329, 1759, 2268, 715]. spectrometry [1532]. spectrum [1594, 1444, 2326]. speed [1436, 455]. spillovers [1736]. Spline [2254, 1919, 2155, 1988, 1778, 82, 590, 722, 1329, 1702, 1794, 1941, 2016, 2015]. Splines [1953, 1092, 365, 852, 1369, 1452, 1320, 1193, 1489, 359, 860, 542, 317, 2160, 494, 1138, 2247, 996, 563, 843]. Split [389, 387, 512, 1554]. split-plot [387]. spreadsheet [471]. spreadsheets [1855]. Springer [1361]. square [411, 702, 2023, 246, 2007, 267, 1882, 1779]. squared [470, 1299, 1488, 1204, 452]. Squares [1596, 132, 1502, 876, 1028, 1721, 73, 1136, 543, 1843, 671, 308, 1043, 1598, 697, 658, 801, 568, 410, 1599, 746, 664, 1440, 850, 625]. SSN [104, 114, 122, 131, 139, 148, 158, 165, 172, 179, 187, 287, 314, 321, 194, 209, 217, 225, 234, 243, 258, 272, 297, 348, 366, 383, 393, 402, 412, 421]. SSN-section [104, 114, 122, 131, 139, 148, 158, 165, 172, 179, 187, 287, 314, 321, 194, 209, 217, 225, 234, 243, 258, 272, 297, 348, 366, 383, 393, 402, 412, 421]. Stability [1106, 853, 2041, 1537, 739]. Stable [632, 310, 1202, 2134, 2133, 905, 1631, 745]. Stacked [1701]. staff [1444]. stage [1092, 1001, 2306, 575, 981, 1677, 2080, 1041, 36]. Stan [438]. Standard [2000, 840, 46, 771, 528, 569, 2164, 1618, 2116]. Standardising [1864]. standardized [1773]. stars [1472]. starting [341]. Stat [2345, 767]. STATA [995]. State [1608, 945, 150, 923, 1183, 1688, 1731, 2302, 1686]. state-dependent [1688]. state-space [945, 150, 923]. States [2195]. stationary [2232, 1348, 1154, 253, 1585, 928]. STATIS [1321, 551]. STATIS-ACT [1321, 551]. Statist [2344, 120, 826, 2067, 939, 2347, 2348, 2346, 277]. statistic [998, 1758, 51, 558, 19, 1331, 534, 2281, 1387, 601, 386, 102, 1769, 1296, 2094, 1716, 2256, 1802]. Statistical [2070, 1726, 1546, 677, 810, 630, 640, 2321, 1564, 976, 1235, 2196, 727, 2236, 487, 401, 2059, 1580, 769, 1461, 1159, 1072, 552, 1636, 259, 461, 392, 18, 190, 533, 1013, 1211, 146, 733, 283, 1015, 728, 200, 1737, 1567, 1733, 1376, 456, 1290, 679, 27, 1354, 1490, 291, 787, 1851, 1828, 1206, 1986, 918, 958, 1243, 1687, 1020, 4, 2313, 184, 1852, 2223]. Statistically [2087]. Statistics [20, 693, 1303, 30, 1361, 1366, 1304, 827, 1301, 2070, 2340, 790, 1574, 1488, 1066, 2027, 193, 1520, 1980, 2237, 404, 2295, 2006, 1900, 2178, 1508, 799, 993, 360, 750, 1865, 1855, 1658, 460, 1158, 9, 2180, 1339, 442, 2184, 1766]. statistics-linear [1574]. status [1828]. stay [324]. Step [1617, 2051, 1720, 2117, 262, 373, 2168, 1686]. step-down [1617]. step-stress [2051, 2168]. Step-up [1617]. Stepwise [553, 1787, 1535]. stereotype [904]. Stochastic [1745, 761, 1231, 196, 40, 257, 613, 1526, 2162, 378, 1419, 946, 2253, 1345, 1728, 1146, 2321, 2126, 949, 1733, 1569, 1169, 2133, 615, 2136, 1730, 2252, 718, 662, 893, 111, 532, 924, 2140, 2141, 1645, 1542, 2197, 1170]. Stock [882, 1872, 1739, 1103]. stopping [198, 772]. strain [798]. strata [717, 2308]. strategies [629, 219, 1196, 228, 846, 580]. strategy [500, 474, 1279, 1703, 551]. stratification [1191, 1475, 2092, 1972]. stratified [1231, 99, 1521, 1814, 1944, 2067, 2174, 986]. stratum [1521]. Strauss [1124]. streaming [2292, 1575]. stress [2051, 2168]. strip [455]. Strong [2021]. Structural [2239, 1875, 1606, 820, 1890, 318, 1942, 2259, 1235, 444, 873, 1873, 1286, 2079, 757, 116, 442, 443, 967]. structure [1841, 2108, 2110, 1392, 2118, 2198, 432, 2128, 1705, 1600, 1457, 1332, 1748, 743]. structured [1932, 852, 118, 966, 591, 1321, 1697]. structures [1556, 208, 302, 912, 2309, 582, 454, 550]. Student [1899, 55, 2]. Student- [1899]. Studentized [1398, 1083]. studies [189, 1143, 2084, 583, 1284, 2092, 2079, 1131, 633, 1457, 496, 1228, 2094, 346, 2286, 641]. study [476, 1502, 1633, 1722, 1720, 556, 1846, 2124, 400, 439, 1263, 2289, 911, 164, 1486, 721, 1290, 1969, 1671, 675, 27, 359, 1975, 1990, 351, 336, 1100, 1236, 568, 171, 886, 1240, 216, 1509, 1977, 1677, 156, 2163, 1790, 103, 522, 1466, 2210, 1507, 2206]. studying [111]. stuttering [2343]. style [619]. Stylized [1168]. sub [117, 585, 1203]. sub-populations [585]. sub-sampling [1203]. sub-set [117]. subject [990, 985, 1471]. subjects [1635]. submodels [2019, 1579]. subpopulations [1258]. Subsampling [1666, 943, 676]. Subset [1782, 1518, 785, 513, 1465]. subspace [16, 2110]. successes [166, 830]. successive [553]. sudden [336]. sufficient [607, 2165, 2301, 1920, 1418, 1796]. Sum [1358, 1066, 1136, 410]. sum-of-squares [410]. Summary [2180, 1920]. Sums [719, 1639, 1478, 994]. supersaturated [1964, 1226, 2148, 805]. Supervised [118, 2049, 2089, 1314]. supplement [2091]. supplier [2339]. Support [1494, 2014, 1023, 1543, 616, 1706, 1435, 2089, 1815, 2088]. SURE [374]. surface [761, 168, 1087, 232, 24, 277]. surgical [1444]. surrogate [1318]. Surveillance [2181, 2233, 1833, 2322]. survey [2159]. surveys [1311, 1231, 839, 536]. Survival [2075, 2008, 1519, 1245, 1988, 2032, 535, 2298, 1984, 578, 861, 191, 2093, 722, 977, 1065, 358, 1837, 1004, 1703, 337, 725, 333, 2328, 63, 1466, 1632, 1505]. survivor [1832]. Susarla [1502]. susceptible [485]. susceptible-infected-susceptible [485]. SVD [1492]. SVD-based [1492]. swapped [1197]. swapping [1750]. swarm [867]. Swedish [2120]. Switching [1736, 1345, 1167, 1737, 2037, 1808]. Symbolic [1883, 1970, 1630]. symmetric [516, 796, 884, 1559]. symmetrical [752, 1029]. symmetrised [1391]. symmetrized [637]. symmetry [34, 19, 1316, 601, 1265, 1977]. syndrome [336, 2228]. Synthesis [1526, 797]. synthesize [1594]. System [306, 1566, 461, 1023, 44, 1824, 973, 597, 939, 460, 1652, 1907]. Systematic [1435, 626, 1591, 1829, 667]. Systemic [1939]. systems [1638, 733, 1110, 1522, 1439, 488, 1582, 1687].

T [1570]. table

[2336, 1353, 1286, 1911, 1321, 888]. tables [642, 608, 514, 465, 2179, 1743, 1884, 2170, 2019, 2023, 99, 389, 689, 1044, 1223, 462, 551, 2207, 2210]. tabu [616]. tail [838, 1, 1485]. tailed [1638, 1205, 2045, 2337, 2071]. tails [1634, 2136, 396]. taper [2326]. TARGET [1620]. task [163]. tasks [1709, 2297]. tau [1447]. Taxonomy [2297]. TCA [1583].

TDT [2102]. Teaching [1855]. technical [1159]. technique [311, 1434]. Techniques [864, 1958, 1310, 305, 1266, 1666, 2322, 1489, 1350, 1277, 460, 982, 1687, 1203]. Technological [617]. teletraffic [262, 489]. tempering [570]. template [996]. Temporal [1472, 1591, 2187, 2194, 432, 2234, 111, 550, 57, 1137]. ten [351]. tensor [1482]. term [2108, 1330, 2110, 2128, 1791, 333, 743, 705]. termination [565]. Terms [2345]. terrorist [1207]. Test [2102, 1044, 1911, 778, 274, 895, 1307, 533, 33, 1676, 2183, 603, 1654, 2052, 1352, 1679, 1789, 973, 1351, 1021, 2129, 896, 2023, 1000, 2260, 1346, 1397, 23, 2033, 1003, 1316, 978, 1135, 2225, 67, 479, 546, 52, 2044, 135, 1398, 386, 1386, 1882, 95, 694, 1628, 784, 827, 2318, 1265, 333, 1652, 1506, 2050, 1151, 857, 2026, 1093, 2276, 1083, 2153, 1715, 915, 442, 1507]. Testing [476, 2188, 608, 1066, 1955, 1165, 588, 1174, 2126, 1230, 477, 1328, 1000, 1323, 1933, 1944, 2067, 754, 153, 1509, 1103, 2337, 2227, 1977, 42, 1700, 156, 1788, 1093, 1622, 2168, 448, 443, 2008, 763, 1606, 387, 2300, 907, 85, 2255, 557, 1408, 1022, 2078, 564, 1776, 1094, 1060, 528, 2281, 580, 1490, 1657, 1699, 2157, 1155, 377, 601, 1503, 1628, 2059, 755, 2040, 325, 1501, 2238, 1295, 1067, 1203, 1465, 986, 1050, 902]. Tests [2162, 879, 950, 250, 726, 327, 2103, 2152, 2051, 129, 1868, 1841, 268, 642, 506, 1340, 1326, 1886, 2113, 1299, 1922, 545, 465, 2287, 2179, 2275, 2316, 2344, 1360, 1252, 284, 26, 229, 1723, 2127, 1533, 1762, 1205, 1255, 736, 1253, 951, 539, 1938, 1812, 897, 1486, 176, 689, 2290, 1060, 1974, 739, 977, 898, 1998, 391, 675, 2131, 252, 1780, 2038, 553, 1286, 626, 241, 1186, 2231, 2291, 409, 1089, 380, 2135, 35, 1236, 885, 1995, 115, 95, 2317, 1839, 2334, 584, 859, 1757, 396, 1187, 643, 1618, 1142, 325, 417, 57, 667, 2100, 1306, 1946, 2241, 2034, 967]. tests [753, 1355, 749]. textile [1781].

Texture [344]. their

[508, 107, 2032, 509, 1393, 2237, 569, 1560]. theoretic [2100]. Theoretical [1108, 206, 573, 1746]. Theory [1361, 112, 306, 1015, 200, 2349, 2336, 754, 381, 887, 1014, 332]. thermodynamic [2296]. thin [1942]. thin-plate [1942]. Third [1251, 742]. Third-order [742]. Thomas [638]. Three [406, 1128, 275, 942, 2251, 891, 879, 2214, 1034, 1967, 2031, 1843, 1060, 781, 586, 525, 1615, 1897, 1414, 2180]. three-level [942]. Three-mode [1128]. three-parameter [2251, 879, 2031, 1843, 781, 525, 1615, 1897]. Three-way [406, 275, 891, 2214, 1414]. Threshold [1095, 614, 1735, 1682, 1634, 2195, 542, 1305, 1821, 192, 1524, 2325]. Threshold-based [1095]. thresholds [239]. throughput [1305]. tickets [845]. ties [386]. Time [1064, 1655, 1154, 1178, 921, 1614, 928, 1868, 943, 1650, 2072, 2181, 141, 328, 1023, 167, 486, 639, 1363, 1168, 947, 2226, 577, 1407, 682, 732, 1271, 929, 406, 881, 1612, 339, 1586, 39, 355, 320, 684, 2052, 2058, 557, 1797, 43, 420, 1595, 1533, 1733, 1294, 1695, 923, 884, 356, 999, 1833, 861, 319, 1348, 737, 1741, 1974, 233, 1701, 898, 1713, 1813, 1074, 675, 1004, 134, 169, 253, 1585, 2079, 1880, 1646, 966, 468, 142, 1474, 154, 2137, 1134, 1759, 374, 1862, 1706, 619, 1265, 1839, 1114, 2035, 1480, 1817, 2337, 1133, 1455, 2147, 2279]. time [1582, 1104, 745, 417, 57, 1593, 1622, 1179, 1834, 1184, 1970, 1185, 934, 1337, 364, 2213]. Time-adaptive [1614]. time-course [2072, 2079, 1706, 1622, 934]. Time-dependent [928]. Time-frequency [1154]. time-series [167, 417]. time-to-event [2035]. time-varying [2226, 1741, 1974, 1701, 374, 1455, 745, 1593, 1184]. timecourse [822]. times [2008, 871, 1065]. timeslot [2322]. titer [1892]. TLS [311, 1609]. tobacco [921]. Toeplitz [1600]. TOF [1835]. tolerance [1483, 1362, 849]. tool [2121, 1022, 933, 550, 1314]. toolbox [1047]. Tools [1018, 1538, 707, 28, 142]. tooth [2032]. Total [1596, 320, 1598, 801, 1599]. totals [2159, 465]. toxicity [1143]. traffic [1009, 856]. trait [1497, 2098, 2082]. traits [2047]. trajectories [406, 881]. TRAMO [920]. TRAMO-SEATS [920]. trans [1649]. trans-dimensional [1649]. transcription [2088]. transcriptional [2080]. transects [455]. transfer [942]. Transformation [419, 1672, 1352, 1288, 185, 1492, 35, 1839, 237, 1324, 2193, 521]. Transformations [390, 456, 1809, 764, 1491, 285, 2204, 1328, 952]. transformed [2252]. transforms [290]. transient [1638]. transition [328, 1390, 1697]. transitional [345]. transitions [1067, 2230]. Transmission [2102, 988, 857]. transmission/disequilibrium [857]. treatment [822, 1967, 1791, 1997, 1475, 1711, 553, 454, 1309]. treatments [725]. Tree [1697, 7, 247, 190, 538, 1620, 779, 860]. Tree-structured [1697]. trees [160, 2041, 765, 2273, 2032, 823, 2150, 777, 836, 1056, 519, 1380, 605, 616, 263, 71, 512, 520, 8, 1554]. tremor [163]. Trend [589, 2213, 895, 35, 2254, 364]. Trend-resistant [589]. Trends [1745, 271, 2187]. trial [1991, 1831, 2174, 1429]. trials [1982, 268, 1220, 284, 155, 1984, 1235, 99, 565, 872, 1994, 517, 1291, 1814, 1944, 2067, 1107, 830, 2001, 2040, 1461, 1318, 1385, 986]. Triangular [1360]. Trimmed [1545, 415, 132, 1667, 1371, 1090, 671, 1078, 697, 2151, 1540]. Trimming [1920, 1541]. triple [981]. triple-goal [981]. trivial [772]. true [1993, 633, 2139]. truncated [593, 1599]. truncation [2328]. Tucker [972]. Tucker- [972]. Tukey [1888, 586]. tumor [189]. tuning [146, 686]. Tweedie [1640]. Two [1326, 1943, 1997, 1594, 800, 575, 759, 2205, 1664, 1125, 1306, 36, 2210, 2002, 219, 506, 2072, 1717, 666, 465, 1488, 1720, 1433, 1654, 2170, 1767, 1789, 717, 574, 466, 2078, 34, 558, 976, 1967, 2336, 564, 1353, 689, 1001, 1060, 1150, 898, 938, 821, 358, 1837, 1354, 1983, 2306, 479, 1680, 660, 754, 376, 725, 414, 980, 1236, 1304, 885, 1629, 981, 254, 2249, 68, 1894, 1511, 1194, 1677, 1450, 1309, 1618, 2080, 1415, 1801, 758, 1159, 2215, 144, 551, 855, 1041, 2030, 270, 646, 1357, 1389, 874]. two-color [1159]. Two-level [1943]. Two-mode [1664, 2215]. two-parameter [1767, 1354, 414, 980, 1304, 1801, 855, 646]. two-part [874]. two-phase [1194]. two-release [358]. Two-sample [1326, 2072, 1654, 1789, 558, 1837, 479, 1389]. two-sided [1967, 270]. Two-stage [575, 36, 1001, 2306, 981, 1677, 2080, 1041]. two-step [1720]. Two-way [1997, 800, 1306, 465, 2336, 564, 1353, 689, 1150, 2215, 144]. Type [2119, 993, 1304, 2152, 2051, 388, 2027, 2078, 1119, 2237, 1812, 1397, 977, 1347, 1780, 2320, 1387, 1640, 683, 1869, 1485, 1821, 1792, 1717, 1811, 937, 980, 2249]. type-I [2051, 1780]. Type-II [1304, 2152, 683, 1717, 1811, 937, 980, 2249]. type-XII [2152]. types [1129, 374, 2205, 2180]. typical [1009].

Uhlenbeck [2142]. UK [1890, 921]. ultra [1171]. ultra-high [1171]. ULV [311]. umbrella [1151]. unbalanced [895, 52, 454]. Unbiased [1056, 1554, 605, 1259, 786]. Uncertainty [1797, 1984, 242, 2243, 1525, 1014]. Unconditional [643, 268, 506, 66, 1894]. uncorrelated [116]. undamped [1152]. underdispersion [452]. unemployment [1890, 379]. Unequal [226, 33, 312, 1446, 241, 52, 1503, 2227]. unfolding [2214, 2215]. unidimensional [1531]. unified [2085]. uniform [1359, 274, 481, 1543, 1221]. unifying [1586, 1557]. Unimodal [1952, 953]. uniqueness [899, 802]. unit [1307, 420, 951, 1938, 2131, 966, 2050, 1067]. United [2195]. units [629]. unity [1195]. Univariate [1020, 290]. unknown [648, 1303, 497, 950, 2309, 1847, 1746, 1147, 549, 767]. unmasking [110]. unnatural [1037]. unobservable [1256]. Unobserved [919, 929, 1074, 1621]. unrelated [312, 439]. unreplicated [108, 822, 41]. unrestricted [1606, 1353]. unsmoothness [2261]. unspecified [329, 1702]. unstructured [1244]. unsupervised [545, 881]. unusual [5]. updated [2150]. Updating [1529, 1954, 2003]. upper [242]. uraemic [2228]. Use [1492, 2345, 2342, 942, 704, 1086, 433, 239, 1478, 509, 86, 1561, 2063, 1525, 2091, 1076]. used [2146, 901]. useful [1670]. usefulness [1919, 1560]. User [426]. Using [906, 2109, 2110, 247, 1208, 1567, 1522, 1837, 429, 69, 1962, 471, 1501, 521, 9, 192, 2042, 22, 943, 930, 25, 2108, 1340, 461, 1497, 1770, 1092, 1886, 2187, 6, 1035, 2115, 1899, 1, 1403, 1066, 1672, 2292, 878, 998, 1735, 193, 1483, 1612, 1372, 355, 125, 1436, 1165, 1174, 400, 1487, 1130, 1856, 949, 1270, 150, 1452, 2232, 1620, 478, 1949, 302, 1055, 1732, 319, 1782, 911, 2031, 721, 1517, 1416, 1129, 2010, 1049, 1074, 1215, 1121, 1221, 1489, 358, 779, 860, 246, 2301, 1951, 1280, 317, 169, 1559, 1075, 1585, 1432, 1646, 1536, 1006, 1749, 1765]. using [2053, 687, 329, 1152, 1904, 1380, 135, 1540, 235, 2049, 1206, 1437, 1945, 1658, 460, 1706, 1787, 216, 381, 1103, 2337, 536, 1870, 1902, 71, 1652, 995, 1953, 1551, 2089, 1815, 2088, 1158, 1854, 2229, 2346, 2097, 857, 2141, 869, 1318, 1622, 1957, 870, 1203, 1834, 2257, 986, 2202, 1996, 265, 1325, 1137, 1786, 1505, 2116, 2075]. usual [5]. Utilizing [2095]. UV [2195].

V4 [1608]. vague [1030]. Valid [1203]. Validating [292]. validation [1690, 1999, 1383, 2267, 510, 329, 2088]. validity [2168]. Vallée [1040]. value [1811, 2124, 1175, 100, 2139, 1529, 489, 1339, 792]. value-at-risk [1175]. valued [1061, 1230, 1793]. values [1770, 2146, 1574, 341, 1211, 94, 949, 951, 342, 1395, 626, 2249, 1498, 102, 1464, 1677, 768, 1625, 1295, 1050, 88, 2293]. VAR [1114]. variability [245]. Variable [1517, 676, 1156, 2280, 520, 157, 1283, 1681, 611, 1578, 1029, 2319, 1472, 1624, 785, 1056, 988, 2169, 2167, 1536, 101, 1658, 1115, 1099, 1560, 956, 1769, 2193, 901, 2172, 1093, 1319, 1844, 2161, 418]. variables [181, 1447, 1932, 1574, 2046, 1369, 98, 1360, 1612, 339, 1558, 2048, 466, 1022, 1346, 1279, 555, 47, 251, 559, 2173, 1215, 515, 444, 2007, 1559, 1018, 1153, 566, 659, 518, 52, 691, 599, 953, 994, 866, 269, 1629, 371, 582, 1961, 1688, 637, 1704, 701, 77, 1596, 1454, 1929, 1597, 1639, 1471]. Variance [2159, 1359, 62, 227, 892, 2074, 648, 1303, 1433, 1310, 588, 2126, 1705, 37, 1481, 992, 978, 1670, 1757, 1659, 1007, 1645, 809, 685, 1098, 1396]. variance-structure [1705]. variances [919, 33, 1433, 1446, 571, 52, 343, 1602, 87]. variant [529]. variants [531, 1665]. variate [797, 80]. variates [2248, 1822]. variation [965, 223, 765]. Variational [1413, 1973, 1578]. VARIMAX [664]. variogram [303]. variograms [320]. various [308, 426]. Varying [1627, 2226, 817, 1196, 382, 1482, 1346, 1741, 1974, 722, 1150, 1701, 374, 1455, 900, 745, 1593, 397, 1184]. Varying-coefficient [1627, 1346]. VDM [1225]. VDM/ECM [1225]. vector [355, 999, 1584, 1782, 1543, 816, 1121, 1494, 1005, 616, 2304, 1706, 1839, 1455, 2089, 2014, 1815, 2088, 1085, 1524, 2213]. vectors [153, 582]. vegetation [2187]. vehicle [70]. Verification [2063, 1052, 885, 1995, 1992]. versatility [1508]. version [1232, 1802]. versus [920, 1727, 1358]. vertical [266]. Very [1073, 1025, 657, 293, 93]. vi [464]. via [2296, 2074, 1370, 698, 1558, 2048, 1682, 2019, 89, 590, 1634, 2095, 357, 1119, 1903, 1543, 1078, 1226, 240, 1186, 659, 487, 126, 616, 2304, 1114, 824, 1464, 983, 1418, 1952, 1911, 1283, 1440, 1512, 2285, 1802, 1339, 1940, 1076]. vibration [1749]. Vieu [1361]. view [1376, 1542]. Viewing [738]. views [116]. vine [1221]. vine-copula [1221]. violations [569]. viral [1936]. virological [1791]. virus [485]. visible [9]. ViSta [427]. visual [425, 672, 292]. visualisation [1781]. Visualization [107, 1467, 304, 432, 728, 433, 424, 1201]. Visualizations [2331]. Visualizing [1009, 427, 435, 914]. volatilities [926, 1727]. Volatility [1735, 1736, 1738, 969, 378, 1459, 2114, 1345, 2118, 1728, 2120, 2126, 2127, 1169, 2133, 1734, 1155, 2136, 1730, 2252, 893, 2140, 2141, 747, 1170]. volitional [163]. volume [105, 140, 173]. vs [543, 735, 2231].

wage [2125]. Waiting [1880]. Wald [135]. walk [229, 1984, 1202, 1700]. walks [1491, 1487]. ward [1444]. Waring [1477, 2263]. warning [1110]. waveform [79]. waveforms [1372]. Wavelet [266, 598, 693, 405, 1532, 1739, 1455, 931, 238, 703, 1467, 963, 1098]. Wavelet-based [1532, 931]. wavelets [930, 1165, 1585]. way [25, 476, 891, 465, 406, 2214, 2336, 564, 1812, 1353, 689, 1150, 1997, 1215, 800, 52, 275, 1306, 2215, 1414, 144]. weak [130]. web [207, 461, 302]. Web-based [461]. Webmining [207]. Weibull [124, 1966, 1810, 407, 1843, 1969, 2018, 360, 2299, 742, 2333, 683, 1158, 1415, 419, 1636]. weighing [1453]. weight [838, 495, 1860]. Weighted [1393, 237, 325, 1502, 1633, 32, 883, 2048, 975, 2129, 681, 1331, 1843, 1082, 1508, 1837, 801, 212, 2089, 1008]. weighting [2074, 17, 2030]. weights [1189]. Weinberg [2086]. Weiß [386]. West [2232]. Where [2070]. which [634, 7]. white [1746]. Whitney [1331]. Whittaker [1590]. whose [2096]. Wichmann [1853]. wide [1999, 207, 2094, 1041]. width [1359]. Wilcoxon [1066, 1331]. Wild [627, 735, 736, 1938]. Wilk [1397, 2281]. Wilk-type [1397]. Williams [895]. wind [1436]. Wisconsin [2309]. wise [1537, 801, 1953]. Wishart [2233]. within [112, 1980, 150, 1117, 23, 1077, 1316, 957]. within-blocks [957]. without [1976, 400, 1934, 2261]. Woodruff [30]. word [800]. Workarounds [1111]. workflow [1444, 1201]. works [1268]. world [207, 435]. worms [1847]. Worst [1256]. Worst-case [1256]. wrapped [1631].

X [30]. XGobi [424]. XII [2152, 525, 1795]. XP [291]. xx [1361].

Yohai [453]. Youden [267].

Zero [335, 250, 184, 353, 2162, 74, 588, 1235, 1286, 2044, 1757, 2241, 88]. Zero-Inflated [250, 335, 184, 353, 2241]. zero-inflation [2162, 2044]. zero-modified [74]. Zipf [1123]. zonoid [2151].

References

Caers:1998:NTE

[1] Jef Caers and Jozef Van Dyck. Nonparametric tail estimation using a double bootstrap method. Computational Statistics & Data Analysis, 29(2):191– 211, December 28, 1998. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/S0167947398000607.

Li:1998:CNA

[2] Baibing Li and Bart De Moor. A corrected normal approximation for the Student’s t distribution.

Computational Statistics & Data Analysis, 29(2):213–216, December 28, 1998. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947398000656. See

note [55].

Coleman:1999:SCI

[3] Dan Coleman, Xioapeng Dong, Johanna Hardin, David M. Rocke, and David L. Woodruff. Some computational issues in cluster analysis with no a priori metric. Computational Statistics & Data Analysis, 31(1):1– 11, July 28, 1999. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000092. See

comments [30].

Wegman:2000:ECS

[4] Edward J. Wegman. On the eve of the 21st century: statistical science at a crossroads. Computational Statistics & Data Analysis, 32(3–4): 239–243, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739900078X.

Downing:2000:LDS

[5] D. J. Downing, V. V. Fedorov, W. F. Lawkins, M. D. Morris, and G. Ostrouchov. Large data series: modeling the usual to identify the unusual. Computational Statistics & Data Analysis, 32(3–4):245–258, January 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000791.

Bonetti:2000:NGA

[6] Marco Bonetti. A new geometric approach to data analysis using the Minkowski polytope. Computational Statistics & Data Analysis, 32(3–4): 259–271, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000808.

Berry:2000:CTW

[7] V. Berry, O. Gascuel, and G. Caraux. Choosing the tree which actually best explains the data: another look at the bootstrap in phylogenetic reconstruction. Computational Statistics & Data Analysis, 32(3–4): 273–283, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739900081X.

Siciliano:2000:MDA

[8] Roberta Siciliano and Francesco Mola. Multivariate data analysis and modeling through classification and regression trees. Computational Statistics & Data Analysis, 32(3–4):285– 301, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000821.

Unwin:2000:UYE

[9] Antony Unwin. Using your eyes — making statistics more visible with computers. Computational Statistics & Data Analysis, 32(3–4):303– 312, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000833.

Shortle:2000:NMD

[10] John F. Shortle and Max B. Mendel. A new method of drawing probability densities. Computational Statistics & Data Analysis, 32(3–4):313– 321, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000845.

Hudson:2000:MFS

[11] H. Malcolm Hudson and Craig Walsh. Model fitting for sequences of images. Computational Statistics & Data Analysis, 32(3–4):323–335, January 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000857.

Clarkson:2000:REI

[12] Douglas B. Clarkson. A random effects individual difference multidimensional scaling model. Computational Statistics & Data Analysis, 32(3–4): 337–347, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000870.

Wijsman:2000:MCM

[13] Ellen M. Wijsman. Monte Carlo Markov chain methods and model selection in genetic epidemiology. Computational Statistics & Data Analysis, 32(3–4):349–360, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947399000882.

Mazzuchi:2000:CME

[14] T. A. Mazzuchi, E. S. Soofi, and R. Soyer. Computation of maximum entropy Dirichlet for modeling lifetime data. Computational Statistics & Data Analysis, 32(3–4):361– 378, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000900.

Sato-Ilic:2000:AAO

[15] Mika Sato-Ilic and Yoshiharu Sato. Asymmetric aggregation operator and its application to fuzzy clustering model. Computational Statistics & Data Analysis, 32(3–4):379–394, January 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000912.

Balbi:2000:RCA

[16] S. Balbi and V. Esposito. Rotated canonical analysis onto a reference subspace. Computational Statistics & Data Analysis, 32(3–4):395– 410, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000924.

Michailidis:2000:MHA

[17] George Michailidis and Jan de Leeuw. Multilevel homogeneity analysis with differential weighting. Computational Statistics & Data Analysis, 32(3–4): 411–442, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000869.

Capobianco:2000:NNS

[18] Enrico Capobianco. Neural networks and statistical inference: seeking robust and efficient learning. Computational Statistics & Data Analysis, 32(3– 4):443–454, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000894.

Hilton:2000:NAD

[19] Joan F. Hilton. A new asymptotic distribution for Hollander’s bivariate symmetry statistic. Computational Statistics & Data Analysis, 32(3–4): 455–463, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000936.

Bentley:2000:SAI

[20] Jim Bentley and Tammi J. Schneider. Statistics and archaeology in Israel. Computational Statistics & Data Analysis, 32(3–4):465–483, January 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000948.

Anonymous:2000:Ia

[21] Anonymous. Index. Computational Statistics & Data Analysis, 32(3–4): 511–512, January 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001188.

Adatia:2000:EPH

[22] A. Adatia. Estimation of parameters of the half-logistic distribution using generalized ranked set sampling. Computational Statistics & Data Analysis, 33(1):1–13, March 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000353.

Kang:2000:ELR

[23] Seung-Ho Kang and Soo Min Park. Exact likelihood ratio test of independence of binary responses within clusters. Computational Statistics & Data Analysis, 33(1):15–23, March 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000444.

Trinca:2000:AAR

[24] Luzia A. Trinca and Steven G. Gilmour. An algorithm for arranging response surface designs in small blocks. Computational Statistics & Data Analysis, 33(1):25–43, March 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739900033X. See

erratum [277].

Anderson-Cook:2000:IEU

[25] Christine M. Anderson-Cook. An industrial example using one-way analysis of circular-linear data. Computational Statistics & Data Analysis, 33 (1):45–57, March 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000365.

Famoye:2000:GFT

[26] Felix Famoye. Goodness-of-fit tests for generalized logarithmic series distribution. Computational Statistics & Data Analysis, 33(1):59–67, March 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000481.

Lai:2000:DCL

[27] Dejian Lai and Guanrong Chen. Distribution of controlled Lyapunov exponents: a statistical simulation study. Computational Statistics & Data Analysis, 33(1):69–77, March 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739900047X.

Lindsey:2000:DTR

[28] P. J. Lindsey and J. K. Lindsey. Diagnostic tools for random effects in the repeated measures growth curve model. Computational Statistics & Data Analysis, 33(1):79–100, March 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000493.

Soler:2000:OCA

[29] Júlia M. P. Soler and Julio M. Singer. Optimal covariance adjustment in growth curve models. Computational Statistics & Data Analysis, 33(1):101– 110, March 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000341.

Fraley:2000:CDC

[30] C. Fraley. Comments on: D. Coleman, X. Dong, J. Hardin, D. M. Rocke, D. L. Woodruff, “Some computational issues in cluster analysis with no a priori metric”, Computational Statistics & Data Analysis 31:1–12 (July 1999). Computational Statistics & Data Analysis, 33 (2):131–133, April 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001061. See

[3].

Rocke:2000:ARC

[31] David M. Rocke and David L. Woodruff. Author’s response: Comments on “Some computational issues in cluster analysis with no a priori metric”. Computational Statistics & Data Analysis, 33(2):135– 136, April 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739900105X. See

[3, 30].

Chen:2000:EOD

[32] Ray-Bing Chen and Mong-Na Lo Huang. Exact D-optimal designs for weighted polynomial regression model. Computational Statistics & Data Analysis, 33(2):137– 149, April 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000547.

Chen:2000:ATE

[33] Shun-Yi Chen, Hubert J. Chen, and Cherng G. Ding. An ANOVA test for the equivalency of means under unequal variances. Computational Statistics & Data Analysis, 33(2):151– 175, April 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000535.

Gower:2000:ROR

[34] John C. Gower. Rank-one and ranktwo departures from symmetry. Computational Statistics & Data Analysis, 33(2):177–188, April 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739900050X.

Neuhauser:2000:PLS

[35] Markus Neuhäuser and Ludwig A. Hothorn. Parametric location-scale and scale trend tests based on Levene’s transformation. Computational Statistics & Data Analysis, 33(2):189– 200, April 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000511.

Wu:2000:TSM

[36] Shu-Fei Wu and Hubert J. Chen. Two-stage multiple comparisons with the average for normal distributions under heteroscedasticity. Computational Statistics & Data Analysis, 33 (2):201–213, April 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000523.

Karlsson:2000:CED

[37] Sune Karlsson and Mickael Löthgren. Computationally efficient double bootstrap variance estimation. Computational Statistics & Data Analysis, 33 (3):237–247, May 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000663.

Carling:2000:ROR

[38] Kenneth Carling. Resistant outlier rules and the non-Gaussian case. Computational Statistics & Data Analysis, 33(3):249–258, May 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000572.

DeGooijer:2000:NCP

[39] Jan G. De Gooijer and Ali Gannoun. Nonparametric conditional predictive regions for time series. Computational Statistics & Data Analysis, 33 (3):259–275, May 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000560.

Krivy:2000:SAN

[40] Ivan Křivý, Josef Tvrdík, and Radek Krpec. Stochastic algorithms in nonlinear regression. Computational Statistics & Data Analysis, 33(3):277– 290, May 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000596.

Liao:2000:IDE

[41] C. T. Liao. Identification of dispersion effects from unreplicated 2^{n−k} fractional factorial designs. Computational Statistics & Data Analysis, 33 (3):291–298, May 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000584.

Sanchez-Meca:2000:TSC

[42] Julio Sánchez-Meca and Fulgencio Marín-Martínez. Testing the significance of a common risk difference in meta-analysis. Computational Statistics & Data Analysis, 33 (3):299–313, May 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000559.

Ferreira:2000:SAS

[43] Eva Ferreira, Vicente Núñez-Antón, and Juan Rodríguez-Póo. Semiparametric approaches to signal extraction problems in economic time series. Computational Statistics & Data Analysis, 33(3):315–333, May 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000626.

Chen:2000:CNB

[44] Z. Y. Chen and Y. C. Chou. Computing the noncentral beta distribution with S-system. Computational Statistics & Data Analysis, 33(4):343– 360, June 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000675.

Atkinson:2000:RLR

[45] A. C. Atkinson and Tsung-Chi Cheng. On robust linear regression with incomplete data. Computational Statistics & Data Analysis, 33(4):361– 380, June 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000614.

Fan:2000:RBD

[46] Tsai-Hung Fan and James O. Berger. Robust Bayesian displays for standard inferences concerning a normal mean. Computational Statistics & Data Analysis, 33(4):381– 399, June 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000602.

Kim:2000:HEL

[47] Sung-Ho Kim. Hyper-EM for large recursive models of categorical variables. Computational Statistics & Data Analysis, 33(4):401– 424, June 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794739900064X.

Nicole:2000:FNN

[48] Sandro Nicole. Feedforward neural networks for principal components extraction. Computational Statistics & Data Analysis, 33(4):425– 437, June 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000638.

Spiess:2000:CDM

[49] Martin Spiess and Alfred Hamerle. A comparison of different methods for the estimation of regression models with correlated binary responses. Computational Statistics & Data Analysis, 33 (4):439–455, June 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000651.

Anonymous:2000:Ib

[50] Anonymous. Index. Computational Statistics & Data Analysis, 33(4):473– 474, June 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730000013X.

Drew:2000:CCD

[51] John H. Drew, Andrew G. Glen, and Lawrence M. Leemis. Computing the cumulative distribution function of the Kolmogorov–Smirnov statistic. Computational Statistics & Data Analysis, 34(1):1–15, July 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000699.

Miwa:2000:CLP

[52] Tetsuhisa Miwa, A. J. Hayter, and Wei Liu. Calculations of level probabilities for normal random variables with unequal variances with applications to Bartholomew’s test in unbalanced one-way models. Computational Statistics & Data Analysis, 34 (1):17–32, July 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000754.

Natarajan:2000:MCE

[53] Ranjini Natarajan, Charles E. Mc-Culloch, and Nicholas M. Kiefer. A Monte Carlo EM method for estimating multinomial probit models. Computational Statistics & Data Analysis, 34(1):33–50, July 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000730.

Clemons:2000:NMO

[54] Traci E. Clemons and Edwin L. Bradley, Jr. A nonparametric measure of the overlapping coefficient. Computational Statistics & Data Analysis, 34(1):51–61, July 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000742. See

erratum [120].

Gleason:2000:NPS

[55] John R. Gleason. A note on a proposed Student t approximation. Computational Statistics & Data Analysis, 34(1):63–66, July 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000705. See

[2].

Visek:2000:DE

[56] Jan Ámos Víšek. On the diversity of estimates. Computational Statistics & Data Analysis, 34(1):67– 89, July 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000687.

Teles:2000:ETA

[57] Paulo Teles and William W. S. Wei. The effects of temporal aggregation on tests of linearity of a time series. Computational Statistics & Data Analysis, 34(1):91–103, July 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000729.

Mulekar:2000:CIE

[58] Madhuri S. Mulekar and Satya N. Mishra. Confidence interval estimation of overlap: equal means case. Computational Statistics & Data Analysis, 34 (2):121–137, August 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000961.

Hartford:2000:CMA

[59] Alan Hartford and Marie Davidian. Consequences of misspecifying assumptions in nonlinear mixed effects models. Computational Statistics & Data Analysis, 34(2):139–164, August 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399000766.

Lee:2000:NRS

[60] Sauchi Stephen Lee. Noisy replication in skewed binary classification. Computational Statistics & Data Analysis, 34 (2):165–191, August 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794739900095X.

Piegorsch:2000:ACB

[61] Walter W. Piegorsch, R. Webster West, Obaid M. Al-Saidy, and Kelly D. Bradley. Asymmetric confidence bands for simple linear regression over bounded intervals. Computational Statistics & Data Analysis, 34(2):193– 217, August 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000973.

Borkowf:2000:NNM

[62] Craig B. Borkowf. A new nonparametric method for variance estimation and confidence interval construction for Spearman’s rank correlation. Computational Statistics & Data Analysis, 34 (2):219–241, August 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000778.

Xiang:2000:CPN

[63] Anny Xiang, Pablo Lapuerta, Alex Ryutov, Jonathan Buckley, and Stanley Azen. Comparison of the performance of neural network methods and Cox regression for censored survival data. Computational Statistics & Data Analysis, 34(2):243–257, August 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000985.

Billor:2000:BBA

[64] Nedret Billor, Ali S. Hadi, and Paul F. Velleman. BACON: blocked adaptive computationally efficient outlier nominators. Computational Statistics & Data Analysis, 34(3):279–298, September 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001012.

DeOliveira:2000:BPC

[65] Victor De Oliveira. Bayesian prediction of clipped Gaussian random fields. Computational Statistics & Data Analysis, 34(3):299–314, September 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001036.

Jandhyala:2000:CUC

[66] Venkata K. Jandhyala, Stergios B. Fotopoulos, and Nicholas E. Evaggelopoulos. A comparison of unconditional and conditional solutions to the maximum likelihood estimation of a change-point. Computational Statistics & Data Analysis, 34(3):315–334, September 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001024.

Lui:2000:ACT

[67] Kung-Jong Lui. Asymptotic conditional test procedures for relative difference under inverse sampling. Computational Statistics & Data Analysis, 34(3):335–343, September 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001048.

Price:2000:ERT

[68] Robert M. Price and Douglas G. Bonett. Estimating the ratio of two Poisson rates. Computational Statistics & Data Analysis, 34(3):345–356, September 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001000.

Shen:2000:UOP

[69] Larry Z. Shen and John O’Quigley. Using a one-parameter model to sequentially estimate the root of a regression function. Computational Statistics & Data Analysis, 34(3):357–369, September 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001073.

Kuhnert:2000:CNP

[70] Petra M. Kuhnert, Kim-Anh Do, and Rod McClure. Combining nonparametric models with logistic regression: an application to motor vehicle injury data. Computational Statistics & Data Analysis, 34(3):371– 386, September 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399000997.

Salmeron:2000:ISB

[71] Antonio Salmerón, Andrés Cano, and Serafín Moral. Importance sampling in Bayesian networks using probability trees. Computational Statistics & Data Analysis, 34(4):387–413, October 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001103.

Struyf:2000:HDC

[72] Anja Struyf and Peter J. Rousseeuw. High-dimensional computation of the deepest location. Computational Statistics & Data Analysis, 34(4):415– 426, October 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001127.

DUrso:2000:LSA

[73] Pierpaolo D’Urso and Tommaso Gastaldi. A least-squares approach to fuzzy linear regression analysis. Computational Statistics & Data Analysis, 34(4):427–440, October 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001097.

Dietz:2000:EPP

[74] Ekkehart Dietz and Dankmar Böhning. On estimation of the Poisson parameter in zero-modified Poisson models. Computational Statistics & Data Analysis, 34(4):441–459, October 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947399001115.

Mittlbock:2000:AMP

[75] M. Mittlböck and T. Waldhör. Adjustments for R^{2}-measures for Poisson regression models. Computational Statistics & Data Analysis, 34(4):461– 472, October 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001139.

Orre:2000:BNN

[76] R. Orre, A. Lansner, A. Bate, and M. Lindquist. Bayesian neural networks with confidence estimations applied to data mining. Computational Statistics & Data Analysis, 34(4):473– 493, October 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001140.

Tanizaki:2000:BCO

[77] Hisashi Tanizaki. Bias correction of OLSE in the regression model with lagged dependent variables. Computational Statistics & Data Analysis, 34(4):495–511, October 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001085.

Anonymous:2000:Ic

[78] Anonymous. Index. Computational Statistics & Data Analysis, 34(4):547– 548, October 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730000044X.

Shimansky:2000:CMS

[79] Yury P. Shimansky. Continuous measure of significant linear dimensionality of a waveform set. Computational Statistics & Data Analysis, 35 (1):1–10, November 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000050.

Landim:2000:DHM

[80] Flávia Landim and Dani Gamerman. Dynamic hierarchical models: an extension to matrix-variate observations. Computational Statistics & Data Analysis, 35(1):11–42, November 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000049.

Priebe:2000:AKM

[81] Carey E. Priebe and David J. Marchette. Alternating kernel and mixture density estimates. Computational Statistics & Data Analysis, 35(1):43– 65, November 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000037.

Doksum:2000:SEP

[82] Kjell Doksum and Ja-Yong Koo. On spline estimators and prediction intervals in nonparametric regression. Computational Statistics & Data Analysis, 35(1):67–82, November 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001164.

Jhun:2000:ABM

[83] Myoungshic Jhun and Hyeong-Chul Jeong. Applications of bootstrap methods for categorical data analysis. Computational Statistics & Data Analysis, 35(1):83–91, November 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001152.

Oliveira:2000:EIB

[84] Aurelio R. L. Oliveira, Mário A. Nascimento, and Christiano Lyra. Efficient implementation and benchmark of interior point methods for the polynomial L_{1} fitting problem. Computational Statistics & Data Analysis, 35(2):119– 135, December 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000062.

Childs:2000:SAM

[85] Aaron Childs and N. Balakrishnan. Some approximations to the multivariate hypergeometric distribution with applications to hypothesis testing. Computational Statistics & Data Analysis, 35(2):137–154, December 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000074.

Neath:2000:RMS

[86] Andrew A. Neath and Joseph E. Cavanaugh. A regression model selection criterion based on bootstrap bumping for use with resistant fitting. Computational Statistics & Data Analysis, 35(2):155–169, December 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000104.

Thorpe:2000:SMC

[87] David P. Thorpe and Burt Holland. Some multiple comparison procedures for variances from non-normal populations. Computational Statistics & Data Analysis, 35(2):171–199, December 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000086.

Zhou:2000:IER

[88] Xiao-Hua Zhou and Wanzhu Tu. Interval estimation for the ratio in means of log-normally distributed medical costs with zero values. Computational Statistics & Data Analysis, 35(2):201– 210, December 28, 2000. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000098.

Fernandez:2000:RCL

[89] J. M. Vilar Fernández and W. González Manteiga. Resampling for checking linear regression models via non-parametric regression estimation. Computational Statistics & Data Analysis, 35(2):211–231, December 28, 2000. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947399001176.

Choulakian:2001:AND

[90] V. Choulakian and J. Almhana. An algorithm for nonmetric discriminant analysis. Computational Statistics & Data Analysis, 35(3):253–264, January 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000190.

Galecki:2001:SFA

[91] Andrzej T. Galecki, Thomas R. Ten Have, and Geert Molenberghs. A simple and fast alternative to the EM algorithm for incomplete categorical data and latent class models. Computational Statistics & Data Analysis, 35(3):265– 281, January 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000153.

Higdon:2001:MLC

[92] Roger Higdon and Daniel W. Schafer. Maximum likelihood computations for regression with measurement error. Computational Statistics & Data Analysis, 35(3):283–299, January 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000141.

Smirnov:2001:FML

[93] Oleg Smirnov and Luc Anselin. Fast maximum likelihood estimation of very large spatial autoregressive models: a characteristic polynomial approach. Computational Statistics & Data Analysis, 35(3):301–319, January 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000189.

Dupuis:2001:FLM

[94] D. J. Dupuis. Fitting log- F models robustly, with an application to the analysis of extreme values. Computational Statistics & Data Analysis, 35(3):321– 333, January 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000128.

OGorman:2001:APT

[95] Thomas W. O’Gorman. An adaptive permutation test procedure for several common tests of significance. Computational Statistics & Data Analysis, 35(3):335–350, January 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000116.

Ritter:2001:PFA

[96] Gunter Ritter and Christoph Pesch. Polarity-free automatic classification of chromosomes. Computational Statistics & Data Analysis, 35(3):351–372, January 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000220.

Gauderman:2001:MCN

[97] W. James Gauderman and William Navidi. A Monte Carlo Newton– Raphson procedure for maximizing complex likelihoods on pedigree data. Computational Statistics & Data Analysis, 35(4):395–415, February 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000219.

Casin:2001:GPC

[98] Ph. Casin. A generalization of principal component analysis to K sets of variables. Computational Statistics & Data Analysis, 35(4):417–428, February 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000244.

Hartzel:2001:DHE

[99] Jonathan Hartzel, I-Ming Liu, and Alan Agresti. Describing heterogeneous effects in stratified ordinal contingency tables, with application to multicenter clinical trials. Computational Statistics & Data Analysis, 35(4):429– 449, February 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000207.

Heinzl:2001:CNS

[100] Harald Heinzl and Clemens Tempfer. A cautionary note on segmenting a cyclical covariate by minimum P - value search. Computational Statistics & Data Analysis, 35(4):451–461, February 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000232.

Montanari:2001:PPA

[101] Angela Montanari and Laura Lizzani. A projection pursuit approach to variable selection. Computational Statistics & Data Analysis, 35(4):463– 473, February 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000268.

Priebe:2001:ISS

[102] Carey E. Priebe, Daniel Q. Naiman, and Leslie M. Cope. Importance sampling for spatial scan analysis: computing scan statistic p-values for marked point processes. Computational Statistics & Data Analysis, 35(4):475–485, February 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000177.

VanKeilegom:2001:ECD

[103] Ingrid Van Keilegom, Michael G. Akritas, and Noël Veraverbeke. Estimation of the conditional distribution in regression with censored data: a comparative study. Computational Statistics & Data Analysis, 35(4):487– 500, February 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000256.

Anonymous:2001:SSa

[104] Anonymous. SSN-section. Computational Statistics & Data Analysis, 35(4):501–512, February 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947300000608.

Anonymous:2001:AIVa

[105] Anonymous. Author index volume 35 (2001). Computational Statistics & Data Analysis, 35(4):513–514, February 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000591.

Chan:2001:EIB

[106] Kenny Y. F. Chan and Stephen M. S. Lee. An exact iterated bootstrap algorithm for small-sample bias reduction. Computational Statistics & Data Analysis, 36(1):1–13, March 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000293.

Dzemyda:2001:VSP

[107] Gintautas Dzemyda. Visualization of a set of parameters characterized by their correlation matrix. Computational Statistics & Data Analysis, 36 (1):15–30, March 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730000030X.

Aboukalam:2001:RAU

[108] M. A. F. Aboukalam and A. A. Al-Shiha. A robust analysis for unreplicated factorial experiments. Computational Statistics & Data Analysis, 36 (1):31–46, March 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730000027X.

Heo:2001:FGD

[109] Moonseong Heo and K. Ruben Gabriel. The fit of graphical displays to patterns of expectations. Computational Statistics & Data Analysis, 36(1):47– 67, March 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000165.

Justel:2001:BUL

[110] Ana Justel and Daniel Peña. Bayesian unmasking in linear models. Computational Statistics & Data Analysis, 36 (1):69–84, March 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000335.

Renshaw:2001:GPP

[111] Eric Renshaw and Aila Särkkä. Gibbs point processes for studying the development of spatial-temporal stochastic processes. Computational Statistics & Data Analysis, 36(1):85–105, March 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000281.

Bartolucci:2001:DMC

[112] F. Bartolucci. Developments of the Markov chain approach within the distribution theory of runs. Computational Statistics & Data Analysis, 36 (1):107–118, March 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000311.

Becker:2001:LNO

[113] Claudia Becker and Ursula Gather. The largest nonidentifiable outlier: a comparison of multivariate simultaneous outlier identification rules. Computational Statistics & Data Analysis, 36 (1):119–127, March 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000323.

Anonymous:2001:SSb

[114] Anonymous. SSN-section. Computational Statistics & Data Analysis, 36 (1):129–143, March 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000032.

Noubiap:2001:EAC

[115] Roger Fandom Noubiap and Wilfried Seidel. An efficient algorithm for constructing Γ-minimax tests for finite parameter spaces. Computational Statistics & Data Analysis, 36 (2):145–161, April 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000347.

Scrucca:2001:RCC

[116] Luca Scrucca. A review and computer code for assessing the structural dimension of a regression model: uncorrelated 2D views. Computational Statistics & Data Analysis, 36 (2):163–177, April 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000359.

Glendinning:2001:SSS

[117] Richard H. Glendinning. Selecting sub-set autoregressions from outlier contaminated data. Computational Statistics & Data Analysis, 36(2):179– 207, April 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000384.

Hand:2001:SCS

[118] David J. Hand, Hua Gui Li, and Niall M. Adams. Supervised classification with structured class definitions. Computational Statistics & Data Analysis, 36(2):209– 225, April 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000372.

Velilla:2001:BMR

[119] Santiago Velilla. On the bootstrap in misspecified regression models. Computational Statistics & Data Analysis, 36(2):227–242, April 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000396.

Clemons:2001:ENM

[120] T. E. Clemons. Erratum to “A nonparametric measure of the overlapping coefficient” [comput. statist. data anal. 34 (2000) 51–61]. Computational Statistics & Data Analysis, 36 (2):243, April 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000475. See

[54].

Kundu:2001:ENS

[121] Debasis Kundu and Amit Mitra. Estimating the number of signals of the damped exponential models. Computational Statistics & Data Analysis, 36 (2):245–256, April 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000360.

Anonymous:2001:SSc

[122] Anonymous. SSN-section. Computational Statistics & Data Analysis, 36 (2):257–278, April 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000044.

Bashtannyk:2001:BSK

[123] David M. Bashtannyk and Rob J. Hyndman. Bandwidth selection for kernel conditional density estimation. Computational Statistics & Data Analysis, 36(3):279–298, May 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000463.

Bowman:2001:WDW

[124] K. O. Bowman and L. R. Shenton. Weibull distributions when the shape parameter is defined. Computational Statistics & Data Analysis, 36(3):299– 310, May 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000487.

Devidas:2001:LDE

[125] Meenakshi Devidas and E. Olusȩgun George. Low-dose extrapolation using the power family of response functions. Computational Statistics & Data Analysis, 36(3):311– 317, May 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000402.

Ogliari:2001:ALD

[126] Paulo J. Ogliari and Dalton F. Andrade. Analysing longitudinal data via nonlinear models in randomized block designs. Computational Statistics & Data Analysis, 36(3):319– 332, May 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000499.

Polansky:2001:BSS

[127] Alan M. Polansky. Bandwidth selection for the smoothed bootstrap percentile method. Computational Statistics & Data Analysis, 36(3):333– 349, May 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000451.

Wisnowski:2001:CAM

[128] James W. Wisnowski, Douglas C. Montgomery, and James R. Simpson. A comparative analysis of multiple outlier detection procedures in the linear regression model. Computational Statistics & Data Analysis, 36 (3):351–382, May 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000426.

Aerts:2001:BTM

[129] Marc Aerts and Gerda Claeskens. Bootstrap tests for misspecified models, with application to clustered binary data. Computational Statistics & Data Analysis, 36(3):383– 401, May 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000517.

Ahmad:2001:MCK

[130] Ibrahim Ahmad, Xiaohong Chen, and Qi Li. Model check by kernel methods under weak moment conditions. Computational Statistics & Data Analysis, 36(3):403–409, May 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000438.

Anonymous:2001:SSd

[131] Anonymous. SSN-section. Computational Statistics & Data Analysis, 36 (3):411–423, May 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000238.

Agullo:2001:NAC

[132] José Agulló. New algorithms for computing the least trimmed squares regression estimator. Computational Statistics & Data Analysis, 36(4):425– 439, June 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000566.

Cuevas:2001:CAF

[133] Antonio Cuevas, Manuel Febrero, and Ricardo Fraiman. Cluster analysis: a further approach based on density estimation. Computational Statistics & Data Analysis, 36(4):441– 459, June 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000529.

Liu:2001:BMD

[134] Shu-Ing Liu. Bayesian model determination for binary-time-series data with applications. Computational Statistics & Data Analysis, 36(4):461– 473, June 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000530.

Neto:2001:WOS

[135] José Cardoso Neto and Gilberto A. Paula. Wald one-sided test using generalized estimating equations approach. Computational Statistics & Data Analysis, 36(4):475– 495, June 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000414.

Cao:2001:RDE

[136] Ricardo Cao, Paul Janssen, and Noël Veraverbeke. Relative density estimation and local bandwidth selection for censored data. Computational Statistics & Data Analysis, 36 (4):497–510, June 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000554.

Furno:2001:LER

[137] Marilena Furno. LAD estimation with random coefficient autocorrelated errors. Computational Statistics & Data Analysis, 36(4):511– 523, June 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000505.

Huang:2001:IEE

[138] Yangxin Huang. Interval estimation of the ED_{50} when a logistic dose–response curve is incorrectly assumed. Computational Statistics & Data Analysis, 36 (4):525–537, June 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000542.

Anonymous:2001:SSe

[139] Anonymous. SSN-section. Computational Statistics & Data Analysis, 36 (4):539–560, June 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000366.

Anonymous:2001:AIVb

[140] Anonymous. Author index volume 36 (2001). Computational Statistics & Data Analysis, 36(4):561– 562, June 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000378.

Baragona:2001:GAI

[141] Roberto Baragona, Francesco Battaglia, and Claudio Calzini. Genetic algorithms for the identification of additive and innovation outliers in time series. Computational Statistics & Data Analysis, 37(1):1–12, July 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730000058X.

Nielsen:2001:GSC

[142] Henrik Aa. Nielsen and Henrik Madsen. A generalization of some classical time series tools. Computational Statistics & Data Analysis, 37 (1):13–31, July 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730000061X.

Soong:2001:ESC

[143] W. C. Soong. Exact simultaneous confidence intervals for multiple comparisons with the mean. Computational Statistics & Data Analysis, 37 (1):33–47, July 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000621.

Vichi:2001:FMA

[144] Maurizio Vichi and Henk A. L. Kiers. Factorial k-means analysis for two-way data. Computational Statistics & Data Analysis, 37(1):49– 64, July 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000645.

Christmann:2001:MOB

[145] Andreas Christmann and Peter J. Rousseeuw. Measuring overlap in binary regression. Computational Statistics & Data Analysis, 37(1):65– 75, July 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000633.

Cox:2001:SMT

[146] Dennis D. Cox, Jeong-Soo Park, and Clifford E. Singer. A statistical method for tuning a computer code to a data base. Computational Statistics & Data Analysis, 37 (1):77–92, July 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000578.

Molenberghs:2001:IAA

[147] Geert Molenberghs, Geert Verbeke, Herbert Thijs, Emmanuel Lesaffre, and Michael G. Kenward. Influence analysis to assess sensitivity of the dropout process. Computational Statistics & Data Analysis, 37(1):93– 113, July 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000657.

Anonymous:2001:SSf

[148] Anonymous. SSN-section. Computational Statistics & Data Analysis, 37 (1):115–134, July 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730100038X.

Choulakian:2001:RMP

[149] V. Choulakian. Robust Q-mode principal component analysis in L_{1}. Computational Statistics & Data Analysis, 37 (2):135–150, August 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000056.

Geweke:2001:BES

[150] John Geweke and Hisashi Tanizaki. Bayesian estimation of state-space models using the Metropolis–Hastings algorithm within Gibbs sampling. Computational Statistics & Data Analysis, 37(2):151–170, August 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947301000093.

Hummer:2001:ELR

[151] Amanda Hummer and Song Yang. On the extended log-rank estimating function for the censored regression model. Computational Statistics & Data Analysis, 37(2):171–180, August 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000694.

Kowalczyk:2001:AMK

[152] T. Kowalczyk and M. Niewiadomska-Bugaj. An algorithm for maximizing Kendall’s τ. Computational Statistics & Data Analysis, 37(2):181–193, August 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947300000670.

Nettleton:2001:TED

[153] Dan Nettleton and T. Banerjee. Testing the equality of distributions of random vectors with categorical components. Computational Statistics & Data Analysis, 37(2):195–208, August 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000159.

Nottingham:2001:LLR

[154] Quinton J. Nottingham and Deborah F. Cook. Local linear regression for estimating time series data. Computational Statistics & Data Analysis, 37 (2):209–217, August 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000068.

Friede:2001:NCP

[155] T. Friede, F. Miller, W. Bischoff, and M. Kieser. A note on change point estimation in dose–response trials. Computational Statistics & Data Analysis, 37 (2):219–232, August 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000669.

Shin:2001:TOS

[156] Dong Wan Shin, Chul Gyu Park, and Taesung Park. Testing for onesided group effects in repeated measures study. Computational Statistics & Data Analysis, 37(2):233–247, August 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000081.

Steel:2001:VSD

[157] S. J. Steel and N. Louw. Variable selection in discriminant analysis: measuring the influence of individual cases. Computational Statistics & Data Analysis, 37(2):249–260, August 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000123.

Anonymous:2001:SSg

[158] Anonymous. SSN-section. Computational Statistics & Data Analysis, 37 (2):261–274, August 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000500.

Angelis:2001:OEE

[159] L. Angelis, E. Bora-Senta, and C. Moyssiadis. Optimal exact experimental designs with correlated errors through a simulated annealing algorithm. Computational Statistics & Data Analysis, 37(3):275–296, September 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000111.

Badsberg:2001:IIP

[160] J. H. Badsberg and F. M. Malvestuto. An implementation of the iterative proportional fitting procedure by propagation trees. Computational Statistics & Data Analysis, 37(3):297– 322, September 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000135.

Hawkins:2001:FMC

[161] Douglas M. Hawkins. Fitting multiple change-point models to data. Computational Statistics & Data Analysis, 37(3):323–341, September 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947300000682.

Kellison:2001:ACT

[162] F. W. Kellison and D. R. Halverson. Applications of curvature toward the measurement of robustness for data processors. Computational Statistics & Data Analysis, 37(3):343– 362, September 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730100007X.

Eakin:2001:DTV

[163] T. Eakin, K. Francis, and W. Spirduso. Decoupling tremor and volitional force control factors in analyses of motor task performance. Computational Statistics & Data Analysis, 37(3):363– 372, September 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000196.

Intrator:2001:INN

[164] Orna Intrator and Nathan Intrator. Interpreting neural-network results: a simulation study. Computational Statistics & Data Analysis, 37(3):373– 393, September 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000160.

Anonymous:2001:SSh

[165] Anonymous. SSN-section. Computational Statistics & Data Analysis, 37(3):395–404, September 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947301000603.

Martin:2001:AAD

[166] Donald E. K. Martin. Applications of an algorithm for the distribution of the number of successes in fourth-order Markovian sequences. Computational Statistics & Data Analysis, 37(4):405– 418, October 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000226.

Bickel:2001:ADT

[167] David R. Bickel and Dejian Lai. Asymptotic distribution of timeseries intermittency estimates: applications to economic and clinical data. Computational Statistics & Data Analysis, 37(4):419–431, October 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000214.

Goos:2001:ORS

[168] Peter Goos and Martina Vandebroek. D-optimal response surface designs in the presence of random block effects. Computational Statistics & Data Analysis, 37(4):433–453, October 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730100010X.

Liu:2001:DMT

[169] Lon-Mu Liu, Siddhartha Bhattacharyya, Stanley L. Sclove, Rong Chen, and William J. Lattyak. Data mining on time series: an illustration using fast-food restaurant franchise data. Computational Statistics & Data Analysis, 37(4):455–476, October 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000147.

Huang:2001:DFQ

[170] Mei Ling Huang. On a distributionfree quantile estimator. Computational Statistics & Data Analysis, 37(4):477– 486, October 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000202.

Nur:2001:PRN

[171] Darfiana Nur, Rodney C. Wolff, and Kerrie L. Mengersen. Phase randomisation: numerical study of higher cumulants behaviour. Computational Statistics & Data Analysis, 37(4):487– 513, October 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000184.

Anonymous:2001:SSi

[172] Anonymous. SSN-section. Computational Statistics & Data Analysis, 37(4):515–542, October 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947301000792.

Anonymous:2001:AIVc

[173] Anonymous. Author index volume 37 (2001). Computational Statistics & Data Analysis, 37(4):543–544, October 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000809.

Mazucheli:2001:BIP

[174] Josmar Mazucheli, Francisco Louzada-Neto, and Jorge A. Achcar. Bayesian inference for polyhazard models in the presence of covariates. Computational Statistics & Data Analysis, 38(1):1– 14, November 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000275.

Hawkins:2001:DNC

[175] Dollena S. Hawkins, David M. Allen, and Arnold J. Stromberg. Determining the number of components in mixtures of linear models. Computational Statistics & Data Analysis, 38(1):15– 48, November 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000172.

Jeong:2001:BTA

[176] Jinook Jeong and Seoung Chung. Bootstrap tests for autocorrelation. Computational Statistics & Data Analysis, 38(1):49–69, November 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000317.

Ostermark:2001:GMM

[177] Ralf Östermark. Genetic modelling of multivariate EGARCHXprocesses: evidence on the international asset return signal response mechanism. Computational Statistics & Data Analysis, 38(1):71–93, November 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000287.

Wang:2001:DES

[178] P. C. Wang and D. F. Lin. Dispersion effects in signal-response data from fractional factorial experiments. Computational Statistics & Data Analysis, 38(1):95–111, November 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000457.

Anonymous:2001:SSj

[179] Anonymous. SSN-section. Computational Statistics & Data Analysis, 38(1):113–123, November 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016794730100086X.

Mojirsheibani:2001:ICR

[180] Majid Mojirsheibani. An iterated classification rule based on auxiliary pseudo-predictors. Computational Statistics & Data Analysis, 38(2):125– 138, December 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730100024X.

Asparoukhov:2001:CDP

[181] Ognian K. Asparoukhov and Wojtek J. Krzanowski. A comparison of discriminant procedures for binary variables. Computational Statistics & Data Analysis, 38(2):139–160, December 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000329.

Kiang:2001:EKS

[182] Melody Y. Kiang. Extending the Kohonen self-organizing map networks for clustering analysis. Computational Statistics & Data Analysis, 38(2):161– 180, December 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000408.

Wu:2001:SSL

[183] Jianrong Wu and Guoyong Jiang. Small sample likelihood inference for the ratio of means. Computational Statistics & Data Analysis, 38(2):181– 190, December 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000251.

Xie:2001:ZIP

[184] M. Xie, B. He, and T. N. Goh. Zero-inflated Poisson model in statistical process control. Computational Statistics & Data Analysis, 38(2):191– 201, December 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000330.

Headrick:2001:IRT

[185] Todd C. Headrick and Ourania Rotou. An investigation of the rank transformation in multiple regression. Computational Statistics & Data Analysis, 38(2):203–215, December 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000342.

Meulders:2001:PMD

[186] Michel Meulders, Paul De Boeck, and Iven Van Mechelen. Probability matrix decomposition models and main-effects generalized linear models for the analysis of replicated binary associations. Computational Statistics & Data Analysis, 38(2):217–233, December 28, 2001. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000299.

Anonymous:2001:SSk

[187] Anonymous. SSN-section. Computational Statistics & Data Analysis, 38(2):235–248, December 28, 2001. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947301000895.

Luceno:2002:AAC

[188] Alberto Luceño and Jaime Puig-Pey. An accurate algorithm to compute the run length probability distribution, and its convolutions, for a Cusum chart to control normal mean. Computational Statistics & Data Analysis, 38(3):249– 261, January 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000433.

Ahn:2002:NBA

[189] Hongshik Ahn, Hojin Moon, Sunyoung Kim, and Ralph L. Kodell. A Newton-based approach for attributing tumor lethality in animal carcinogenicity studies. Computational Statistics & Data Analysis, 38(3):263– 283, January 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730100041X.

Cappelli:2002:SAG

[190] Carmela Cappelli, Francesco Mola, and Roberta Siciliano. A statistical approach to growing a reliable honest tree. Computational Statistics & Data Analysis, 38(3):285–299, January 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000445.

Howlader:2002:BSE

[191] Hatem A. Howlader and Anwar M. Hossain. Bayesian survival estimation of Pareto distribution of the second kind based on failurecensored data. Computational Statistics & Data Analysis, 38(3):301–314, January 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000391.

Wu:2002:UGA

[192] Berlin Wu and Chih-Li Chang. Using genetic algorithms to parameters (d, r) estimation for threshold autoregressive models. Computational Statistics & Data Analysis, 38(3):315– 330, January 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000305.

Childs:2002:SAM

[193] Aaron Childs and N. Balakrishnan. Series approximations for moments of order statistics using MAPLE. Computational Statistics & Data Analysis, 38(3):331–347, January 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000421.

Anonymous:2002:SSa

[194] Anonymous. SSN-section. Computational Statistics & Data Analysis, 38(3):349–364, January 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947301001013.

Barra:2002:P

[195] Simone Barra and Agostino Di Ciaccio. Preface. Computational Statistics & Data Analysis, 38(4):365–366, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000640.

Friedman:2002:SGB

[196] Jerome H. Friedman. Stochastic gradient boosting. Computational Statistics & Data Analysis, 38(4):367–378, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000652.

Ridgeway:2002:LLB

[197] Greg Ridgeway. Looking for lumps: boosting and bagging for density estimation. Computational Statistics & Data Analysis, 38(4):379–392, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000664.

Drucker:2002:EPE

[198] Harris Drucker. Effect of pruning and early stopping on performance of a boosting ensemble. Computational Statistics & Data Analysis, 38(4):393– 406, February 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000676.

Borra:2002:INR

[199] Simone Borra and Agostino Di Ciaccio. Improving nonparametric regression methods by bagging and boosting. Computational Statistics & Data Analysis, 38(4):407–420, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000688.

Evgeniou:2002:RSL

[200] Theodoros Evgeniou, Tomaso Poggio, Massimiliano Pontil, and Alessandro Verri. Regularization and statistical learning theory for data analysis. Computational Statistics & Data Analysis, 38(4):421–432, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730100069X.

Peracchi:2002:ECQ

[201] Franco Peracchi. On estimating conditional quantiles and distribution functions. Computational Statistics & Data Analysis, 38(4):433–447, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000706.

Borgelt:2002:LID

[202] Christian Borgelt and Rudolf Kruse. Learning from imprecise data: possibilistic graphical models. Computational Statistics & Data Analysis, 38(4):449–463, February 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947301000718.

Saporta:2002:DFD

[203] Gilbert Saporta. Data fusion and data grafting. Computational Statistics & Data Analysis, 38(4):465–473, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730100072X.

Denison:2002:BPM

[204] D. G. T. Denison, N. M. Adams, C. C. Holmes, and D. J. Hand. Bayesian partition modelling. Computational Statistics & Data Analysis, 38(4):475– 485, February 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000731.

Conversano:2002:GAM

[205] Claudio Conversano, Roberta Siciliano, and Francesco Mola. Generalized additive multi-mixture model for data mining. Computational Statistics & Data Analysis, 38(4):487–500, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000743.

Coppi:2002:TFD

[206] Renato Coppi. A theoretical framework for data mining: the “Informational Paradigm”. Computational Statistics & Data Analysis, 38(4):501– 515, February 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000755.

Larsen:2002:WLW

[207] Jan Larsen, Lars Kai Hansen, Anna Szymkowiak Have, Torben Christiansen, and Thomas Kolenda. Webmining: learning from the world wide web. Computational Statistics & Data Analysis, 38(4):517–532, February 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000767.

Giudici:2002:DMA

[208] Paolo Giudici and Gianluca Passerone. Data mining of association structures to model consumer behaviour. Computational Statistics & Data Analysis, 38(4):533–541, February 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947301000779.

Anonymous:2002:SSb

[209] Anonymous. SSN-section. Computational Statistics & Data Analysis, 38(4):543–557, February 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947302000038.

Anonymous:2002:AIa

[210] Anonymous. Author index. Computational Statistics & Data Analysis, 38(4):559–560, February 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016794730200004X.

Allison:2002:MMA

[211] David B. Allison, Gary L. Gadbury, Moonseong Heo, José R. Fernández, Cheol-Koo Lee, Tomas A. Prolla, and Richard Weindruch. A mixture model approach for the analysis of microarray gene expression data. Computational Statistics & Data Analysis, 39 (1):1–20, March 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000469.

Park:2002:RAF

[212] Chanseok Park, Ayanendranath Basu, and Bruce G. Lindsay. The residual adjustment function and weighted likelihood: a graphical interpretation of robustness of minimum disparity estimators. Computational Statistics & Data Analysis, 39(1):21– 33, March 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000470.

Shi:2002:EAD

[213] Mingren Shi and Mark A. Lukas. An L_{1} estimation algorithm with degeneracy and linear constraints. Computational Statistics & Data Analysis, 39 (1):35–55, March 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000494.

Bull:2002:MSF

[214] Shelley B. Bull, Carmen Mak, and Celia M. T. Greenwood. A modified score function estimator for multinomial logistic regression in small samples. Computational Statistics & Data Analysis, 39(1):57–74, March 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000482.

Modarres:2002:ENE

[215] Reza Modarres. Efficient nonparametric estimation of a distribution function. Computational Statistics & Data Analysis, 39(1):75–95, March 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000354.

Raqab:2002:BLI

[216] Mohammad Z. Raqab, Elies Kouider, and Qasim M. Al-Shboul. Best linear invariant estimators using ranked set sampling procedure: comparative study. Computational Statistics & Data Analysis, 39(1):97–105, March 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000512.

Anonymous:2002:SSc

[217] Anonymous. SSN-section. Computational Statistics & Data Analysis, 39 (1):107–126, March 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200035X.

Agresti:2002:MRM

[218] Alan Agresti and Brian Caffo. Measures of relative model fit. Computational Statistics & Data Analysis, 39 (2):127–136, April 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000548.

Altaleb:2002:BAL

[219] Anas Altaleb and Didier Chauveau. Bayesian analysis of the logit model and comparison of two Metropolis– Hastings strategies. Computational Statistics & Data Analysis, 39(2):137– 152, April 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730100055X.

Bickel:2002:REM

[220] David R. Bickel. Robust estimators of the mode and skewness of continuous data. Computational Statistics & Data Analysis, 39(2):153– 163, April 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000573.

Sain:2002:MLA

[221] Stephan R. Sain. Multivariate locally adaptive density estimation. Computational Statistics & Data Analysis, 39 (2):165–186, April 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000536.

Concordet:2002:SPM

[222] Didier Concordet and Olivier G. Nunez. A simulated pseudo-maximum likelihood estimator for nonlinear mixed models. Computational Statistics & Data Analysis, 39(2):187– 201, April 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000524.

Browne:2002:BLM

[223] William J. Browne, David Draper, Harvey Goldstein, and Jon Rasbash. Bayesian and likelihood methods for fitting multilevel models with complex level-1 variation. Computational Statistics & Data Analysis, 39(2):203– 225, April 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000585.

Lawson:2002:RAE

[224] John Lawson. Regression analysis of experiments with complex confounding patterns guided by the alias matrix. Computational Statistics & Data Analysis, 39(2):227– 241, April 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000561.

Anonymous:2002:SSd

[225] Anonymous. SSN-section. Computational Statistics & Data Analysis, 39 (2):243–260, April 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000828.

Booth:2002:USM

[226] James G. Booth and Brian S. Caffo. Unequal sampling for Monte Carlo EM algorithms. Computational Statistics & Data Analysis, 39(3):261– 270, May 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000627.

Borkowf:2002:CNA

[227] Craig B. Borkowf. Computing the nonnull asymptotic variance and the asymptotic relative efficiency of Spearman’s rank correlation. Computational Statistics & Data Analysis, 39 (3):271–286, May 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000810.

Dean:2002:CGS

[228] A. M. Dean and S. M. Lewis. Comparison of group screening strategies for factorial experiments. Computational Statistics & Data Analysis, 39 (3):287–297, May 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000822.

Fang:2002:CRR

[229] Yue Fang. The compass rose and random walk tests. Computational Statistics & Data Analysis, 39(3):299– 310, May 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000639.

Liao:2002:BMC

[230] Tim Futing Liao. Bayesian model comparison in generalized linear models across multiple groups. Computational Statistics & Data Analysis, 39 (3):311–327, May 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000615.

Lee:2002:CRM

[231] A. J. Lee, S. O. Nyangoma, and G. A. F. Seber. Confidence regions for multinomial parameters. Computational Statistics & Data Analysis, 39 (3):329–342, May 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000834.

Jang:2002:GME

[232] Dae-Heung Jang. A graphical method for evaluating slope-rotatability in axial directions for second order response surface designs. Computational Statistics & Data Analysis, 39 (3):343–349, May 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000597.

Karioti:2002:DOS

[233] V. Karioti and C. Caroni. Detecting outlying series in sets of short time series. Computational Statistics & Data Analysis, 39(3):351– 364, May 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000780.

Anonymous:2002:SSe

[234] Anonymous. SSN-section. Computational Statistics & Data Analysis, 39 (3):365–370, May 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000993.

Ng:2002:EPP

[235] H. K. T. Ng, P. S. Chan, and N. Balakrishnan. Estimation of parameters from progressively censored data using EM algorithm. Computational Statistics & Data Analysis, 39(4):371– 386, June 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000913.

Oh:2002:BAR

[236] Man-Suk Oh, Dong Wan Shin, and Han Joon Kim. Bayesian analysis of regression models with spatially correlated errors and missing observations. Computational Statistics & Data Analysis, 39(4):387– 400, June 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000846.

Rayner:2002:WQB

[237] G. D. Rayner and H. L. MacGillivray. Weighted quantile-based estimation for a class of transformation distributions. Computational Statistics & Data Analysis, 39(4):401– 433, June 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000901.

Abramovich:2002:EBA

[238] Felix Abramovich, Panagiotis Besbeas, and Theofanis Sapatinas. Empirical Bayes approach to block wavelet function estimation. Computational Statistics & Data Analysis, 39(4):435– 451, June 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000858.

El-Aroui:2002:UPT

[239] Mhamed-Ali El-Aroui and Jean Diebolt. On the use of the peaks over thresholds method for estimating out-of-sample quantiles. Computational Statistics & Data Analysis, 39 (4):453–475, June 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000871.

Loschi:2002:AIS

[240] R. H. Loschi and F. R. B. Cruz. An analysis of the influence of some prior specifications in the identification of change points via product partition model. Computational Statistics & Data Analysis, 39(4):477– 501, June 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000950.

Marazzi:2002:BTR

[241] A. Marazzi. Bootstrap tests for robust means of asymmetric distributions with unequal shapes. Computational Statistics & Data Analysis, 39 (4):503–528, June 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000883.

Modarres:2002:EUQ

[242] Reza Modarres, Tapan K. Nayak, and Joseph L. Gastwirth. Estimation of upper quantiles under model and parameter uncertainty. Computational Statistics & Data Analysis, 39 (4):529–554, June 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000949.

Anonymous:2002:SSf

[243] Anonymous. SSN-section. Computational Statistics & Data Analysis, 39 (4):555–573, June 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001263.

Anonymous:2002:AIb

[244] Anonymous. Author index. Computational Statistics & Data Analysis, 39 (4):575–576, June 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001135.

Jandhyala:2002:DEA

[245] Venkata K. Jandhyala, Stergios B. Fotopoulos, and Douglas M. Hawkins. Detection and estimation of abrupt changes in the variability of a process. Computational Statistics & Data Analysis, 40(1):1–19, July 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301001086.

Li:2002:ADU

[246] Baibing Li and Elaine B. Martin. An approximation to the F distribution using the chi-square distribution. Computational Statistics & Data Analysis, 40(1):21–26, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301000974.

Buttrey:2002:UNN

[247] Samuel E. Buttrey and Ciril Karo. Using k-nearest-neighbor classification in the leaves of a tree. Computational Statistics & Data Analysis, 40 (1):27–37, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000986.

Dittrich:2002:MDP

[248] R. Dittrich, R. Hatzinger, and W. Katzenbeisser. Modelling dependencies in paired comparison data: a log-linear approach. Computational Statistics & Data Analysis, 40(1):39– 57, July 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301001062.

Huang:2002:MFI

[249] Yangxin Huang, Peter Harris, Simon P. J. Kirby, and John C. Dearden. A modified Fieller interval for the interval estimation of effective doses for a logistic dose–response curve. Computational Statistics & Data Analysis, 40(1):59–74, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301001025.

Jansakul:2002:STZ

[250] N. Jansakul and J. P. Hinde. Score tests for zero-inflated Poisson models. Computational Statistics & Data Analysis, 40(1):75–96, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301001049.

Kim:2002:CIE

[251] Sung-Ho Kim. Calibrated initials for an EM applied to recursive models of categorical variables. Computational Statistics & Data Analysis, 40 (1):97–110, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301001050.

Li:2002:RTA

[252] Ju Li and Daniel Zelterman. Rank tests of association for exchangeable paired data. Computational Statistics & Data Analysis, 40(1):111– 129, July 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000962.

Maharaj:2002:CNS

[253] Elizabeth Ann Maharaj. Comparison of non-stationary time series in the frequency domain. Computational Statistics & Data Analysis, 40 (1):131–141, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301001001.

Pan:2002:ACI

[254] Wei Pan. Approximate confidence intervals for one proportion and difference of two proportions. Computational Statistics & Data Analysis, 40 (1):143–157, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301001074.

Viljoen:2002:IMD

[255] H. Viljoen and J. H. Venter. Identifying multivariate discordant observations: a computer-intensive approach. Computational Statistics & Data Analysis, 40 (1):159–172, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301001037.

Vu:2002:HME

[256] Hien T. V. Vu and Matthew W. Knuiman. A hybrid ML-EM algorithm for calculation of maximum likelihood estimates in semiparametric shared frailty models. Computational Statistics & Data Analysis, 40 (1):173–187, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000998.

Noble:2002:SRJ

[257] Chris Noble and Steve Sugden. Stochastic recurrences of jackpot keno. Computational Statistics & Data Analysis, 40(1):189–205, July 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000925.

Anonymous:2002:SSg

[258] Anonymous. SSN-section. Computational Statistics & Data Analysis, 40 (1):207–222, July 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001561.

Aloupis:2002:LBC

[259] Greg Aloupis, Carmen Cortés, Francisco Gómez, Michael Soss, and Godfried Toussaint. Lower bounds for computing statistical depth. Computational Statistics & Data Analysis, 40 (2):223–229, August 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000324.

Arnold:2002:ENC

[260] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Exact and near compatibility of discrete conditional distributions. Computational Statistics & Data Analysis, 40(2):231– 252, August 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301001116.

Hawkins:2002:FAR

[261] Douglas M. Hawkins and Xiangrong Yin. A faster algorithm for ridge regression of reduced rank data. Computational Statistics & Data Analysis, 40 (2):253–262, August 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000348.

Hsieh:2002:EFS

[262] Ping-Hung Hsieh. An exploratory first step in teletraffic data modeling: evaluation of long-run performance of parameter estimators. Computational Statistics & Data Analysis, 40 (2):263–283, August 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000336.

Pupko:2002:FCM

[263] T. Pupko and D. Graur. Fast computation of maximum likelihood trees by numerical approximation of amino acid replacement probabilities. Computational Statistics & Data Analysis, 40 (2):285–291, August 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000087.

Racine:2002:PDK

[264] Jeff Racine. Parallel distributed kernel estimation. Computational Statistics & Data Analysis, 40(2):293–302, August 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301001098.

Yu:2002:QRU

[265] Keming Yu. Quantile regression using RJMCMC algorithm. Computational Statistics & Data Analysis, 40 (2):303–315, August 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947301000937.

Chang:2002:WEB

[266] Woojin Chang and Brani Vidakovic. Wavelet estimation of a base-line signal from repeated noisy measurements by vertical block shrinkage. Computational Statistics & Data Analysis, 40 (2):317–328, August 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000531.

Mansson:2002:MOY

[267] Ralph Mansson and Philip Prescott. Missing observations in Youden square designs. Computational Statistics & Data Analysis, 40(2):329–338, August 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000099.

Andres:2002:ATC

[268] A. Martín Andrés, M. J. Sánchez Quevedo, and A. Silva Mato. Asymptotical tests in 2 × 2 comparative trials: (unconditional approach). Computational Statistics & Data Analysis, 40 (2):339–354, August 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000051.

Narula:2002:SAP

[269] Subhash C. Narula and John F. Wellington. Sensitivity analysis for predictor variables in the MSAE regression. Computational Statistics & Data Analysis, 40(2):355–373, August 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947301001104.

Wu:2002:TSS

[270] Shu-Fei Wu and Mao-Yi Cheng. A two-sided sequential screening procedure based on individual misclassification error. Computational Statistics & Data Analysis, 40(2):375–391, August 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000063.

Beran:2002:SMS

[271] Jan Beran and Yuanhua Feng. SEMI-FAR models — a semiparametric approach to modelling trends, long-range dependence and nonstationarity. Computational Statistics & Data Analysis, 40(2):393–419, August 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000075.

Anonymous:2002:SSh

[272] Anonymous. SSN-section. Computational Statistics & Data Analysis, 40 (2):421–433, August 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002189.

Anonymous:2002:EBa

[273] Anonymous. Editorial board. Computational Statistics & Data Analysis, 40 (3):iv–v, September 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002396.

Bonett:2002:TNH

[274] Douglas G. Bonett and Edith Seier. A test of normality with high uniform power. Computational Statistics & Data Analysis, 40(3):435–445, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000749.

Timmerman:2002:TWC

[275] Marieke E. Timmerman and Henk A. L. Kiers. Three-way component analysis with smoothness constraints. Computational Statistics & Data Analysis, 40(3):447–470, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000592.

Li:2002:PCA

[276] Baibing Li, Elaine B. Martin, and A. Julian Morris. On principal component analysis in L_{1}. Computational Statistics & Data Analysis, 40(3):471– 474, September 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000762.

Trinca:2002:EAA

[277] L. A. Trinca and S. G. Gilmour. Erratum to “An algorithm for arranging response surface designs in small blocks” [comput. statist. data anal. 33 (2000) 25–43]. Computational Statistics & Data Analysis, 40 (3):475, September 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0925527302001548. See

[24].

Fok:2002:OLA

[278] Dennis Fok and Philip Hans Franses. Ordered logit analysis for selectively sampled data. Computational Statistics & Data Analysis, 40(3):477–497, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000634.

Gediga:2002:AQS

[279] Günther Gediga and Ivo Düntsch. Approximation quality for sorting rules. Computational Statistics & Data Analysis, 40(3):499–526, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000658.

Huang:2002:LSP

[280] Yangxin Huang. On large-sample parametric approaches for interval estimation of the ED_{90}. Computational Statistics & Data Analysis, 40(3):527– 537, September 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000646.

Song:2002:BMS

[281] Xin-Yuan Song and Sik-Yum Lee. A Bayesian model selection method with applications. Computational Statistics & Data Analysis, 40(3):539–557, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000737.

Tack:2002:AAO

[282] Lieven Tack and Martina Vandebroek. An adjustment algorithm for optimal run orders in design of experiments. Computational Statistics & Data Analysis, 40(3):559–577, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000567.

Dippon:2002:SAC

[283] J. Dippon, P. Fritz, and M. Kohler. A statistical approach to case based reasoning, with application to breast cancer data. Computational Statistics & Data Analysis, 40(3):579–602, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000580.

Duchateau:2002:SFM

[284] Luc Duchateau, Paul Janssen, Patrick Lindsey, Catherine Legrand, Rosemary Nguti, and Richard Sylvester. The shared frailty model and the power for heterogeneity tests in multicenter trials. Computational Statistics & Data Analysis, 40(3):603–620, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000579.

Franses:2002:DTE

[285] Philip Hans Franses and Paul de Bruin. On data transformations and evidence of nonlinearity. Computational Statistics & Data Analysis, 40(3):621–632, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000543.

Rotondi:2002:IPD

[286] R. Rotondi. On the influence of the proposal distributions on a reversible jump MCMC algorithm applied to the detection of multiple change-points. Computational Statistics & Data Analysis, 40(3):633–653, September 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000555.

Anonymous:2002:SSPa

[287] Anonymous. SSN-section, PS-file follows with cover proofs. Computational Statistics & Data Analysis, 40(3):655– 664, September 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002591.

Anonymous:2002:EBb

[288] Anonymous. Editorial board. Computational Statistics & Data Analysis, 40 (4):iv–v, October 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002682.

Buhlmann:2002:ANG

[289] Peter Bühlmann and Alexander J. Mc-Neil. An algorithm for nonparametric GARCH modelling. Computational Statistics & Data Analysis, 40(4):665– 683, October 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000804.

Headrick:2002:FFO

[290] Todd C. Headrick. Fast fifth-order polynomial transforms for generating univariate and multivariate nonnormal distributions. Computational Statistics & Data Analysis, 40(4):685– 711, October 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000725.

McCullough:2002:ASP

[291] B. D. McCullough and Berry Wilson. On the accuracy of statistical procedures in Microsoft Excel 2000 and Excel XP. Computational Statistics & Data Analysis, 40(4):713– 721, October 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000956.

Hennig:2002:VVC

[292] Christian Hennig and Norbert Christlieb. Validating visual clusters in large datasets: fixed point clusters of spectral features. Computational Statistics & Data Analysis, 40(4):723–739, October 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000774.

Rousseeuw:2002:REV

[293] Peter J. Rousseeuw and Sabine Verboven. Robust estimation in very small samples. Computational Statistics & Data Analysis, 40(4):741–758, October 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000786.

Xiang:2002:IDG

[294] Liming Xiang, Siu-Keung Tse, and Andy H. Lee. Influence diagnostics for generalized linear mixed models: applications to clustered data. Computational Statistics & Data Analysis, 40(4):759–774, October 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000750.

Yi:2002:IHO

[295] Grace Y. Yi, Jianrong Wu, and Ying Liu. Implementation of higher-order asymptotics to S-plus. Computational Statistics & Data Analysis, 40(4):775– 800, October 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000968.

Cribari-Neto:2002:IEC

[296] Francisco Cribari-Neto, Alejandro C. Frery, and Michel F. Silva. Improved estimation of clutter properties in speckled imagery. Computational Statistics & Data Analysis, 40(4):801– 824, October 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001020.

Anonymous:2002:SSi

[297] Anonymous. SSN-section. Computational Statistics & Data Analysis, 40(4):825–844, October 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947302002840.

Anonymous:2002:AIc

[298] Anonymous. Author index. Computational Statistics & Data Analysis, 40(4):845–846, October 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947302002785.

Anonymous:2002:EBc

[299] Anonymous. Editorial board. Computational Statistics & Data Analysis, 41(1):viii–ix, November 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947302003110.

Barlow:2002:P

[300] Jesse L. Barlow, Michael W. Berry, Axel Ruhe, and Hongyuan Zha.

Preface. Computational Statistics & Data Analysis, 41(1):1–2, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002049.

Kukush:2002:CFM

[301] A. Kukush, I. Markovsky, and S. Van Huffel. Consistent fundamental matrix estimation in a quadratic measurement error model arising in motion analysis. Computational Statistics & Data Analysis, 41(1):3–18, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000683.

He:2002:WDC

[302] Xiaofeng He, Hongyuan Zha, Chris H. Q. Ding, and Horst D. Simon. Web document clustering using hyperlink structures. Computational Statistics & Data Analysis, 41(1):19– 45, November 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000701.

Genton:2002:NVC

[303] Marc G. Genton and David J. Gorsich. Nonparametric variogram and covariogram estimation with Fourier– Bessel matrices. Computational Statistics & Data Analysis, 41(1):47–57, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000622.

Dhillon:2002:CVH

[304] Inderjit S. Dhillon, Dharmendra S. Modha, and W. Scott Spangler. Class visualization of high-dimensional data with applications. Computational Statistics & Data Analysis, 41(1):59– 90, November 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001445.

Fodor:2002:DRT

[305] Imola K. Fodor and Chandrika Kamath. Dimension reduction techniques and the classification of bent double galaxies. Computational Statistics & Data Analysis, 41(1):91–122, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000610.

Chang:2002:NLA

[306] Xiao-Wen Chang and Christopher C. Paige. Numerical linear algebra in the integrity theory of the global positioning system. Computational Statistics & Data Analysis, 41(1):123–142, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000609.

Malone:2002:BIC

[307] Samuel W. Malone, Pablo Tarazaga, and Michael W. Trosset. Better initial configurations for metric multidimensional scaling. Computational Statistics & Data Analysis, 41(1):143– 156, November 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001457.

Kiers:2002:SAL

[308] Henk A. L. Kiers. Setting up alternating least squares and iterative majorization algorithms for solving various matrix optimization problems. Computational Statistics & Data Analysis, 41(1):157–170, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001421.

Georgiou:2002:GMS

[309] S. Georgiou, C. Koukouvinos, and S. Stylianou. On good matrices, skew Hadamard matrices and optimal designs. Computational Statistics & Data Analysis, 41(1):171–184, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000671.

Doornik:2002:NSC

[310] Jurgen A. Doornik and R. J. O’Brien. Numerically stable cointegration analysis. Computational Statistics & Data Analysis, 41(1):185–193, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001433.

Erbay:2002:MGS

[311] Hasan Erbay, Jesse L. Barlow, and Zhenyue Zhang. A modified Gram–Schmidt-based downdating technique for ULV decompositions with applications to recursive TLS problems. Computational Statistics & Data Analysis, 41(1):195–209, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000695.

Foschi:2002:SUR

[312] Paolo Foschi and Erricos J. Kontoghiorghes. Seemingly unrelated regression model with unequal size observations: computational aspects. Computational Statistics & Data Analysis, 41(1):211–229, November 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001469.

Vigneau:2002:NAL

[313] Evelyne Vigneau and El Mostafa Qannari. A new algorithm for latent root regression analysis. Computational Statistics & Data Analysis, 41(1):231– 242, November 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302000713.

Anonymous:2002:SSPb

[314] Anonymous. SSN-section, PS-file follows with cover proofs. Computational Statistics & Data Analysis, 41(1):243– 254, November 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003286.

Anonymous:2002:EBd

[315] Anonymous. Editorial board. Computational Statistics & Data Analysis, 41 (2):iv–v, December 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003341.

Anonymous:2002:EBe

[316] Anonymous. Editorial board. Computational Statistics & Data Analysis, 41 (2):iv–v, December 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003341.

Lindstrom:2002:BEF

[317] Mary J. Lindstrom. Bayesian estimation of free-knot splines using reversible jumps. Computational Statistics & Data Analysis, 41(2):255–269, December 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200066X.

Chou:2002:MMS

[318] Chih-Ping Chou and P. M. Bentler. Model modification in structural equation modeling by imposing constraints. Computational Statistics & Data Analysis, 41(2):271–287, December 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200097X.

Holmstrom:2002:MIA

[319] Lasse Holmström and Panu Erästö. Making inferences about past environmental change using smoothing in multiple time scales. Computational Statistics & Data Analysis, 41(2):289– 309, December 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000798.

DeIaco:2002:STV

[320] S. De Iaco, D. E. Myers, and D. Posa. Space–time variograms and a functional form for total air pollution measurements. Computational Statistics & Data Analysis, 41(2):311–328, December 28, 2002. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302000816.

Anonymous:2002:SSPc

[321] Anonymous. SSN-section, PS-file follows with cover proofs. Computational Statistics & Data Analysis, 41(2):329– 348, December 28, 2002. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200347X.

Anonymous:2003:EBa

[322] Anonymous. Editorial board. Computational Statistics & Data Analysis, 41 (3–4):iv–v, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003560.

Bohning:2003:ERD

[323] Dankmar Böhning and Wilfried Seidel. Editorial: recent developments in mixture models. Computational Statistics & Data Analysis, 41(3–4): 349–357, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001615.

Yau:2003:FMR

[324] Kelvin K. W. Yau, Andy H. Lee, and Angus S. K. Ng. Finite mixture regression model with random effects: application to neonatal hospital length of stay. Computational Statistics & Data Analysis, 41(3–4): 359–366, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001809.

Susko:2003:WTH

[325] Edward Susko. Weighted tests of homogeneity for testing the number of components in a mixture. Computational Statistics & Data Analysis, 41(3–4): 367–378, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001627.

McLachlan:2003:MHD

[326] G. J. McLachlan, D. Peel, and R. W. Bean. Modelling highdimensional data by mixtures of factor analyzers. Computational Statistics & Data Analysis, 41(3–4):379– 388, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001834.

Mao:2003:TDH

[327] Chang Xuan Mao and Bruce G. Lindsay. Tests and diagnostics for heterogeneity in the species problem. Computational Statistics & Data Analysis, 41(3–4):389–398, January 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001640.

Berchtold:2003:MTD

[328] André Berchtold. Mixture transition distribution (MTD) modeling of heteroscedastic time series. Computational Statistics & Data Analysis, 41(3–4): 399–411, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001913.

Miloslavsky:2003:FMU

[329] Maja Miloslavsky and Mark J. van der Laan. Fitting of mixtures with unspecified number of components using cross validation distance estimate. Computational Statistics & Data Analysis, 41(3– 4):413–428, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001664.

Hunt:2003:MMC

[330] Lynette Hunt and Murray Jorgensen. Mixture model clustering for mixed data with missing information. Computational Statistics & Data Analysis, 41(3–4):429–440, January 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001901.

Schlattmann:2003:ENC

[331] Peter Schlattmann. Estimating the number of components in a finite mixture model: the special case of homogeneity. Computational Statistics & Data Analysis, 41(3–4):441– 451, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001731.

vandeGeer:2003:ATM

[332] Sara van de Geer. Asymptotic theory for maximum likelihood in nonparametric mixture models. Computational Statistics & Data Analysis, 41(3–4): 453–464, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001883.

Pons:2003:ETL

[333] O. Pons and M. Lemdani. Estimation and test in long-term survival mixture models. Computational Statistics & Data Analysis, 41(3–4):465– 479, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001780.

Peng:2003:FSC

[334] Yingwei Peng. Fitting semiparametric cure models. Computational Statistics & Data Analysis, 41(3–4):481– 490, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001846.

Dalrymple:2003:FMZ

[335] M. L. Dalrymple, I. L. Hudson, and R. P. K. Ford. Finite mixture, zeroinflated Poisson and hurdle models with application to SIDS. Computational Statistics & Data Analysis, 41(3– 4):491–504, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001871.

Mooney:2003:FMM

[336] Jennifer A. Mooney, Peter J. Helms, and Ian T. Jolliffe. Fitting mixtures of von Mises distributions: a case study involving sudden infant death syndrome. Computational Statistics & Data Analysis, 41(3–4):505– 513, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001810.

Morbiducci:2003:CCI

[337] Marta Morbiducci, Alessandra Nardi, and Carla Rossi. Classification of “cured” individuals in survival analysis: the mixture approach to the diagnostic–prognostic problem. Computational Statistics & Data Analysis, 41(3–4):515–529, January 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001858.

Vermunt:2003:LCM

[338] Jeroen K. Vermunt and Jay Magidson. Latent class models for classification. Computational Statistics & Data Analysis, 41(3–4):531–537, January 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001792.

Croft:2003:DMB

[339] J. Croft and J. Q. Smith. Discrete mixtures in Bayesian networks with hidden variables: a latent time budget example. Computational Statistics & Data Analysis, 41(3–4):539– 547, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001676.

Formann:2003:LCM

[340] Anton K. Formann. Latent class model diagnostics — a review and some proposals. Computational Statistics & Data Analysis, 41(3–4):549– 559, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200186X.

Biernacki:2003:CSV

[341] Christophe Biernacki, Gilles Celeux, and Gérard Govaert. Choosing starting values for the EM algorithm for getting the highest likelihood in multivariate Gaussian mixture models. Computational Statistics & Data Analysis, 41(3– 4):561–575, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001639.

Karlis:2003:CIV

[342] Dimitris Karlis and Evdokia Xekalaki. Choosing initial values for the EM algorithm for finite mixtures. Computational Statistics & Data Analysis, 41(3– 4):577–590, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001779.

Nityasuddhi:2003:APE

[343] Dechavudh Nityasuddhi and Dankmar Böhning. Asymptotic properties of the EM algorithm estimate for normal mixture models with component specific variances. Computational Statistics & Data Analysis, 41(3–4):591– 601, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001767.

Grim:2003:TMD

[344] J. Grim and M. Haindl. Texture modelling by discrete distribution mixtures. Computational Statistics & Data Analysis, 41(3–4):603– 615, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001743.

Biggeri:2003:TNP

[345] A. Biggeri, E. Dreassi, C. Lagazio, and D. Böhning. A transitional non-parametric maximum pseudolikelihood estimator for disease mapping. Computational Statistics & Data Analysis, 41(3–4):617–629, January 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001895.

Viwatwongkasem:2003:CRD

[346] Chukiat Viwatwongkasem and Walailuck Böhning. A comparison of risk difference estimators in multi-center studies under baseline-risk heterogeneity. Computational Statistics & Data Analysis, 41(3–4):631–644, January 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001755.

Murphy:2003:MDB

[347] Thomas Brendan Murphy and Donal Martin. Mixtures of distance-based models for ranking data. Computational Statistics & Data Analysis, 41(3– 4):645–655, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001652.

Anonymous:2003:SS

[348] Anonymous. SSN-section. Computational Statistics & Data Analysis, 41(3– 4):657–676, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302003699.

Anonymous:2003:AIa

[349] Anonymous. Author index. Computational Statistics & Data Analysis, 41(3– 4):677–678, January 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302003614.

Anonymous:2003:EBb

[350] Anonymous. Editorial board. Computational Statistics & Data Analysis, 42(1–2):viii–ix, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947302003754.

Masse:2003:MCS

[351] Jean-Claude Massé and Jean-François Plante. A Monte Carlo study of the accuracy and robustness of ten bivariate location estimators. Computational Statistics & Data Analysis, 42(1–2): 1–26, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001032.

Albers:2003:EDA

[352] C. J. Albers and W. Schaafsma. Estimating a density by adapting an initial guess. Computational Statistics & Data Analysis, 42(1–2):27–36, February 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200124X.

Angers:2003:BAZ

[353] Jean-François Angers and Atanu Biswas. A Bayesian analysis of zeroinflated generalized Poisson model. Computational Statistics & Data Analysis, 42(1–2):37–46, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947302001548.

DUrso:2003:LRA

[354] Pierpaolo D’Urso. Linear regression analysis for fuzzy/crisp input and fuzzy/crisp output data. Computational Statistics & Data Analysis, 42(1– 2):47–72, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001172.

DeGooijer:2003:MVN

[355] Jan G. De Gooijer and Bonnie K. Ray. Modeling vector nonlinear time series using POLYMARS. Computational Statistics & Data Analysis, 42(1–2): 73–90, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001238.

Hall:2003:BPL

[356] Charles B. Hall, Jun Ying, Lynn Kuo, and Richard B. Lipton. Bayesian and profile likelihood change point methods for modeling cognitive function over time. Computational Statistics & Data Analysis, 42(1–2):91– 109, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001482.

Hazan:2003:RME

[357] Alon Hazan, Zinoviy Landsman, and Udi E. Makov. Robustness via a mixture of exponential power distributions. Computational Statistics & Data Analysis, 42(1–2):111– 121, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001536.

Lee:2003:ESR

[358] Ching-Yung Lee, Shen-Ming Lee, and Mei-Jih Gee. Estimating survival rates using an extended Ricker’s tworelease method. Computational Statistics & Data Analysis, 42(1–2):123– 137, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001500.

Lee:2003:SPS

[359] Thomas C. M. Lee. Smoothing parameter selection for smoothing splines: a simulation study. Computational Statistics & Data Analysis, 42(1–2):

139–148, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001597.

Mahmoud:2003:OSI

[360] M. A. W. Mahmoud, K. S. Sultan, and S. M. Amer. Order statistics from inverse Weibull distribution and associated inference. Computational Statistics & Data Analysis, 42(1–2): 149–163, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001512.

Ortega:2003:IDG

[361] Edwin M. M. Ortega, Heleno Bolfarine, and Gilberto A. Paula. Influence diagnostics in generalized log-gamma regression models. Computational Statistics & Data Analysis, 42(1–2):165– 186, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001044.

Peng:2003:EBD

[362] Yingwei Peng. Estimating baseline distribution in proportional hazards cure models. Computational Statistics & Data Analysis, 42(1–2):187– 201, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001585.

Wu:2003:FER

[363] Hsien-Chung Wu. Fuzzy estimates of regression parameters in linear regression models for imprecise input and output data. Computational Statistics & Data Analysis, 42(1–2):203– 217, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001160.

Zhao:2003:JPT

[364] Shan Zhao and G. W. Wei. Jump process for the trend estimation of time series. Computational Statistics & Data Analysis, 42(1–2):219– 241, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001251.

Boracchi:2003:JMC

[365] Patrizia Boracchi, Elia Biganzoli, and Ettore Marubini. Joint modelling of cause-specific hazard functions with cubic splines: an application to a large series of breast cancer patients. Computational Statistics & Data Analysis, 42(1–2):243– 262, February 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001226.

Anonymous:2003:SSPa

[366] Anonymous. SSN-section, PS-file follows. Computational Statistics & Data Analysis, 42(1–2):263–276, February 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000045.

Anonymous:2003:EBc

[367] Anonymous. Editorial board. Computational Statistics & Data Analysis, 42(3):x–xi, March 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000136.

Belsley:2003:E

[368] David A. Belsley and Erricos John Kontoghiorghes. Editorial. Computational Statistics & Data Analysis, 42 (3):277–278, March 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200350X.

Pollock:2003:IFS

[369] D. S. G. Pollock. Improved frequency selective filters. Computational Statistics & Data Analysis, 42 (3):279–297, March 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002281.

Gilli:2003:GOH

[370] M. Gilli and P. Winker. A global optimization heuristic for estimating agent based models. Computational Statistics & Data Analysis, 42(3):299– 312, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002141.

Paolella:2003:CMR

[371] Marc S. Paolella. Computing moments of ratios of quadratic forms in normal variables. Computational Statistics & Data Analysis, 42(3):313– 331, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730200213X.

Doornik:2003:CAM

[372] Jurgen A. Doornik and Marius Ooms. Computational aspects of maximum likelihood estimation of autoregressive fractionally integrated moving average models. Computational Statistics & Data Analysis, 42(3):333– 348, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002128.

Skouras:2003:ACE

[373] Spyros Skouras. An algorithm for computing estimators that optimize step functions. Computational Statistics & Data Analysis, 42(3):349– 361, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002104.

Orbe:2003:AET

[374] Susan Orbe, Eva Ferreira, and Juan Rodriguez-Poo. An algorithm to estimate time-varying parameter SURE models under different types of restriction. Computational Statistics & Data Analysis, 42(3):363– 383, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002293.

Capobianco:2003:IMC

[375] Enrico Capobianco. Independent multiresolution component analysis and matching pursuit. Computational Statistics & Data Analysis, 42(3):385– 402, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002177.

Munkin:2003:MSE

[376] Murat K. Munkin. The MCMC and SML estimation of a self-selection model with two outcomes. Computational Statistics & Data Analysis, 42 (3):403–424, March 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002098.

Monfardini:2003:ICN

[377] Chiara Monfardini. An illustration of Cox’s non-nested testing procedure for logit and probit models. Computational Statistics & Data Analysis, 42 (3):425–444, March 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002165.

Bartolucci:2003:LBI

[378] F. Bartolucci and G. De Luca. Likelihood-based inference for asymmetric stochastic volatility models.

Computational Statistics & Data Analysis, 42(3):445–449, March 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947302002153.

Proietti:2003:FUU

[379] Tommaso Proietti. Forecasting the US unemployment rate. Computational Statistics & Data Analysis, 42 (3):451–476, March 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730200230X.

Miles:2003:PNS

[380] Daniel Miles and Juan Mora. On the performance of nonparametric specification tests in regression models. Computational Statistics & Data Analysis, 42(3):477– 490, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730200227X.

Reagle:2003:INT

[381] Derrick P. Reagle and H. D. Vinod. Inference for negativist theory using numerically computed rejection regions. Computational Statistics & Data Analysis, 42(3):491–512, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002311.

Gamerman:2003:SVR

[382] Dani Gamerman, Ajax R. B. Moreira, and Håvard Rue. Space-varying regression models: specifications and simulation. Computational Statistics & Data Analysis, 42(3):513– 533, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002116.

Anonymous:2003:SSPb

[383] Anonymous. SSN-section, PS-file follows. Computational Statistics & Data Analysis, 42(3):535–544, March 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000392.

Anonymous:2003:EBd

[384] Anonymous. Editorial board. Computational Statistics & Data Analysis, 42(4):iv–v, April 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000471.

Bowman:2003:CAN

[385] A. W. Bowman and A. Azzalini. Computational aspects of nonparametric smoothing with illustrations from the SM library. Computational Statistics & Data Analysis, 42(4):545– 560, April 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001184.

Neuhauser:2003:NET

[386] Markus Neuhäuser. A note on the exact test based on the Baumgartner– Weiß–Schindler statistic in the presence of ties. Computational Statistics & Data Analysis, 42(4):561– 568, April 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001214.

Beasley:2003:CAF

[387] T. Mark Beasley and Bruno D. Zumbo. Comparison of aligned Friedman rank and parametric methods for testing interactions in split-plot designs. Computational Statistics & Data Analysis, 42(4):569–593, April 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001470.

Beirlant:2003:RRD

[388] Jan Beirlant and Yuri Goegebeur. Regression with response distributions of Pareto-type. Computational Statistics & Data Analysis, 42(4):595– 619, April 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001202.

Hojsgaard:2003:SMC

[389] Søren Højsgaard. Split models for contingency tables. Computational Statistics & Data Analysis, 42(4):621– 645, April 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001196.

Huang:2003:TRG

[390] Yufen Huang and Norman R. Draper. Transformations, regression geometry and R^{2}. Computational Statistics & Data Analysis, 42(4):647– 664, April 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001494.

Kulldorff:2003:PCD

[391] Martin Kulldorff, Toshiro Tango, and Peter J. Park. Power comparisons for disease clustering tests. Computational Statistics & Data Analysis, 42 (4):665–684, April 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001603.

Bose:2003:MSC

[392] Smarajit Bose. Multilayer statistical classifiers. Computational Statistics & Data Analysis, 42(4):685– 701, April 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001718.

Anonymous:2003:SSPc

[393] Anonymous. SSN-section, PS-file follows with cover proofs. Computational Statistics & Data Analysis, 42 (4):703–723, April 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000616.

Anonymous:2003:AIb

[394] Anonymous. Author index. Computational Statistics & Data Analysis, 42 (4):725–726, April 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000537.

Anonymous:2003:EBe

[395] Anonymous. Editorial board. Computational Statistics & Data Analysis, 43(1):viii–ix, May 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000756.

Schmid:2003:STP

[396] Friedrich Schmid and Mark Trede. Simple tests for peakedness, fat tails and leptokurtosis based on quantiles. Computational Statistics & Data Analysis, 43(1):1–12, May 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001706.

Tutz:2003:GLR

[397] Gerhard Tutz and Göran Kauermann. Generalized linear random effects models with varying coefficients. Computational Statistics & Data Analysis, 43(1):13–28, May 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002050.

Hou:2003:FSC

[398] Chia-Ding Hou, Jengtung Chiang, and John Jen Tai. A family of simultaneous confidence intervals for multinomial proportions. Computational Statistics & Data Analysis, 43(1):29– 45, May 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200169X.

Robinson:2003:QDG

[399] Kevin S. Robinson and André I. Khuri. Quantile dispersion graphs for evaluating and comparing designs for logistic regression models. Computational Statistics & Data Analysis, 43 (1):47–62, May 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001822.

Feijoo:2003:MQD

[400] Santiago Rodríguez Feijoó, Alejandro Rodríguez Caro, and Delia Dávila Quintana. Methods for quarterly disaggregation without indicators; a comparative study using simulation. Computational Statistics & Data Analysis, 43(1):63–78, May 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302001688.

Pardalos:2003:SIA

[401] Panos M. Pardalos, J. Chris Sackellares, Leonidas D. Iasemidis, Vitality Yatsenko, Mark C. K. Yang, Deng-Shan Shiau, and Wanpracha Chaovalitwongse. Statistical information approaches for the modelling of the epileptic brain. Computational Statistics & Data Analysis, 43 (1):79–108, May 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302001524.

Anonymous:2003:SSPd

[402] Anonymous. SSN-section, PS-file follows with cover proofs. Computational Statistics & Data Analysis, 43 (1):109–119, May 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000902.

Anonymous:2003:EBf

[403] Anonymous. Editorial board. Computational Statistics & Data Analysis, 43(2):iv–v, June 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001026.

Hothorn:2003:EDM

[404] Torsten Hothorn and Berthold Lausen. On the exact distribution of maximally selected rank statistics. Computational Statistics & Data Analysis, 43 (2):121–137, June 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002256.

Charles:2003:WDP

[405] C. Charles and J. P. Rasson. Wavelet denoising of Poisson-distributed data and applications. Computational Statistics & Data Analysis, 43(2):139– 148, June 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002086.

Coppi:2003:TWF

[406] Renato Coppi and Pierpaolo D’Urso. Three-way fuzzy clustering models for LR fuzzy time trajectories. Computational Statistics & Data Analysis, 43 (2):149–177, June 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002268.

Gupta:2003:DBW

[407] Rameshwar D. Gupta and Debasis Kundu. Discriminating between Weibull and generalized exponential distributions. Computational Statistics & Data Analysis, 43(2):179– 196, June 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002062.

Lian:2003:RDG

[408] Ie-Bin Lian. Reducing over-dispersion by generalized degree of freedom and propensity score. Computational Statistics & Data Analysis, 43(2):197– 214, June 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002232.

Menendez:2003:THB

[409] M. L. Menéndez, J. A. Pardo, L. Pardo, and K. Zografos. On tests of homogeneity based on minimum ϕ-divergence estimator with constraints. Computational Statistics & Data Analysis, 43 (2):215–234, June 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002074.

Pacheco:2003:DHM

[410] Joaquín Pacheco and Olga Valencia. Design of hybrids for the minimum sum-of-squares clustering problem. Computational Statistics & Data Analysis, 43(2):235– 248, June 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002244.

Benton:2003:CDM

[411] Denise Benton and K. Krishnamoorthy. Computing discrete mixtures of continuous distributions: noncentral chisquare, noncentral t and the distribution of the square of the sample multiple correlation coefficient. Computational Statistics & Data Analysis, 43 (2):249–267, June 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002839.

Anonymous:2003:SSPe

[412] Anonymous. SSN-section, PS-file follows with cover proofs. Computational Statistics & Data Analysis, 43 (2):269–282, June 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001178.

Anonymous:2003:EBg

[413] Anonymous. Editorial board. Computational Statistics & Data Analysis, 43(3):iv–v, July 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001245.

Ng:2003:MME

[414] H. K. T. Ng, D. Kundu, and N. Balakrishnan. Modified moment estimation for the two-parameter Birnbaum– Saunders distribution. Computational Statistics & Data Analysis, 43(3):283– 298, July 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002542.

Elamir:2003:TM

[415] Elsayed A. H. Elamir and Allan H. Seheult. Trimmed L-moments. Computational Statistics & Data Analysis, 43 (3):299–314, July 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002505.

Rousseeuw:2003:RAS

[416] Peter J. Rousseeuw and Andreas Christmann. Robustness against separation and outliers in logistic regression. Computational Statistics & Data Analysis, 43(3):315– 332, July 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003043.

Swanepoel:2003:NGF

[417] C. J. Swanepoel and W. O. Doku. New goodness-of-fit tests for the error distribution of autoregressive timeseries models. Computational Statistics & Data Analysis, 43(3):333– 340, July 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002335.

Wisnowski:2003:RMV

[418] James W. Wisnowski, James R. Simpson, Douglas C. Montgomery, and George C. Runger. Resampling methods for variable selection in robust regression. Computational Statistics & Data Analysis, 43(3):341– 355, July 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002359.

Yang:2003:TAC

[419] Zhenlin Yang, Stanley P. See, and M. Xie. Transformation approaches for the construction of Weibull prediction interval. Computational Statistics & Data Analysis, 43(3):357– 368, July 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002323.

Fong:2003:TSR

[420] Pak Wing Fong and Wai Keung Li. On time series with randomized unit root and randomized seasonal unit root. Computational Statistics & Data Analysis, 43(3):369– 395, July 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002980.

Anonymous:2003:SSPf

[421] Anonymous. SSN-section, PS-file follows with cover proofs. Computational Statistics & Data Analysis, 43 (3):397–420, July 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001397.

Anonymous:2003:EBh

[422] Anonymous. Editorial board. Computational Statistics & Data Analysis, 43(4):iv–v, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001579.

Wegman:2003:P

[423] Edward J. Wegman, Wendy L. Martinez, and Jeffrey L. Solka. Preface. Computational Statistics & Data Analysis, 43(4):421–422, August 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002852.

Swayne:2003:GEX

[424] Deborah F. Swayne, Duncan Temple Lang, Andreas Buja, and Dianne Cook. GGobi: evolving from XGobi into an extensible framework for interactive data visualization. Computational Statistics & Data Analysis, 43 (4):423–444, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002864.

Bohlen:2003:VDM

[425] Michael Böhlen, Linas Bukauskas, Poul Svante Eriksen, Steffen Lilholt Lauritzen, Artūras Mažeika, Peter Musaeus, and Peer Mylov. 3D visual data mining — goals and experiences. Computational Statistics & Data Analysis, 43(4):445–469, August 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002876.

Wilhelm:2003:UIV

[426] Adalbert Wilhelm. User interaction at various levels of data displays. Computational Statistics & Data Analysis, 43 (4):471–494, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002888.

Valero-Mora:2003:VCD

[427] Pedro M. Valero-Mora, Forrest W. Young, and Michael Friendly. Visualizing categorical data in ViSta. Computational Statistics & Data Analysis, 43 (4):495–508, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200289X.

Friendly:2003:EOD

[428] Michael Friendly and Ernest Kwan. Effect ordering for data displays. Computational Statistics & Data Analysis, 43 (4):509–539, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002906.

Marchette:2003:UDI

[429] David J. Marchette and Jeffrey L. Solka. Using data images for outlier detection. Computational Statistics & Data Analysis, 43(4):541–552, August 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002918.

Unwin:2003:PCE

[430] Antony Unwin, Chris Volinsky, and Sylvia Winkler. Parallel coordinates for exploratory modelling analysis. Computational Statistics & Data Analysis, 43(4):553–564, August 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200292X.

Hofmann:2003:CRM

[431] Heike Hofmann. Constructing and reading mosaicplots. Computational Statistics & Data Analysis, 43(4):565– 580, August 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002931.

Dykes:2003:SSR

[432] J. A. Dykes and D. M. Mountain. Seeking structure in records of spatiotemporal behaviour: visualization issues, efforts and applications. Computational Statistics & Data Analysis, 43 (4):581–603, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002943.

Edsall:2003:PCP

[433] Robert M. Edsall. The parallel coordinate plot in action: design and use for geographic visualization. Computational Statistics & Data Analysis, 43 (4):605–619, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002955.

Priebe:2003:CCC

[434] Carey E. Priebe, Jeffrey L. Solka, David J. Marchette, and B. Ted Clark. Class cover catch digraphs for latent class discovery in gene expression monitoring by DNA microarrays. Computational Statistics & Data Analysis, 43 (4):621–632, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002967.

Wegman:2003:VCW

[435] Edward J. Wegman and Alan Dorfman. Visualizing cereal world. Computational Statistics & Data Analysis, 43 (4):633–649, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002979.

Anonymous:2003:AIc

[436] Anonymous. Author index. Computational Statistics & Data Analysis, 43 (4):651–652, August 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001609.

Anonymous:2003:EBi

[437] Anonymous. Editorial board. Computational Statistics & Data Analysis, 44(1–2):viii–ix, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947303001920.

Niland:2003:SIH

[438] Joyce C. Niland, Abdelmonem A. Afifi, and Erricos John Kontoghiorghes. Special issue in honour of Stan Azen: a birthday celebration. Computational Statistics & Data Analysis, 44 (1–2):1–2, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001786.

Foschi:2003:CSA

[439] Paolo Foschi, David A. Belsley, and Erricos J. Kontoghiorghes. A comparative study of algorithms for solving seemingly unrelated regressions models. Computational Statistics & Data Analysis, 44(1–2):3–35, October 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000288.

Pollock:2003:REE

[440] D. S. G. Pollock. Recursive estimation in econometrics. Computational Statistics & Data Analysis, 44(1–2): 37–75, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001506.

Jhun:2003:BNP

[441] Myoungshic Jhun, Seuck Heun Song, and Byoung Cheol Jung. BLUP in the nested panel regression model with serially correlated errors. Computational Statistics & Data Analysis, 44(1– 2):77–88, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003481.

Yuan:2003:ETS

[442] Ke-Hai Yuan and Peter M. Bentler. Eight test statistics for multilevel structural equation models. Computational Statistics & Data Analysis, 44(1–2): 89–107, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003493.

Zeileis:2003:TDS

[443] Achim Zeileis, Christian Kleiber, Walter Krämer, and Kurt Hornik. Testing and dating of structural changes in practice. Computational Statistics & Data Analysis, 44(1–2):109– 123, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000306.

Lee:2003:MLE

[444] Sik-Yum Lee and Xin-Yuan Song. Maximum likelihood estimation and model comparison of nonlinear structural equation models with continuous and polytomous variables. Computational Statistics & Data Analysis, 44(1– 2):125–142, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003055.

Tanaka:2003:LIP

[445] Yutaka Tanaka, Fanghong Zhang, and Yuichi Mori. Local influence in principal component analysis: relationship between the local influence and influence function approaches revisited. Computational Statistics & Data Analysis, 44(1–2):143–160, October 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003444.

Ferraty:2003:CDN

[446] F. Ferraty and P. Vieu. Curves discrimination: a nonparametric functional approach. Computational Statistics & Data Analysis, 44(1–2):161– 173, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300032X.

Geng:2003:BML

[447] Zhi Geng, Yang-Bo He, Xue-Li Wang, and Qiang Zhao. Bayesian method for learning graphical models with incompletely categorical data. Computational Statistics & Data Analysis, 44(1–2): 175–192, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000665.

Woodward:2003:TOM

[448] Wayne A. Woodward and Stephan R. Sain. Testing for outliers from a mixture distribution when some data are missing. Computational Statistics & Data Analysis, 44(1–2):193– 210, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000082.

Lee:2003:MAC

[449] Seung-Chun Lee and Moon Yul Huh. A measure of association for complex data. Computational Statistics & Data Analysis, 44(1–2):211– 222, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000318.

Wilcox:2003:IBM

[450] Rand R. Wilcox. Inferences based on multiple skipped correlations. Computational Statistics & Data Analysis, 44(1–2):223–236, October 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000434.

Caussinus:2003:MDM

[451] H. Caussinus, M. Fekri, S. Hakam, and A. Ruiz-Gazen. A monitoring display of multivariate outliers. Computational Statistics & Data Analysis, 44(1–2): 237–252, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000598.

Heinzl:2003:PSM

[452] Harald Heinzl and Martina Mittlböck. Pseudo R-squared measures for Poisson regression models with over- or underdispersion. Computational Statistics & Data Analysis, 44(1–2):253– 271, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000628.

Croux:2003:IBY

[453] Christophe Croux and Gentiane Haesbroeck. Implementing the Bianco and Yohai estimator for logistic regression. Computational Statistics & Data Analysis, 44(1–2):273–295, October 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000422.

Payne:2003:GBL

[454] Roger W. Payne. General balance, large data sets and extensions to unbalanced treatment structures. Computational Statistics & Data Analysis, 44(1– 2):297–304, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730300094X.

Trenkel:2003:DES

[455] Verena M. Trenkel. Density estimator for strip transects when animals show directional movement and observation speed is slow. Computational Statistics & Data Analysis, 44(1–2): 305–312, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000604.

Kafadar:2003:TBE

[456] Karen Kafadar and Tzulip Phang. Transformations, background estimation, and process effects in the statistical analysis of microarrays. Computational Statistics & Data Analysis, 44(1– 2):313–338, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000690.

Husing:2003:QSN

[457] Johannes Hüsing, Irina Gana Dresen, and Karl-Heinz Jöckel. Quest for a sensible null distribution in longitudinal microarray experiments. Computational Statistics & Data Analysis, 44(1– 2):339–347, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000653.

Basak:2003:REU

[458] Indrani Basak and N. Balakrishnan. Robust estimation under progressive censoring. Computational Statistics & Data Analysis, 44(1–2):349–376, October 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300029X.

Zakovic:2003:PAS

[459] S. Žaković, B. Rustem, and S. P. Asprey. A parallel algorithm for semiinfinite programming. Computational Statistics & Data Analysis, 44(1–2): 377–390, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300152X.

Pardalos:2003:AED

[460] Panos M. Pardalos, Vitaliy Yatsenko, J. Chris Sackellares, Deng-Shan Shiau, Wanpracha Chaovalitwongse, and Leonidas D. Iasemidis. Analysis of EEG data using optimization, statistics, and dynamical system techniques. Computational Statistics & Data Analysis, 44(1–2):391– 408, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000276.

Asano:2003:WBS

[461] Chooichiro Asano and Akinobu Takeuchi. Web-based statistical system by using the DLL. Computational Statistics & Data Analysis, 44(1–2):409– 418, October 28, 2003. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001440.

Victor:2003:CFA

[462] N. Victor and M. Kieser. Configural frequency analysis and association analyses in contingency tables. Computational Statistics & Data Analysis, 44(1–2):419–429, October 28, 2003. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000380.

Anonymous:2004:EBa

[463] Anonymous. Editorial board. Computational Statistics & Data Analysis, 44(3):iv–v, January 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002159.

Anonymous:2004:PNV

[464] Anonymous. Publisher’s note vi. Computational Statistics & Data Analysis, 44(3):vi, January 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730300224X.

Borkowf:2004:EAG

[465] Craig B. Borkowf. An efficient algorithm for generating two-way contingency tables with fixed marginal totals and arbitrary mean proportions, with applications to permutation tests. Computational Statistics & Data Analysis, 44(3):431–449, January 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002530.

Glen:2004:CDP

[466] Andrew G. Glen, Lawrence M. Leemis, and John H. Drew. Computing the distribution of the product of two continuous random variables. Computational Statistics & Data Analysis, 44 (3):451–464, January 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002347.

Lopez-Fidalgo:2004:CMS

[467] Jesús López-Fidalgo and Chiara Tommasi. Construction of MV- and SMV-optimum designs for binary response models. Computational Statistics & Data Analysis, 44(3):465–475, January 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002566.

Mira:2004:PDF

[468] José Mira and María Jesús Sánchez. Prediction of deterministic functions: an application of a Gaussian kriging model to a time series outlier problem. Computational Statistics & Data Analysis, 44(3):477–491, January 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002554.

Wilkinson:2004:SMA

[469] Darren J. Wilkinson and Stephen KH Yeung. A sparse matrix approach to Bayesian computation in large linear models. Computational Statistics & Data Analysis, 44(3):493–516, January 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002529.

Besbeas:2004:ISE

[470] Panagiotis Besbeas and Byron J. T. Morgan. Integrated squared error estimation of normal mixtures. Computational Statistics & Data Analysis, 44 (3):517–526, January 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002517.

Stokes:2004:USS

[471] Lynne Stokes and John Plummer. Using spreadsheet solvers in sample design. Computational Statistics & Data Analysis, 44(3):527–546, January 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003225.

Anonymous:2004:EBb

[472] Anonymous. Editorial board. Computational Statistics & Data Analysis, 44 (4):iv–v, January 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002524.

Gaudart:2004:CPM

[473] Jean Gaudart, Bernard Giusiano, and Laetitia Huiart. Comparison of the performance of multi-layer perceptron and linear regression for epidemiological data. Computational Statistics & Data Analysis, 44(4):547–570, January 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002578.

Gilmour:2004:ECS

[474] Arthur Gilmour, Brian Cullis, Sue Welham, Beverley Gogel, and Robin Thompson. An efficient computing strategy for prediction in mixed linear models. Computational Statistics & Data Analysis, 44(4):571–586, January 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730200258X.

Hwang:2004:PBR

[475] Jinsoo Hwang, Hongsuk Jorn, and Jeankyung Kim. On the performance of bivariate robust location estimators under contamination. Computational Statistics & Data Analysis, 44(4):587– 601, January 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002827.

Argac:2004:THG

[476] Doğan Argaç. Testing for homogeneity in a general one-way classification with fixed effects: power simulations and comparative study. Computational Statistics & Data Analysis, 44(4):603– 612, January 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002645.

Guo:2004:THR

[477] Jianhua Guo, Yanping Ma, Ning-Zhong Shi, and Tai Shing Lau. Testing for homogeneity of relative differences under inverse sampling. Computational Statistics & Data Analysis, 44(4):613– 624, January 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002621.

Hardin:2004:ODM

[478] Johanna Hardin and David M. Rocke. Outlier detection in the multiple cluster setting using the minimum covariance determinant estimator. Computational Statistics & Data Analysis, 44(4):625– 638, January 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302002803.

Marozzi:2004:BANa

[479] Marco Marozzi. A bi-aspect nonparametric test for the two-sample location problem. Computational Statistics & Data Analysis, 44(4):639–648, January 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002797.

Renard:2004:PLA

[480] Didier Renard, Geert Molenberghs, and Helena Geys. A pairwise likelihood approach to estimation in multilevel probit models. Computational Statistics & Data Analysis, 44(4):649– 667, January 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302002633.

Huang:2004:BSP

[481] Wen-Tao Huang and Yu-Pin Lin. Bayesian sampling plans for exponential distribution based on uniform random censored data. Computational Statistics & Data Analysis, 44(4):669– 691, January 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003304.

Anonymous:2004:AIa

[482] Anonymous. Author index. Computational Statistics & Data Analysis, 44(4):693–694, January 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947303002573.

Anonymous:2004:EBc

[483] Anonymous. Editorial board. Computational Statistics & Data Analysis, 45 (1):viii–ix, February 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303003116.

Solka:2004:E

[484] Jeffrey L. Solka and Edward J. Wegman. Editorial. Computational Statistics & Data Analysis, 45(1):1– 2, February 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001129.

Wierman:2004:MCV

[485] John C. Wierman and David J. Marchette. Modeling computer virus prevalence with a susceptibleinfected-susceptible model with reintroduction. Computational Statistics & Data Analysis, 45(1):3–23, February 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001130.

Bohacek:2004:DMR

[486] Stephan Bohacek and Boris Rozovskii. A diffusion model of roundtrip time. Computational Statistics & Data Analysis, 45(1):25–50, February 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001142.

Naiman:2004:SAD

[487] Daniel Q. Naiman. Statistical anomaly detection via httpd data analysis. Computational Statistics & Data Analysis, 45(1):51–67, February 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001154.

Scott:2004:BPD

[488] Steven L. Scott. A Bayesian paradigm for designing intrusion detection systems. Computational Statistics & Data Analysis, 45(1):69–83, February 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001774.

Tsourti:2004:EVA

[489] Zoi Tsourti and John Panaretos. Extreme-value analysis of teletraffic data. Computational Statistics & Data Analysis, 45(1):85–103, February 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001166.

Anonymous:2004:EBd

[490] Anonymous. Editorial board. Computational Statistics & Data Analysis, 45(2):iv–v, March 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000088.

Basu:2004:IRE

[491] Ayanendranath Basu and Bruce G. Lindsay. The iteratively reweighted estimating equation in minimum distance problems. Computational Statistics & Data Analysis, 45(2):105– 124, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003262.

Handley:2004:CEA

[492] John C. Handley. Computationally efficient approximate likelihood procedures for the Boolean model. Computational Statistics & Data Analysis, 45 (2):125–136, March 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003250.

Jamshidian:2004:ARM

[493] Mortaza Jamshidian. On algorithms for restricted maximum likelihood estimation. Computational Statistics & Data Analysis, 45(2):137– 157, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003456.

Molinari:2004:BOK

[494] Nicolas Molinari, Jean-François Durand, and Robert Sabatier. Bounded optimal knots for regression splines. Computational Statistics & Data Analysis, 45(2):159–178, March 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947302003432.

Pace:2004:CAL

[495] R. Kelley Pace and James P. LeSage. Chebyshev approximation of log-determinants of spatial weight matrices. Computational Statistics & Data Analysis, 45(2):179– 196, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003213.

Paoletti:2004:DED

[496] Xavier Paoletti, John O’Quigley, and Jean Maccario. Design efficiency in dose finding studies. Computational Statistics & Data Analysis, 45(2):197– 214, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003237.

Cribari-Neto:2004:AIU

[497] Francisco Cribari-Neto. Asymptotic inference under heteroskedasticity of unknown form. Computational Statistics & Data Analysis, 45(2):215– 233, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003663.

Dean:2004:PQL

[498] C. B. Dean, M. D. Ugarte, and A. F. Militino. Penalized quasi-likelihood with spatially correlated data. Computational Statistics & Data Analysis, 45(2):235–248, March 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302003249.

Delaigle:2004:PBS

[499] A. Delaigle and I. Gijbels. Practical bandwidth selection in deconvolution kernel density estimation. Computational Statistics & Data Analysis, 45 (2):249–267, March 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302003298.

Duczmal:2004:SAS

[500] Luiz Duczmal and Renato Assunção. A simulated annealing strategy for the detection of arbitrarily shaped spatial clusters. Computational Statistics & Data Analysis, 45(2):269– 286, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730200302X.

Gupta:2004:NGN

[501] Ramesh C. Gupta and S. H. Ong. A new generalization of the negative binomial distribution. Computational Statistics & Data Analysis, 45 (2):287–300, March 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302003018.

Hubert:2004:FRD

[502] Mia Hubert and Katrien Van Driessen. Fast and robust discriminant analysis. Computational Statistics & Data Analysis, 45(2):301–320, March 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947302002992.

Lee:2004:IAN

[503] Sik-Yum Lee and Liang Xu. Influence analyses of nonlinear mixedeffects models. Computational Statistics & Data Analysis, 45(2):321– 341, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003031.

Zhang:2004:NMC

[504] Li-Chun Zhang. Nonparametric Markov chain bootstrap for multiple imputation. Computational Statistics & Data Analysis, 45(2):343– 353, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003006.

Hurlimann:2004:FBC

[505] Werner Hürlimann. Fitting bivariate cumulative returns with copulas. Computational Statistics & Data Analysis, 45(2):355–372, March 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003468.

Andres:2004:EUN

[506] A. Martín Andrés and I. Herranz Tejedor. Exact unconditional nonclassical tests on the difference of two proportions. Computational Statistics & Data Analysis, 45(2):373– 388, March 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003511.

Anonymous:2004:EBe

[507] Anonymous. Editorial board. Computational Statistics & Data Analysis, 45(3):iv–v, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000416.

Ait-Sidi-Allal:2004:NAE

[508] M. L. Aït-Sidi-Allal, A. Baccini, and A. M. Mondot. A new algorithm for estimating the parameters and their asymptotic covariance in correlation and association models. Computational Statistics & Data Analysis, 45 (3):389–421, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000355.

Georgiou:2004:CNS

[509] S. Georgiou, C. Koukouvinos, and S. Stylianou. Construction of new skew Hadamard matrices and their use in screening experiments. Computational Statistics & Data Analysis, 45 (3):423–429, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000586.

Kozek:2004:GQP

[510] A. S. Kozek and J. Yin. On Gauss quadrature and partial cross validation. Computational Statistics & Data Analysis, 45(3):431– 448, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000409.

Price:2004:ICI

[511] Robert M. Price and Douglas G. Bonett. An improved confidence interval for a linear function of binomial proportions. Computational Statistics & Data Analysis, 45(3):449– 456, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000070.

Shih:2004:NSS

[512] Y.-S. Shih. A note on split selection bias in classification trees. Computational Statistics & Data Analysis, 45 (3):457–466, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000641.

Stamey:2004:PSS

[513] James D. Stamey, Tom L. Bratcher, and Dean M. Young. Parameter subset selection and multiple comparisons of Poisson rate parameters with misclassification. Computational Statistics & Data Analysis, 45(3):467– 479, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000689.

Becue-Bertaut:2004:PAM

[514] Mónica Bécue-Bertaut and Jérôme Pagès. A principal axes method for comparing contingency tables: MFACT. Computational Statistics & Data Analysis, 45(3):481– 503, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000033.

Lautre:2004:MML

[515] Ignacio García Lautre and Elena Abascal Fernández. A methodology for measuring latent variables based on multiple factor analysis. Computational Statistics & Data Analysis, 45 (3):505–517, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000379.

Giordani:2004:PCA

[516] Paolo Giordani and Henk A. L. Kiers. Principal component analysis of symmetric fuzzy data. Computational Statistics & Data Analysis, 45(3):519– 548, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302003523.

Lai:2004:EHE

[517] Dejian Lai. Estimating the Hurst effect and its application in monitoring clinical trials. Computational Statistics & Data Analysis, 45(3):549– 562, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000859.

Merbouha:2004:RLM

[518] A. Merbouha and A. Mkhadri. Regularization of the location model in discrimination with mixed discrete and continuous variables. Computational Statistics & Data Analysis, 45 (3):563–576, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000677.

Miglio:2004:CBC

[519] Rossella Miglio and Gabriele Soffritti. The comparison between classification trees through proximity measures. Computational Statistics & Data Analysis, 45(3):577– 593, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300063X.

Shih:2004:VSB

[520] Yu-Shan Shih and Hsin-Wen Tsai. Variable selection bias in regression trees with constant fits. Computational Statistics & Data Analysis, 45 (3):595–607, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000367.

Tan:2004:UNQ

[521] W. D. Tan, F. F. Gan, and T. C. Chang. Using normal quantile plot to select an appropriate transformation to achieve normality. Computational Statistics & Data Analysis, 45 (3):609–619, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000094.

Vu:2004:ESC

[522] Hien T. V. Vu. Estimation in semiparametric conditional shared frailty models with events before study entry. Computational Statistics & Data Analysis, 45(3):621– 637, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947302003687.

Yun:2004:CHM

[523] Sungcheol Yun and Youngjo Lee. Comparison of hierarchical and marginal likelihood estimators for binary outcomes. Computational Statistics & Data Analysis, 45(3):639–650, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000331.

Halekoh:2004:BAS

[524] U. Halekoh and W. Vach. A Bayesian approach to seriation problems in archaeology. Computational Statistics & Data Analysis, 45(3):651– 673, April 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000343.

Shao:2004:NML

[525] Quanxi Shao. Notes on maximum likelihood estimation for the threeparameter Burr XII distribution. Computational Statistics & Data Analysis, 45(3):675–687, April 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003675.

Anonymous:2004:EBf

[526] Anonymous. Editorial board. Computational Statistics & Data Analysis, 45(4):iv–v, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400057X.

Ahmed:2004:BES

[527] S. E. Ahmed and W. J. Krzanowski. Biased estimation in a simple multivariate regression model. Computational Statistics & Data Analysis, 45 (4):689–696, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000884.

Krishnamoorthy:2004:EMT

[528] K. Krishnamoorthy, Jessica Thomson, and Yong Cai. An exact method of testing equality of several binomial proportions to a specified standard. Computational Statistics & Data Analysis, 45 (4):697–707, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000951.

Favre:2004:VCA

[529] Anne-Catherine Favre, Alina Matei, and Yves Tillé. A variant of the Cox algorithm for the imputation of nonresponse of qualitative data. Computational Statistics & Data Analysis, 45 (4):709–719, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000938.

Sulieman:2004:PBA

[530] H. Sulieman, P. J. McLellan, and D. W. Bacon. A profile-based approach to parametric sensitivity in multiresponse regression models. Computational Statistics & Data Analysis, 45 (4):721–740, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000860.

Basu:2004:SVM

[531] Ayanendranath Basu, Chanseok Park, Bruce G. Lindsay, and Haihong Li. Some variants of minimum disparity estimation. Computational Statistics & Data Analysis, 45(4):741– 763, May 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000987.

Renshaw:2004:MHS

[532] Eric Renshaw. Metropolis–Hastings from a stochastic population dynamics perspective. Computational Statistics & Data Analysis, 45(4):765– 786, May 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300118X.

Chaouche:2004:STP

[533] Ali Chaouche and Jean-Noël Bacro. A statistical test procedure for the shape parameter of a generalized Pareto distribution. Computational Statistics & Data Analysis, 45(4):787– 803, May 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000872.

Liang:2004:CCB

[534] Hua Liang. Comparison of curves based on a Cramér–von Mises statistic. Computational Statistics & Data Analysis, 45(4):805–812, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001749.

Gao:2004:IPF

[535] Feng Gao, Amita K. Manatunga, and Shande Chen. Identification of prognostic factors with multivariate survival data. Computational Statistics & Data Analysis, 45(4):813– 824, May 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000896.

Rueda:2004:SIE

[536] M. M. Rueda, A. Arcos, M. D. Martínez-Miranda, and Y. Román. Some improved estimators of finite population quantile using auxiliary information in sample surveys. Computational Statistics & Data Analysis, 45 (4):825–848, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000975.

Song:2004:RMC

[537] James X. Song, James T. Wassell, and Asha Kapadia. Relative mortality for correlated lifetime data. Computational Statistics & Data Analysis, 45 (4):849–864, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000963.

Fu:2004:CLM

[538] Chong Yau Fu. Combining loglinear model with classification and regression tree (CART): an application to birth data. Computational Statistics & Data Analysis, 45(4):865– 874, May 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303000926.

Hennig:2004:DBP

[539] Christian Hennig and Bernhard Hausdorf. Distance-based parametric bootstrap tests for clustering of species ranges. Computational Statistics & Data Analysis, 45(4):875– 895, May 10, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303000914.

Anonymous:2004:AIb

[540] Anonymous. Author index. Computational Statistics & Data Analysis, 45 (4):897–899, May 10, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000623.

Anonymous:2004:EBg

[541] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 46(1):ix–x, May 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000726.

Li:2004:RST

[542] Chin-Shang Li and Daniel Hunt. Regression splines for threshold selection with application to a random-effects logistic dose-response model. Computational Statistics & Data Analysis, 46(1):1–9, May 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001361.

Elden:2004:PLS

[543] Lars Eldén. Partial least-squares vs. Lanczos bidiagonalization — i: analysis of a projection method for multiple regression. Computational Statistics & Data Analysis, 46(1):11– 31, May 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001385.

Huo:2004:NFA

[544] Xiaoming Huo and Jye-Chyi Lu. A network flow approach in finding maximum likelihood estimate of high concentration regions. Computational Statistics & Data Analysis, 46(1):33– 56, May 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001348.

Bolton:2004:STU

[545] Richard J. Bolton, David J. Hand, and Martin Crowder. Significance tests for unsupervised pattern discovery in large continuous multivariate data sets. Computational Statistics & Data Analysis, 46(1):57– 79, May 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001427.

Marozzi:2004:BANb

[546] Marco Marozzi. A bi-aspect nonparametric test for the multi-sample location problem. Computational Statistics & Data Analysis, 46(1):81– 92, May 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001439.

Olive:2004:REM

[547] David J. Olive. A resistant estimator of multivariate location and dispersion. Computational Statistics & Data Analysis, 46(1):93– 102, May 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001191.

Gannoun:2004:SIR

[548] Ali Gannoun, Stéphane Girard, Christiane Guinot, and Jérôme Saracco. Sliced inverse regression in reference curves estimation. Computational Statistics & Data Analysis, 46(1):103– 122, May 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001415.

Sherman:2004:CIB

[549] Michael Sherman and Edward Carlstein. Confidence intervals based on estimators with unknown rates of convergence. Computational Statistics & Data Analysis, 46(1):123– 139, May 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001452.

Taylor:2004:CPT

[550] C. C. Taylor and A. Zempléni. Chain plot: a tool for exploiting bivariate temporal structures. Computational Statistics & Data Analysis, 46 (1):141–153, May 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001208.

Vivien:2004:GSA

[551] Myrtille Vivien and Robert Sabatier. A generalization of STATIS-ACT strategy: DO-ACT for two multiblocks tables. Computational Statistics & Data Analysis, 46(1):155– 171, May 28, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001464.

Walde:2004:SAM

[552] J. F. Walde, G. Tappeiner, U. Tappeiner, E. Tasser, and H. W. Holub. Statistical aspects of multilayer perceptrons under data limitations. Computational Statistics & Data Analysis, 46 (1):173–188, May 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001403.

Liu:2004:SMT

[553] W. Liu and P. N. Somerville. Stepwise multiple tests for successive comparisons of treatment effects. Computational Statistics & Data Analysis, 46 (1):189–199, May 28, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001373.

Anonymous:2004:EBh

[554] Anonymous. Editorial board. Computational Statistics & Data Analysis, 46(2):iv–v, June 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000891.

Kawai:2004:ASM

[555] Shinichi Kawai. Approximation of some moments in eigenvalue problem including some random variables and its application. Computational Statistics & Data Analysis, 46(2):201– 212, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001518.

Chou:2004:PGC

[556] Chih-Ping Chou, Dongyun Yang, Mary Ann Pentz, and Yih-Ing Hser. Piecewise growth curve modeling approach for longitudinal prevention study. Computational Statistics & Data Analysis, 46(2):213– 225, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300149X.

Duchesne:2004:RTC

[557] Pierre Duchesne. On robust testing for conditional heteroscedasticity in time series models. Computational Statistics & Data Analysis, 46 (2):227–256, June 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001737.

Greenwell:2004:RRP

[558] Raymond N. Greenwell and Stephen J. Finch. Randomized rejection procedure for the two-sample Kolmogorov– Smirnov statistic. Computational Statistics & Data Analysis, 46(2):257– 267, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001488.

Kojadinovic:2004:AHC

[559] Ivan Kojadinovic. Agglomerative hierarchical clustering of continuous variables based on mutual information. Computational Statistics & Data Analysis, 46(2):269– 294, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001531.

Pardoe:2004:MAP

[560] Iain Pardoe. Model assessment plots for multilevel logistic regression. Computational Statistics & Data Analysis, 46(2):295–307, June 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001476.

Piegorsch:2004:SSI

[561] Walter W. Piegorsch. Sample sizes for improved binomial confidence intervals. Computational Statistics & Data Analysis, 46(2):309– 316, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002354.

Shao:2004:DED

[562] Quanxi Shao, Wai-Cheung Ip, and Heung Wong. Determination of embedded distributions. Computational Statistics & Data Analysis, 46(2):317– 334, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001713.

Wei:2004:DDR

[563] Wen Hsiang Wei. Derivatives diagnostics and robustness for smoothing splines. Computational Statistics & Data Analysis, 46(2):335– 356, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001701.

Horn:2004:RBG

[564] Paul S. Horn and Douglas A. Wolfe. A rank-based goodness-of-fit approach to testing for non-additivity in the twoway layout with no replications. Computational Statistics & Data Analysis, 46(2):357–376, June 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947302003274.

Hogel:2004:CCP

[565] Josef Högel and Martina Kron. Comparison of conditional power criteria for early termination of clinical trials. Computational Statistics & Data Analysis, 46(2):377– 395, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001725.

Magnussen:2004:AGP

[566] Steen Magnussen. An algorithm for generating positively correlated betadistributed random variables with known marginal distributions and a specified correlation. Computational Statistics & Data Analysis, 46(2):397– 406, June 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001695.

Anonymous:2004:EBi

[567] Anonymous. Editorial board. Computational Statistics & Data Analysis, 46(3):v–vi, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001306.

Nguyen:2004:PLS

[568] Danh V. Nguyen and David M. Rocke. On partial least squares dimension reduction for microarray-based classification: a simulation study. Computational Statistics & Data Analysis, 46 (3):407–425, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001798.

Maas:2004:IVA

[569] Cora J. M. Maas and Joop J. Hox. The influence of violations of assumptions on multilevel parameter estimates and their standard errors. Computational Statistics & Data Analysis, 46 (3):427–440, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001816.

Kozumi:2004:PAL

[570] Hideo Kozumi. Posterior analysis of latent competing risk models by parallel tempering. Computational Statistics & Data Analysis, 46(3):441– 458, June 15, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001828.

Mak:2004:EVA

[571] Tak K. Mak. Estimating variances for all sample sizes by the bootstrap. Computational Statistics & Data Analysis, 46(3):459–467, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300183X.

Munoz-Pichardo:2004:FMM

[572] J. M. Muñoz-Pichardo, A. Enguix-González, J. Muñoz-García, and A. Pascual-Acosta. The Fréchet’s metric as a measure of influence in multivariate linear models with random errors elliptically distributed. Computational Statistics & Data Analysis, 46 (3):469–491, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001841.

Mercatanti:2004:ARE

[573] Andrea Mercatanti. Analyzing a randomized experiment with imperfect compliance and ignorable conditions for missing data: theoretical and computational issues. Computational Statistics & Data Analysis, 46 (3):493–509, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001853.

Giampaoli:2004:CTN

[574] Viviana Giampaoli and Julio da Motta Singer. Comparison of two normal populations with restricted means. Computational Statistics & Data Analysis, 46 (3):511–529, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001865.

Montepiedra:2004:TSD

[575] Grace Montepiedra and Arthur B. Yeh. Two-stage designs for identification and estimation of polynomial models. Computational Statistics & Data Analysis, 46(3):531– 546, June 15, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002081.

Famoye:2004:CGP

[576] Felix Famoye and Weiren Wang. Censored generalized Poisson regression model. Computational Statistics & Data Analysis, 46(3):547– 560, June 15, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002093.

Caroni:2004:DIO

[577] C. Caroni and V. Karioti. Detecting an innovative outlier in a set of time series. Computational Statistics & Data Analysis, 46(3):561– 570, June 15, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002111.

Govindarajulu:2004:SMC

[578] Usha Govindarajulu, Lisa Sullivan, and Ralph B. D’Agostino. SAS macros for confidence interval estimation for CHD risk prediction by different survival models. Computational Statistics & Data Analysis, 46 (3):571–592, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001750.

Mansouri:2004:NMC

[579] Hossein Mansouri and Carrie Shaw. Nonparametric multiple comparison procedures for ordered parameters in balanced incomplete blocks. Computational Statistics & Data Analysis, 46 (3):593–604, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303001762.

Maiste:2004:OTS

[580] Paul J. Maiste and Bruce S. Weir. Optimal testing strategies for large, sparse multinomial models. Computational Statistics & Data Analysis, 46 (3):605–620, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303001804.

Anonymous:2004:EBj

[581] Anonymous. Editorial board. Computational Statistics & Data Analysis, 46(4):iv–v, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001884.

Park:2004:CRV

[582] Chul Gyu Park. Construction of random vectors of heterogeneous component variables under specified correlation structures. Computational Statistics & Data Analysis, 46(4):621– 630, July 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002457.

Kim:2004:SNM

[583] Inyoung Kim and Noah Cohen. Semiparametric and nonparametric modeling for effect modification in matched studies. Computational Statistics & Data Analysis, 46(4):631– 643, July 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730300210X.

Rodel:2004:LRT

[584] Egmar Rödel and Wolfgang Kössler. Linear rank tests for independence in bivariate distributions — power comparisons by simulation. Computational Statistics & Data Analysis, 46 (4):645–660, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002251.

Tao:2004:DRA

[585] Jian Tao, Ning-Zhong Shi, and Sik-Yum Lee. Drug risk assessment with determining the number of subpopulations under finite mixture normal models. Computational Statistics & Data Analysis, 46(4):661– 676, July 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002263.

Randal:2004:MLE

[586] John A. Randal and Peter J. Thomson. Maximum likelihood estimation for Tukey’s three corners. Computational Statistics & Data Analysis, 46 (4):677–687, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002299.

Mevik:2004:APC

[587] Bjørn-Helge Mevik. Assessing the performance of classifiers when classes arise from a continuum. Computational Statistics & Data Analysis, 46 (4):689–705, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002305.

El-Bassiouni:2004:TNV

[588] M. Y. El-Bassiouni and H. A. Charif. Testing a null variance ratio in mixed models with zero degrees of freedom for error. Computational Statistics & Data Analysis, 46(4):707– 719, July 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002317.

Tack:2004:TRC

[589] Lieven Tack and Martina Vandebroek. Trend-resistant and cost-efficient crossover designs for mixed models. Computational Statistics & Data Analysis, 46(4):721–746, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002329.

Ge:2004:ALA

[590] Zhiyu Ge, Peter J. Bickel, and John A. Rice. An approximate likelihood approach to nonlinear mixed effects models via spline approximation. Computational Statistics & Data Analysis, 46 (4):747–776, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002330.

Tutz:2004:GSS

[591] Gerhard Tutz. Generalized semiparametrically structured mixed models. Computational Statistics & Data Analysis, 46(4):777–800, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947303002342.

Manor:2004:SSI

[592] Orly Manor and David M. Zucker. Small sample inference for the fixed effects in the mixed linear model. Computational Statistics & Data Analysis, 46(4):801–817, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300238X.

BuHamra:2004:ICQ

[593] Sana S. BuHamra, Noriah M. Al-Kandari, and S. E. Ahmed. Inference concerning quantile for left truncated and right censored data. Computational Statistics & Data Analysis, 46 (4):819–831, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300241X.

Anonymous:2004:AIc

[594] Anonymous. Author index. Computational Statistics & Data Analysis, 46 (4):833–835, July 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001926.

Anonymous:2004:EBk

[595] Anonymous. Editorial board. Computational Statistics & Data Analysis, 47 (1):viii–ix, August 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002129.

Anonymous:2004:CP

[596] Anonymous. Call for papers. Computational Statistics & Data Analysis, 47(1):xi, August 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400218X.

Nagahara:2004:MSM

[597] Yuichi Nagahara. A method of simulating multivariate nonnormal distributions by the Pearson distribution system and estimation. Computational Statistics & Data Analysis, 47 (1):1–29, August 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002421. See

erratum [939].

Chang:2004:WEP

[598] Xiao-Wen Chang and Leming Qu. Wavelet estimation of partially linear models. Computational Statistics & Data Analysis, 47(1):31–48, August 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002445. See

erratum [693].

Mugdadi:2004:BSK

[599] A. R. Mugdadi and Ibrahim A. Ahmad. A bandwidth selection for kernel density estimation of functions of random variables. Computational Statistics & Data Analysis, 47(1):49– 62, August 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002470.

Clark:2004:ENP

[600] Allan B. Clark and Andrew B. Lawson. An evaluation of nonparametric relative risk estimators for disease maps. Computational Statistics & Data Analysis, 47(1):63– 78, August 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002482.

Nagar:2004:DPP

[601] Daya K. Nagar, Jie Chen, and Arjun K. Gupta. Distribution and percentage points of the likelihood ratio statistic for testing circular symmetry. Computational Statistics & Data Analysis, 47 (1):79–89, August 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002639.

Prescott:2004:RDC

[602] Philip Prescott and Ralph Mansson. Robustness of diallel cross designs to the loss of one or more observations. Computational Statistics & Data Analysis, 47(1):91–109, August 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002688.

Cuevas:2004:ATF

[603] Antonio Cuevas, Manuel Febrero, and Ricardo Fraiman. An ANOVA test for functional data. Computational Statistics & Data Analysis, 47(1):111– 122, August 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730300269X.

Kukush:2004:CEI

[604] Alexander Kukush, Ivan Markovsky, and Sabine Van Huffel. Consistent estimation in an implicit quadratic measurement error model. Computational Statistics & Data Analysis, 47 (1):123–147, August 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002706.

Noh:2004:UMC

[605] Hyun Gon Noh, Moon Sup Song, and Sung Hyun Park. An unbiased method for constructing multilabel classification trees. Computational Statistics & Data Analysis, 47 (1):149–164, August 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002433.

Schwertman:2004:SMG

[606] Neil C. Schwertman, Margaret Ann Owens, and Robiah Adnan. A simple more general boxplot method for identifying outliers. Computational Statistics & Data Analysis, 47(1):165– 174, August 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002469.

Li:2004:CBE

[607] Lexin Li, R. Dennis Cook, and Christopher J. Nachtsheim. Clusterbased estimation for sufficient dimension reduction. Computational Statistics & Data Analysis, 47(1):175– 193, August 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002640.

Bartolucci:2004:TPA

[608] F. Bartolucci and L. Scaccia. Testing for positive association in contingency tables with fixed margins. Computational Statistics & Data Analysis, 47 (1):195–210, August 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002676.

Anonymous:2004:EBl

[609] Anonymous. Editorial board. Computational Statistics & Data Analysis, 47 (2):viii–ix, September 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002580.

Winker:2004:AOH

[610] Peter Winker and Manfred Gilli. Applications of optimization heuristics to estimation and modelling problems. Computational Statistics & Data Analysis, 47(2):211–223, September 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002718.

Cadima:2004:CAA

[611] Jorge Cadima, J. Orestes Cerdeira, and Manuel Minhoto. Computational aspects of algorithms for variable selection in the context of principal components. Computational Statistics & Data Analysis, 47(2):225–236, September 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300272X.

Woodruff:2004:EAM

[612] David L. Woodruff and Torsten Reiners. Experiments with, and on, algorithms for maximum likelihood clustering. Computational Statistics & Data Analysis, 47(2):237–253, September 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002731.

Alcock:2004:GEA

[613] Jamie Alcock and Kevin Burrage. A genetic estimation algorithm for parameters of stochastic ordinary differential equations. Computational Statistics & Data Analysis, 47(2):255– 275, September 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002743.

Baragona:2004:FPL

[614] R. Baragona, F. Battaglia, and D. Cucina. Fitting piecewise linear threshold autoregressive models by means of genetic algorithms. Computational Statistics & Data Analysis, 47(2):277–295, September 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002755.

Luebke:2004:GPO

[615] Karsten Luebke and Claus Weihs. Generation of prediction optimal projection on latent factors by a stochastic search algorithm. Computational Statistics & Data Analysis, 47(2):297– 310, September 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002767.

Orsenigo:2004:DSV

[616] Carlotta Orsenigo and Carlo Vercellis. Discrete support vector decision trees via tabu search. Computational Statistics & Data Analysis, 47(2):311– 322, September 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002779.

Roverato:2004:TMG

[617] Alberto Roverato and Sandra Paterlini. Technological modelling for graphical models: an approach based on genetic algorithms. Computational Statistics & Data Analysis, 47(2):323– 337, September 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002792.

Maringer:2004:FRR

[618] Dietmar G. Maringer. Finding the relevant risk factors for asset pricing. Computational Statistics & Data Analysis, 47(2):339–352, September 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002809.

Pattarin:2004:CFT

[619] Francesco Pattarin, Sandra Paterlini, and Tommaso Minerva. Clustering financial time series: an application to mutual funds style analysis. Computational Statistics & Data Analysis, 47(2):353–372, September 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002780.

Schlottmann:2004:HHA

[620] Frank Schlottmann and Detlef Seese. A hybrid heuristic approach to discrete multi-objective optimization of credit portfolios. Computational Statistics & Data Analysis, 47(2):373–399, September 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002810.

Anonymous:2004:EBm

[621] Anonymous. Editorial board. Computational Statistics & Data Analysis, 47 (3):viii–ix, October 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002804.

Salvador:2004:AMI

[622] Manuel Salvador and Pilar Gargallo. Automatic monitoring and intervention in multivariate dynamic linear models. Computational Statistics & Data Analysis, 47(3):401–431, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002925.

Tirler:2004:EKR

[623] Günter Tirler, Peter Dalgaard, Wolfgang Hörmann, and Josef Leydold. An error in the Kinderman–Ramage method and how to fix it. Computational Statistics & Data Analysis, 47 (3):433–440, October 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002937.

Lee:2004:NAG

[624] Jaeyong Lee and Yongdai Kim. A new algorithm to generate beta processes. Computational Statistics & Data Analysis, 47(3):441–453, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002986.

Yeredor:2004:HLS

[625] Arie Yeredor and Bart De Moor. On homogeneous least-squares problems and the inconsistency introduced by mis-constraining. Computational Statistics & Data Analysis, 47(3):455– 465, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303003001.

Loughin:2004:SCM

[626] Thomas M. Loughin. A systematic comparison of methods for combining p-values from independent tests. Computational Statistics & Data Analysis, 47(3):467–485, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002950.

Manteiga:2004:CSP

[627] W. González Manteiga, M. D. Martínez Miranda, and A. Pérez González. The choice of smoothing parameter in nonparametric regression through wild bootstrap. Computational Statistics & Data Analysis, 47(3):487–515, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002998.

Bhattacharjee:2004:EHR

[628] Arnab Bhattacharjee. Estimation in hazard regression models under ordered departures from proportionality. Computational Statistics & Data Analysis, 47(3):517–536, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303003013.

Abrahantes:2004:CUA

[629] José Cortiñas Abrahantes, Geert Molenberghs, Tomasz Burzykowski, Ziv Shkedy, Ariel Alonso Abad, and Didier Renard. Choice of units of analysis and modeling strategies in multilevel hierarchical models. Computational Statistics & Data Analysis, 47 (3):537–563, October 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303003025.

Chan:2004:SIG

[630] Jennifer S. K. Chan, Yeh Lam, and Doris Y. P. Leung. Statistical inference for geometric processes with gamma distributions. Computational Statistics & Data Analysis, 47(3):565– 581, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303003037.

Mendes:2004:DDE

[631] Beatriz Vaz de Melo Mendes and Hedibert Freitas Lopes. Data driven estimates for mixtures. Computational Statistics & Data Analysis, 47(3):583– 598, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303003050.

Li:2004:SCA

[632] Chin-Shang Li and Cheng Cheng. Stable classification with applications to microarray data. Computational Statistics & Data Analysis, 47(3):599– 609, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303003207.

Nguyen:2004:EPT

[633] Danh V. Nguyen. On estimating the proportion of true null hypotheses for false discovery rate controlling procedures in exploratory DNA microarray studies. Computational Statistics & Data Analysis, 47(3):611– 637, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000040.

Agresti:2004:EWM

[634] Alan Agresti, Brian Caffo, and Pamela Ohman-Strickland. Examples in which misspecification of a random effects distribution reduces efficiency, and possible remedies. Computational Statistics & Data Analysis, 47(3):639– 653, October 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303003190.

Anonymous:2004:EBn

[635] Anonymous. Editorial board. Computational Statistics & Data Analysis, 47 (4):vi–vii, November 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400307X.

Ruggoo:2004:BSO

[636] Arvind Ruggoo and Martina Vandebroek. Bayesian sequential D- D optimal model-robust designs. Computational Statistics & Data Analysis, 47(4):655–673, November 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947303002664.

Polzehl:2004:SSE

[637] Jörg Polzehl and Silvelyn Zwanzig. On a symmetrized simulation extrapolation estimator in linear errorsin-variables models. Computational Statistics & Data Analysis, 47(4):675– 688, November 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002846.

Neves:2004:RTA

[638] Cláudia Neves and M. I. Fraga Alves. Reiss and Thomas’ automatic selection of the number of extremes. Computational Statistics & Data Analysis, 47(4):689–704, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002858.

Brahim-Belhouari:2004:GPN

[639] Sofiane Brahim-Belhouari and Amine Bermak. Gaussian process for nonstationary time series prediction. Computational Statistics & Data Analysis, 47(4):705–712, November 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947304000301.

Chernov:2004:SEC

[640] N. Chernov and C. Lesort. Statistical efficiency of curve fitting algorithms. Computational Statistics & Data Analysis, 47(4):713–728, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002822.

Zwane:2004:SMC

[641] E. N. Zwane and P. G. M. van der Heijden. Semiparametric models for capture–recapture studies with covariates. Computational Statistics & Data Analysis, 47(4):729–743, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002834.

Andres:2004:CAP

[642] A. Martín Andrés, A. Silva Mato, J. M. Tapia García, and M. J. Sánchez Quevedo. Comparing the asymptotic power of exact tests in 2 × 2 tables. Computational Statistics & Data Analysis, 47(4):745–756, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730300286X.

Skipka:2004:UET

[643] G. Skipka, A. Munk, and G. Freitag. Unconditional exact tests for the difference of binomial probabilities — contrasted and compared. Computational Statistics & Data Analysis, 47(4):757– 773, November 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002871.

Naik:2004:ISI

[644] Prasad A. Naik and Chih-Ling Tsai. Isotonic single-index model for high-dimensional database marketing. Computational Statistics & Data Analysis, 47(4):775–790, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002883.

Adem:2004:ACM

[645] Jan Adem and Willy Gochet. Aggregating classifiers with mathematical programming. Computational Statistics & Data Analysis, 47(4):791– 807, November 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002901.

Wu:2004:IIE

[646] Jianrong Wu and A. C. M. Wong. Improved interval estimation for the twoparameter Birnbaum–Saunders distribution. Computational Statistics & Data Analysis, 47(4):809–821, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002913.

Boumaza:2004:DAI

[647] Rachid Boumaza. Discriminant analysis with independently repeated multivariate measurements: an L^{2} approach. Computational Statistics & Data Analysis, 47(4):823–843, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000167.

Chen:2004:OCI

[648] Hubert J. Chen and Shun-Yi Chen. Optimal confidence interval for the largest normal mean with unknown variance. Computational Statistics & Data Analysis, 47(4):845–866, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000179. See

erratum [1303].

Theobald:2004:BSF

[649] Chris M. Theobald and Mike Talbot. Bayesian selection of fertilizer level when crop price depends on quality. Computational Statistics & Data Analysis, 47(4):867–880, November 1, 2004. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002895.

Anonymous:2004:AId

[650] Anonymous. Author index. Computational Statistics & Data Analysis, 47(4):881–883, November 1, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947304003123.

Anonymous:2005:EBa

[651] Anonymous. Editorial board. Computational Statistics & Data Analysis, 48 (1):vii–viii, January 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003317.

Vinzi:2005:E

[652] Vincenzo Esposito Vinzi and Carlo Lauro. Editorial. Computational Statistics & Data Analysis, 48(1):1– 4, January 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001124.

Abraham:2005:DRA

[653] Bovas Abraham and Giovanni Merola. Dimensionality reduction approach to multivariate prediction. Computational Statistics & Data Analysis, 48 (1):5–16, January 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002962.

Bastien:2005:PGL

[654] Philippe Bastien, Vincenzo Esposito Vinzi, and Michel Tenenhaus. PLS generalised linear regression. Computational Statistics & Data Analysis, 48 (1):17–46, January 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000271.

Dijksterhuis:2005:CPA

[655] Garmt Dijksterhuis, Harald Martens, and Magni Martens. Combined Procrustes analysis and PLSR for internal and external mapping of data from multiple sources. Computational Statistics & Data Analysis, 48 (1):47–62, January 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002652.

Hanafi:2005:AAP

[656] M. Hanafi and E. M. Qannari. An alternative algorithm to the PLS B problem. Computational Statistics & Data Analysis, 48(1):63–67, January 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002275.

Kettaneh:2005:PPV

[657] Nouna Kettaneh, Anders Berglund, and Svante Wold. PCA and PLS with very large data sets. Computational Statistics & Data Analysis, 48 (1):69–85, January 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002949.

Li:2005:NLN

[658] Baibing Li, Per A. Hassel, A. Julian Morris, and Elaine B. Martin. A non-linear nested partial leastsquares algorithm. Computational Statistics & Data Analysis, 48(1):87– 101, January 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002391.

Martens:2005:RDM

[659] Harald Martens, Endre Anderssen, Arnar Flatberg, Lars Halvor Gidskehaug, Martin Høy, Frank Westad, Anette Thybo, and Magni Martens. Regression of a data matrix on descriptors of both its rows and of its columns via latent variables: L-PLSR. Computational Statistics & Data Analysis, 48 (1):103–123, January 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002378.

Milidiu:2005:DPT

[660] Ruy L. Milidiú and Raúl P. Rentería. DPLS and PPLS: two PLS algorithms for large data sets. Computational Statistics & Data Analysis, 48(1):125– 138, January 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002408.

Nocairi:2005:DLC

[661] Hicham Nocairi, El Mostafa Qannari, Evelyne Vigneau, and Dominique Bertrand. Discrimination on latent components with respect to patterns. application to multicollinear data. Computational Statistics & Data Analysis, 48(1):139–147, January 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002287.

Preda:2005:PRS

[662] C. Preda and G. Saporta. PLS regression on a stochastic process. Computational Statistics & Data Analysis, 48 (1):149–158, January 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303002366.

Tenenhaus:2005:PPM

[663] Michel Tenenhaus, Vincenzo Esposito Vinzi, Yves-Marie Chatelin, and Carlo Lauro. PLS path modeling. Computational Statistics & Data Analysis, 48 (1):159–205, January 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000519.

Wang:2005:IPL

[664] Huiwen Wang, Qiang Liu, and Yongping Tu. Interpretation of partial least-squares regression models with VARIMAX rotation. Computational Statistics & Data Analysis, 48(1):207– 219, January 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947303003049.

Anonymous:2005:EBb

[665] Anonymous. Editorial board. Computational Statistics & Data Analysis, 48 (2):viii–ix, February 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003457.

Boone:2005:ATA

[666] Edward L. Boone, Keying Ye, and Eric P. Smith. Assessment of two approximation methods for computing posterior model probabilities. Computational Statistics & Data Analysis, 48(2):221–234, February 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000210.

Turner:2005:IBM

[667] Heather Turner, Trevor Bailey, and Wojtek Krzanowski. Improved biclustering of microarray data demonstrated through systematic performance tests. Computational Statistics & Data Analysis, 48(2):235–254, February 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000295.

Loschi:2005:EPP

[668] R. H. Loschi and F. R. B. Cruz. Extension to the product partition model: computing the probability of a change. Computational Statistics & Data Analysis, 48(2):255–268, February 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000490.

Kim:2005:BAP

[669] Hea-Jung Kim. A Bayesian approach to paired comparison rankings based on a graphical model. Computational Statistics & Data Analysis, 48(2):269– 290, February 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000180.

Nashimoto:2005:MCP

[670] Kane Nashimoto and F. T. Wright. Multiple comparison procedures for detecting differences in simply ordered means. Computational Statistics & Data Analysis, 48(2):291–306, February 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000222.

Jung:2005:MLT

[671] Kang-Mo Jung. Multivariate leasttrimmed squares regression estimator. Computational Statistics & Data Analysis, 48(2):307–316, February 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000234.

Godtliebsen:2005:VDD

[672] Fred Godtliebsen and Tor Arne Øigård. A visual display device for significant features in complicated signals. Computational Statistics & Data Analysis, 48(2):317–343, February 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000246.

Sutradhar:2005:LIB

[673] Brajendra C. Sutradhar and Rahul Mukerjee. On likelihood inference in binary mixed model with an application to COPD data. Computational Statistics & Data Analysis, 48(2):345– 361, February 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400026X.

Levine:2005:ICH

[674] Richard A. Levine, Zhaoxia Yu, William G. Hanley, and John J. Nitao. Implementing componentwise Hastings algorithms. Computational Statistics & Data Analysis, 48(2):363– 389, February 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000283.

Kwan:2005:CSF

[675] Andy C. C. Kwan, Ah-Boon Sim, and Yangru Wu. A comparative study of the finite-sample performance of some portmanteau tests for randomness of a time series. Computational Statistics & Data Analysis, 48(2):391– 413, February 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000192.

LaRocca:2005:VSN

[676] Michele La Rocca and Cira Perna. Variable selection in neural network regression models with dependent data: a subsampling approach. Computational Statistics & Data Analysis, 48(2):415– 429, February 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000209.

Boginski:2005:SAF

[677] Vladimir Boginski, Sergiy Butenko, and Panos M. Pardalos. Statistical analysis of financial networks. Computational Statistics & Data Analysis, 48(2):431–443, February 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000258.

Anonymous:2005:EBc

[678] Anonymous. Editorial board. Computational Statistics & Data Analysis, 48(3):viii–ix, March 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003615.

Knusel:2005:ASD

[679] Leo Knüsel. On the accuracy of statistical distributions in Microsoft Excel 2003. Computational Statistics & Data Analysis, 48(3):445– 449, March 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000337.

Zhou:2005:GML

[680] Xiuqing Zhou and Jinde Wang. A genetic method of LAD estimation for models with censored data. Computational Statistics & Data Analysis, 48 (3):451–466, March 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000362.

Hens:2005:KWI

[681] Niel Hens, Marc Aerts, Geert Molenberghs, Herbert Thijs, and Geert Verbeke. Kernel weighted influence measures. Computational Statistics & Data Analysis, 48(3):467– 487, March 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000313.

Chiou:2005:SED

[682] Paul Chiou and Weiwen Miao. Shrinkage estimation for the difference between exponential guarantee time parameters. Computational Statistics & Data Analysis, 48(3):489– 507, March 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000325.

Singh:2005:EPE

[683] Umesh Singh, Pramod K. Gupta, and S. K. Upadhyay. Estimation of parameters for exponentiated-Weibull family under type-II censoring scheme. Computational Statistics & Data Analysis, 48(3):509–523, March 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000349.

DeIaco:2005:MPM

[684] S. De Iaco, M. Palma, and D. Posa. Modeling and prediction of multivariate space–time random fields. Computational Statistics & Data Analysis, 48 (3):525–547, March 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000350.

Zhang:2005:GNM

[685] Nansong Zhang, Bo cheng Wei, and Jin guan Lin. Generalized nonlinear models and variance function estimation. Computational Statistics & Data Analysis, 48(3):549– 570, March 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000507.

Warwick:2005:DBM

[686] J. Warwick. A data-based method for selecting tuning parameters in minimum distance estimators. Computational Statistics & Data Analysis, 48 (3):571–585, March 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000672.

Michailidis:2005:HAU

[687] George Michailidis and Jan De Leeuw. Homogeneity analysis using absolute deviations. Computational Statistics & Data Analysis, 48(3):587– 603, March 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000684.

Cai:2005:SII

[688] Yong Cai and K. Krishnamoorthy. A simple improved inferential method for some discrete distributions. Computational Statistics & Data Analysis, 48 (3):605–621, March 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000829.

Jeong:2005:BTI

[689] Hyeong Chul Jeong, Myoungshic Jhun, and Daehak Kim. Bootstrap tests for independence in two-way ordinal contingency tables. Computational Statistics & Data Analysis, 48 (3):623–631, March 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000830.

Hutter:2005:DMI

[690] Marcus Hutter and Marco Zaffalon. Distribution of mutual information from complete and incomplete data. Computational Statistics & Data Analysis, 48(3):633– 657, March 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000842.

Moustaki:2005:LCM

[691] Irini Moustaki and Ioulia Papageorgiou. Latent class models for mixed variables with applications in archaeometry. Computational Statistics & Data Analysis, 48(3):659– 675, March 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000374.

Anonymous:2005:EBd

[692] Anonymous. Editorial board. Computational Statistics & Data Analysis, 48(4):viii–ix, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003780.

Chang:2005:EWE

[693] Xiao-Wen Chang and Leming Qu. Erratum to “Wavelet estimation of partially linear models” [computational statistics and data analysis 47/1 (2004) 31–48]. Computational Statistics & Data Analysis, 48(4): 677, April 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002981. See

[598].

Ozonoff:2005:PCI

[694] Al Ozonoff, Marco Bonetti, Laura Forsberg, and Marcello Pagano. Power comparisons for an improved disease clustering test. Computational Statistics & Data Analysis, 48(4):679– 684, April 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000994.

Jank:2005:QMC

[695] Wolfgang Jank. Quasi-Monte Carlo sampling to improve the efficiency of Monte Carlo EM. Computational Statistics & Data Analysis, 48(4):685– 701, April 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001033.

Willems:2005:FRB

[696] Gert Willems and Stefan Van Aelst. Fast and robust bootstrap for LTS. Computational Statistics & Data Analysis, 48(4):703–715, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947304001057.

Li:2005:ACE

[697] Lei M. Li. An algorithm for computing exact least-trimmed squares estimate of simple linear regression with constraints. Computational Statistics & Data Analysis, 48(4):717– 734, April 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001082.

Congdon:2005:BPM

[698] P. Congdon. Bayesian predictive model comparison via parallel sampling. Computational Statistics & Data Analysis, 48(4):735– 753, April 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304000982.

Preisser:2005:ACF

[699] John S. Preisser and Daniel I. Garcia. Alternative computational formulae for generalized linear model diagnostics: identifying influential observations with SAS software. Computational Statistics & Data Analysis, 48 (4):755–764, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001008.

Moerbeek:2005:RPO

[700] M. Moerbeek. Robustness properties of A-, D-, and E-optimal designs for polynomial growth models with autocorrelated errors. Computational Statistics & Data Analysis, 48 (4):765–778, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001021.

Shi:2005:SAC

[701] Mingren Shi and Mark A. Lukas. Sensitivity analysis of constrained linear L_{1} regression: perturbations to response and predictor variables. Computational Statistics & Data Analysis, 48(4):779–802, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001045.

Canal:2005:NAC

[702] Luisa Canal. A normal approximation for the chi-square distribution. Computational Statistics & Data Analysis, 48 (4):803–808, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001069.

Oh:2005:HLP

[703] Hee-Seok Oh and Thomas C. M. Lee. Hybrid local polynomial wavelet shrinkage: wavelet regression with automatic boundary adjustment. Computational Statistics & Data Analysis, 48 (4):809–819, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001070.

Bartolucci:2005:UMD

[704] F. Bartolucci and L. Scaccia. The use of mixtures for dealing with nonnormal regression errors. Computational Statistics & Data Analysis, 48 (4):821–834, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001100.

Wust:2005:ILS

[705] Kirsten Wüst and Meinhard Kieser. Including long- and short-term data in blinded sample size recalculation for binary endpoints. Computational Statistics & Data Analysis, 48(4):835– 855, April 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001112.

Groenewald:2005:BCL

[706] Pieter C. N. Groenewald and Lucky Mokgatlhe. Bayesian computation for logistic regression. Computational Statistics & Data Analysis, 48(4):857– 868, April 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001148.

Lee:2005:ECR

[707] Jae Won Lee, Jung Bok Lee, Mira Park, and Seuck Heun Song. An extensive comparison of recent classification tools applied to microarray data. Computational Statistics & Data Analysis, 48 (4):869–885, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400101X.

Voutilainen:2005:SMC

[708] Arto Voutilainen and Jari P. Kaipio. Sequential Monte Carlo estimation of aerosol size distributions. Computational Statistics & Data Analysis, 48 (4):887–908, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304000854.

Anonymous:2005:AIa

[709] Anonymous. Author index. Computational Statistics & Data Analysis, 48 (4):909–911, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003834.

Anonymous:2005:EBe

[710] Anonymous. Editorial board. Computational Statistics & Data Analysis, 49(1):viii–ix, April 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000149.

Bognar:2005:BIS

[711] Matthew A. Bognar. Bayesian inference for spatially inhomogeneous pairwise interacting point processes. Computational Statistics & Data Analysis, 49(1):1–18, April 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001136.

Edlund:2005:CCE

[712] Ove Edlund and Håkan Ekblom. Computing the constrained M-estimates for regression. Computational Statistics & Data Analysis, 49(1):19– 32, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001173.

Neerchal:2005:IMC

[713] Nagaraj K. Neerchal and Jorge G. Morel. An improved method for the computation of maximum likeliood estimates for multinomial overdispersion models. Computational Statistics & Data Analysis, 49(1):33– 43, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001409.

Rodeiro:2005:EEE

[714] Carmen L. Vidal Rodeiro and Andrew B. Lawson. An evaluation of the edge effects in disease map modelling. Computational Statistics & Data Analysis, 49(1):45– 62, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001434.

Tritchler:2005:SCM

[715] David Tritchler, Shafagh Fallah, and Joseph Beyene. A spectral clustering method for microarray data. Computational Statistics & Data Analysis, 49 (1):63–76, April 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400115X.

Gao:2005:ECP

[716] Wei Gao and Ning-Zhong Shi. Estimating cell probabilities under orderrestricted odds ratios. Computational Statistics & Data Analysis, 49(1):77– 84, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001161.

Fu:2005:NAS

[717] Chong Yau Fu, Jeng-Hsiu Hung, Shih-Hua Liu, and Yung-Lin Chien. A new algorithm for solving binary discrimination in conditional logistic regression, with two choices of strata. Computational Statistics & Data Analysis, 49(1):85–97, April 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001185.

Preda:2005:CPR

[718] C. Preda and G. Saporta. Clusterwise PLS regression on a stochastic process. Computational Statistics & Data Analysis, 49(1):99– 108, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001203.

Nadarajah:2005:SPR

[719] Saralees Nadarajah. Sums, products, and ratios for the bivariate Lomax distribution. Computational Statistics & Data Analysis, 49(1):109– 129, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001215.

Huang:2005:FIE

[720] Yangxin Huang. On a family of interval estimators of effective doses. Computational Statistics & Data Analysis, 49 (1):131–146, April 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001227.

Jorgensen:2005:MML

[721] Murray Jorgensen. Minimum message length estimation using EM methods: a case study. Computational Statistics & Data Analysis, 49(1):147– 167, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001239.

Kauermann:2005:PSS

[722] Göran Kauermann. Penalized spline smoothing in multivariable survival models with varying coefficients. Computational Statistics & Data Analysis, 49(1):169–186, April 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001240.

Kundu:2005:GRD

[723] Debasis Kundu and Mohammad Z. Raqab. Generalized Rayleigh distribution: different methods of estimations. Computational Statistics & Data Analysis, 49(1):187– 200, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001410.

Goos:2005:MRM

[724] Peter Goos, André Kobilinsky, Timothy E. O’Brien, and Martina Vandebroek. Model-robust and modelsensitive designs. Computational Statistics & Data Analysis, 49(1):201– 216, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001422.

Nam:2005:ETT

[725] Jun mo Nam, Jinheum Kim, and Seungyeoun Lee. Equivalence of two treatments and sample size determination under exponential survival model with censoring. Computational Statistics & Data Analysis, 49(1):217– 226, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001446.

Klar:2005:TNM

[726] Bernhard Klar and Simos G. Meintanis. Tests for normal mixtures based on the empirical characteristic function. Computational Statistics & Data Analysis, 49(1):227– 242, April 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400146X.

Karavasilis:2005:SMS

[727] G. J. Karavasilis, V. K. Kotti, D. S. Tsitsis, V. G. Vassiliadis, and A. G. Rigas. Statistical methods and software for risk assessment: applications to a neurophysiological data set. Computational Statistics & Data Analysis, 49(1):243–263, April 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001458.

Dzemyda:2005:MDV

[728] Gintautas Dzemyda. Multidimensional data visualization in the statistical analysis of curricula. Computational Statistics & Data Analysis, 49 (1):265–281, April 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001197.

Anonymous:2005:EBf

[729] Anonymous. Editorial board. Computational Statistics & Data Analysis, 49 (2):vii–viii, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000575.

Belsley:2005:SSI

[730] David A. Belsley and Erricos John Kontoghiorghes. Second special issue on computational econometrics. Computational Statistics & Data Analysis, 49(2):283–285, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001471.

Barone-Adesi:2005:OPF

[731] Giovanni Barone-Adesi, Henrik Rasmussen, and Claudia Ravanelli. An option pricing formula for the GARCH diffusion model. Computational Statistics & Data Analysis, 49(2):287– 310, April 30, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001483.

Chipman:2005:OAL

[732] J. Chipman and P. Winker. Optimal aggregation of linear time series models. Computational Statistics & Data Analysis, 49(2):311– 331, April 30, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001495.

Cubadda:2005:SSI

[733] Gianluca Cubadda and Pieter Omtzigt. Small-sample improvements in the statistical analysis of seasonally cointegrated systems. Computational Statistics & Data Analysis, 49(2):333– 348, April 30, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001501.

Eklund:2005:ECR

[734] Bruno Eklund. Estimating confidence regions over bounded domains. Computational Statistics & Data Analysis, 49(2):349–360, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001525.

Flachaire:2005:BHR

[735] Emmanuel Flachaire. Bootstrapping heteroskedastic regression models: wild bootstrap vs. pairs bootstrap. Computational Statistics & Data Analysis, 49 (2):361–376, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001537.

Godfrey:2005:WBH

[736] L. G. Godfrey and A. R. Tremayne. The wild bootstrap and heteroskedasticityrobust tests for serial correlation in dynamic regression models. Computational Statistics & Data Analysis, 49 (2):377–395, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001549.

Jewitt:2005:CEC

[737] Giles Jewitt and J. Roderick Mc-Crorie. Computing estimates of continuous time macroeconometric models on the basis of discrete data. Computational Statistics & Data Analysis, 49 (2):397–416, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001550.

Joseph:2005:VRE

[738] Agnes S. Joseph and Jan F. Kiviet. Viewing the relative efficiency of IV estimators in models with lagged and instantaneous feedbacks. Computational Statistics & Data Analysis, 49 (2):417–444, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001562.

Khalaf:2005:ETS

[739] Lynda Khalaf and Maral Kichian. Exact tests of the stability of the Phillips curve: the Canadian case. Computational Statistics & Data Analysis, 49 (2):445–460, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001574.

Nankervis:2005:CAD

[740] John C. Nankervis. Computational algorithms for double bootstrap confidence intervals. Computational Statistics & Data Analysis, 49(2):461– 475, April 30, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001586.

Proietti:2005:NAD

[741] Tommaso Proietti. New algorithms for dating the business cycle. Computational Statistics & Data Analysis, 49 (2):477–498, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001598.

Rekkas:2005:TOI

[742] M. Rekkas and A. Wong. Thirdorder inference for the Weibull distribution. Computational Statistics & Data Analysis, 49(2):499– 525, April 30, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001604.

Sanford:2005:SBB

[743] Andrew D. Sanford and Gael M. Martin. Simulation-based Bayesian estimation of an affine term structure model. Computational Statistics & Data Analysis, 49(2):527– 554, April 30, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001616.

Stewart:2005:CSE

[744] Mark B. Stewart. A comparison of semiparametric estimators for the ordered response model. Computational Statistics & Data Analysis, 49 (2):555–573, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001628.

Swamy:2005:HSM

[745] P. A. V. B. Swamy, George S. Tavlas, and I-Lok Chang. How stable are monetary policy rules: estimating the timevarying coefficients in monetary policy reaction function for the US. Computational Statistics & Data Analysis, 49 (2):575–590, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400163X.

Vassiliou:2005:AAL

[746] E. Vassiliou and I. C. Demetriou. An adaptive algorithm for least squares piecewise monotonic data fitting. Computational Statistics & Data Analysis, 49(2):591–609, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001641.

Xekalaki:2005:EVF

[747] Evdokia Xekalaki and Stavros Degiannakis. Evaluating volatility forecasts in option pricing in the context of a simulated options market. Computational Statistics & Data Analysis, 49 (2):611–629, April 30, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001653.

Anonymous:2005:EBg

[748] Anonymous. Editorial board. Computational Statistics & Data Analysis, 49(3):ix–x, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000897.

Zhao:2005:CNA

[749] Yanli Zhao, Patricia M. Grambsch, and James D. Neaton. Comparison of numerical algorithms for bivariate sequential tests based on marginal criteria. Computational Statistics & Data Analysis, 49(3):631– 641, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001513.

Marhuenda:2005:CBR

[750] Y. Marhuenda, D. Morales, J. A. Pardo, and M. C. Pardo. Choosing the best Rukhin goodness-of-fit statistics. Computational Statistics & Data Analysis, 49(3):643– 662, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001756.

Guedon:2005:HHM

[751] Yann Guédon. Hidden hybrid Markov/ semi-Markov chains. Computational Statistics & Data Analysis, 49(3):663– 688, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001665.

Cysneiros:2005:RMS

[752] Francisco José A. Cysneiros and Gilberto A. Paula. Restricted methods in symmetrical linear regression models. Computational Statistics & Data Analysis, 49(3):689– 708, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001689.

Zhang:2005:LRT

[753] Jin Zhang and Yuehua Wu. Likelihoodratio tests for normality. Computational Statistics & Data Analysis, 49 (3):709–721, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001690.

Munk:2005:TGH

[754] A. Munk, G. Skipka, and B. Stratmann. Testing general hypotheses under binomial sampling: the two sample case — asymptotic theory and exact procedures. Computational Statistics & Data Analysis, 49 (3):723–739, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001707.

Sigut:2005:LDA

[755] J. Sigut, J. Piñeiro, L. Moreno, J. Estevez, R. Aguilar, and R. Marichal. A large deviation approach to normality testing. Computational Statistics & Data Analysis, 49(3):741– 756, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001720.

Kim:2005:IBM

[756] Yuwon Kim and Ja-Yong Koo. Inverse boosting for monotone regression functions. Computational Statistics & Data Analysis, 49(3):757– 770, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001732. See

erratum [826].

Ogasawara:2005:ARA

[757] Haruhiko Ogasawara. Asymptotic robustness of the asymptotic biases in structural equation modeling. Computational Statistics & Data Analysis, 49 (3):771–783, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001744.

Takane:2005:ERA

[758] Yoshio Takane and Heungsun Hwang. An extended redundancy analysis and its applications to two practical examples. Computational Statistics & Data Analysis, 49(3):785– 808, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001768.

Olive:2005:TSR

[759] David J. Olive. Two simple resistant regression estimators. Computational Statistics & Data Analysis, 49 (3):809–819, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400177X.

Austin:2005:BMM

[760] Peter C. Austin and Michael D. Escobar. Bayesian modeling of missing data in clinical research. Computational Statistics & Data Analysis, 49 (3):821–836, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304001781.

Diaz-Garcia:2005:SPM

[761] José A. Díaz-García, Rogelio Ramos-Quiroga, and Enrique Cabrera-Vicencio. Stochastic programming methods in the response surface methodology. Computational Statistics & Data Analysis, 49(3):837– 848, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001793.

Peters:2005:GIC

[762] A. Peters, T. Hothorn, and B. Lausen. Generalised indirect classifiers. Computational Statistics & Data Analysis, 49(3):849–861, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400180X.

Aslam:2005:RTP

[763] Shagufta Aslam and David M. Rocke. A robust testing procedure for the equality of covariance matrices. Computational Statistics & Data Analysis, 49(3):863–874, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001811.

Cheng:2005:RRD

[764] Tsung-Chi Cheng. Robust regression diagnostics with data transformations. Computational Statistics & Data Analysis, 49(3):875– 891, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001823.

Choi:2005:RTA

[765] Yunhee Choi, Hongshik Ahn, and James J. Chen. Regression trees for analysis of count data with extra Poisson variation. Computational Statistics & Data Analysis, 49(3):893– 915, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001975.

DElia:2005:MMP

[766] Angela D’Elia and Domenico Piccolo. A mixture model for preferences data analysis. Computational Statistics & Data Analysis, 49(3):917– 934, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001987.

Sherman:2005:ACI

[767] Michael Sherman and Edward Carlstein. Addendum to “Confidence intervals based on estimators with unknown rates of convergence” [comp. Stat. and data 46/1 (2003) 123–139]. Computational Statistics & Data Analysis, 49(3):935, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002002.

Sartori:2005:MIM

[768] Nicola Sartori, Alberto Salvan, and Karl Thomaseth. Multiple imputation of missing values in a cancer mortality analysis with estimated exposure dose. Computational Statistics & Data Analysis, 49(3):937– 953, June 1, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001999.

Silva:2005:SAN

[769] M. Eduarda Silva, Teresa Mendonça, Isabel Silva, and Hugo Magalhães. Statistical analysis of neuromuscular blockade response: contributions to an automatic controller calibration. Computational Statistics & Data Analysis, 49(3):955–968, June 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304001719.

Anonymous:2005:EBh

[770] Anonymous. Editorial board. Computational Statistics & Data Analysis, 49(4):viii–ix, June 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001076.

Hand:2005:OMC

[771] David J. Hand and Wojtek J. Krzanowski. Optimising k-means clustering results with standard software packages. Computational Statistics & Data Analysis, 49(4):969– 973, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002038.

Peres-Neto:2005:HMP

[772] Pedro R. Peres-Neto, Donald A. Jackson, and Keith M. Somers. How many principal components? stopping rules for determining the number of non-trivial axes revisited. Computational Statistics & Data Analysis, 49 (4):974–997, June 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002014.

Egmont-Petersen:2005:CIP

[773] M. Egmont-Petersen, A. Feelders, and B. Baesens. Confidence intervals for probabilistic network classifiers. Computational Statistics & Data Analysis, 49(4):998–1019, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400204X.

Kuhn:2005:MLE

[774] E. Kuhn and M. Lavielle. Maximum likelihood estimation in nonlinear mixed effects models. Computational Statistics & Data Analysis, 49 (4):1020–1038, June 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002221.

Liang:2005:EME

[775] Faming Liang and Chuanhai Liu. Efficient MCMC estimation of discrete distributions. Computational Statistics & Data Analysis, 49(4):1039– 1052, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002415.

Christophi:2005:ADF

[776] Costas A. Christophi and Reza Modarres. Approximating the distribution function of risk. Computational Statistics & Data Analysis, 49(4):1053– 1067, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002452.

Hothorn:2005:BCB

[777] Torsten Hothorn and Berthold Lausen. Bundling classifiers by bagging trees. Computational Statistics & Data Analysis, 49(4):1068–1078, June 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947304002051.

Begum:2005:MMP

[778] Nelufa Begum and Maxwell L. King. Most mean powerful test of a composite null against a composite alternative. Computational Statistics & Data Analysis, 49(4):1079– 1104, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400221X.

Lee:2005:GMD

[779] Seong Keon Lee. On generalized multivariate decision tree by using GEE. Computational Statistics & Data Analysis, 49(4):1105– 1119, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002063.

Paulino:2005:BAC

[780] Carlos Daniel Paulino, Giovani Silva, and Jorge Alberto Achcar. Bayesian analysis of correlated misclassified binary data. Computational Statistics & Data Analysis, 49(4):1120– 1131, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002233.

Koutrouvelis:2005:ETP

[781] Ioannis A. Koutrouvelis, George C. Canavos, and Simos G. Meintanis. Estimation in the three-parameter inverse Gaussian distribution. Computational Statistics & Data Analysis, 49 (4):1132–1147, June 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002245.

Shore:2005:RMM

[782] Haim Shore. Response modeling methodology (RMM) — maximum likelihood estimation procedures. Computational Statistics & Data Analysis, 49(4):1148–1172, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002257.

Varin:2005:PLI

[783] Cristiano Varin, Gudmund Høst, and Øivind Skare. Pairwise likelihood inference in spatial generalized linear mixed models. Computational Statistics & Data Analysis, 49(4):1173– 1191, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002403.

Park:2005:BTA

[784] Soo Jung Park, Dong Wan Shin, Byeong Uk Park, Woo Chul Kim, and Man-Suk Oh. Bayesian test for asymmetry and nonstationarity in MTAR model with possibly incomplete data. Computational Statistics & Data Analysis, 49(4):1192– 1204, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002427. See

corrigendum [827].

Kojadinovic:2005:RMS

[785] Ivan Kojadinovic. Relevance measures for subset variable selection in regression problems based on k-additive mutual information. Computational Statistics & Data Analysis, 49(4):1205– 1227, June 15, 2005. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002439.

Varma:2005:AUP

[786] Vishal Varma and Don R. Halverson. Applications of unbiased perturbations towards quantifying robustness with pragmatic geometric methods. Computational Statistics & Data Analysis, 49 (4):1228–1243, June 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002440.

McCullough:2005:ASP

[787] B. D. McCullough and Berry Wilson. On the accuracy of statistical procedures in Microsoft Excel 2003. Computational Statistics & Data Analysis, 49 (4):1244–1252, June 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002026.

Anonymous:2005:AIb

[788] Anonymous. Author index. Computational Statistics & Data Analysis, 49 (4):1253–1255, June 15, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001118.

Anonymous:2006:EBa

[789] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50 (1):vii–viii, January 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001817.

Barlow:2006:SIM

[790] Jesse L. Barlow, Haesun Park, Patrick J. F. Groenen, and Hongyuan Zha. 2nd special issue on matrix computations and statistics. Computational Statistics & Data Analysis, 50(1):1– 4, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002488.

Chang:2006:CHE

[791] Xiao-Wen Chang. Computation of Huber’s M-estimates for a block-angular regression problem. Computational Statistics & Data Analysis, 50(1):5– 20, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002397.

deLeeuw:2006:PCA

[792] Jan de Leeuw. Principal component analysis of binary data by iterated singular value decomposition. Computational Statistics & Data Analysis, 50 (1):21–39, January 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002300.

Dax:2006:SLI

[793] Achiya Dax. The ℓ_{1} solution of linear inequalities. Computational Statistics & Data Analysis, 50(1):40–60, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002269.

Eilers:2006:FCS

[794] Paul H. C. Eilers, Iain D. Currie, and Maria Durbán. Fast and compact smoothing on large multidimensional grids. Computational Statistics & Data Analysis, 50(1):61–76, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002270.

Evangelaras:2006:CBG

[795] H. Evangelaras and C. Koukouvinos. A comparison between the Gröbner bases approach and hidden projection properties in factorial designs. Computational Statistics & Data Analysis, 50 (1):77–88, January 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947303002974.

Gower:2006:AML

[796] John C. Gower. An application of the modified Leverrier–Faddeev algorithm to the spectral decomposition of symmetric block-circulant matrices. Computational Statistics & Data Analysis, 50(1):89–106, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002294.

Gardner:2006:SCV

[797] Sugnet Gardner, John C. Gower, and N. J. le Roux. A synthesis of canonical variate analysis, generalised canonical correlation and Procrustes analysis. Computational Statistics & Data Analysis, 50(1):107–134, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002282.

Lewis:2006:SAS

[798] R. M. Lewis and M. W. Trosset. Sensitivity analysis of the strain criterion for multidimensional scaling. Computational Statistics & Data Analysis, 50(1):135–153, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002312.

Li:2006:SEA

[799] Baibing Li. Sign eigenanalysis and its applications to optimization problems and robust statistics. Computational Statistics & Data Analysis, 50(1):154– 162, January 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002324.

Li:2006:TWP

[800] Jia Li and Hongyuan Zha. Twoway Poisson mixture models for simultaneous document classification and word clustering. Computational Statistics & Data Analysis, 50(1):163– 180, January 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002336.

Markovsky:2006:EWW

[801] Ivan Markovsky, Maria Luisa Rastello, Amedeo Premoli, Alexander Kukush, and Sabine Van Huffel. The elementwise weighted total least-squares problem. Computational Statistics & Data Analysis, 50(1):181–209, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002348.

Stegeman:2006:KCU

[802] Alwin Stegeman and Jos M. F. Ten Berge. Kruskal’s condition for uniqueness in Candecomp/Parafac when ranks and k-ranks coincide. Computational Statistics & Data Analysis, 50(1):210–220, January 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794730400235X.

Takane:2006:IMG

[803] Yoshio Takane, Haruo Yanai, and Heungsun Hwang. An improved method for generalized constrained canonical correlation analysis. Computational Statistics & Data Analysis, 50(1):221– 241, January 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002361.

Trendafilov:2006:PGA

[804] Nickolay T. Trendafilov and Ian T. Jolliffe. Projected gradient approach to the numerical solution of the SCoTLASS. Computational Statistics & Data Analysis, 50(1):242–253, January 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002373.

Yamada:2006:GCM

[805] Shu Yamada, Michiyo Matsui, Tomomi Matsui, Dennis K. J. Lin, and Takenori Takahashi. A general construction method for mixed-level supersaturated design. Computational Statistics & Data Analysis, 50(1):254– 265, January 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002385.

Anonymous:2006:AR

[806] Anonymous. Acknowledgement referees of 2004. Computational Statistics & Data Analysis, 50(2):I–III, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002033.

Anonymous:2006:EBb

[807] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50 (2):viii–ix, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001933.

Aldrin:2006:IPP

[808] Magne Aldrin. Improved predictions penalizing both slope and curvature in additive models. Computational Statistics & Data Analysis, 50(2):267– 284, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002476.

Wu:2006:RAD

[809] Jixiang Wu, Dongfeng Wu, Johnie N. Jenkins, and Jack C. McCarty, Jr. A recursive approach to detect multivariable conditional variance components and conditional random effects. Computational Statistics & Data Analysis, 50(2):285–300, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002518.

Bustos:2006:SFP

[810] Oscar H. Bustos and Alejandro C. Frery. Statistical functions and procedures in IDL 5.6 and 6.0. Computational Statistics & Data Analysis, 50(2):301–310, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794730400266X.

Ip:2006:ACN

[811] W. C. Ip, Heung Wong, J. Z. Pan, and D. F. Li. The asymptotic convexity of the negative likelihood function of GARCH models. Computational Statistics & Data Analysis, 50(2):311– 331, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002671.

Lim:2006:PPC

[812] Johan Lim. Permutation procedures with censored data. Computational Statistics & Data Analysis, 50(2):332– 345, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002725.

Congdon:2006:BMC

[813] Peter Congdon. Bayesian model choice based on Monte Carlo estimates of posterior model probabilities. Computational Statistics & Data Analysis, 50(2):346–357, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002464.

Husson:2006:MGI

[814] F. Husson and J. Pagès. INDSCAL model: geometrical interpretation and methodology. Computational Statistics & Data Analysis, 50(2):358–378, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002506.

Demetrescu:2006:EGN

[815] Matei Demetrescu. An extension of the Gauss–Newton algorithm for estimation under asymmetric loss. Computational Statistics & Data Analysis, 50(2):379–401, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400252X.

Khan:2006:CIR

[816] B. U. Khan and S. E. Ahmed. Comparisons of improved risk estimators of the multivariate mean vector. Computational Statistics & Data Analysis, 50(2):402–421, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002531.

Congdon:2006:MNP

[817] Peter Congdon. A model for nonparametric spatially varying regression effects. Computational Statistics & Data Analysis, 50(2):422–445, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002634.

vanderKooij:2006:LMC

[818] Anita J. van der Kooij, Jacqueline J. Meulman, and Willem J. Heiser. Local minima in categorical multiple regression. Computational Statistics & Data Analysis, 50(2):446–462, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002646.

Yu:2006:ESG

[819] Binbing Yu. Estimation of shared gamma frailty models by a modified EM algorithm. Computational Statistics & Data Analysis, 50(2):463– 474, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002658.

Brookhart:2006:SMS

[820] M. Alan Brookhart and Mark J. van der Laan. A semiparametric model selection criterion with applications to the marginal structural model. Computational Statistics & Data Analysis, 50(2):475–498, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002683.

Landgrebe:2006:EDA

[821] Jobst Landgrebe, Frank Bretz, and Edgar Brunner. Efficient design and analysis of two colour factorial microarray experiments. Computational Statistics & Data Analysis, 50(2):499– 517, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002695.

DeCook:2006:IDE

[822] Rhonda DeCook, Dan Nettleton, Carol Foster, and Eve S. Wurtele. Identifying differentially expressed genes in unreplicated multiple-treatment microarray timecourse experiments. Computational Statistics & Data Analysis, 50(2):518–532, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947304002701.

Gey:2006:BIR

[823] Servane Gey and Jean-Michel Poggi. Boosting and instability for regression trees. Computational Statistics & Data Analysis, 50(2):533–550, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002713.

Priebe:2006:SMR

[824] Carey E. Priebe, Michael I. Miller, and J. Tilak Ratnanather. Segmenting magnetic resonance images via hierarchical mixture modelling. Computational Statistics & Data Analysis, 50(2):551–567, January 30, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002737.

Reisen:2006:ESF

[825] Valderio Anselmo Reisen, Alexandre L. Rodrigues, and Wilfredo Palma. Estimation of seasonal fractionally integrated processes. Computational Statistics & Data Analysis, 50(2):568– 582, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400249X.

Kim:2006:EIB

[826] Yuwon Kim and Ja-Yong Koo. Erratum to “Inverse boosting for monotone regression functions” [comput. statist. data anal. 49 (2005) 757–770]. Computational Statistics & Data Analysis, 50(2):583, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001490. See

[756].

Park:2006:CBT

[827] Soo Jung Park, Dong Wan Shin, Byeong Uk Park, Woo Chul Kim, and Man-Suk Oh. Corrigendum to “Bayesian test for asymmetry and nonstationarity in MTAR model with possibly incomplete data”: [computational statistics and data analysis 49/4 (2005) 1192–1204]. Computational Statistics & Data Analysis, 50 (2):584, January 30, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001611. See

[784].

Anonymous:2006:EBc

[828] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50 (3):viii–ix, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002161.

Qiu:2006:SIP

[829] Weiliang Qiu and Harry Joe. Separation index and partial membership for clustering. Computational Statistics & Data Analysis, 50(3):585– 603, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002749.

Martin:2006:RAC

[830] Donald E. K. Martin. A recursive algorithm for computing the distribution of the number of successes in higherorder Markovian trials. Computational Statistics & Data Analysis, 50(3):604– 610, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002890.

Kim:2006:DCA

[831] Seong-Ho Kim and Sung-Ho Kim. A divide-and-conquer approach in applying EM for large recursive models with incomplete categorical data. Computational Statistics & Data Analysis, 50(3):611–641, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002907.

Turlach:2006:EFA

[832] Berwin A. Turlach. An even faster algorithm for ridge regression of reduced rank data. Computational Statistics & Data Analysis, 50(3):642–658, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002919.

Kuo:2006:ICG

[833] Lynn Kuo and Tae Young Yang. An improved collapsed Gibbs sampler for Dirichlet process mixing models. Computational Statistics & Data Analysis, 50(3):659–674, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002968.

Liang:2006:EPL

[834] Hua Liang. Estimation in partially linear models and numerical comparisons. Computational Statistics & Data Analysis, 50(3):675–687, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003172.

Pena:2006:BCE

[835] Daniel Peña and Dolores Redondas. Bayesian curve estimation by model averaging. Computational Statistics & Data Analysis, 50(3):688–709, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002920.

Kim:2006:MPP

[836] Jinseog Kim and Yongdai Kim. Maximum a posteriori pruning on decision trees and its application to bootstrap BUMPing. Computational Statistics & Data Analysis, 50(3):710– 719, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304002932.

Lindsey:2006:MDC

[837] J. K. Lindsey and P. J. Lindsey. Multivariate distributions with correlation matrices for nonlinear repeated measurements. Computational Statistics & Data Analysis, 50(3):720–732, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002944.

Brys:2006:RMT

[838] Guy Brys, Mia Hubert, and Anja Struyf. Robust measures of tail weight. Computational Statistics & Data Analysis, 50(3):733–759, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002956.

Jang:2006:NDE

[839] Woncheol Jang. Nonparametric density estimation and clustering in astronomical sky surveys. Computational Statistics & Data Analysis, 50(3):760– 774, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400297X.

Bonett:2006:ACI

[840] Douglas G. Bonett. Approximate confidence interval for standard deviation of nonnormal distributions. Computational Statistics & Data Analysis, 50(3):775–782, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304002993.

Boulesteix:2006:IIP

[841] Anne-Laure Boulesteix and Gerhard Tutz. Identification of interaction patterns and classification with applications to microarray data. Computational Statistics & Data Analysis, 50(3):783–802, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003007.

Lim:2006:RED

[842] Dong Hoon Lim. Robust edge detection in noisy images. Computational Statistics & Data Analysis, 50(3):803– 812, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003019.

Yuan:2006:GQS

[843] Ming Yuan. GACV for quantile smoothing splines. Computational Statistics & Data Analysis, 50(3):813– 829, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003184.

Jansen:2006:NSM

[844] Ivy Jansen, Niel Hens, Geert Molenberghs, Marc Aerts, Geert Verbeke, and Michael G. Kenward. The nature of sensitivity in monotone missing not at random models. Computational Statistics & Data Analysis, 50(3):830– 858, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003196.

Hartley:2006:HDR

[845] Roger Hartley and Gauthier Lanot. Heterogeneous demand responses to discrete price changes: an application to the purchase of lottery tickets. Computational Statistics & Data Analysis, 50(3):859–877, February 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003020.

Lafratta:2006:EEM

[846] Giovanni Lafratta. Efficiency evaluation of MEV spatial sampling strategies: a scenario analysis. Computational Statistics & Data Analysis, 50(3):878–890, February 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947304003202.

Anonymous:2006:EBd

[847] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50 (4):viii–ix, February 24, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002410.

Ogasawara:2006:AES

[848] Haruhiko Ogasawara. Asymptotic expansion of the sample correlation coefficient under nonnormality. Computational Statistics & Data Analysis, 50(4):891–910, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003251.

Lin:2006:ETI

[849] Tsai-Yu Lin and Chen-Tuo Liao. A β-expectation tolerance interval for general balanced mixed linear models. Computational Statistics & Data Analysis, 50(4):911–925, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003688.

Wasito:2006:NNL

[850] Ito Wasito and Boris Mirkin. Nearest neighbours in least-squares data imputation algorithms with different missing patterns. Computational Statistics & Data Analysis, 50(4):926–949, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003706.

Nadarajah:2006:IML

[851] Saralees Nadarajah. Information matrices for Laplace and Pareto mixtures. Computational Statistics & Data Analysis, 50(4):950–966, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003718. See

[1213].

Brezger:2006:GSA

[852] Andreas Brezger and Stefan Lang. Generalized structured additive regression based on Bayesian P - splines. Computational Statistics & Data Analysis, 50(4):967–991, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003214.

Bertrand:2006:LMR

[853] P. Bertrand and G. Bel Mufti. Loevinger’s measures of rule quality for assessing cluster stability. Computational Statistics & Data Analysis, 50(4):992–1015, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003226.

Ayala:2006:CSP

[854] G. Ayala, I. Epifanio, A. Simó, and V. Zapater. Clustering of spatial point patterns. Computational Statistics & Data Analysis, 50(4):1016– 1032, February 24, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003238.

Wang:2006:MCM

[855] Zhihui Wang, A. F. Desmond, and Xuewen Lu. Modified censored moment estimation for the twoparameter Birnbaum–Saunders distribution. Computational Statistics & Data Analysis, 50(4):1033–1051, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400324X.

Liu:2006:OEA

[856] Z. Liu, J. Almhana, V. Choulakian, and R. McGorman. Online EM algorithm for mixture with application to Internet traffic modeling. Computational Statistics & Data Analysis, 50(4):1052–1071, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003263.

Tai:2006:CTD

[857] John Jen Tai and Chia-Ding Hou. On the combination of transmission/ disequilibrium test and mean test for linkage detection using affected sib pairs. Computational Statistics & Data Analysis, 50(4):1072–1089, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400338X.

Yang:2006:ELC

[858] Chih-Chien Yang. Evaluating latent class analysis models in qualitative phenotype identification. Computational Statistics & Data Analysis, 50(4):1090–1104, February 24, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947304003391.

Salmaso:2006:NIC

[859] Luigi Salmaso and Aldo Solari. Nonparametric iterated combined tests for genetic differentiation. Computational Statistics & Data Analysis, 50(4):1105– 1112, February 24, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003548.

Lee:2006:MCC

[860] Tian-Shyug Lee, Chih-Chou Chiu, Yu-Chao Chou, and Chi-Jie Lu. Mining the customer credit using classification and regression tree and multivariate adaptive regression splines. Computational Statistics & Data Analysis, 50(4):1113–1130, February 24, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400355X.

Hollander:2006:EFF

[861] Norbert Holländer and Martin Schumacher. Estimating the functional form of a continuous covariate’s effect on survival time. Computational Statistics & Data Analysis, 50(4):1131– 1151, February 24, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730400369X.

Anonymous:2006:EBe

[862] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50(5):viii–ix, March 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002641.

Abellana:2006:ICR

[863] Rosa M. Abellana, Josep L. Carrasco, Lluís Jover, and Carlos Ascaso. Improving the convergence rate in conditional autoregressive models. Computational Statistics & Data Analysis, 50 (5):1153–1163, March 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003901.

Du:2006:TGP

[864] C. Du, D. Kurowicka, and R. M. Cooke. Techniques for generic probabilistic inversion. Computational Statistics & Data Analysis, 50(5):1164– 1187, March 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000034.

Hu:2006:HEA

[865] Tianming Hu and Sam Yuan Sung. A hybrid EM approach to spatial clustering. Computational Statistics & Data Analysis, 50(5):1188– 1205, March 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003974.

Nadarajah:2006:RLR

[866] Saralees Nadarajah and Arjun K. Gupta. On the ratio of logistic random variables. Computational Statistics & Data Analysis, 50(5):1206– 1219, March 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003950.

Paterlini:2006:DEP

[867] Sandra Paterlini and Thiemo Krink. Differential evolution and particle swarm optimisation in partitional clustering. Computational Statistics & Data Analysis, 50(5):1220– 1247, March 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003962.

Seidel:2006:ECN

[868] Wilfried Seidel, Krunoslav Sever, and Hana Ševčiková. Efficient calculation of the NPMLE of a mixing distribution for mixtures of exponentials. Computational Statistics & Data Analysis, 50 (5):1248–1271, March 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003925.

Tan:2006:HMR

[869] Ming Tan, Guo-Liang Tian, and Kai Wang Ng. Hierarchical models for repeated binary data using the IBF sampler. Computational Statistics & Data Analysis, 50(5):1272– 1286, March 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003986.

VanAelst:2006:LGU

[870] Stefan Van Aelst, Xiaogang (Steven) Wang, Ruben H. Zamar, and Rong Zhu. Linear grouping using orthogonal regression. Computational Statistics & Data Analysis, 50(5):1287– 1312, March 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730400372X.

Basak:2006:SPT

[871] Indrani Basak, Prasanta Basak, and N. Balakrishnan. On some predictors of times to failure of censored items in progressively censored samples. Computational Statistics & Data Analysis, 50(5):1313– 1337, March 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000265.

Kehl:2006:RIC

[872] Victoria Kehl and Kurt Ulm. Responder identification in clinical trials with censored data. Computational Statistics & Data Analysis, 50 (5):1338–1355, March 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003913.

Lee:2006:ALI

[873] Sik-Yum Lee, Bin Lu, and Xin-Yuan Song. Assessing local influence for nonlinear structural equation models with ignorable missing data. Computational Statistics & Data Analysis, 50 (5):1356–1377, March 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003883.

Zhou:2006:SPS

[874] Xiao-Hua Zhou and Hua Liang. Semiparametric single-index two-part regression models. Computational Statistics & Data Analysis, 50(5):1378– 1390, March 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003937.

Anonymous:2006:EBf

[875] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50 (6):viii–ix, March 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002859.

Barreto:2006:LMS

[876] Humberto Barreto and David Maharry. Least median of squares and regression through the origin. Computational Statistics & Data Analysis, 50(6):1391– 1397, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000046.

Brodin:2006:QEB

[877] Erik Brodin. On quantile estimation by bootstrap. Computational Statistics & Data Analysis, 50(6):1398– 1406, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001738.

Chae:2006:EUP

[878] Seong S. Chae and William D. Warde. Effect of using principal coordinates and principal components on retrieval of clusters. Computational Statistics & Data Analysis, 50(6):1407– 1417, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000289.

Chen:2006:TFT

[879] Colin Chen. Tests of fit for the three-parameter lognormal distribution. Computational Statistics & Data Analysis, 50(6):1418–1440, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000277.

Choulakian:2006:NPP

[880] V. Choulakian. L_{1}-norm projection pursuit principal component analysis. Computational Statistics & Data Analysis, 50(6):1441–1451, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000241.

Coppi:2006:FUC

[881] Renato Coppi and Pierpaolo D’Urso. Fuzzy unsupervised classification of multivariate time trajectories with the Shannon entropy regularization. Computational Statistics & Data Analysis, 50(6):1452–1477, March 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794730500023X.

DellAquila:2006:SBR

[882] Rosario Dell’Aquila and Elvezio Ronchetti. Stock and bond return predictability: the discrimination power of model selection criteria. Computational Statistics & Data Analysis, 50(6):1478– 1495, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000022.

DUrso:2006:WFM

[883] Pierpaolo D’Urso and Paolo Giordani. A weighted fuzzy c-means clustering model for fuzzy data. Computational Statistics & Data Analysis, 50(6):1496– 1523, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304003949.

Hafidi:2006:CAC

[884] B. Hafidi and A. Mkhadri. A corrected Akaike criterion based on Kullback’s symmetric divergence: applications in time series, multiple and multivariate regression. Computational Statistics & Data Analysis, 50(6):1524– 1550, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000083.

Nofuentes:2006:CTB

[885] José Antonio Roldán Nofuentes and Juan de Dios Luna del Castillo. Comparing two binary diagnostic tests in the presence of verification bias. Computational Statistics & Data Analysis, 50(6):1551– 1564, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000071.

Orbe:2006:AAS

[886] Jesus Orbe and Vicente Núñez-Antón. Alternative approaches to study lifetime data under different scenarios: from the PH to the modified semiparametric AFT model. Computational Statistics & Data Analysis, 50(6):1565– 1582, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000253.

Schmid:2006:NEC

[887] Friedrich Schmid and Axel Schmidt. Nonparametric estimation of the coefficient of overlapping — theory and empirical application. Computational Statistics & Data Analysis, 50(6):1583– 1596, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000290.

Wang:2006:ARD

[888] Shun-Fang Wang, Nian-Sheng Tang, and Xue-Ren Wang. Analysis of the risk difference of marginal and conditional probabilities in an incomplete correlated 2 × 2 table. Computational Statistics & Data Analysis, 50(6):1597– 1614, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000058.

Hu:2006:APD

[889] Xiaomi Hu. Approximation of partial distribution in renewal function calculation. Computational Statistics & Data Analysis, 50(6):1615– 1624, March 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730500006X.

Anonymous:2006:EBg

[890] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50(7):viii–ix, April 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305003385.

Bocci:2006:MMC

[891] Laura Bocci, Donatella Vicari, and Maurizio Vichi. A mixture model for the classification of three-way proximity data. Computational Statistics & Data Analysis, 50(7):1625– 1654, April 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000368.

Browne:2006:MAC

[892] William J. Browne. MCMC algorithms for constrained variance matrices. Computational Statistics & Data Analysis, 50(7):1655– 1677, April 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500037X.

Raggi:2006:CSV

[893] Davide Raggi and Silvano Bordignon. Comparing stochastic volatility models through Monte Carlo simulations. Computational Statistics & Data Analysis, 50(7):1678– 1699, April 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000332.

Tomasi:2006:CAF

[894] Giorgio Tomasi and Rasmus Bro. A comparison of algorithms for fitting the PARAFAC model. Computational Statistics & Data Analysis, 50 (7):1700–1734, April 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003895.

Bretz:2006:EWT

[895] Frank Bretz. An extension of the Williams trend test to general unbalanced linear models. Computational Statistics & Data Analysis, 50 (7):1735–1748, April 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000344.

Gupta:2006:PPD

[896] A. K. Gupta, T. Nguyen, and L. Pardo. Preliminary phi-divergence test estimator for multinomial probabilities. Computational Statistics & Data Analysis, 50(7):1749– 1773, April 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000356.

James:2006:PHT

[897] Gareth M. James and Ashish Sood. Performing hypothesis tests on the shape of functional data. Computational Statistics & Data Analysis, 50 (7):1774–1792, April 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000393.

Kim:2006:STB

[898] Myung Suk Kim and Suojin Wang. Sizes of two bootstrap-based nonparametric specification tests for the drift function in continuous time models. Computational Statistics & Data Analysis, 50(7):1793– 1806, April 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000381.

Li:2006:EUR

[899] Baibing Li and Xiangning Huang. Existence and uniqueness of relative incidence estimates in caseseries analysis. Computational Statistics & Data Analysis, 50(7):1807– 1817, April 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000319.

Staniswalis:2006:FGN

[900] Joan G. Staniswalis. On fitting generalized non-linear models with varying coefficients. Computational Statistics & Data Analysis, 50(7):1818– 1839, April 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500040X.

Steel:2006:IDC

[901] S. J. Steel and D. W. Uys. Influential data cases when the C_{p} criterion is used for variable selection in multiple linear regression. Computational Statistics & Data Analysis, 50 (7):1840–1854, April 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000320.

Zhang:2006:SHT

[902] Biao Zhang. A semiparametric hypothesis testing procedure for the ROC curve area under a density ratio model. Computational Statistics & Data Analysis, 50(7):1855– 1876, April 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000307.

Anonymous:2006:EBh

[903] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50(8):viii–ix, April 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000302.

Kuss:2006:ESR

[904] Oliver Kuss. On the estimation of the stereotype regression model. Computational Statistics & Data Analysis, 50 (8):1877–1890, April 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000423.

Menn:2006:CFB

[905] Christian Menn and Svetlozar T. Rachev. Calibrated FFT-based density approximations for α-stable distributions. Computational Statistics & Data Analysis, 50(8):1891– 1904, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000460.

Aguilera:2006:UPC

[906] Ana M. Aguilera, Manuel Escabias, and Mariano J. Valderrama. Using principal components for estimating logistic regression with high-dimensional multicollinear data. Computational Statistics & Data Analysis, 50(8):1905– 1924, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000630.

Ceyhan:2006:RDR

[907] Elvan Ceyhan, Carey E. Priebe, and John C. Wierman. Relative density of the random r-factor proximity catch digraph for testing spatial patterns of segregation and association. Computational Statistics & Data Analysis, 50 (8):1925–1964, April 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000447.

Delaigle:2006:DDB

[908] A. Delaigle and I. Gijbels. Datadriven boundary estimation in deconvolution problems. Computational Statistics & Data Analysis, 50(8):1965– 1994, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000411.

Deschepper:2006:RRP

[909] E. Deschepper, O. Thas, and J. P. Ottoy. Regional residual plots for assessing the fit of linear regression models. Computational Statistics & Data Analysis, 50(8):1995– 2013, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000502.

Fernandez:2006:BML

[910] Arturo J. Fernández. Bounding maximum likelihood estimates based on incomplete ordered data. Computational Statistics & Data Analysis, 50 (8):2014–2027, April 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000459.

Iglesias:2006:DAU

[911] Pilar Iglesias, Hector Jorquera, and Wilfredo Palma. Data analysis using regression models with missing observations and long-memory: an application study. Computational Statistics & Data Analysis, 50(8):2028– 2043, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000496.

Kim:2006:CLL

[912] Sung-Ho Kim. Conditional loglinear structures for log-linear modelling. Computational Statistics & Data Analysis, 50(8):2044–2064, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000514.

Schmidt:2006:MDM

[913] Rafael Schmidt, Tomas Hrycej, and Eric Stützle. Multivariate distribution models with generalized hyperbolic margins. Computational Statistics & Data Analysis, 50(8):2065– 2096, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000629.

Simila:2006:SOM

[914] Timo Similä. Self-organizing map visualizing conditional quantile functions with multidimensional covariates. Computational Statistics & Data Analysis, 50(8):2097–2110, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000435.

Yamada:2006:PTC

[915] Tomoya Yamada and Takakazu Sugiyama. On the permutation test in canonical correlation analysis. Computational Statistics & Data Analysis, 50 (8):2111–2123, April 10, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000484.

DiIorio:2006:DIE

[916] Francesca Di Iorio and Giorgio Calzolari. Discontinuities in indirect estimation: an application to EAR models. Computational Statistics & Data Analysis, 50(8):2124–2136, April 10, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000472.

Anonymous:2006:EBi

[917] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50(9):ix–x, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000582.

Pollock:2006:ISI

[918] D. S. G. Pollock. Introduction to the special issue on statistical signal extraction and filtering. Computational Statistics & Data Analysis, 50 (9):2137–2145, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001568.

Broto:2006:UCM

[919] Carmen Broto and Esther Ruiz. Unobserved component models with asymmetric conditional variances. Computational Statistics & Data Analysis, 50 (9):2146–2166, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304004013.

Maravall:2006:ATS

[920] Agustín Maravall. An application of the TRAMO-SEATS automatic procedure; direct versus indirect adjustment. Computational Statistics & Data Analysis, 50(9):2167– 2190, May 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001532.

Mazzocchi:2006:TPU

[921] Mario Mazzocchi. Time patterns in UK demand for alcohol and tobacco: an application of the EM algorithm. Computational Statistics & Data Analysis, 50 (9):2191–2205, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304004098.

McElroy:2006:IPA

[922] Tucker McElroy and Andrew Sutcliffe. An iterated parametric approach to nonstationary signal extraction. Computational Statistics & Data Analysis, 50(9):2206– 2231, May 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001556.

Godolphin:2006:DTS

[923] E. J. Godolphin and Kostas Triantafyllopoulos. Decomposition of time series models in state-space form. Computational Statistics & Data Analysis, 50 (9):2232–2246, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304004049.

Strickland:2006:BAS

[924] Chris M. Strickland, Catherine S. Forbes, and Gael M. Martin. Bayesian analysis of the stochastic conditional duration model. Computational Statistics & Data Analysis, 50(9):2247– 2267, May 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001520.

Pollock:2006:EMS

[925] D. S. G. Pollock. Econometric methods of signal extraction. Computational Statistics & Data Analysis, 50 (9):2268–2292, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500157X.

Pascual:2006:BPR

[926] Lorenzo Pascual, Juan Romo, and Esther Ruiz. Bootstrap prediction for returns and volatilities in GARCH models. Computational Statistics & Data Analysis, 50(9): 2293–2312, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304004001.

Fried:2006:RMH

[927] Roland Fried, Thorsten Bernholt, and Ursula Gather. Repeated median and hybrid filters. Computational Statistics & Data Analysis, 50(9):2313– 2338, May 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304004050.

Ombao:2006:TDF

[928] Hernando Ombao and Moon ho Ringo Ho. Time-dependent frequency domain principal components analysis of multichannel non-stationary signals. Computational Statistics & Data Analysis, 50(9):2339–2360, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947304004037.

Coakley:2006:UHP

[929] Jerry Coakley, Ana-María Fuertes, and Ron Smith. Unobserved heterogeneity in panel time series models. Computational Statistics & Data Analysis, 50 (9):2361–2380, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304004074.

Aminghafari:2006:MDU

[930] Mina Aminghafari, Nathalie Cheze, and Jean-Michel Poggi. Multivariate denoising using wavelets and principal component analysis. Computational Statistics & Data Analysis, 50 (9):2381–2398, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304004025.

Whitcher:2006:WBB

[931] Brandon Whitcher. Wavelet-based bootstrapping of spatial patterns on a finite lattice. Computational Statistics & Data Analysis, 50(9):2399– 2421, May 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304004086.

Amato:2006:DRF

[932] U. Amato, A. Antoniadis, and I. De Feis. Dimension reduction in functional regression with applications. Computational Statistics & Data Analysis, 50 (9):2422–2446, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947304003998.

Stoev:2006:LTL

[933] Stilian Stoev, Murad S. Taqqu, Cheolwoo Park, George Michailidis, and J. S. Marron. LASS: a tool for the local analysis of self-similarity. Computational Statistics & Data Analysis, 50 (9):2447–2471, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947304004062.

Zeng:2006:NHB

[934] Yujing Zeng and Javier Garcia-Frias. A novel HMM-based clustering algorithm for the analysis of gene expression time-course data. Computational Statistics & Data Analysis, 50 (9):2472–2494, May 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001544.

Lehtinen:2006:SES

[935] Aki Lehtinen. Signal extraction for simulated games with a large number of players. Computational Statistics & Data Analysis, 50(9):2495– 2507, May 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001519.

Anonymous:2006:EBj

[936] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50 (10):viii–ix, June 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000995.

Kundu:2006:ATI

[937] Debasis Kundu and Avijit Joarder. Analysis of Type-II progressively hybrid censored data. Computational Statistics & Data Analysis, 50(10): 2509–2528, June 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001210.

Lafosse:2006:SCA

[938] Roger Lafosse and Jos M. F. ten Berge. A simultaneous CONCOR algorithm for the analysis of two partitioned matrices. Computational Statistics & Data Analysis, 50(10):2529– 2535, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000654.

Nagahara:2006:EMS

[939] Yuichi Nagahara. Erratum to “A method of simulating multivariate nonnormal distributions by the Pearson distribution system and estimation” [comput. statist. data anal. 47 (2004) 1–29]. Computational Statistics & Data Analysis, 50(10): 2536, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000642. See

[597].

Purczynki:2006:FGF

[940] Jan Purczyńki and Przemysław Włodarski. On fast generation of fractional Gaussian noise. Computational Statistics & Data Analysis, 50(10):2537– 2551, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001003.

Ainsworth:2006:AID

[941] L. M. Ainsworth and C. B. Dean. Approximate inference for disease mapping. Computational Statistics & Data Analysis, 50(10):2552– 2570, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001209.

Aitken:2006:ETE

[942] C. G. G. Aitken, D. Lucy, G. Zadora, and J. M. Curran. Evaluation of transfer evidence for three-level multivariate data with the use of graphical models. Computational Statistics & Data Analysis, 50(10):2571– 2588, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730500068X.

Alonso:2006:CTS

[943] Andrés M. Alonso and Elizabeth A. Maharaj. Comparison of time series using subsampling. Computational Statistics & Data Analysis, 50(10): 2589–2599, June 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000733.

Batsidis:2006:EDM

[944] A. Batsidis, K. Zografos, and S. Loukas. Errors in discrimination with monotone missing data from multivariate normal populations. Computational Statistics & Data Analysis, 50(10): 2600–2634, June 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000708.

Bengtsson:2006:IAI

[945] Thomas Bengtsson and Joseph E. Cavanaugh. An improved Akaike information criterion for state-space model selection. Computational Statistics & Data Analysis, 50(10):2635– 2654, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001222.

Bouzas:2006:MMD

[946] P. R. Bouzas, M. J. Valderrama, A. M. Aguilera, and N. Ruiz-Fuentes. Modelling the mean of a doubly stochastic Poisson process by functional data analysis. Computational Statistics & Data Analysis, 50(10): 2655–2667, June 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000812.

Caiado:2006:PBM

[947] Jorge Caiado, Nuno Crato, and Daniel Peña. A periodogrambased metric for time series classification. Computational Statistics & Data Analysis, 50(10):2668– 2684, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000770.

ElAdlouni:2006:CMA

[948] Salaheddine El Adlouni, Anne-Catherine Favre, and Bernard Bobée. Comparison of methodologies to assess the convergence of Markov chain Monte Carlo methods. Computational Statistics & Data Analysis, 50(10):2685– 2701, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000836.

Gad:2006:ALD

[949] Ahmed M. Gad and Abeer S. Ahmed. Analysis of longitudinal data with intermittent missing values using the stochastic EM algorithm. Computational Statistics & Data Analysis, 50(10):2702– 2714, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000678.

Godfrey:2006:TRM

[950] L. G. Godfrey. Tests for regression models with heteroskedasticity of unknown form. Computational Statistics & Data Analysis, 50(10):2715– 2733, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000691.

Harvey:2006:SSL

[951] David I. Harvey and Dick van Dijk. Sample size, lag order and critical values of seasonal unit root tests. Computational Statistics & Data Analysis, 50(10):2734– 2751, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000745.

Marazzi:2006:RBC

[952] Alfio Marazzi and Victor J. Yohai. Robust Box–Cox transformations based on minimum residual autocorrelation. Computational Statistics & Data Analysis, 50(10):2752– 2768, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500071X.

Muino:2006:GDN

[953] J. M. Muiño, E. O. Voit, and A. Sorribas. GS-distributions: a new family of distributions for continuous unimodal variables. Computational Statistics & Data Analysis, 50(10):2769– 2798, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000800.

Nyangoma:2006:IIM

[954] S. O. Nyangoma, W.-K. Fung, and R. C. Jansen. Identifying influential multinomial observations by perturbation. Computational Statistics & Data Analysis, 50(10):2799– 2821, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001180.

Rasekh:2006:LIM

[955] A. R. Rasekh. Local influence in measurement error models with ridge estimate. Computational Statistics & Data Analysis, 50(10):2822– 2834, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001015.

Sessions:2006:IOV

[956] David N. Sessions and Lonnie K. Stevans. Investigating omitted variable bias in regression parameter estimation: a genetic algorithm approach. Computational Statistics & Data Analysis, 50(10):2835–2854, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000666.

Sethuraman:2006:OFD

[957] Venkat S. Sethuraman, Damaraju Raghavarao, and Bikas K. Sinha. Optimal s^{n} factorial designs when observations within-blocks are correlated. Computational Statistics & Data Analysis, 50(10):2855–2862, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730500099X.

Shimansky:2006:CSL

[958] Yury P. Shimansky. Continuous significant linear dimensionality: Geometric interpretation and statistical characteristics. Computational Statistics & Data Analysis, 50(10):2863– 2877, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001192.

Tutz:2006:RSE

[959] Gerhard Tutz and Florian Leitenstorfer. Response shrinkage estimators in binary regression. Computational Statistics & Data Analysis, 50(10): 2878–2901, June 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000721.

VanLandeghem:2006:SFP

[960] Georges Van Landeghem, Patrick Onghena, and Jan Van Damme. Separability in the fixed part of multilevel models. Computational Statistics & Data Analysis, 50(10):2902– 2919, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000794.

Walters:2006:CGL

[961] Eli Walters, Naomi S. Altman, and Laura Elnitski. Clustering of gene locations. Computational Statistics & Data Analysis, 50(10):2920– 2932, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305000988.

Wilcox:2006:PCD

[962] Rand R. Wilcox. Pairwise comparisons of dependent groups based on medians. Computational Statistics & Data Analysis, 50(10):2933– 2941, June 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000824.

Tian:2006:MMR

[963] Yongge Tian and Agnes M. Herzberg. A-minimax and D-minimax robust optimal designs for approximately linear Haar-wavelet models. Computational Statistics & Data Analysis, 50(10): 2942–2951, June 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305000782.

Anonymous:2006:EBk

[964] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50 (11):viii–ix, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001460.

Bonett:2006:CIC

[965] Douglas G. Bonett. Confidence interval for a coefficient of quartile variation. Computational Statistics & Data Analysis, 50(11):2953– 2957, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001271.

Melard:2006:EML

[966] Guy Mélard, Roch Roy, and Abdessamad Saidi. Exact maximum likelihood estimation of structured or unit root multivariate time series models. Computational Statistics & Data Analysis, 50(11):2958– 2986, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001398.

Zeileis:2006:ICS

[967] Achim Zeileis. Implementing a class of structural change tests: an econometric computing approach. Computational Statistics & Data Analysis, 50 (11):2987–3008, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001477.

Zhang:2006:BAB

[968] Xibin Zhang, Maxwell L. King, and Rob J. Hyndman. A Bayesian approach to bandwidth selection for multivariate kernel density estimation. Computational Statistics & Data Analysis, 50 (11):3009–3031, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001362.

Audrino:2006:IGN

[969] Francesco Audrino. The impact of general non-parametric volatility functions in multivariate GARCH models. Computational Statistics & Data Analysis, 50(11):3032–3052, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001313.

Azari:2006:LDM

[970] Rahman Azari, Lexin Li, and Chih-Ling Tsai. Longitudinal data model selection. Computational Statistics & Data Analysis, 50(11):3053– 3066, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001283.

Benhenni:2006:NER

[971] K. Benhenni and M. Rachdi. Nonparametric estimation of the regression function from quantized observations. Computational Statistics & Data Analysis, 50(11):3067– 3085, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001337.

Dahl:2006:BBT

[972] Tobias Dahl and Tormod Næs. A bridge between Tucker-1 and Carroll’s generalized canonical analysis. Computational Statistics & Data Analysis, 50 (11):3086–3098, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001465.

Feyzioglu:2006:DOC

[973] Orhan Feyzioğlu, İ. Kuban Altınel, and Süleyman Özekici. The design of optimum component test plans for system reliability. Computational Statistics & Data Analysis, 50(11):3099– 3112, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001453.

Ghosh:2006:OCN

[974] Anil K. Ghosh. On optimum choice of k in nearest neighbor classification. Computational Statistics & Data Analysis, 50(11):3113– 3123, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001374.

Giloni:2006:RWL

[975] Avi Giloni, Jeffrey S. Simonoff, and Bhaskar Sengupta. Robust weighted LAD regression. Computational Statistics & Data Analysis, 50(11):3124– 3140, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001325.

Gupta:2006:SIC

[976] Ramesh C. Gupta and Xue Li. Statistical inference for the common mean of two log-normal distributions and some applications in reliability. Computational Statistics & Data Analysis, 50 (11):3141–3164, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001246.

Kim:2006:LTT

[977] Jinheum Kim, Dae Ryong Kang, and Chung Mo Nam. Logrank-type tests for comparing survival curves with interval-censored data. Computational Statistics & Data Analysis, 50(11): 3165–3178, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001441.

Lin:2006:APS

[978] Jin-Guan Lin and Bo cheng Wei. Approximate power of score test for variance heterogeneity under local alternatives in nonlinear models. Computational Statistics & Data Analysis, 50 (11):3179–3198, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730500143X.

Munoz-Garcia:2006:CRP

[979] J. Muñoz-García, J. M. Muñoz-Pichardo, and L. Pardo. Cressie and read power-divergences as influence measures for logistic regression models. Computational Statistics & Data Analysis, 50(11):3199– 3221, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001350.

Ng:2006:PIE

[980] H. K. T. Ng, D. Kundu, and N. Balakrishnan. Point and interval estimation for the two-parameter Birnbaum– Saunders distribution based on Type-II censored samples. Computational Statistics & Data Analysis, 50(11): 3222–3242, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001301. See

erratum [1304].

Paddock:2006:FDT

[981] Susan M. Paddock, Greg Ridgeway, Rongheng Lin, and Thomas A. Louis. Flexible distributions for triplegoal estimates in two-stage hierarchical models. Computational Statistics & Data Analysis, 50(11):3243– 3262, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500126X.

Rueda:2006:MEC

[982] M. Rueda, S. Martínez, H. Martínez, and A. Arcos. Mean estimation with calibration techniques in presence of missing data. Computational Statistics & Data Analysis, 50(11):3263– 3277, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001349.

Shimizu:2006:FCO

[983] Shohei Shimizu, Aapo Hyvärinen, Patrik O. Hoyer, and Yutaka Kano. Finding a causal ordering via independent component analysis. Computational Statistics & Data Analysis, 50(11): 3278–3293, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001234.

Senturk:2006:ECA

[984] Damla Şentürk and Danh V. Nguyen. Estimation in covariate-adjusted regression. Computational Statistics & Data Analysis, 50(11):3294– 3310, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001295.

vandenHout:2006:ELR

[985] Ardo van den Hout and Peter Kooiman. Estimating the linear regression model with categorical covariates subject to randomized response. Computational Statistics & Data Analysis, 50(11):3311– 3323, July 20, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001386.

Wu:2006:MTP

[986] K. H. Wu, M. L. Tang, and K. M. Lee. Multiple testing procedures for analyzing stratified comparative clinical trials using odds ratios. Computational Statistics & Data Analysis, 50 (11):3324–3342, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001258.

Headrick:2006:SMN

[987] Todd C. Headrick and Abdel Mugdadi. On simulating multivariate nonnormal distributions from the generalized lambda distribution. Computational Statistics & Data Analysis, 50 (11):3343–3353, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001404.

Li:2006:LVM

[988] Ning Li, Guoqi Qian, and Richard Huggins. A latent variable model for estimating disease transmission rate from data on household outbreaks. Computational Statistics & Data Analysis, 50 (11):3354–3368, July 20, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001416.

Anonymous:2006:EBl

[989] Anonymous. Editorial board. Computational Statistics & Data Analysis, 50(12):ix–x, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001642.

Boese:2006:CIB

[990] Doyle H. Boese, Dean M. Young, and James D. Stamey. Confidence intervals for a binomial parameter based on binary data subject to falsepositive misclassification. Computational Statistics & Data Analysis, 50 (12):3369–3385, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001866.

Celeux:2006:IIS

[991] Gilles Celeux, Jean-Michel Marin, and Christian P. Robert. Iterated importance sampling in missing data problems. Computational Statistics & Data Analysis, 50(12):3386– 3404, August 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001659.

Levine:2006:BSC

[992] M. Levine. Bandwidth selection for a class of difference-based variance estimators in the nonparametric regression: a possible approach. Computational Statistics & Data Analysis, 50 (12):3405–3431, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001702.

Lin:2006:PTE

[993] Hui-Yi Lin and Leann Myers. Power and type I error rates of goodnessof-fit statistics for binomial generalized estimating equations (GEE) models. Computational Statistics & Data Analysis, 50(12):3432– 3448, August 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001647.

Nadarajah:2006:DSP

[994] Saralees Nadarajah and M. Masoom Ali. The distribution of sums, products and ratios for Lawrance and Lewis’s bivariate exponential random variables. Computational Statistics & Data Analysis, 50(12):3449– 3463, August 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001696.

Sauerbrei:2006:MRM

[995] W. Sauerbrei, C. Meier-Hirmer, A. Benner, and P. Royston. Multivariable regression model building by using fractional polynomials: Description of SAS, STATA and R programs. Computational Statistics & Data Analysis, 50 (12):3464–3485, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001623.

Sima:2006:CTS

[996] Diana M. Sima and Sabine Van Huffel. A class of template splines. Computational Statistics & Data Analysis, 50 (12):3486–3499, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500160X.

Bickel:2006:FRE

[997] David R. Bickel and Rudolf Frühwirth. On a fast, robust estimator of the mode: Comparisons to other robust estimators with applications. Computational Statistics & Data Analysis, 50 (12):3500–3530, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001581.

Chae:2006:MPN

[998] Seong S. Chae, Janice L. DuBien, and William D. Warde. A method of predicting the number of clusters using Rand’s statistic. Computational Statistics & Data Analysis, 50(12): 3531–3546, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001751.

Harvill:2006:FCA

[999] Jane L. Harvill and Bonnie K. Ray. Functional coefficient autoregressive models for vector time series. Computational Statistics & Data Analysis, 50 (12):3547–3566, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001635.

Herwartz:2006:TRE

[1000] Helmut Herwartz. Testing for random effects in panel data under cross sectional error correlation — a bootstrap approach to the Breusch Pagan test. Computational Statistics & Data Analysis, 50(12):3567– 3591, August 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001726.

Ji:2006:TSE

[1001] Yuan Ji, Kam-Wah Tsui, and Kyung-Mann Kim. A two-stage empirical Bayes method for identifying differentially expressed genes. Computational Statistics & Data Analysis, 50 (12):3592–3604, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001672.

Kanamori:2006:CME

[1002] Takafumi Kanamori and Ichiro Takeuchi.

Conditional mean estimation under asymmetric and heteroscedastic error by linear combination of quantile regressions. Computational Statistics & Data Analysis, 50(12):3605– 3618, August 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001684.

Klebanov:2006:PTM

[1003] L. Klebanov, A. Gordon, Y. Xiao, H. Land, and A. Yakovlev. A permutation test motivated by microarray data analysis. Computational Statistics & Data Analysis, 50(12): 3619–3628, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500174X.

Li:2006:EKE

[1004] Xiaochun Li and Ronghui Xu. Empirical and kernel estimation of covariate distribution conditional on survival time. Computational Statistics & Data Analysis, 50(12):3629– 3643, August 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001714.

Mauricio:2006:EML

[1005] José Alberto Mauricio. Exact maximum likelihood estimation of partially nonstationary vector ARMA models. Computational Statistics & Data Analysis, 50(12):3644– 3662, August 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001593.

McMahon:2006:GMM

[1006] James M. McMahon, Enrique R. Pouget, and Stephanie Tortu. A guide for multilevel modeling of dyadic data with binary outcomes using SAS PROC NLMIXED. Computational Statistics & Data Analysis, 50(12): 3663–3680, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001878.

Sidik:2006:RVE

[1007] Kurex Sidik and Jeffrey N. Jonkman. Robust variance estimation for random effects meta-analysis. Computational Statistics & Data Analysis, 50 (12):3681–3701, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305001660.

Triantafyllopoulos:2006:MDW

[1008] Kostas Triantafyllopoulos. Multivariate discount weighted regression and local level models. Computational Statistics & Data Analysis, 50(12): 3702–3720, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001507.

Kafadar:2006:VTE

[1009] Karen Kafadar and Edward J. Wegman. Visualizing “typical” and “exotic” Internet traffic data. Computational Statistics & Data Analysis, 50 (12):3721–3743, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305001489.

Anonymous:2006:AI

[1010] Anonymous. Author index. Computational Statistics & Data Analysis, 50 (12):3744–3751, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001691.

Anonymous:2006:EBm

[1011] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51 (1):v–vii, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003008.

Anonymous:2006:Ca

[1012] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (1):viii–x, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600301X.

Coppi:2006:FAS

[1013] Renato Coppi, Maria A. Gil, and Henk A. L. Kiers. The fuzzy approach to statistical analysis. Computational Statistics & Data Analysis, 51 (1):1–14, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001769.

Zadeh:2006:GTU

[1014] Lotfi A. Zadeh. Generalized theory of uncertainty (GTU) — principal concepts and ideas. Computational Statistics & Data Analysis, 51 (1):15–46, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001290.

Dubois:2006:PTS

[1015] Didier Dubois. Possibility theory and statistical reasoning. Computational Statistics & Data Analysis, 51 (1):47–69, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001149.

Nguyen:2006:RFS

[1016] Hung T. Nguyen and Berlin Wu. Random and fuzzy sets in coarse data analysis. Computational Statistics & Data Analysis, 51(1):70–85, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001150.

Baudrit:2006:PRI

[1017] C. Baudrit and D. Dubois. Practical representations of incomplete probabilistic knowledge. Computational Statistics & Data Analysis, 51(1):86– 108, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000454.

Lopez-Diaz:2006:TFR

[1018] Miguel López-Díaz and Dan A.

Ralescu. Tools for fuzzy random variables: Embeddings and measurabilities. Computational Statistics & Data Analysis, 51(1):109–114, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001162.

Coletti:2006:CPF

[1019] Giulianella Coletti and Romano Scozzafava. Conditional probability and fuzzy information. Computational Statistics & Data Analysis, 51(1):115– 132, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001289.

Viertl:2006:USA

[1020] Reinhard Viertl. Univariate statistical analysis with fuzzy data. Computational Statistics & Data Analysis, 51(1):133–147, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306000880.

Gil:2006:BAM

[1021] María Ángeles Gil, Manuel Montenegro, Gil González-Rodríguez, Ana Colubi, and María Rosa Casals. Bootstrap approach to the multisample test of means with imprecise data. Computational Statistics & Data Analysis, 51(1):148–162, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001174.

Gonzalez-Rodriguez:2006:FRR

[1022] Gil González-Rodríguez, Ana Colubi, and María Ángeles Gil. A fuzzy representation of random variables: an operational tool in exploratory analysis and hypothesis testing. Computational Statistics & Data Analysis, 51(1):163– 176, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000922.

Bermudez:2006:DSS

[1023] J. D. Bermúdez, J. V. Segura, and E. Vercher. A decision support system methodology for forecasting of time series based on soft computing. Computational Statistics & Data Analysis, 51(1):177–191, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000466.

Doring:2006:DAF

[1024] Christian Döring, Marie-Jeanne Lesot, and Rudolf Kruse. Data analysis with fuzzy clustering methods. Computational Statistics & Data Analysis, 51(1):192–214, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306001307.

Hathaway:2006:EFP

[1025] Richard J. Hathaway and James C.

Bezdek. Extending fuzzy and probabilistic clustering to very large data sets. Computational Statistics & Data Analysis, 51(1):215–234, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000442.

Nather:2006:RFR

[1026] Wolfgang Näther. Regression with fuzzy random data. Computational Statistics & Data Analysis, 51(1):235– 252, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000661.

Guo:2006:DMP

[1027] Peijun Guo and Hideo Tanaka.

Dual models for possibilistic regression analysis. Computational Statistics & Data Analysis, 51(1):253–266, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000910.

Coppi:2006:LSE

[1028] Renato Coppi, Pierpaolo D’Urso, Paolo Giordani, and Adriana Santoro. Least squares estimation of a linear regression model with LR fuzzy response. Computational Statistics & Data Analysis, 51(1):267–286, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001332.

DUrso:2006:FCL

[1029] Pierpaolo D’Urso and Adriana Santoro. Fuzzy clusterwise linear regression analysis with symmetrical fuzzy output variable. Computational Statistics & Data Analysis, 51(1):287–313, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001915.

Grzegorzewski:2006:CCV

[1030] Przemysław Grzegorzewski. The coefficient of concordance for vague data. Computational Statistics & Data Analysis, 51(1):314–322, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001277.

Hryniewicz:2006:GKM

[1031] Olgierd Hryniewicz. Goodman– Kruskal γ measure of dependence for fuzzy ordered categorical data. Computational Statistics & Data Analysis, 51(1):323–334, November 1, 2006.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306001125.

Hebert:2006:FMS

[1032] Pierre-Alexandre Hébert, Marie-Hélène Masson, and Thierry Denœux. Fuzzy multidimensional scaling. Computational Statistics & Data Analysis, 51(1):335–359, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016794730600065X.

Groenen:2006:SMS

[1033] P. J. F. Groenen, S. Winsberg, O. Rodríguez, and E. Diday. I-Scal: Multidimensional scaling of interval dissimilarities. Computational Statistics & Data Analysis, 51(1):360– 378, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000892.

Giordani:2006:CTM

[1034] Paolo Giordani and Henk A. L. Kiers.

A comparison of three methods for principal component analysis of fuzzy interval data. Computational Statistics & Data Analysis, 51(1):379–397, November 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000648.

Bonissone:2006:DLF

[1035] Piero P. Bonissone, Anil Varma, Kareem S. Aggour, and Feng Xue. Design of local fuzzy models using evolutionary algorithms. Computational Statistics & Data Analysis, 51(1):398– 416, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001113.

Shepherd:2006:FME

[1036] David Shepherd and Francis K. C. Shi. Fuzzy modelling and estimation of economic relationships. Computational Statistics & Data Analysis, 51(1):417– 433, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000909.

Gulbay:2006:DFP

[1037] Murat Gülbay and Cengiz Kahraman. Development of fuzzy process control charts and fuzzy unnatural pattern analyses. Computational Statistics & Data Analysis, 51(1):434– 451, November 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001137.

Anonymous:2006:EBn

[1038] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51 (2):iii–v, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003124.

Anonymous:2006:Cb

[1039] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (2):vi–xi, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003136.

Farebrother:2006:LPP

[1040] Richard William Farebrother. A linear programming procedure based on de la Vallée Poussin’s minimax estimation procedure. Computational Statistics & Data Analysis, 51(2):453– 456, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002574.

Wang:2006:OTS

[1041] Hansong Wang and Daniel O. Stram.

Optimal two-stage genome-wide association designs based on false discovery rate. Computational Statistics & Data Analysis, 51(2):457–465, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001344.

Same:2006:CEA

[1042] Allou Samé, Christophe Ambroise, and Gérard Govaert. A classification EM algorithm for binned data. Computational Statistics & Data Analysis, 51(2):466–480, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002082.

Krijnen:2006:CSP

[1043] Wim P. Krijnen. Convergence of the sequence of parameters generated by alternating least squares algorithms. Computational Statistics & Data Analysis, 51(2):481–489, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002112.

Requena:2006:MIN

[1044] F. Requena and N. Martín Ciudad. A major improvement to the Network Algorithm for Fisher’s exact test in 2×c contingency tables. Computational Statistics & Data Analysis, 51(2):490– 498, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002173.

Knoth:2006:CAC

[1045] Sven Knoth. Computation of the ARL for CUSUM–S^{2} schemes. Computational Statistics & Data Analysis, 51(2):499–512, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002525.

Tran:2006:KKD

[1046] Thanh N. Tran, Ron Wehrens, and Lutgarde M. C. Buydens. KNNkernel density-based clustering for high-dimensional multivariate data.

Computational Statistics & Data Analysis, 51(2):513–525, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002537.

Leisch:2006:TCC

[1047] Friedrich Leisch. A toolbox for Kcentroids cluster analysis. Computational Statistics & Data Analysis, 51(2):526–544, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947305002586.

Cizek:2006:RED

[1048] P. Čižek and W. Härdle. Robust estimation of dimension reduction space. Computational Statistics & Data Analysis, 51(2):545–555, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002938.

Kondylis:2006:DCR

[1049] Athanassios Kondylis and Ali S. Hadi.

Derived components regression using the BACON algorithm. Computational Statistics & Data Analysis, 51(2):556– 569, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002963.

Xiang:2006:IEF

[1050] Qinfang Xiang, Jode Edwards, and Gary L. Gadbury. Interval estimation in a finite mixture model: Modeling P -values in multiple testing applications. Computational Statistics & Data Analysis, 51(2):570–586, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500304X.

Biernacki:2006:MBC

[1051] Christophe Biernacki, Gilles Celeux, Gérard Govaert, and Florent Langrognet. Model-based cluster and discriminant analysis with the MIX-MOD software. Computational Statistics & Data Analysis, 51(2):587–600, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003300.

Martinez:2006:ESS

[1052] Edson Zangiacomi Martinez, Jorge Alberto Achcar, and Francisco Louzada-Neto. Estimators of sensitivity and specificity in the presence of verification bias: a Bayesian approach. Computational Statistics & Data Analysis, 51(2):601–611, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000041.

Kapetanios:2006:COS

[1053] George Kapetanios. Choosing the optimal set of instruments from large instrument sets. Computational Statistics & Data Analysis, 51(2):612–620, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000107.

Bayarri:2006:MMA

[1054] M. J. Bayarri, M. E. Castellanos, and J. Morales. MCMC methods to approximate conditional predictive distributions. Computational Statistics & Data Analysis, 51(2):621–640, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000235.

He:2006:CAU

[1055] Yi He, Wei Pan, and Jizhen Lin. Cluster analysis using multivariate normal mixture models to detect differential gene expression with microarray data. Computational Statistics & Data Analysis, 51(2):641–658, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600048X.

Lee:2006:UVS

[1056] Tzu-Haw Lee and Yu-Shan Shih. Unbiased variable selection for classification trees with multivariate responses. Computational Statistics & Data Analysis, 51(2):659–667, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600051X.

Kokolakis:2006:BDP

[1057] G. Kokolakis, Ph. Nanopoulos, and D. Fouskakis. Bregman divergences in the m × k-partitioning problem. Computational Statistics & Data Analysis, 51(2):668–678, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000521.

vanLieshout:2006:PSM

[1058] M. N. M. van Lieshout and R. S. Stoica. Perfect simulation for marked point processes. Computational Statistics & Data Analysis, 51(2):679–698, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000703.

Skaug:2006:AAM

[1059] Hans J. Skaug and David A. Fournier. Automatic approximation of the marginal likelihood in non-Gaussian hierarchical models. Computational Statistics & Data Analysis, 51(2):699– 709, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000764.

Kang:2006:STP

[1060] Seung-Ho Kang, Yonghee Lee, and EunSug Park. The sizes of the three popular asymptotic tests for testing homogeneity of two binomial proportions. Computational Statistics & Data Analysis, 51(2):710–722, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000776.

Adewale:2006:NCR

[1061] Adeniyi J. Adewale and Douglas P.

Wiens. New criteria for robust integer-valued designs in linear models. Computational Statistics & Data Analysis, 51(2):723–736, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000806.

Choulakian:2006:RCM

[1062] V. Choulakian, J. Allard, and J. Almhana. Robust centroid method. Computational Statistics & Data Analysis, 51(2):737–746, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000831.

Jank:2006:AEF

[1063] Wolfgang Jank. Ascent EM for fast and global solutions to finite mixtures: an application to curve-clustering of online auctions. Computational Statistics & Data Analysis, 51(2):747–761, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000843.

Alonso:2006:TSC

[1064] A. M. Alonso, J. R. Berrendero, A. Hernández, and A. Justel. Time series clustering based on forecast densities. Computational Statistics & Data Analysis, 51(2):762–776, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001186.

Kneib:2006:MMB

[1065] Thomas Kneib. Mixed model-based inference in geoadditive hazard regression for interval-censored survival times. Computational Statistics & Data Analysis, 51(2):777–792, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600212X.

Capanu:2006:TPU

[1066] Marinela Capanu, Gregory A. Jones, and Ronald H. Randles. Testing for preference using a sum of Wilcoxon signed rank statistics. Computational Statistics & Data Analysis, 51(2):793– 796, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003324.

Vougas:2006:URT

[1067] Dimitrios V. Vougas. On unit root testing with smooth transitions. Computational Statistics & Data Analysis, 51(2):797–800, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002295.

Zheng:2006:MHA

[1068] Guoqing Zheng and Pingjian Zhang.

Meta-heuristic algorithms for parameter estimation of semi-parametric linear regression models. Computational Statistics & Data Analysis, 51(2):801– 808, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002070.

Huard:2006:BCS

[1069] David Huard, Guillaume Évin, and Anne-Catherine Favre. Bayesian copula selection. Computational Statistics & Data Analysis, 51(2):809–822, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002094.

Perez:2006:MBL

[1070] C. J. Pérez, J. Martín, and M. J.

Rufo. MCMC-based local parametric sensitivity estimations. Computational Statistics & Data Analysis, 51(2):823– 835, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002203.

Hinkle:2006:RC

[1071] John E. Hinkle and William S. Rayens.

Reciprocal curves. Computational Statistics & Data Analysis, 51(2):836– 858, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002215.

Tscheschel:2006:SRR

[1072] A. Tscheschel and D. Stoyan. Statistical reconstruction of random point patterns. Computational Statistics & Data Analysis, 51(2):859–871, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002306.

Gilardoni:2006:VAP

[1073] G. L. Gilardoni. Very accurate posterior approximations based on finite mixtures of the hyperparameters conditionals. Computational Statistics & Data Analysis, 51(2):872–884, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002318.

Koopman:2006:FDT

[1074] Siem Jan Koopman and Marius Ooms. Forecasting daily time series using periodic unobserved components time series models. Computational Statistics & Data Analysis, 51(2):885–903, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730500232X.

Luceno:2006:FGP

[1075] Alberto Luceño. Fitting the generalized Pareto distribution to data using maximum goodness-of-fit estimators. Computational Statistics & Data Analysis, 51(2):904–917, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002331.

Zhu:2006:ADS

[1076] Mu Zhu and Ali Ghodsi. Automatic dimensionality selection from the scree plot via the use of profile likelihood. Computational Statistics & Data Analysis, 51(2):918–930, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002343.

Kohn:2006:CID

[1077] Hans-Friedrich Köhn. Combinatorial individual differences scaling within the city-block metric. Computational Statistics & Data Analysis, 51(2):931– 946, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002501.

Karvanen:2006:EQM

[1078] Juha Karvanen. Estimation of quantile mixtures via L-moments and trimmed L-moments. Computational Statistics & Data Analysis, 51(2):947–959, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002513.

Ospina:2006:IPI

[1079] Raydonal Ospina, Francisco Cribari-Neto, and Klaus L. P. Vasconcellos. Improved point and interval estimation for a beta regression model. Computational Statistics & Data Analysis, 51(2):960–981, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002549. See

erratum [2347].

Chen:2006:OCI

[1080] Hubert J. Chen and Miin-Jye Wen. Optimal confidence interval for the largest normal mean under heteroscedasticity. Computational Statistics & Data Analysis, 51(2):982–1001, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002562.

Silva:2006:LBA

[1081] E. M. Silva, G. C. Franco, V. A. Reisen, and F. R. B. Cruz. Local bootstrap approaches for fractional differential parameter estimation in ARFIMA models. Computational Statistics & Data Analysis, 51(2):1002– 1011, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002744.

Lee:2006:IEB

[1082] Seung-Chun Lee. Interval estimation of binomial proportions based on weighted Polya posterior. Computational Statistics & Data Analysis, 51(2):1012–1021, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002756.

Wen:2006:SRT

[1083] Miin-Jye Wen and Hubert J. Chen. A Studentized range test for the equivalency of normal means under heteroscedasticity. Computational Statistics & Data Analysis, 51(2):1022– 1038, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002768.

Modarres:2006:RMR

[1084] Reza Modarres, Terrence P. Hui, and Gang Zheng. Resampling methods for ranked set samples. Computational Statistics & Data Analysis, 51(2):1039– 1050, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500277X.

Xu:2006:ICR

[1085] Jin Xu and Arjun K. Gupta. Improved confidence regions for a mean vector under general conditions. Computational Statistics & Data Analysis, 51(2):1051–1062, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002781.

Cuevas:2006:UBE

[1086] Antonio Cuevas, Manuel Febrero, and Ricardo Fraiman. On the use of the bootstrap for estimating functions with functional data. Computational Statistics & Data Analysis, 51(2):1063– 1074, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002793.

Goos:2006:BRS

[1087] P. Goos and A. N. Donev. Blocking response surface designs. Computational Statistics & Data Analysis, 51(2):1075– 1088, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002951.

Ruggoo:2006:MSS

[1088] Arvind Ruggoo and Martina Vandebroek. Model-sensitive sequential optimal designs. Computational Statistics & Data Analysis, 51(2):1089– 1099, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305002975.

Menendez:2006:TIB

[1089] M. L. Menéndez, J. A. Pardo, L. Pardo, and K. Zografos. On tests of independence based on minimum ϕ-divergence estimator with constraints: an application to modeling DNA. Computational Statistics & Data Analysis, 51(2):1100– 1118, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002987.

Fernandez:2006:BEB

[1090] Arturo J. Fernández. Bayesian estimation based on trimmed samples from Pareto populations. Computational Statistics & Data Analysis, 51(2):1119– 1130, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003014.

Xu:2006:DMG

[1091] Liang Xu, Sik-Yum Lee, and Wai-Yin Poon. Deletion measures for generalized linear mixed effects models. Computational Statistics & Data Analysis, 51(2):1131–1146, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003026.

Banerjee:2006:CPS

[1092] Tathagata Banerjee, Tapabrata Maiti, and Pushpal Mukhopadhyay. Classification of pathological stage of prostate cancer patients using penalized splines. Computational Statistics & Data Analysis, 51(2):1147– 1155, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003051.

Tian:2006:TEI

[1093] Lili Tian. Testing equality of inverse Gaussian means under heterogeneity, based on generalized test variable. Computational Statistics & Data Analysis, 51(2):1156–1162, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003063.

Hwang:2006:NMT

[1094] Yi-Ting Hwang and Peir Feng Wei. A novel method for testing normality in a mixed model of a nested classification. Computational Statistics & Data Analysis, 51(2):1163– 1183, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003075.

Nikulin:2006:TBC

[1095] V. Nikulin. Threshold-based clustering with merging and regularization in application to network intrusion detection. Computational Statistics & Data Analysis, 51(2):1184– 1196, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003087.

Hammill:2006:SIS

[1096] Bradley G. Hammill and John S. Preisser. A SAS/IML software program for GEE and regression diagnostics. Computational Statistics & Data Analysis, 51(2):1197–1212, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003099.

Lukas:2006:SAC

[1097] Mark A. Lukas and Mingren Shi. Sensitivity analysis of constrained linear L_{1} regression: Perturbations to constraints, addition and deletion of observations. Computational Statistics & Data Analysis, 51(2):1213– 1231, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002064.

terBraak:2006:BSS

[1098] Cajo J. F. ter Braak. Bayesian sigmoid shrinkage with improper variance priors and an application to wavelet denoising. Computational Statistics & Data Analysis, 51(2):1232– 1242, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002015.

Qian:2006:SBA

[1099] Zhiguang Qian and Alexander Shapiro.

Simulation-based approach to estimation of latent variable models. Computational Statistics & Data Analysis, 51(2):1243–1259, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000533.

NGuessan:2006:EMA

[1100] Assi N’Guessan, Azeddine Essai, and Modeste N’Zi. An estimation method of the average effect and the different accident risks when modelling a road safety measure: a simulation study. Computational Statistics & Data Analysis, 51(2):1260– 1277, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002100.

Ouali:2006:DMB

[1101] Abdelaziz Ouali, Amar Ramdane Cherif, and Marie-Odile Krebs. Data mining based Bayesian networks for best classification. Computational Statistics & Data Analysis, 51(2):1278– 1292, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002355.

Ane:2006:AFA

[1102] Thierry Ané. An analysis of the flexibility of asymmetric power GARCH models. Computational Statistics & Data Analysis, 51(2):1293–1311, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500294X.

Roch:2006:TBD

[1103] Oriol Roch and Antonio Alegre. Testing the bivariate distribution of daily equity returns using copulas. an application to the Spanish stock market. Computational Statistics & Data Analysis, 51(2):1312–1329, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305002999.

Stern:2006:AOT

[1104] David I. Stern. An atmosphere– ocean time series model of global climate change. Computational Statistics & Data Analysis, 51(2):1330– 1346, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003038.

Arato:2006:HBM

[1105] N. Miklós Arató, Ian L. Dryden, and Charles C. Taylor. Hierarchical Bayesian modelling of spatial agedependent mortality. Computational Statistics & Data Analysis, 51(2):1347– 1363, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000429.

Longford:2006:SHI

[1106] N. T. Longford and M. G. Pittau. Stability of household income in European countries in the 1990s. Computational Statistics & Data Analysis, 51(2):1364– 1383, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000478.

Mady:2006:SEL

[1107] Afaf M. Mady. Some extensions of Langenberg model for clinical trials with delayed observations normally distributed responses. Computational Statistics & Data Analysis, 51(2):1384– 1392, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000508.

Sicard:2006:TFL

[1108] E. Sicard and R. Sabatier. Theoretical framework for local PLS1 regression, and application to a rainfall data set. Computational Statistics & Data Analysis, 51(2):1393– 1410, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001368.

Chen:2006:EAR

[1109] Haiying Chen, Elizabeth A. Stasny, and Douglas A. Wolfe. An empirical assessment of ranking accuracy in ranked set sampling. Computational Statistics & Data Analysis, 51(2):1411– 1419, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002350.

Fuertes:2006:EWS

[1110] Ana-Maria Fuertes and Elena Kalotychou. Early warning systems for sovereign debt crises: the role of heterogeneity. Computational Statistics & Data Analysis, 51(2):1420– 1441, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002842.

Anonymous:2006:MEF

[1111] Anonymous. Microsoft Excel 2000 and 2003 faults, problems, workarounds and fixes. Computational Statistics & Data Analysis, 51(2):1442–1443, November 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000028.

Anonymous:2006:EBo

[1112] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51 (3):iii–v, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004117.

Anonymous:2006:Cc

[1113] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (3):vi–ix, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004129.

Prado:2006:MTS

[1114] Raquel Prado, Francisco Molina, and Gabriel Huerta. Multivariate time series modeling and classification via hierarchical VAR mixtures. Computational Statistics & Data Analysis, 51(3):1445–1462, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000739.

Pacheco:2006:ANV

[1115] Joaquín Pacheco, Silvia Casado, Laura Núñez, and Olga Gómez. Analysis of new variable selection methods for discriminant analysis. Computational Statistics & Data Analysis, 51(3):1463– 1478, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001198.

Garlipp:2006:DLC

[1116] T. Garlipp and C. H. Müller. Detection of linear and circular shapes in image analysis. Computational Statistics & Data Analysis, 51(3):1479– 1490, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001216.

Hanafi:2006:ASD

[1117] Mohamed Hanafi and Henk A. L. Kiers. Analysis of K sets of data, with differential emphasis on agreement between and within sets. Computational Statistics & Data Analysis, 51(3):1491– 1508, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600123X.

Jokinen:2006:FEA

[1118] Jukka Jokinen. Fast estimation algorithm for likelihood-based analysis of repeated categorical responses. Computational Statistics & Data Analysis, 51(3):1509–1522, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001241.

Helu:2006:EIC

[1119] Amal Helu and Dayanand N. Naik. Estimation of interclass correlation via a kotz-type distribution. Computational Statistics & Data Analysis, 51(3):1523– 1534, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001356.

Gonzalez:2006:NIL

[1120] Jorge González, Francis Tuerlinckx, Paul De Boeck, and Ronald Cools. Numerical integration in logistic-normal models. Computational Statistics & Data Analysis, 51(3):1535–1548, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600137X.

Kuroda:2006:ACE

[1121] Masahiro Kuroda and Michio Sakakihara. Accelerating the convergence of the EM algorithm using the vector ϵ algorithm. Computational Statistics & Data Analysis, 51(3):1549– 1561, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001381.

Samson:2006:ESA

[1122] Adeline Samson, Marc Lavielle, and France Mentré. Extension of the SAEM algorithm to leftcensored data in nonlinear mixedeffects model: Application to HIV dynamics model. Computational Statistics & Data Analysis, 51(3):1562– 1574, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001393.

Izsak:2006:MLE

[1123] F. Izsák. Maximum likelihood estimation for constrained parameters of multinomial distributions — application to Zipf–Mandelbrot models. Computational Statistics & Data Analysis, 51(3):1575–1583, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600154X.

Nadarajah:2006:FAS

[1124] Saralees Nadarajah. FIM for Arnold and Strauss’s bivariate gamma distribution. Computational Statistics & Data Analysis, 51(3):1584–1590, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001563.

Torsney:2006:TCM

[1125] B. Torsney and S. Mandal. Two classes of multiplicative algorithms for constructing optimizing distributions. Computational Statistics & Data Analysis, 51(3):1591–1601, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001782.

Guolo:2006:IIS

[1126] Annamaria Guolo, Alessandra R. Brazzale, and Alessandra Salvan. Improved inference on a scalar fixed effect of interest in nonlinear mixedeffects models. Computational Statistics & Data Analysis, 51(3):1602– 1613, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001812.

Wichmann:2006:GGP

[1127] B. A. Wichmann and I. D. Hill. Generating good pseudo-random numbers. Computational Statistics & Data Analysis, 51(3):1614–1622, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001836.

Schepers:2006:TMP

[1128] Jan Schepers, Iven van Mechelen, and Eva Ceulemans. Three-mode partitioning. Computational Statistics & Data Analysis, 51(3):1623– 1642, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001927.

Kim:2006:MSL

[1129] Yongdai Kim, Sunghoon Kwon, and Seuck Heun Song. Multiclass sparse logistic regression for classification of multiple cancer types using gene expression data. Computational Statistics & Data Analysis, 51(3):1643– 1655, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001976.

Frommlet:2006:PAD

[1130] F. Frommlet, M. Bogdan, and A. Futschik. Power analysis of database search using multiple scoring matrices. Computational Statistics & Data Analysis, 51(3):1656– 1663, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600199X.

Neugebauer:2006:CEL

[1131] Romain Neugebauer and Mark J. van der Laan. Causal effects in longitudinal studies: Definition and maximum likelihood estimation. Computational Statistics & Data Analysis, 51(3):1664– 1675, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002040.

Neugebauer:2006:GCE

[1132] Romain Neugebauer and Mark J. van der Laan. G-computation estimation for causal inference with complex longitudinal data. Computational Statistics & Data Analysis, 51(3):1676– 1697, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002076.

Sarkka:2006:AMP

[1133] Aila Särkkä and Eric Renshaw.

The analysis of marked point patterns evolving through space and time. Computational Statistics & Data Analysis, 51(3):1698–1718, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002222.

Oigaard:2006:BMA

[1134] Tor Arne Øigård, Håvard Rue, and Fred Godtliebsen. Bayesian multiscale analysis for time series data. Computational Statistics & Data Analysis, 51(3):1719–1730, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002520.

Lin:2006:IPR

[1135] Jen-Wen Lin and A. Ian McLeod.

Improved Peña–Rodriguez portmanteau test. Computational Statistics & Data Analysis, 51(3):1731–1738, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002003.

Demidenko:2006:CGM

[1136] Eugene Demidenko. Criteria for global minimum of sum of squares in nonlinear regression. Computational Statistics & Data Analysis, 51(3):1739– 1753, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002027.

Yuan:2006:FTE

[1137] Ming Yuan. Flexible temporal expression profile modelling using the Gaussian process. Computational Statistics & Data Analysis, 51(3):1754– 1764, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003105.

Reynes:2006:CBS

[1138] Christelle Reynès, Robert Sabatier, and Nicolas Molinari. Choice of Bsplines with free parameters in the flexible discriminant analysis context. Computational Statistics & Data Analysis, 51(3):1765–1778, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003117.

Fong:2006:SMA

[1139] P. W. Fong, W. K. Li, and Hong-Zhi An. A simple multivariate ARCH model specified by random coefficients. Computational Statistics & Data Analysis, 51(3):1779– 1802, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003129.

Storti:2006:MDE

[1140] G. Storti. Minimum distance estimation of GARCH(1, 1) models. Computational Statistics & Data Analysis, 51(3):1803–1821, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003130.

Yu:2006:CBD

[1141] Lining Yu and Eberhard O. Voit. Construction of bivariate S-distributions with copulas. Computational Statistics & Data Analysis, 51(3):1822– 1839, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003142.

Stover:2006:ASP

[1142] Jason H. Stover. An algorithm for simulating permutation tests of graph-intersection measures of association. Computational Statistics & Data Analysis, 51(3):1840–1847, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003154.

Faes:2006:HMA

[1143] Christel Faes, Helena Geys, Marc Aerts, and Geert Molenberghs. A hierarchical modeling approach for risk assessment in developmental toxicity studies. Computational Statistics & Data Analysis, 51(3):1848– 1861, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305003166.

Godolphin:2006:SRR

[1144] J. D. Godolphin. The specification of rank reducing observation sets in experimental design. Computational Statistics & Data Analysis, 51(3):1862– 1874, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305003178.

Adimari:2006:PPI

[1145] Gianfranco Adimari and Monica Chiogna. Partially parametric interval estimation of Pr{Y > X}. Computational Statistics & Data Analysis, 51(3):1875–1891, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730500318X.

DeGooijer:2006:DCP

[1146] Jan G. De Gooijer. Detecting changepoints in multidimensional stochastic processes. Computational Statistics & Data Analysis, 51(3):1892– 1903, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947305003191.

Philippe:2006:BAA

[1147] Anne Philippe. Bayesian analysis of autoregressive moving average processes with unknown orders. Computational Statistics & Data Analysis, 51(3):1904– 1923, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003208.

Zhao:2006:SNA

[1148] Qiang Zhao and Jianguo Sun. Semiparametric and nonparametric analysis of recurrent events with observation gaps. Computational Statistics & Data Analysis, 51(3):1924– 1933, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730500321X.

Wilcox:2006:CM

[1149] Rand R. Wilcox. Comparing medians. Computational Statistics & Data Analysis, 51(3):1934–1943, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003221.

Kauermann:2006:ATW

[1150] Göran Kauermann and Pavel Khomski.

Additive two-way hazards model with varying coefficients. Computational Statistics & Data Analysis, 51(3):1944– 1956, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003233.

Singh:2006:TAU

[1151] Parminder Singh and Wei Liu. A test against an umbrella ordered alternative. Computational Statistics & Data Analysis, 51(3):1957–1964, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003245.

Mitra:2006:FEU

[1152] Amit Mitra, Debasis Kundu, and Gunjan Agrawal. Frequency estimation of undamped exponential signals using genetic algorithms. Computational Statistics & Data Analysis, 51(3):1965– 1985, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003257.

Lu:2006:NEP

[1153] Zeng-Hua Lu. The numerical evaluation of the probability density function of a quadratic form in normal variables. Computational Statistics & Data Analysis, 51(3):1986– 1996, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003269.

Jiang:2006:TFA

[1154] Huiping Jiang, Henry L. Gray, and Wayne A. Woodward. Timefrequency analysis — G(λ)-stationary processes. Computational Statistics & Data Analysis, 51(3):1997–2028, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003270.

Miazhynskaia:2006:BTN

[1155] Tatiana Miazhynskaia, Sylvia Frühwirth-Schnatter, and Georg Dorffner. Bayesian testing for non-linearity in volatility modeling. Computational Statistics & Data Analysis, 51(3):2029– 2042, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003282.

Louw:2006:VSK

[1156] N. Louw and S. J. Steel. Variable selection in kernel Fisher discriminant analysis by means of recursive feature elimination. Computational Statistics & Data Analysis, 51(3):2043– 2055, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003294.

Seong:2006:ASB

[1157] Byeongchan Seong, A. K. M. Mahbub Morshed, and Sung K. Ahn. Additional sources of bias in half-life estimation. Computational Statistics & Data Analysis, 51(3):2056– 2064, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947305003312.

Soliman:2006:CEU

[1158] Ahmed A. Soliman, A. H. Abd Ellah, and K. S. Sultan. Comparison of estimates using record statistics from Weibull model: Bayesian and non-Bayesian approaches. Computational Statistics & Data Analysis, 51(3):2065– 2077, December 1, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600003X.

Tsai:2006:SDT

[1159] Shin-Fu Tsai, Chen-Tuo Liao, and Feng-Shun Chai. Statistical designs for two-color microarray experiments involving technical replication. Computational Statistics & Data Analysis, 51(3):2078–2090, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001824.

Lambert-Lacroix:2006:LLR

[1160] Sophie Lambert-Lacroix and Julie Peyre. Local likelihood regression in generalized linear single-index models with applications to microarray data. Computational Statistics & Data Analysis, 51(3):2091–2113, December 1, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002143.

Anonymous:2006:EBp

[1161] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51 (4):iii–v, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004038.

Anonymous:2006:Cd

[1162] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (4):vi–vii, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600404X.

Amendola:2006:SIN

[1163] Alessandra Amendola, Christian Francq, and Siem Jan Koopman. Special issue on nonlinear modelling and financial econometrics. Computational Statistics & Data Analysis, 51(4):2115– 2117, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003513.

Arteche:2006:SEP

[1164] J. Arteche. Semiparametric estimation in perturbed long memory series. Computational Statistics & Data Analysis, 51(4):2118–2141, December 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002398.

Duchesne:2006:TMA

[1165] Pierre Duchesne. Testing for multivariate autoregressive conditional heteroskedasticity using wavelets. Computational Statistics & Data Analysis, 51(4):2142–2163, December 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002386.

Chen:2006:CNA

[1166] Cathy W. S. Chen, Richard Gerlach, and Mike K. P. So. Comparison of nonnested asymmetric heteroskedastic models. Computational Statistics & Data Analysis, 51(4):2164– 2178, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002428.

DeLuca:2006:RSP

[1167] Giovanni De Luca and Paola Zuccolotto. Regime-switching Pareto distributions for ACD models. Computational Statistics & Data Analysis, 51(4):2179–2191, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947306002751.

Bulla:2006:SFF

[1168] Jan Bulla and Ingo Bulla. Stylized facts of financial time series and hidden semi-Markov models. Computational Statistics & Data Analysis, 51(4):2192– 2209, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002374.

Kim:2006:ASV

[1169] Myung Suk Kim and Suojin Wang. On the applicability of stochastic volatility models. Computational Statistics & Data Analysis, 51(4):2210– 2217, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600260X.

Yu:2006:CNS

[1170] Jun Yu, Zhenlin Yang, and Xibin Zhang. A class of nonlinear stochastic volatility models and its implications for pricing currency options. Computational Statistics & Data Analysis, 51(4):2218–2231, December 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002854.

Brownlees:2006:FEA

[1171] C. T. Brownlees and G. M. Gallo. Financial econometric analysis at ultrahigh frequency: Data handling concerns. Computational Statistics & Data Analysis, 51(4):2232–2245, December 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003458.

Frijns:2006:NDN

[1172] Bart Frijns and Peter C. Schotman.

Nonlinear dynamics in Nasdaq dealer quotes. Computational Statistics & Data Analysis, 51(4):2246–2266, December 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003355.

Audrino:2006:DME

[1173] Francesco Audrino and Giovanni Barone-Adesi. A dynamic model of expected bond returns: a functional gradient descent approach. Computational Statistics & Data Analysis, 51(4):2267– 2277, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002404.

Escanciano:2006:TMD

[1174] J. Carlos Escanciano and Carlos Velasco. Testing the martingale difference hypothesis using integrated regression functions. Computational Statistics & Data Analysis, 51(4):2278– 2294, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002416.

Hartz:2006:AVR

[1175] Christoph Hartz, Stefan Mittnik, and Marc Paolella. Accurate value-at-risk forecasting based on the normal-GARCH model. Computational Statistics & Data Analysis, 51(4):2295– 2312, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003367.

Zhang:2006:EFR

[1176] Zhengjun Zhang and James Huang. Extremal financial risk models and portfolio evaluation. Computational Statistics & Data Analysis, 51(4):2313– 2338, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003392.

Hlavka:2006:FAN

[1177] Zdeněk Hlávka. Fast algorithm for nonparametric arbitrage-free SPD estimation. Computational Statistics & Data Analysis, 51(4):2339–2349, December 15, 2006. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002660.

Jung:2006:TSC

[1178] Robert C. Jung, Martin Kukuk, and Roman Liesenfeld. Time series of count data: modeling, estimation and diagnostics. Computational Statistics & Data Analysis, 51(4):2350– 2364, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002581.

Varin:2006:PLI

[1179] Cristiano Varin and Paolo Vidoni. Pairwise likelihood inference for ordinal categorical time series. Computational Statistics & Data Analysis, 51(4):2365– 2373, December 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600332X.

Anonymous:2007:EBa

[1180] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51 (5):iv–vi, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004415.

Anonymous:2007:Ca

[1181] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (5):vii–viii, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004427.

Davis:2007:GK

[1182] Matthew K. Davis and Gemai Chen. Graphing Kendall’s τ. Computational Statistics & Data Analysis, 51(5):2375– 2378, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730600329X.

Guedon:2007:ESS

[1183] Yann Guédon. Exploring the state sequence space for hidden Markov and semi-Markov chains. Computational Statistics & Data Analysis, 51(5):2379– 2409, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002118.

Wu:2007:CEM

[1184] Lang Wu. A computationally efficient method for nonlinear mixedeffects models with nonignorable missing data in time-varying covariates. Computational Statistics & Data Analysis, 51(5):2410–2419, February 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002556.

Yu:2007:EJC

[1185] Binbing Yu, Michael J. Barrett, Hyune-Ju Kim, and Eric J. Feuer. Estimating joinpoints in continuous time scale for multiple change-point models. Computational Statistics & Data Analysis, 51(5):2420–2427, February 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002623.

Marsh:2007:GFT

[1186] Patrick Marsh. Goodness of fit tests via exponential series density estimation. Computational Statistics & Data Analysis, 51(5):2428– 2441, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002830.

Siani:2007:APB

[1187] Carole Siani and Christian de Peretti.

Analysing the performance of bootstrap neural tests for conditional heteroskedasticity in ARCH-M models. Computational Statistics & Data Analysis, 51(5):2442–2460, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306002866.

Mount:2007:PAA

[1188] David M. Mount, Nathan S. Netanyahu, Kathleen Romanik, Ruth Silverman, and Angela Y. Wu. A practical approximation algorithm for the LMS line estimator. Computational Statistics & Data Analysis, 51(5):2461– 2486, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002921.

Merler:2007:PAW

[1189] Stefano Merler, Bruno Caprile, and Cesare Furlanello. Parallelizing AdaBoost by weights dynamics. Computational Statistics & Data Analysis, 51(5):2487– 2498, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003239.

Hyvarinen:2007:SES

[1190] Aapo Hyvärinen. Some extensions of score matching. Computational Statistics & Data Analysis, 51(5):2499– 2512, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003264.

Abellanas:2007:PSS

[1191] Manuel Abellanas, Mercè Claverol, and Ferran Hurtado. Point set stratification and Delaunay depth. Computational Statistics & Data Analysis, 51(5):2513– 2530, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003276.

Bellio:2007:ABI

[1192] Ruggero Bellio. Algorithms for bounded-influence estimation. Computational Statistics & Data Analysis, 51(5):2531–2541, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306003379.

Jullion:2007:RSR

[1193] Astrid Jullion and Philippe Lambert. Robust specification of the roughness penalty prior distribution in spatially adaptive Bayesian P - splines models. Computational Statistics & Data Analysis, 51(5):2542– 2558, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003549.

Rueda:2007:QET

[1194] María del Mar Rueda, Antonio Arcos, Juan Francisco Muñoz, and Sarjinder Singh. Quantile estimation in two-phase sampling. Computational Statistics & Data Analysis, 51(5):2559– 2572, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000053.

DiZio:2007:MMM

[1195] Marco Di Zio, Ugo Guarnera, and Roberto Rocci. A mixture of mixture models for a classification problem: the unity measure error. Computational Statistics & Data Analysis, 51(5):2573– 2585, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000065.

Congdon:2007:BMS

[1196] P. Congdon. Bayesian modelling strategies for spatially varying regression coefficients: a multivariate perspective for multiple outcomes. Computational Statistics & Data Analysis, 51(5):2586– 2601, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000077.

Lian:2007:IDE

[1197] I. B. Lian, C. J. Chang, Y. J. Liang, M. J. Yang, and C. S. J. Fann. Identifying differentially expressed genes in dye-swapped microarray experiments of small sample size. Computational Statistics & Data Analysis, 51(5):2602– 2620, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000089.

Zhu:2007:HMI

[1198] Li-Xing Zhu, Megu Ohtaki, and Yingxing Li. On hybrid methods of inverse regression-based algorithms. Computational Statistics & Data Analysis, 51(5):2621–2635, February 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000090.

Ausin:2007:BEG

[1199] María Concepción Ausín and Pedro Galeano. Bayesian estimation of the Gaussian mixture GARCH model. Computational Statistics & Data Analysis, 51(5):2636–2652, February 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000119.

deUna-Alvarez:2007:NEL

[1200] Jacobo de Uña-Álvarez and Alberto Rodríguez-Casal. Nonparametric estimation from length-biased data under competing risks. Computational Statistics & Data Analysis, 51(5):2653– 2669, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000120.

Zhang:2007:DVP

[1201] Ping Zhang and Nicoleta Serban. Discovery, visualization and performance analysis of enterprise workflow. Computational Statistics & Data Analysis, 51(5):2670–2687, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306000132.

Lombardi:2007:BIS

[1202] Marco J. Lombardi. Bayesian inference for α-stable distributions: a random walk MCMC approach. Computational Statistics & Data Analysis, 51(5):2688– 2700, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000144.

Walde:2007:VHT

[1203] Janette F. Walde. Valid hypothesis testing in face of spatially dependent data using multi-layer perceptrons and sub-sampling techniques. Computational Statistics & Data Analysis, 51(5):2701–2719, February 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000156.

Gonzalez-Manteiga:2007:EMS

[1204] W. González-Manteiga, M. J. Lombardía, I. Molina, D. Morales, and L. Santamaría. Estimation of the mean squared error of predictors of small area linear parameters under a logistic mixed model. Computational Statistics & Data Analysis, 51(5):2720– 2733, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600017X.

Gel:2007:RDT

[1205] Yulia R. Gel, Weiwen Miao, and Joseph L. Gastwirth. Robust directed tests of normality against heavy-tailed alternatives. Computational Statistics & Data Analysis, 51(5):2734– 2746, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002805.

Novikov:2007:OLS

[1206] Ilya Novikov and Bernice Oberman. Optimization of large simulations using statistical software. Computational Statistics & Data Analysis, 51(5):2747– 2752, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600209X.

Porter:2007:DLR

[1207] Michael D. Porter and Donald E.

Brown. Detecting local regions of change in high-dimensional criminal or terrorist point processes. Computational Statistics & Data Analysis, 51(5):2753–2768, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306002179.

Contreras-Cristan:2007:UEA

[1208] Alberto Contreras-Cristán. Using the EM algorithm for inference in a mixture of distributions with censored but partially identifiable data. Computational Statistics & Data Analysis, 51(5):2769– 2781, February 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002519.

Anonymous:2007:EBb

[1209] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51(6):iii–v, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000436.

Anonymous:2007:Cb

[1210] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (6):vi–viii, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000448.

Cordeiro:2007:LOS

[1211] José Antonio Cordeiro. Logical operators of some statistical computing packages and missing values. Computational Statistics & Data Analysis, 51 (6):2783–2787, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004579.

Berger:2007:NOP

[1212] Roger L. Berger. Nonstandard operator precedence in Excel. Computational Statistics & Data Analysis, 51 (6):2788–2791, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003641.

Arslan:2007:CIM

[1213] Olcay Arslan. Comment on “Information matrices for Laplace and Pareto mixtures” by S. Nadarajah. Computational Statistics & Data Analysis, 51 (6):2792–2793, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002106. See

[851] and reply [1214].

Nadarajah:2007:RCA

[1214] Saralees Nadarajah. Reply to the comment by O. Arslan on “Information matrices for Laplace and Pareto mixtures”. Computational Statistics & Data Analysis, 51(6):2794– 2795, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003483. See

[1213].

Kundu:2007:CWG

[1215] Debasis Kundu and Rameshwar D.

Gupta. A convenient way of generating gamma random variables using generalized exponential distribution. Computational Statistics & Data Analysis, 51 (6):2796–2802, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003616.

Wang:2007:NIM

[1216] Hansheng Wang. A note on iterative marginal optimization: a simple algorithm for maximum rank correlation estimation. Computational Statistics & Data Analysis, 51(6):2803– 2812, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003665.

Fournier:2007:EPG

[1217] B. Fournier, N. Rupin, M. Bigerelle, D. Najjar, A. Iost, and R. Wilcox. Estimating the parameters of a generalized lambda distribution. Computational Statistics & Data Analysis, 51 (6):2813–2835, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003677.

Kim:2007:CSP

[1218] Gunky Kim, Mervyn J. Silvapulle, and Paramsothy Silvapulle. Comparison of semiparametric and parametric methods for estimating copulas. Computational Statistics & Data Analysis, 51 (6):2836–2850, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003690.

Avalos:2007:PAM

[1219] Marta Avalos, Yves Grandvalet, and Christophe Ambroise. Parsimonious additive models. Computational Statistics & Data Analysis, 51(6):2851– 2870, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003732.

Diaz:2007:CPL

[1220] Rafael E. Diaz. Comparison of PQL and Laplace 6 estimates of hierarchical linear models when comparing groups of small incident rates in cluster randomised trials. Computational Statistics & Data Analysis, 51 (6):2871–2888, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003744.

Kurowicka:2007:SAG

[1221] D. Kurowicka and R. M. Cooke. Sampling algorithms for generating joint uniform distributions using the vinecopula method. Computational Statistics & Data Analysis, 51(6):2889– 2906, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004269.

Broda:2007:SAD

[1222] Simon Broda and Marc S. Paolella.

Saddlepoint approximations for the doubly noncentral t distribution. Computational Statistics & Data Analysis, 51(6):2907–2918, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306004609.

Tang:2007:IEA

[1223] Man-Lai Tang, Kai Wang Ng, Guo-Liang Tian, and Ming Tan. On improved EM algorithm and confidence interval construction for incomplete r × c tables. Computational Statistics & Data Analysis, 51(6):2919– 2933, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004634.

Nandram:2007:CBI

[1224] Balgobin Nandram and Daniel Zelterman. Computational Bayesian inference for estimating the size of a finite population. Computational Statistics & Data Analysis, 51(6):2934– 2945, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004683.

Wang:2007:LPV

[1225] Ji-Ping Wang. A linearization procedure and a VDM/ECM algorithm for penalized and constrained nonparametric maximum likelihood estimation for mixture models. Computational Statistics & Data Analysis, 51 (6):2946–2957, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004695.

Koukouvinos:2007:ECS

[1226] Christos Koukouvinos, Kalliopi Mylona, and Dimitris E. Simos. Exploring k-circulant supersaturated designs via genetic algorithms. Computational Statistics & Data Analysis, 51 (6):2958–2968, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004750.

Steinsland:2007:PES

[1227] Ingelin Steinsland. Parallel exact sampling and evaluation of Gaussian Markov random fields. Computational Statistics & Data Analysis, 51 (6):2969–2981, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000181.

Stamey:2007:BSS

[1228] James Stamey and Richard Gerlach.

Bayesian sample size determination for case-control studies with misclassification. Computational Statistics & Data Analysis, 51(6):2982– 2992, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000193.

Karasoy:2007:NBE

[1229] Durdu Sertkaya Karasoy and Cem Kadilar. A new Bayes estimate of the change point in the hazard function. Computational Statistics & Data Analysis, 51(6):2993– 3001, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002210.

Gil:2007:TLI

[1230] María Ángeles Gil, Gil González-Rodríguez, Ana Colubi, and Manuel Montenegro. Testing linear independence in linear models with intervalvalued data. Computational Statistics & Data Analysis, 51(6):3002– 3015, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600020X.

Diaz-Garcia:2007:OAS

[1231] José A. Díaz-García and Ma. Magdalena Garay-Tápia. Optimum allocation in stratified surveys: Stochastic programming. Computational Statistics & Data Analysis, 51(6):3016– 3026, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000211.

Candel:2007:EBE

[1232] Math J. J. M. Candel. Empirical Bayes estimators of the random intercept in multilevel analysis: Performance of the classical, Morris and Rao version. Computational Statistics & Data Analysis, 51(6):3027– 3040, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000223.

Fotopoulos:2007:NLP

[1233] Stergios B. Fotopoulos, Venkata K. Jandhyala, and Kim-Heng Chen. Nonlinear properties of conditional returns under scale mixtures. Computational Statistics & Data Analysis, 51 (6):3041–3056, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000247.

Hazelton:2007:RKD

[1234] Martin L. Hazelton and Berwin A. Turlach. Reweighted kernel density estimation. Computational Statistics & Data Analysis, 51(6):3057– 3069, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000375.

Gupta:2007:SIR

[1235] Ramesh C. Gupta and Suzhong Tian.

Statistical inference for the risk ratio in 2 × 2 binomial trials with structural zero. Computational Statistics & Data Analysis, 51(6):3070– 3084, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000387.

Ng:2007:CST

[1236] H. K. T. Ng, K. Gu, and M. L. Tang. A comparative study of tests for the difference of two Poisson means. Computational Statistics & Data Analysis, 51 (6):3085–3099, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000399.

Haldrup:2007:EFI

[1237] Niels Haldrup and Morten Ørregaard Nielsen. Estimation of fractional integration in the presence of data noise. Computational Statistics & Data Analysis, 51(6):3100– 3114, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000405.

Olive:2007:PIR

[1238] David J. Olive. Prediction intervals for regression models. Computational Statistics & Data Analysis, 51 (6):3115–3122, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000417.

Renshaw:2007:DMP

[1239] Eric Renshaw, Jorge Mateu, and Fuensanta Saura. Disentangling mark/ point interaction in marked-point processes. Computational Statistics & Data Analysis, 51(6):3123– 3144, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002532.

Park:2007:CAA

[1240] Mira Park, Jae Won Lee, and Choongrak Kim. Correspondence analysis approach for finding allele associations in population genetic study. Computational Statistics & Data Analysis, 51(6):3145– 3155, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003240.

Cai:2007:CFS

[1241] D. Michael Cai, Maya Gokhale, and James Theiler. Comparison of feature selection and classification algorithms in identifying malicious executables. Computational Statistics & Data Analysis, 51(6):3156– 3172, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003288.

Li:2007:CEM

[1242] Yan Li, Julia L. Bienias, and David A.

Bennett. Confounding in the estimation of mediation effects. Computational Statistics & Data Analysis, 51 (6):3173–3186, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600377X.

Tran:2007:FOC

[1243] Nam H. Tran. Fracture orientation characterization: Minimizing statistical modelling errors. Computational Statistics & Data Analysis, 51(6):3187– 3196, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003938.

Congdon:2007:MSU

[1244] Peter Congdon. Mixtures of spatial and unstructured effects for spatially discontinuous health outcomes. Computational Statistics & Data Analysis, 51 (6):3197–3212, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004531.

Chen:2007:IME

[1245] Yuh-Ing Chen and Yu-Mei Chang.

Identification of the minimum effective dose for right-censored survival data. Computational Statistics & Data Analysis, 51(6):3213– 3222, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004555.

Garcia-Zattera:2007:CIM

[1246] María José García-Zattera, Alejandro Jara, Emmanuel Lesaffre, and Dominique Declerck. Conditional independence of multivariate binary data with an application in caries research. Computational Statistics & Data Analysis, 51(6):3223– 3234, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004580.

Liao:2007:RBM

[1247] Chen-Tuo Liao and Tsai-Yu Lin. Robust balanced measurement designs when errors are serially correlated. Computational Statistics & Data Analysis, 51(6):3235–3243, March 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004646.

Wang:2007:EAN

[1248] Jing Wang. EM algorithms for nonlinear mixed effects models. Computational Statistics & Data Analysis, 51 (6):3244–3256, March 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004592.

Anonymous:2007:EBc

[1249] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51(7):iii–v, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000813.

Anonymous:2007:Cc

[1250] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (7):vi–viii, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000825.

Belsley:2007:TSI

[1251] David A. Belsley, Erricos John Kontoghiorghes, and Jan R. Magnus. The third special issue on computational econometrics. Computational Statistics & Data Analysis, 51(7): 3258, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000588.

Davidson:2007:IRB

[1252] Russell Davidson and James G. MacKinnon. Improving the reliability of bootstrap tests with the fast double bootstrap. Computational Statistics & Data Analysis, 51(7):3259– 3281, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000879.

Godfrey:2007:AAI

[1253] L. G. Godfrey. Alternative approaches to implementing Lagrange multiplier tests for serial correlation in dynamic regression models. Computational Statistics & Data Analysis, 51 (7):3282–3295, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001848.

Kiviet:2007:AFS

[1254] Jan F. Kiviet and Jerzy Niemczyk. The asymptotic and finite sample distributions of OLS and simple IV in simultaneous equations. Computational Statistics & Data Analysis, 51 (7):3296–3318, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003446.

Gelper:2007:MST

[1255] Sarah Gelper and Christophe Croux. Multivariate out-of-sample tests for Granger causality. Computational Statistics & Data Analysis, 51(7):3319– 3329, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003501.

Esteban-Bravo:2007:WCE

[1256] Mercedes Esteban-Bravo and Jose M. Vidal-Sanz. Worst-case estimation for econometric models with unobservable components. Computational Statistics & Data Analysis, 51(7):3330– 3354, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003434.

Broda:2007:BAE

[1257] Simon Broda, Kai Carstensen, and Marc S. Paolella. Bias-adjusted estimation in the ARX(1) model. Computational Statistics & Data Analysis, 51 (7):3355–3367, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002234.

Flachaire:2007:EID

[1258] Emmanuel Flachaire and Olivier Nuñez. Estimation of the income distribution and detection of subpopulations: an explanatory model. Computational Statistics & Data Analysis, 51 (7):3368–3380, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002192.

Swamy:2007:EBL

[1259] P. A. V. B. Swamy, Wisam Yaghi, Jatinder S. Mehta, and I-Lok Chang. Empirical best linear unbiased prediction in misspecified and improved panel data models with an application to gasoline demand. Computational Statistics & Data Analysis, 51 (7):3381–3392, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000867.

Detemple:2007:MCM

[1260] Jérôme Detemple and Marcel Rindisbacher. Monte Carlo methods for derivatives of options with discontinuous payoffs. Computational Statistics & Data Analysis, 51(7):3393– 3417, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600288X.

Kim:2007:HLE

[1261] Jae H. Kim, Param Silvapulle, and Rob J. Hyndman. Half-life estimation based on the bias-corrected bootstrap: a highest density region approach. Computational Statistics & Data Analysis, 51(7):3418– 3432, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002209.

Trindade:2007:ADE

[1262] A. Alexandre Trindade and Yun Zhu. Approximating the distributions of estimators of financial risk under an asymmetric Laplace law. Computational Statistics & Data Analysis, 51 (7):3433–3447, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002647.

Fuertes:2007:SCM

[1263] Ana-Maria Fuertes and Elena Kalotychou. On sovereign credit migration: a study of alternative estimators and rating dynamics. Computational Statistics & Data Analysis, 51 (7):3448–3469, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002180.

Bartolucci:2007:MLE

[1264] Francesco Bartolucci and Valentina Nigro. Maximum likelihood estimation of an extended latent Markov model for clustered binary panel data. Computational Statistics & Data Analysis, 51 (7):3470–3483, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600466X.

Perez-Alonso:2007:BAT

[1265] Alicia Pérez-Alonso. A bootstrap approach to test the conditional symmetry in time series models. Computational Statistics & Data Analysis, 51 (7):3484–3504, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004841.

Geweke:2007:CTA

[1266] John Geweke, Patrick J. F. Groenen, Richard Paap, and Herman K. van Dijk. Computational techniques for applied econometric analysis of macroeconomic and financial processes. Computational Statistics & Data Analysis, 51 (7):3506–3508, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600449X.

Fruhwirth-Schnatter:2007:AMS

[1267] Sylvia Frühwirth-Schnatter and Rudolf Frühwirth. Auxiliary mixture sampling with applications to logistic models. Computational Statistics & Data Analysis, 51(7):3509– 3528, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003720.

Geweke:2007:IIM

[1268] John Geweke. Interpretation and inference in mixture models: Simple MCMC works. Computational Statistics & Data Analysis, 51(7):3529– 3550, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004506.

Bauwens:2007:MMN

[1269] L. Bauwens, C. M. Hafner, and J. V. K. Rombouts. Multivariate mixed normal conditional heteroskedasticity. Computational Statistics & Data Analysis, 51 (7):3551–3566, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003811.

Gelfand:2007:MMU

[1270] Alan E. Gelfand, Sudipto Banerjee, C. F. Sirmans, Yong Tu, and Seow Eng Ong. Multilevel modeling using spatial processes: Application to the Singapore housing market. Computational Statistics & Data Analysis, 51 (7):3567–3579, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004300.

Clements:2007:BPI

[1271] Michael P. Clements and Jae H. Kim.

Bootstrap prediction intervals for autoregressive time series. Computational Statistics & Data Analysis, 51 (7):3580–3594, April 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003331.

Gulpinar:2007:ROD

[1272] Nalan Gülpınar and Berc Rustem. Robust optimal decisions with imprecise forecasts. Computational Statistics & Data Analysis, 51(7):3595– 3611, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004543.

Heij:2007:FCP

[1273] Christiaan Heij, Patrick J. F. Groenen, and Dick van Dijk. Forecast comparison of principal component regression and principal covariate regression. Computational Statistics & Data Analysis, 51(7):3612– 3625, April 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003860.

Anonymous:2007:EBd

[1274] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51(8):iii–v, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001259.

Anonymous:2007:Cd

[1275] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (8):vi–viii, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001260.

Bertram:2007:EMP

[1276] William K. Bertram and M. Shelton Peiris. An example of a misclassification problem applied to Australian equity data. Computational Statistics & Data Analysis, 51(8): 3627–3630, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004828.

Paciorek:2007:CTS

[1277] Christopher J. Paciorek. Computational techniques for spatial logistic regression with large data sets. Computational Statistics & Data Analysis, 51 (8):3631–3653, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004294.

Drignei:2007:COD

[1278] Dorin Drignei. The computational order of a DACE dynamical model. Computational Statistics & Data Analysis, 51(8):3654– 3665, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004762.

Jakobowicz:2007:MPP

[1279] Emmanuel Jakobowicz and Christian Derquenne. A modified PLS path modeling algorithm handling reflective categorical variables and a new model building strategy. Computational Statistics & Data Analysis, 51 (8):3666–3678, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004804.

Lim:2007:AEE

[1280] Johan Lim, Xinlei Wang, and Michael Sherman. An adjustment for edge effects using an augmented neighborhood model in the spatial autologistic model. Computational Statistics & Data Analysis, 51(8):3679– 3688, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004816.

Berlinet:2007:ASA

[1281] A. Berlinet and Ch. Roland. Acceleration schemes with application to the EM algorithm. Computational Statistics & Data Analysis, 51 (8):3689–3702, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004919.

Matei:2007:CAO

[1282] Alina Matei and Yves Tillé. Computational aspects of order πps sampling schemes. Computational Statistics & Data Analysis, 51(8):3703– 3717, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004968.

Trendafilov:2007:DVS

[1283] Nickolay T. Trendafilov and Ian T. Jolliffe. DALASS: Variable selection in discriminant analysis via the LASSO. Computational Statistics & Data Analysis, 51(8):3718– 3736, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306005032.

Langholz:2007:CMC

[1284] Bryan Langholz and Jenny Jiao. Computational methods for casecohort studies. Computational Statistics & Data Analysis, 51(8):3737– 3748, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306005068.

Gilmour:2007:MMR

[1285] A. R. Gilmour. Mixed model regression mapping for QTL detection in experimental crosses. Computational Statistics & Data Analysis, 51 (8):3749–3764, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306005111.

Lloyd:2007:EET

[1286] Chris J. Lloyd. Efficient and exact tests of the risk ratio in a correlated 2 × 2 table with structural zero. Computational Statistics & Data Analysis, 51 (8):3765–3775, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005123.

Wang:2007:MLC

[1287] Yong Wang. Maximum likelihood computation based on the Fisher scoring and Gauss–Newton quadratic approximations. Computational Statistics & Data Analysis, 51(8):3776– 3787, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005147.

Hamasaki:2007:BCP

[1288] Toshimitsu Hamasaki and Seo Young Kim. Box and Cox power-transformation to confined and censored nonnormal responses in regression. Computational Statistics & Data Analysis, 51 (8):3788–3799, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004890.

Schwertman:2007:IOS

[1289] Neil C. Schwertman and Rapti de Silva.

Identifying outliers with sequential fences. Computational Statistics & Data Analysis, 51(8):3800– 3810, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000430.

Keeling:2007:CSR

[1290] Kellie B. Keeling and Robert J. Pavur. A comparative study of the reliability of nine statistical software packages. Computational Statistics & Data Analysis, 51(8):3811– 3831, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000491.

Lui:2007:SSC

[1291] Kung-Jong Lui. Sample size calculation for non-compliance randomized trials with repeated measurements in binary data. Computational Statistics & Data Analysis, 51 (8):3832–3843, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000697.

Huang:2007:IFL

[1292] Yufen Huang, Tzu-Ling Kao, and Tai-Ho Wang. Influence functions and local influence in linear discriminant analysis. Computational Statistics & Data Analysis, 51(8):3844– 3861, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000715.

Jarrett:2007:QCC

[1293] Jeffrey E. Jarrett and Xia Pan. The quality control chart for monitoring multivariate autocorrelated processes. Computational Statistics & Data Analysis, 51(8):3862– 3870, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000727.

Giordano:2007:FNT

[1294] Francesco Giordano, Michele La Rocca, and Cira Perna. Forecasting nonlinear time series with neural network sieve bootstrap. Computational Statistics & Data Analysis, 51(8): 3871–3884, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000740.

Tsai:2007:KEA

[1295] Chen-An Tsai and James J. Chen. Kernel estimation for adjusted p-values in multiple testing. Computational Statistics & Data Analysis, 51(8): 3885–3897, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000788.

Sturdivant:2007:SRB

[1296] Rodney X. Sturdivant and David W.

Hosmer, Jr. A smoothed residual based goodness-of-fit statistic for logistic hierarchical regression models. Computational Statistics & Data Analysis, 51 (8):3898–3912, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600079X.

Abrahantes:2007:CDE

[1297] José Cortiñas Abrahantes, Catherine Legrand, Tomasz Burzykowski, Paul Janssen, Vincent Ducrocq, and Luc Duchateau. Comparison of different estimation procedures for proportional hazards model with random effects. Computational Statistics & Data Analysis, 51(8):3913– 3930, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000818.

Demattei:2007:ASM

[1298] Christophe Dematteï, Nicolas Molinari, and Jean-Pierre Daurès. Arbitrarily shaped multiple spatial cluster detection for case event data. Computational Statistics & Data Analysis, 51 (8):3931–3945, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730600082X.

Best:2007:CSC

[1299] D. J. Best and J. C. W. Rayner. Chisquared components for tests of fit and improved models for the grouped exponential distribution. Computational Statistics & Data Analysis, 51 (8):3946–3954, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306000855.

Mohr:2007:BIC

[1300] Donna L. Mohr. Bayesian identification of clustered outliers in multiple regression. Computational Statistics & Data Analysis, 51(8):3955– 3967, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306000934.

Rayward-Smith:2007:SMC

[1301] V. J. Rayward-Smith. Statistics to measure correlation for data mining applications. Computational Statistics & Data Analysis, 51(8):3968– 3982, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001897.

Su:2007:NML

[1302] Steve Su. Numerical maximum log likelihood estimation for generalized lambda distributions. Computational Statistics & Data Analysis, 51(8): 3983–3998, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001988.

Chen:2007:EOC

[1303] Hubert J. Chen and Shun-Yi Chen.

Erratum to “The optimal confidence interval for the largest normal mean with unknown variance” [computational statistics and data analysis 47 (2004) 845–866]. Computational Statistics & Data Analysis, 51(8):3999– 4000, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/

article/pii/S0167947306000752. See

[648].

Ng:2007:EPI

[1304] H. K. T. Ng, D. Kundu, and N. Balakrishnan. Erratum to “Point and interval estimation for the two-parameter Birnbaum–Saunders distribution based on Type-II censored samples”: [computational statistics & data analysis 50 (2006) 3222–3242]. Computational Statistics & Data Analysis, 51(8):4001, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002325. See

[980].

Tong:2007:AMM

[1305] Donald D. M. Tong, Stephen Buxser, and Thomas J. Vidmar. Application of a mixture model for determining the cutoff threshold for activity in high-throughput screening. Computational Statistics & Data Analysis, 51 (8):4002–4012, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004889.

VanGinkel:2007:TWI

[1306] Joost R. Van Ginkel, L. Andries Van der Ark, Klaas Sijtsma, and Jeroen K. Vermunt. Two-way imputation: a Bayesian method for estimating missing scores in tests and questionnaires, and an accurate approximation. Computational Statistics & Data Analysis, 51(8):4013– 4027, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004993.

Cerrato:2007:BPU

[1307] Mario Cerrato and Nicholas Sarantis.

A bootstrap panel unit root test under cross-sectional dependence, with an application to PPP. Computational Statistics & Data Analysis, 51 (8):4028–4037, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005020.

Gutierrez:2007:AGC

[1308] R. Gutiérrez, N. Rico, P. Román, and F. Torres. Approximate and generalized confidence bands for the mean and mode functions of the lognormal diffusion process. Computational Statistics & Data Analysis, 51 (8):4038–4053, May 1, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005056.

Sauerbrei:2007:DIB

[1309] Willi Sauerbrei, Patrick Royston, and Karina Zapien. Detecting an interaction between treatment and a continuous covariate: a comparison of two approaches. Computational Statistics & Data Analysis, 51(8):4054– 4063, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005172.

Demirtas:2007:PBR

[1310] Hakan Demirtas, Lester M. Arguelles, Hwan Chung, and Donald Hedeker. On the performance of bias–reduction techniques for variance estimation in approximate Bayesian bootstrap imputation. Computational Statistics & Data Analysis, 51(8):4064– 4068, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000023.

Arbia:2007:SSP

[1311] Giuseppe Arbia, Giovanni Lafratta, and Carla Simeoni. Spatial sampling plans to monitor the 3-D spatial distribution of extremes in soil pollution surveys. Computational Statistics & Data Analysis, 51(8):4069– 4082, May 1, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000072.

Anonymous:2007:EBe

[1312] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51(9):iii–v, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001478.

Anonymous:2007:Ce

[1313] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (9):vi–ix, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700148X.

Tenenhaus:2007:KLP

[1314] Arthur Tenenhaus, Alain Giron, Emmanuel Viennet, Michel Béra, Gilbert Saporta, and Bernard Fertil. Kernel logistic PLS: a tool for supervised nonlinear dimensionality reduction and binary classification. Computational Statistics & Data Analysis, 51 (9):4083–4100, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000114.

Smidl:2007:BPC

[1315] Václav Šmídl and Anthony Quinn.

On Bayesian principal component analysis. Computational Statistics & Data Analysis, 51(9):4101– 4123, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000151.

Krampe:2007:BTS

[1316] Anne Krampe and Sonja Kuhnt.

Bowker’s test for symmetry and modifications within the algebraic framework. Computational Statistics & Data Analysis, 51(9):4124– 4142, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000229.

Wang:2007:PEW

[1317] P. C. Wang. Planning experiments when some specified interactions are investigated. Computational Statistics & Data Analysis, 51(9):4143– 4151, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700062X.

Tilahun:2007:FSM

[1318] Abel Tilahun, Assam Pryseley, Ariel Alonso, and Geert Molenberghs. Flexible surrogate marker evaluation from several randomized clinical trials with continuous endpoints, using R and SAS. Computational Statistics & Data Analysis, 51(9):4152– 4163, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000679.

Tsonaka:2007:PCG

[1319] R. Tsonaka and I. Moustaki. Parameter constraints in generalized linear latent variable models. Computational Statistics & Data Analysis, 51 (9):4164–4177, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001204.

Johnson:2007:MDI

[1320] Matthew S. Johnson. Modeling dichotomous item responses with freeknot splines. Computational Statistics & Data Analysis, 51(9):4178– 4192, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001228.

Vallejo-Arboleda:2007:CSB

[1321] Amparo Vallejo-Arboleda, José L.

Vicente-Villardón, and M. P. Galindo-Villardón. Canonical STATIS: Biplot analysis of multi-table group structured data based on STATIS-ACT methodology. Computational Statistics & Data Analysis, 51(9):4193– 4205, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001319.

vanDijk:2007:ASN

[1322] Dick van Dijk, Philip Hans Franses, and H. Peter Boswijk. Absorption of shocks in nonlinear autoregressive models. Computational Statistics & Data Analysis, 51(9):4206– 4226, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001320.

Jamshidian:2007:TEC

[1323] Mortaza Jamshidian and James R.

Schott. Testing equality of covariance matrices when data are incomplete. Computational Statistics & Data Analysis, 51(9):4227– 4239, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600140X.

Royston:2007:IRF

[1324] P. Royston and W. Sauerbrei. Improving the robustness of fractional polynomial models by preliminary covariate transformation: a pragmatic approach. Computational Statistics & Data Analysis, 51(9):4240– 4253, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001551.

Yu:2007:PAN

[1325] Jun-Wu Yu, Guo-Liang Tian, and Man-Lai Tang. Predictive analyses for nonhomogeneous Poisson processes with power law using Bayesian approach. Computational Statistics & Data Analysis, 51(9):4254– 4268, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001575.

Bajorunaite:2007:TST

[1326] Ruta Bajorunaite and John P. Klein. Two-sample tests of the equality of two cumulative incidence functions. Computational Statistics & Data Analysis, 51(9):4269– 4281, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001587.

Alston:2007:BMM

[1327] C. L. Alston, K. L. Mengersen, C. P. Robert, J. M. Thompson, P. J. Littlefield, D. Perry, and A. J. Ball. Bayesian mixture models in a longitudinal setting for analysing sheep CAT scan images. Computational Statistics & Data Analysis, 51(9):4282– 4296, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001770.

Gurka:2007:TTL

[1328] Matthew J. Gurka, Lloyd J. Edwards, and Leena Nylander-French. Testing transformations for the linear mixed model. Computational Statistics & Data Analysis, 51(9):4297– 4307, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001794.

Mante:2007:ARL

[1329] Claude Manté. Application of resampling and linear spline methods to spectral and dispersional analyses of long-memory processes. Computational Statistics & Data Analysis, 51 (9):4308–4323, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001800.

Baccini:2007:PSP

[1330] Michela Baccini, Annibale Biggeri, Corrado Lagazio, Aitana Lertxundi, and Marc Saez. Parametric and semiparametric approaches in the analysis of short-term effects of air pollution on health. Computational Statistics & Data Analysis, 51(9):4324– 4336, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001903.

John:2007:DAM

[1331] Majnu John and Carey E. Priebe. A data-adaptive methodology for finding an optimal weighted generalized Mann–Whitney–Wilcoxon statistic. Computational Statistics & Data Analysis, 51(9):4337–4353, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947306001939.

Osorio:2007:ALI

[1332] Felipe Osorio, Gilberto A. Paula, and Manuel Galea. Assessment of local influence in elliptical linear models with longitudinal structure. Computational Statistics & Data Analysis, 51 (9):4354–4368, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001940.

Al-Saleh:2007:FMG

[1333] Jamal A. Al-Saleh and Satish K. Agarwal. Finite mixture of gamma distributions: a conjugate prior. Computational Statistics & Data Analysis, 51 (9):4369–4378, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001952.

Woo:2007:REM

[1334] Mi-Ja Woo and T. N. Sriram. Robust estimation of mixture complexity for count data. Computational Statistics & Data Analysis, 51(9):4379– 4392, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001964.

Hazelton:2007:BRK

[1335] Martin L. Hazelton. Bias reduction in kernel binary regression. Computational Statistics & Data Analysis, 51 (9):4393–4402, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002039.

Garcia-Escudero:2007:ISH

[1336] L. A. García-Escudero and A. Gordaliza. The importance of the scales in heterogeneous robust clustering. Computational Statistics & Data Analysis, 51 (9):4403–4412, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002052.

Zhang:2007:AEM

[1337] Jiajia Zhang and Yingwei Peng. An alternative estimation method for the accelerated failure time frailty model. Computational Statistics & Data Analysis, 51(9):4413–4423, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002088.

Tang:2007:CNB

[1338] Yongqiang Tang and Subhashis Ghosal.

A consistent nonparametric Bayesian procedure for estimating autoregressive conditional densities. Computational Statistics & Data Analysis, 51 (9):4424–4437, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002131.

Yang:2007:BNN

[1339] Tae Young Yang and Jae Chang Lee. Bayesian nearest-neighbor analysis via record value statistics and nonhomogeneous spatial Poisson processes. Computational Statistics & Data Analysis, 51(9):4438– 4449, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002155.

Archer:2007:GFT

[1340] Kellie J. Archer, Stanley Lemeshow, and David W. Hosmer. Goodness-of-fit tests for logistic regression models when data are collected using a complex sampling design. Computational Statistics & Data Analysis, 51 (9):4450–4464, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002167.

Schweinberger:2007:MMD

[1341] Michael Schweinberger and Tom A. B. Snijders. Markov models for digraph panel data: Monte Carlo-based derivative estimation. Computational Statistics & Data Analysis, 51(9):4465– 4483, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002283.

Asquith:2007:MMG

[1342] William H. Asquith. L-moments and T L-moments of the generalized lambda distribution. Computational Statistics & Data Analysis, 51(9):4484– 4496, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002301.

Kus:2007:NLD

[1343] Coşkun Kuş. A new lifetime distribution. Computational Statistics & Data Analysis, 51(9):4497– 4509, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002313.

Mielniczuk:2007:EHE

[1344] J. Mielniczuk and P. Wojdyłło.

Estimation of Hurst exponent revisited. Computational Statistics & Data Analysis, 51(9):4510–4525, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002337.

Carvalho:2007:SBS

[1345] Carlos M. Carvalho and Hedibert F.

Lopes. Simulation-based sequential analysis of Markov switching stochastic volatility models. Computational Statistics & Data Analysis, 51(9):4526– 4542, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002349.

Ip:2007:GLR

[1346] Wai-Cheung Ip, Heung Wong, and Riquan Zhang. Generalized likelihood ratio test for varying-coefficient models with different smoothing variables. Computational Statistics & Data Analysis, 51(9):4543– 4561, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002441.

Langdon:2007:CTC

[1347] Matthew J. Langdon, Charles C. Taylor, and Robert M. West. Classification of type I-censored bivariate data. Computational Statistics & Data Analysis, 51(9):4562– 4576, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002544.

Huang:2007:MCS

[1348] H.-C. Huang, F. Martinez, J. Mateu, and F. Montes. Model comparison and selection for stationary space–time models. Computational Statistics & Data Analysis, 51(9):4577– 4596, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730600257X.

Chen:2007:BEE

[1349] D. G. Chen. Bootstrapping estimation for estimating relative potency in combinations of bioassays. Computational Statistics & Data Analysis, 51 (9):4597–4604, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002611.

Leitenstorfer:2007:KSB

[1350] Florian Leitenstorfer and Gerhard Tutz. Knot selection by boosting techniques. Computational Statistics & Data Analysis, 51(9):4605– 4621, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002684.

Gerard:2007:EST

[1351] Patrick D. Gerard and William R.

Schucany. An enhanced sign test for dependent binary data with small numbers of clusters. Computational Statistics & Data Analysis, 51(9):4622– 4632, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002714.

Dobric:2007:GFT

[1352] Jadran Dobrić and Friedrich Schmid. A goodness of fit test for copulas based on Rosenblatt’s transformation. Computational Statistics & Data Analysis, 51 (9):4633–4642, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002763.

Iliopoulos:2007:BEU

[1353] G. Iliopoulos, M. Kateri, and I. Ntzoufras. Bayesian estimation of unrestricted and order-restricted association models for a two-way contingency table. Computational Statistics & Data Analysis, 51(9):4643– 4655, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002775.

Lemonte:2007:ISI

[1354] Artur J. Lemonte, Francisco Cribari-Neto, and Klaus L. P. Vasconcellos. Improved statistical inference for the two-parameter Birnbaum– Saunders distribution. Computational Statistics & Data Analysis, 51(9):4656– 4681, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002799.

Zhang:2007:STB

[1355] Jin Zhang and Yuehua Wu. ksample tests based on the likelihood ratio. Computational Statistics & Data Analysis, 51(9):4682– 4691, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002891.

Xie:2007:DAL

[1356] Feng-Chang Xie and Bo-Cheng Wei. Diagnostics analysis for log-Birnbaum– Saunders regression models. Computational Statistics & Data Analysis, 51 (9):4692–4706, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002908.

Yamamoto:2007:CAP

[1357] Michiyo Yamamoto, Takakazu Sugiyama, Hidetoshi Murakami, and Fumitake Sakaori. Correlation analysis of principal components from two populations. Computational Statistics & Data Analysis, 51(9):4707– 4716, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002933.

Urmeneta:2007:RRS

[1358] Henar Urmeneta and Víctor Hernández.

Random Riemann sum estimator versus Monte Carlo. Computational Statistics & Data Analysis, 51(9):4717– 4730, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003252.

Bensic:2007:EWU

[1359] Mirta Benšić and Kristian Sabo. Estimating the width of a uniform distribution when data are measured with additive normal errors with known variance. Computational Statistics & Data Analysis, 51(9):4731– 4741, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003318.

Colubi:2007:TFR

[1360] Ana Colubi and Gil González-Rodríguez. Triangular fuzzification of random variables and power of distribution tests: Empirical discussion. Computational Statistics & Data Analysis, 51(9):4742–4750, May 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004324.

Hall:2007:BRB

[1361] Peter Hall. Book review: Nonparametric functional data analysis: Theory and practice, F. Ferraty, P. Vieu, in: Springer Series in Statistics. Springer, Berlin (2006), xx + 268 pp, ISBN 978-0-387-30369-7. Computational Statistics & Data Analysis, 51 (9):4751–4752, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000941.

Hu:2007:ADP

[1362] Xiaomi Hu. Asymptotical distributions, parameters and coverage probabilities of tolerance limits. Computational Statistics & Data Analysis, 51 (9):4753–4760, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000084.

Breeden:2007:MDM

[1363] Joseph L. Breeden. Modeling data with multiple time dimensions. Computational Statistics & Data Analysis, 51 (9):4761–4785, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000631.

Anonymous:2007:EBf

[1364] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51(10):iii–v, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001740.

Anonymous:2007:Cf

[1365] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (10):vi–viii, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001752.

Manteiga:2007:SFD

[1366] Wenceslao González Manteiga and Philippe Vieu. Statistics for functional data. Computational Statistics & Data Analysis, 51(10):4788– 4792, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003859.

Antoniadis:2007:EIF

[1367] Anestis Antoniadis and Theofanis Sapatinas. Estimation and inference in functional mixed-effects models. Computational Statistics & Data Analysis, 51(10):4793–4813, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003628.

Cuesta-Albertos:2007:RPM

[1368] J. A. Cuesta-Albertos, E. del Barrio, R. Fraiman, and C. Matrán. The random projection method in goodness of fit for functional data. Computational Statistics & Data Analysis, 51(10): 4814–4831, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003306.

Cardot:2007:SSE

[1369] Hervé Cardot, Christophe Crambes, Alois Kneip, and Pascal Sarda. Smoothing splines estimators in functional linear regression with errorsin-variables. Computational Statistics & Data Analysis, 51(10):4832– 4848, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002477.

Chiou:2007:DFR

[1370] Jeng-Min Chiou and Hans-Georg Müller. Diagnostics for functional regression via residual processes. Computational Statistics & Data Analysis, 51(10):4849–4863, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002465.

Cuesta-Albertos:2007:ITM

[1371] Juan Antonio Cuesta-Albertos and Ricardo Fraiman. Impartial trimmed k-means for functional data. Computational Statistics & Data Analysis, 51(10):4864–4877, June 15, 2007.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306002258.

Dabo-Niang:2007:UMC

[1372] Sophie Dabo-Niang, Frédéric Ferraty, and Philippe Vieu. On the using of modal curves for radar waveforms classification. Computational Statistics & Data Analysis, 51(10):4878– 4890, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600226X.

Escabias:2007:FPL

[1373] M. Escabias, A. M. Aguilera, and M. J. Valderrama. Functional PLS logit regression model. Computational Statistics & Data Analysis, 51(10): 4891–4902, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002726.

Ferraty:2007:FBC

[1374] Frédéric Ferraty, Philippe Vieu, and Sylvie Viguier-Pla. Factorbased comparison of groups of curves. Computational Statistics & Data Analysis, 51(10):4903–4910, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003689.

Harezlak:2007:PSF

[1375] Jaroslaw Harezlak, Brent A. Coull, Nan M. Laird, Shannon R. Magari, and David C. Christiani. Penalized solutions to functional regression problems. Computational Statistics & Data Analysis, 51(10):4911– 4925, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600346X.

Hlubinka:2007:CAR

[1376] Daniel Hlubinka and Luboš Prchal. Changes in atmospheric radiation from the statistical point of view. Computational Statistics & Data Analysis, 51(10):4926– 4941, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002489.

Hyndman:2007:RFM

[1377] Rob J. Hyndman and Md. Shahid Ullah. Robust forecasting of mortality and fertility rates: a functional data approach. Computational Statistics & Data Analysis, 51(10):4942– 4956, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002453.

Lopez-Pintado:2007:DBI

[1378] Sara López-Pintado and Juan Romo. Depth-based inference for functional data. Computational Statistics & Data Analysis, 51(10):4957–4968, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003872.

Mante:2007:PCA

[1379] Claude Manté, Anne-Françoise Yao, and Claude Degiovanni. Principal component analysis of measures, with special emphasis on grainsize curves. Computational Statistics & Data Analysis, 51(10):4969– 4983, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002635.

Nerini:2007:CDU

[1380] David Nerini and Badih Ghattas.

Classifying densities using functional regression trees: Applications in oceanology. Computational Statistics & Data Analysis, 51(10):4984– 4993, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003550.

Ruiz-Medina:2007:KFP

[1381] M. D. Ruiz-Medina, R. Salmerón, and J. M. Angulo. Kalman filtering from POP-based diagonalization of ARH(1). Computational Statistics & Data Analysis, 51(10):4994– 5008, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002271.

Wilcox:2007:NRM

[1382] Rand R. Wilcox, Simon Sheather, Edgar Brunner, and Michael G.

Schimek. Nonparametric and robust methods. Computational Statistics & Data Analysis, 51(10):5010– 5012, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003975.

Hubert:2007:FCV

[1383] Mia Hubert and Sanne Engelen. Fast cross-validation of high-breakdown resampling methods for PCA. Computational Statistics & Data Analysis, 51(10):5013–5024, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016794730600291X.

Wellmann:2007:CSD

[1384] R. Wellmann, S. Katina, and Ch. H. Müller. Calculation of simplicial depth estimators for polynomial regression with applications. Computational Statistics & Data Analysis, 51(10): 5025–5040, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003847.

Werner:2007:RMA

[1385] Carola Werner and Edgar Brunner. Rank methods for the analysis of clustered data in diagnostic trials. Computational Statistics & Data Analysis, 51(10):5041– 5054, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001873.

Neuhauser:2007:CLT

[1386] Markus Neuhäuser, Christian Lösch, and Karl-Heinz Jöckel. The Chen– Luo test in case of heteroscedasticity. Computational Statistics & Data Analysis, 51(10):5055–5060, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001253.

Murakami:2007:LTS

[1387] Hidetoshi Murakami. Lepage type statistic based on the modified Baumgartner statistic. Computational Statistics & Data Analysis, 51(10):5061– 5067, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001265.

Nashimoto:2007:NMC

[1388] Kane Nashimoto and F. T. Wright.

Nonparametric multiple-comparison methods for simply ordered medians. Computational Statistics & Data Analysis, 51(10):5068–5076, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306001861.

Zhan:2007:BET

[1389] Xiaojiang Zhan and Thomas P.

Hettmansperger. Bayesian R-estimates in two-sample location models. Computational Statistics & Data Analysis, 51(10):5077–5089, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947306000636.

Zetlaoui:2007:RET

[1390] M. Zetlaoui, N. Picard, and A. Bar-Hen. Robustness of the estimators of transition rates for size-classified matrix models. Computational Statistics & Data Analysis, 51(10):5090– 5102, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002490.

Taskinen:2007:ICA

[1391] S. Taskinen, S. Sirkiä, and H. Oja. Independent component analysis based on symmetrised scatter matrices. Computational Statistics & Data Analysis, 51(10):5103–5111, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947306002246.

Belzunce:2007:QCD

[1392] F. Belzunce, A. Castaño, A. Olvera-Cervantes, and A. Suárez-Llorens. Quantile curves and dependence structure for bivariate distributions. Computational Statistics & Data Analysis, 51(10):5112–5129, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947306002738.

Glenn:2007:WEL

[1393] N. L. Glenn and Yichuan Zhao. Weighted empirical likelihood estimates and their robustness properties. Computational Statistics & Data Analysis, 51(10):5130–5141, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002507.

Jacqmin-Gadda:2007:RLM

[1394] Hélène Jacqmin-Gadda, Solenne Sibillot, Cécile Proust, Jean-Michel Molina, and Rodolphe Thiébaut. Robustness of the linear mixed model to misspecified error distribution. Computational Statistics & Data Analysis, 51(10): 5142–5154, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600185X.

Kovac:2007:SFL

[1395] A. Kovac. Smooth functions and local extreme values. Computational Statistics & Data Analysis, 51(10): 5155–5171, June 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600274X.

Bertaccini:2007:RAV

[1396] Bruno Bertaccini and Roberta Varriale. Robust ANalysis of VAriance: an approach based on the forward search. Computational Statistics & Data Analysis, 51(10):5172– 5183, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002702.

Jureckova:2007:SWT

[1397] Jana Jurečková and Jan Picek. Shapiro–Wilk-type test of normality under nuisance regression and scale. Computational Statistics & Data Analysis, 51(10):5184–5191, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002878.

Neubert:2007:SPT

[1398] Karin Neubert and Edgar Brunner. A Studentized permutation test for the non-parametric Behrens– Fisher problem. Computational Statistics & Data Analysis, 51(10):5192– 5204, June 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306001885.

Anonymous:2007:EBg

[1399] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51(11):iii–v, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001879.

Anonymous:2007:Cg

[1400] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (11):vi–vii, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001880.

Bohning:2007:AMM

[1401] Dankmar Böhning, Wilfried Seidel, Macro Alfó, Bernard Garel, Valentin Patilea, and Günther Walther. Advances in mixture models. Computational Statistics & Data Analysis, 51 (11):5205–5210, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003963.

Mao:2007:EPS

[1402] Chang Xuan Mao. Estimating population sizes for capture–recapture sampling with binomial mixtures. Computational Statistics & Data Analysis, 51 (11):5211–5219, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003525.

Caffo:2007:FRI

[1403] Brian Caffo, Ming-Wen An, and Charles Rohde. Flexible random intercept models for binary outcomes using mixtures of normals. Computational Statistics & Data Analysis, 51 (11):5220–5235, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003574.

Formann:2007:MAM

[1404] Anton K. Formann. Mixture analysis of multivariate categorical data with covariates and missing entries. Computational Statistics & Data Analysis, 51 (11):5236–5246, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002829.

Grun:2007:FFM

[1405] Bettina Grün and Friedrich Leisch. Fitting finite mixtures of generalized linear regressions in R. Computational Statistics & Data Analysis, 51(11): 5247–5252, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002787.

Alfo:2007:RSM

[1406] Marco Alfò, Alessio Farcomeni, and Luca Tardella. Robust semiparametric mixing for detecting differentially expressed genes in microarray experiments. Computational Statistics & Data Analysis, 51(11):5253– 5265, July 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002696.

Carvalho:2007:MNC

[1407] Alexandre X. Carvalho and Martin A. Tanner. Modelling nonlinear count time series with local mixtures of Poisson autoregressions. Computational Statistics & Data Analysis, 51(11): 5266–5294, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003586.

Garel:2007:RAR

[1408] Bernard Garel. Recent asymptotic results in testing for mixtures. Computational Statistics & Data Analysis, 51 (11):5295–5304, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003537.

DiZio:2007:ITF

[1409] Marco Di Zio, Ugo Guarnera, and Orietta Luzi. Imputation through finite Gaussian mixture models. Computational Statistics & Data Analysis, 51 (11):5305–5316, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003707.

Kheiri:2007:BAI

[1410] Soleiman Kheiri, Alan Kimber, and Mohammad Reza Meshkani. Bayesian analysis of an inverse Gaussian correlated frailty model. Computational Statistics & Data Analysis, 51(11): 5317–5326, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003422.

McLachlan:2007:EMF

[1411] G. J. McLachlan, R. W. Bean, and L. Ben-Tovim Jones. Extension of the mixture of factor analyzers model to incorporate the multivariate tdistribution. Computational Statistics & Data Analysis, 51(11):5327– 5338, July 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003410.

Ingrassia:2007:CME

[1412] Salvatore Ingrassia and Roberto Rocci. Constrained monotone EM algorithms for finite mixture of multivariate Gaussians. Computational Statistics & Data Analysis, 51(11):5339– 5351, July 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730600380X.

McGrory:2007:VAB

[1413] C. A. McGrory and D. M. Titterington. Variational approximations in Bayesian model selection for finite mixture distributions. Computational Statistics & Data Analysis, 51(11): 5352–5367, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002362.

Vermunt:2007:HMM

[1414] Jeroen K. Vermunt. A hierarchical mixture model for clustering threeway data sets. Computational Statistics & Data Analysis, 51(11):5368– 5376, July 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306002659.

Sultan:2007:MTI

[1415] K. S. Sultan, M. A. Ismail, and A. S. Al-Moisheer. Mixture of two inverse Weibull distributions: Properties and estimation. Computational Statistics & Data Analysis, 51(11): 5377–5387, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003409.

Karlis:2007:CIH

[1416] Dimitris Karlis and Valentin Patilea.

Confidence intervals of the hazard rate function for discrete distributions using mixtures. Computational Statistics & Data Analysis, 51(11):5388– 5401, July 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002568.

Jara:2007:DPM

[1417] Alejandro Jara, María José García-Zattera, and Emmanuel Lesaffre. A Dirichlet process mixture model for the analysis of correlated binary responses. Computational Statistics & Data Analysis, 51(11):5402– 5415, July 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003343.

Steiner:2007:CLD

[1418] P. M. Steiner and M. Hudec. Classification of large data sets with mixture models via sufficient EM. Computational Statistics & Data Analysis, 51 (11):5416–5428, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003380.

Bordes:2007:SEA

[1419] Laurent Bordes, Didier Chauveau, and Pierre Vandekerkhove. A stochastic EM algorithm for a semiparametric mixture model. Computational Statistics & Data Analysis, 51(11): 5429–5443, July 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002593.

Eilers:2007:NPL

[1420] Paul H. C. Eilers and M. W. Borgdorff. Non-parametric log-concave mixtures. Computational Statistics & Data Analysis, 51(11):5444– 5451, July 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002817.

Rufo:2007:BAF

[1421] M. J. Rufo, C. J. Pérez, and J. Martín. Bayesian analysis of finite mixtures of multinomial and negative-multinomial distributions. Computational Statistics & Data Analysis, 51(11):5452– 5466, July 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600243X.

Anonymous:2007:EBh

[1422] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 51 (12):iii–v, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002769.

Anonymous:2007:Ch

[1423] Anonymous. Contents. Computational Statistics & Data Analysis, 51 (12):vi–xi, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002770.

Rosenthal:2007:ARI

[1424] Jeffrey S. Rosenthal. AMCMC: an R interface for adaptive MCMC. Computational Statistics & Data Analysis, 51(12):5467–5470, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947307000916.

Malefaki:2007:SMD

[1425] Sonia Malefaki and George Iliopoulos. Simulating from a multinomial distribution with large number of categories. Computational Statistics & Data Analysis, 51(12):5471– 5476, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001405.

Park:2007:CAF

[1426] Juhyun Park. On the choice of an auxiliary function in the M/G/∞ estimation. Computational Statistics & Data Analysis, 51(12):5477– 5482, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002253.

Robin:2007:SPA

[1427] Stéphane Robin, Avner Bar-Hen, Jean-Jacques Daudin, and Laurent Pierre. A semi-parametric approach for mixture models: Application to local false discovery rate estimation. Computational Statistics & Data Analysis, 51(12): 5483–5493, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000990.

Qian:2007:MAD

[1428] Dajun Qian, John Kaddis, and Joyce C. Niland. A matching algorithm for the distribution of human pancreatic islets. Computational Statistics & Data Analysis, 51(12): 5494–5506, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001016.

Ren:2007:BAA

[1429] Dianxu Ren and Roslyn A. Stone. A Bayesian approach for analyzing a cluster-randomized trial with adjustment for risk misclassification. Computational Statistics & Data Analysis, 51(12):5507–5518, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001107.

Chen:2007:NSM

[1430] Ray-Bing Chen and Ying Nian Wu. A null space method for over-complete blind source separation. Computational Statistics & Data Analysis, 51(12): 5519–5536, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001120.

Hines:2007:CMS

[1431] R. J. O’Hara Hines and W. G. S.

Hines. Covariance miss-specification and the local influence approach in sensitivity analyses of longitudinal data with drop-outs. Computational Statistics & Data Analysis, 51(12):5537– 5546, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700120X.

Martin-Martin:2007:CMC

[1432] R. Martín-Martín, B. Torsney, and J. López-Fidalgo. Construction of marginally and conditionally restricted designs using multiplicative algorithms. Computational Statistics & Data Analysis, 51(12):5547– 5561, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001387.

Cojbasic:2007:NCI

[1433] Vesna Cojbasic and Andrija Tomovic.

Nonparametric confidence intervals for population variance of one sample and the difference of variances of two samples. Computational Statistics & Data Analysis, 51(12):5562– 5578, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001399.

Xu:2007:EGS

[1434] C. Xu and G. Gertner. Extending a global sensitivity analysis technique to models with correlated parameters. Computational Statistics & Data Analysis, 51(12):5579–5590, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001582.

Pea:2007:SSM

[1435] Johan Pea, Lionel Qualité, and Yves Tillé. Systematic sampling is a minimum support design. Computational Statistics & Data Analysis, 51(12): 5591–5602, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001661.

Dobigeon:2007:JSW

[1436] Nicolas Dobigeon and Jean-Yves Tourneret. Joint segmentation of wind speed and direction using a hierarchical model. Computational Statistics & Data Analysis, 51(12):5603– 5621, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001971.

Obereder:2007:BDE

[1437] Andreas Obereder, Otmar Scherzer, and Arne Kovac. Bivariate density estimation using BV regularisation. Computational Statistics & Data Analysis, 51(12):5622–5634, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001995.

Isaia:2007:QPM

[1438] A. Durio E. D. Isaia. A quick procedure for model selection in the case of mixture of normal densities. Computational Statistics & Data Analysis, 51(12):5635–5643, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002228.

Patan:2007:OED

[1439] Maciej Patan and Barbara Bogacka.

Optimum experimental designs for dynamic systems in the presence of correlated errors. Computational Statistics & Data Analysis, 51(12):5644– 5661, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700223X.

Wang:2007:MDC

[1440] Yong Wang. Minimum disparity computation via the iteratively reweighted least integrated squares algorithms. Computational Statistics & Data Analysis, 51(12):5662– 5672, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002265.

Pal:2007:RCM

[1441] Nabendu Pal, Jyh-Jiuan Lin, Ching-Hui Chang, and Somesh Kumar. A revisit to the common mean problem: Comparing the maximum likelihood estimator with the Graybill– Deal estimator. Computational Statistics & Data Analysis, 51(12):5673– 5681, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001417.

Aitken:2007:EEE

[1442] C. G. G. Aitken, Q. Shen, R. Jensen, and B. Hayes. The evaluation of evidence for exponentially distributed data. Computational Statistics & Data Analysis, 51(12):5682– 5693, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002277.

Leiva:2007:IDL

[1443] Víctor Leiva, Michelli Barros, Gilberto A. Paula, and Manuel Galea. Influence diagnostics in log-Birnbaum– Saunders regression models with censored data. Computational Statistics & Data Analysis, 51(12):5694– 5707, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003495.

Numasaki:2007:NCM

[1444] Hodaka Numasaki, Hajime Harauchi, Yuko Ohno, Kiyonari Inamura, Satoko Kasahara, Morito Monden, and Masato Sakon. New classification of medical staff clinical services for optimal reconstruction of job workflow in a surgical ward: Application of spectrum analysis and sequence relational analysis. Computational Statistics & Data Analysis, 51(12):5708– 5717, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003562.

Vaida:2007:EHE

[1445] Florin Vaida, Anthony P. Fitzgerald, and Victor DeGruttola. Efficient hybrid EM for linear and nonlinear mixed effects models with censored response. Computational Statistics & Data Analysis, 51(12):5718– 5730, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003604.

Krishnamoorthy:2007:PBA

[1446] K. Krishnamoorthy, Fei Lu, and Thomas Mathew. A parametric bootstrap approach for ANOVA with unequal variances: Fixed and random models. Computational Statistics & Data Analysis, 51(12):5731– 5742, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600363X.

Beaudoin:2007:IEK

[1447] David Beaudoin, Thierry Duchesne, and Christian Genest. Improving the estimation of Kendall’s tau when censoring affects only one of the variables. Computational Statistics & Data Analysis, 51(12):5743–5764, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003653.

Pan:2007:QMC

[1448] Jianxin Pan and Robin Thompson. Quasi-Monte Carlo estimation in generalized linear mixed models. Computational Statistics & Data Analysis, 51(12):5765–5775, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003719.

Li:2007:LPL

[1449] Erning Li, Daowen Zhang, and Marie Davidian. Likelihood and pseudolikelihood methods for semiparametric joint models for a primary endpoint and longitudinal data. Computational Statistics & Data Analysis, 51(12): 5776–5790, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003756.

Santner:2007:SSC

[1450] Thomas J. Santner, Vivek Pradhan, Pralay Senchaudhuri, Cyrus R. Mehta, and Ajit Tamhane. Small-sample comparisons of confidence intervals for the difference of two independent binomial proportions. Computational Statistics & Data Analysis, 51(12):

5791–5799, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003768.

Quintela-del-Rio:2007:PBS

[1451] Alejandro Quintela del Río. Plugin bandwidth selection in kernel hazard estimation from dependent data. Computational Statistics & Data Analysis, 51(12):5800–5812, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003793.

Ghosh:2007:SCH

[1452] Joyee Ghosh, Jerome P. Reiter, and Alan F. Karr. Secure computation with horizontally partitioned data using adaptive regression splines. Computational Statistics & Data Analysis, 51(12):5813– 5820, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003823.

Ceranka:2007:COC

[1453] B. Ceranka, M. Graczyk, and K. Katulska. On certain Aoptimal chemical balance weighing designs. Computational Statistics & Data Analysis, 51(12):5821–5827, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003896.

Vittadini:2007:RAL

[1454] Giorgio Vittadini, Simona C. Minotti, Marco Fattore, and Pietro G. Lovaglio. On the relationships among latent variables and residuals in PLS path modeling: the formative-reflective scheme. Computational Statistics & Data Analysis, 51(12):5828–5846, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730600394X.

Sato:2007:WBT

[1455] João R. Sato, Pedro A. Morettin, Paula R. Arantes, and Edson Amaro, Jr. Wavelet based timevarying vector autoregressive modelling. Computational Statistics & Data Analysis, 51(12):5847–5866, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004208.

Agostinelli:2007:REC

[1456] Claudio Agostinelli. Robust estimation for circular data. Computational Statistics & Data Analysis, 51(12): 5867–5875, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730600421X.

Oliveira:2007:MSS

[1457] Manuela M. Oliveira and João T. Mexia. Modelling series of studies with a common structure. Computational Statistics & Data Analysis, 51(12):

5876–5885, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004233.

Huang:2007:PPI

[1458] Yufen Huang, Mei-Ling Kuo, and Tai-Ho Wang. Pair-perturbation influence functions and local influence in PCA. Computational Statistics & Data Analysis, 51(12):5886–5899, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004245.

Bordignon:2007:GLM

[1459] Silvano Bordignon, Massimiliano Caporin, and Francesco Lisi. Generalised long-memory GARCH models for intra-daily volatility. Computational Statistics & Data Analysis, 51(12): 5900–5912, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004257.

Dahl:2007:CRN

[1460] Fredrik A. Dahl. Convergence of random k-nearest-neighbour imputation. Computational Statistics & Data Analysis, 51(12):5913–5917, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004270.

Tang:2007:SIE

[1461] Man-Lai Tang and Wai-Yin Poon. Statistical inference for equivalence trials with ordinal responses: a latent normal distribution approach. Computational Statistics & Data Analysis, 51(12): 5918–5926, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004312.

Diaz-Garcia:2007:NFL

[1462] José A. Díaz-García and José Ramón Domínguez-Molina. A new family of life distributions for dependent data: Estimation. Computational Statistics & Data Analysis, 51(12):5927– 5939, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004336.

Kiang:2007:ESS

[1463] Melody Y. Kiang, Michael Y. Hu, and Dorothy M. Fisher. The effect of sample size on the extended self-organizing map network — a market segmentation application. Computational Statistics & Data Analysis, 51(12):5940– 5948, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004348.

Racine:2007:IKS

[1464] Jeffrey S. Racine and James G. MacKinnon. Inference via kernel smoothing of bootstrap P values. Computational Statistics & Data Analysis, 51(12):

5949–5957, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600435X.

Wen:2007:TSR

[1465] Miin-Jye Wen, Shun-Yi Chen, and Hubert J. Chen. On testing a subset of regression parameters under heteroskedasticity. Computational Statistics & Data Analysis, 51(12):5958– 5976, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600452X.

Xiang:2007:IDR

[1466] Liming Xiang, Kelvin K. W. Yau, S. K. Tse, and Andy H. Lee. Influence diagnostics for random effect survival models: Application to a recurrent infection study for kidney patients on portable dialysis. Computational Statistics & Data Analysis, 51(12): 5977–5993, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004610.

Park:2007:VIB

[1467] Cheolwoo Park, Fred Godtliebsen, Murad Taqqu, Stilian Stoev, and J. S. Marron. Visualization and inference based on wavelet coefficients, SiZer and SiNos. Computational Statistics & Data Analysis, 51(12):5994– 6012, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004701.

Polansky:2007:DCP

[1468] Alan M. Polansky. Detecting changepoints in Markov chains. Computational Statistics & Data Analysis, 51(12):6013–6026, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947306004725.

Lawson:2007:LPC

[1469] Andrew B. Lawson, Silvia Simeon, Martin Kulldorff, Annibale Biggeri, and Corrado Magnani. Line and point cluster models for spatial health data. Computational Statistics & Data Analysis, 51(12):6027– 6043, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004737.

Tutz:2007:BRR

[1470] Gerhard Tutz and Harald Binder. Boosting ridge regression. Computational Statistics & Data Analysis, 51(12):6044–6059, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947306004749.

vandenHout:2007:LRM

[1471] Ardo van den Hout, Peter G. M. van der Heijden, and Robert Gilchrist. The logistic regression model with response variables subject to randomized response. Computational Statistics & Data Analysis, 51(12):6060– 6069, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004774.

Hooper:2007:PAV

[1472] Peter M. Hooper. Period analysis of variable stars: Temporal dependence and local optima. Computational Statistics & Data Analysis, 51(12): 6070–6083, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600483X.

Kumar:2007:CDM

[1473] Mahesh Kumar and Nitin R. Patel. Clustering data with measurement errors. Computational Statistics & Data Analysis, 51(12):6084– 6101, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004920.

Nieto-Barajas:2007:BSP

[1474] Luis E. Nieto-Barajas and Stephen G.

Walker. A Bayesian semi-parametric bivariate failure time model. Computational Statistics & Data Analysis, 51(12):6102–6113, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016794730600497X.

Leon:2007:QSB

[1475] Andrew C. Leon and Donald Hedeker.

Quintile stratification based on a misspecified propensity score in longitudinal treatment effectiveness analyses of ordinal doses. Computational Statistics & Data Analysis, 51(12): 6114–6122, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004981.

Zhang:2007:DMC

[1476] Tonglin Zhang and Ge Lin. A decomposition of Moran’s I for clustering detection. Computational Statistics & Data Analysis, 51(12):6123– 6137, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600507X.

Rodriguez-Avi:2007:NGW

[1477] J. Rodríguez-Avi, A. Conde-Sánchez, A. J. Sáez-Castillo, and M. J. Olmo-Jiménez. A new generalization of the Waring distribution. Computational Statistics & Data Analysis, 51(12): 6138–6150, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005081.

Galeano:2007:UCS

[1478] Pedro Galeano. The use of cumulative sums for detection of changepoints in the rate parameter of a Poisson process. Computational Statistics & Data Analysis, 51(12):6151– 6165, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005184.

Ahn:2007:CER

[1479] Hongshik Ahn, Hojin Moon, Melissa J.

Fazzari, Noha Lim, James J. Chen, and Ralph L. Kodell. Classification by ensembles from random partitions of high-dimensional data. Computational Statistics & Data Analysis, 51(12): 6166–6179, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005196.

Qian:2007:TSM

[1480] Guoqi Qian and Xindong Zhao. On time series model selection involving many candidate ARMA models. Computational Statistics & Data Analysis, 51(12):6180–6196, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005202.

Kuchenhoff:2007:AVE

[1481] Helmut Küchenhoff, Wolfgang Lederer, and Emmanuel Lesaffre. Asymptotic variance estimation for the misclassification SIMEX. Computational Statistics & Data Analysis, 51(12): 6197–6211, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005214.

Heim:2007:SVC

[1482] S. Heim, L. Fahrmeir, P. H. C. Eilers, and B. D. Marx. 3D space-varying coefficient models with application to diffusion tensor imaging. Computational Statistics & Data Analysis, 51(12): 6212–6228, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000035.

Choudhary:2007:SRA

[1483] Pankaj K. Choudhary. Semiparametric regression for assessing agreement using tolerance bands. Computational Statistics & Data Analysis, 51(12): 6229–6241, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000059.

Chang:2007:CML

[1484] George T. Chang and Guenther Walther. Clustering with mixtures of log-concave distributions. Computational Statistics & Data Analysis, 51(12):6242–6251, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947307000060.

Vandewalle:2007:RET

[1485] B. Vandewalle, J. Beirlant, A. Christmann, and M. Hubert. A robust estimator for the tail index of Pareto-type distributions. Computational Statistics & Data Analysis, 51(12):6252– 6268, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000102.

Jamshidian:2007:SPT

[1486] Mortaza Jamshidian, Robert I. Jennrich, and Wei Liu. A study of partial F tests for multiple linear regression models. Computational Statistics & Data Analysis, 51(12):6269– 6284, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000126.

Firat:2007:GCS

[1487] Aykut Firat, Sangit Chatterjee, and Mustafa Yilmaz. Genetic clustering of social networks using random walks. Computational Statistics & Data Analysis, 51(12):6285– 6294, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700014X.

Boulesteix:2007:MSC

[1488] Anne-Laure Boulesteix and Carolin Strobl. Maximally selected chi-squared statistics and non-monotonic associations: an exact approach based on two cutpoints. Computational Statistics & Data Analysis, 51(12):6295– 6306, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000187.

Lambert:2007:ACE

[1489] Philippe Lambert. Archimedean copula estimation using Bayesian splines smoothing techniques. Computational Statistics & Data Analysis, 51(12):

6307–6320, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000217.

Martin:2007:BHT

[1490] Michael A. Martin. Bootstrap hypothesis testing for some common statistical problems: a critical evaluation of size and power properties. Computational Statistics & Data Analysis, 51(12): 6321–6342, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000230.

Dorea:2007:LFA

[1491] Chang C. Y. Dorea, Cira E. Guevara Otiniano, Raul Matsushita, and Pushpa N. Rathie. Lévy flight approximations for scaled transformations of random walks. Computational Statistics & Data Analysis, 51(12): 6343–6354, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000242.

Liang:2007:USB

[1492] Faming Liang. Use of SVD-based probit transformation in clustering gene expression profiles. Computational Statistics & Data Analysis, 51(12): 6355–6366, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700059X.

Klugkist:2007:BFI

[1493] Irene Klugkist and Herbert Hoijtink. The Bayes factor for inequality and about equality constrained models. Computational Statistics & Data Analysis, 51(12):6367–6379, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000606.

Liu:2007:SVM

[1494] Yufeng Liu, Hao Helen Zhang, Cheolwoo Park, and Jeongyoun Ahn. Support vector machines with adaptive L_{q} penalty. Computational Statistics & Data Analysis, 51(12):6380– 6394, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000618.

Benito:2007:DDI

[1495] Mónica Benito and Daniel Peña. Detecting defects with image data. Computational Statistics & Data Analysis, 51(12):6395–6403, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947307000667.

Soubeyrand:2007:RBS

[1496] Samuel Soubeyrand and Joël Chadœuf.

Residual-based specification of a hidden random field included in a hierarchical spatial model. Computational Statistics & Data Analysis, 51(12): 6404–6422, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000680.

Baierl:2007:LMI

[1497] Andreas Baierl, Andreas Futschik, Małgorzata Bogdan, and Przemysław Biecek. Locating multiple interacting quantitative trait loci using robust model selection. Computational Statistics & Data Analysis, 51(12): 6423–6434, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000709.

Phipps:2007:FCI

[1498] Mary C. Phipps and Peter M. Byron. A filter for “Confidence interval P -values”. Computational Statistics & Data Analysis, 51(12):6435– 6446, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000722.

Agresti:2007:NES

[1499] Alan Agresti and Anna Gottard. Nonconservative exact small-sample inference for discrete data. Computational Statistics & Data Analysis, 51(12): 6447–6458, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000953.

Gill:2007:CES

[1500] Ryan Gill, Kiseop Lee, and Seongjoo Song. Computation of estimates in segmented regression and a liquidity effect model. Computational Statistics & Data Analysis, 51(12):6459– 6475, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000965.

Takasaki:2007:URB

[1501] Shigeru Takasaki and Yoshihoro Kawamura. Using radial basis function networks and significance testing to select effective siRNA sequences. Computational Statistics & Data Analysis, 51(12):6476–6487, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000977.

Bao:2007:KSV

[1502] Yanchun Bao, Shuyuan He, and Changlin Mei. The Koul–Susarla– Van Ryzin and weighted least squares estimates for censored linear regression model: a comparative study. Computational Statistics & Data Analysis, 51(12):6488–6497, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000989.

Naik:2007:TEI

[1503] Dayanand N. Naik and Amal Helu.

On testing equality of intraclass correlations under unequal family sizes. Computational Statistics & Data Analysis, 51(12):6498–6510, August 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001004.

Goncalves:2007:LAC

[1504] M. Helena Gonçalves, M. Salomé Cabral, Maria Carme Ruiz de Villa, Eduardo Escrich, and Montse Solanas. Likelihood approach for count data in longitudinal experiments. Computational Statistics & Data Analysis, 51(12):6511–6520, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794730700103X.

da-Silva:2007:BES

[1505] C. Q. da Silva, P. V. S. Gomes, and M. A. Stradioto. Bayesian estimation of survival and capture probabilities using logit link and photoidentification data. Computational Statistics & Data Analysis, 51(12):6521– 6534, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001041.

Schott:2007:TEC

[1506] James R. Schott. A test for the equality of covariance matrices when the dimension is large relative to the sample sizes. Computational Statistics & Data Analysis, 51(12):6535– 6542, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001053.

Zhang:2007:EST

[1507] Chun-Xia Zhang, Chang-Lin Mei, and Jiang-She Zhang. An empirical study of a test for polynomial relationships in randomly right censored regression models. Computational Statistics & Data Analysis, 51(12):6543– 6556, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001065.

Lee:2007:VCW

[1508] Seung-Hwan Lee. On the versatility of the combination of the weighted logrank statistics. Computational Statistics & Data Analysis, 51(12):6557– 6564, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001077.

Roca-Pardinas:2007:TLW

[1509] Javier Roca-Pardiñas and Stefan Sperlich. Testing the link when the index is semiparametric — a comparative study. Computational Statistics & Data Analysis, 51(12):6565– 6581, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001090.

Yang:2007:DAM

[1510] Yang Yang, Ira M. Longini, Jr., and M. Elizabeth Halloran. A dataaugmentation method for infectious disease incidence data from close contact groups. Computational Statistics & Data Analysis, 51(12):6582– 6595, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001119.

Rueda:2007:EQU

[1511] M. Rueda, J. F. Muñoz, S. González, and A. Arcos. Estimating quantiles under sampling on two occasions with arbitrary sample designs. Computational Statistics & Data Analysis, 51(12): 6596–6613, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001132.

Wang:2007:NRE

[1512] Xiaoning Wang, Alan Schumitzky, and David Z. D’Argenio. Nonlinear random effects mixture models: Maximum likelihood estimation via the EM algorithm. Computational Statistics & Data Analysis, 51(12):6614– 6623, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001144.

Gonzalez-Davila:2007:SBD

[1513] Enrique González-Dávila, Roberto Dorta-Guerra, and Josep Ginebra. Simulation-based designs for multiperiod control. Computational Statistics & Data Analysis, 51(12):6624– 6641, August 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001648.

Anonymous:2007:EBi

[1514] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52 (1):v–vii, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003131.

Anonymous:2007:Ci

[1515] Anonymous. Contents. Computational Statistics & Data Analysis, 52(1):viii– xi, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003143.

Gilli:2007:SIA

[1516] Manfred Gilli and Peter Winker. 2nd special issue on applications of optimization heuristics to estimation and modelling problems. Computational Statistics & Data Analysis, 52(1):2– 3, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001946.

Kapetanios:2007:VSR

[1517] George Kapetanios. Variable selection in regression models using nonstandard optimisation of information criteria. Computational Statistics & Data Analysis, 52(1):4–15, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001375.

Hofmann:2007:EAC

[1518] Marc Hofmann, Cristian Gatu, and Erricos John Kontoghiorghes. Efficient algorithms for computing the best subset regression models for largescale problems. Computational Statistics & Data Analysis, 52(1):16–29, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001168.

Ambrogi:2007:SAN

[1519] Federico Ambrogi, Nicola Lama, Patrizia Boracchi, and Elia Biganzoli. Selection of artificial neural network models for survival analysis with genetic algorithms. Computational Statistics & Data Analysis, 52(1):30–42, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700196X.

Duczmal:2007:GAI

[1520] Luiz Duczmal, André L. F. Cançado, Ricardo H. C. Takahashi, and Lupércio F. Bessegato. A genetic algorithm for irregularly shaped spatial scan statistics. Computational Statistics & Data Analysis, 52(1):43– 52, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000199.

Keskinturk:2007:GAA

[1521] Timur Keskintürk and Şebnem Er. A genetic algorithm approach to determine stratum boundaries and sample sizes of each stratum in stratified sampling. Computational Statistics & Data Analysis, 52(1):53–67, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001363.

Krink:2007:UDE

[1522] Thiemo Krink, Sandra Paterlini, and Andrea Resti. Using differential evolution to improve the accuracy of bank rating systems. Computational Statistics & Data Analysis, 52(1):68– 87, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000539.

Fitzenberger:2007:ICC

[1523] Bernd Fitzenberger and Peter Winker.

Improving the computation of censored quantile regressions. Computational Statistics & Data Analysis, 52(1):88– 108, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000175.

Yang:2007:GBM

[1524] Zheng Yang, Zheng Tian, and Zixia Yuan. GSA-based maximum likelihood estimation for threshold vector error correction model. Computational Statistics & Data Analysis, 52(1):109– 120, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002393.

Staszewska:2007:RUA

[1525] Anna Staszewska. Representing uncertainty about response paths: the use of heuristic optimisation methods. Computational Statistics & Data Analysis, 52(1):121–132, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306005007.

Ali:2007:SDA

[1526] M. M. Ali. Synthesis of the βdistribution as an aid to stochastic global optimization. Computational Statistics & Data Analysis, 52(1):133– 149, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004907.

Croux:2007:MLR

[1527] Christophe Croux, Efstratios Gallopoulos, Stefan Van Aelst, and Hongyuan Zha. Machine learning and robust data mining. Computational Statistics & Data Analysis, 52(1):151– 154, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002538.

Berry:2007:AAA

[1528] Michael W. Berry, Murray Browne, Amy N. Langville, V. Paul Pauca, and Robert J. Plemmons. Algorithms and applications for approximate nonnegative matrix factorization. Computational Statistics & Data Analysis, 52(1):155–173, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004191.

Tougas:2007:UPS

[1529] Jane E. Tougas and Raymond J. Spiteri. Updating the partial singular value decomposition in latent semantic indexing. Computational Statistics & Data Analysis, 52(1):174–183, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004944.

Zha:2007:CIM

[1530] Hongyuan Zha and Zhenyue Zhang.

Continuum isomap for manifold learnings. Computational Statistics & Data Analysis, 52(1):184–200, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004518.

Ao:2007:CUS

[1531] S. I. Ao, Michael K. Ng, and Pak C. Sham. Constrained unidimensional scaling with application to genomics. Computational Statistics & Data Analysis, 52(1):201–210, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004282.

Chen:2007:WBP

[1532] Shuo Chen, Don Hong, and Yu Shyr. Wavelet-based procedures for proteomic mass spectrometry data processing. Computational Statistics & Data Analysis, 52(1):211–220, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000576.

Fried:2007:RTS

[1533] Roland Fried and Ursula Gather. On rank tests for shift detection in time series. Computational Statistics & Data Analysis, 52(1):221–233, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004932.

Bednarski:2007:RMB

[1534] Tadeusz Bednarski. On a robust modification of Breslow’s cumulated hazard estimator. Computational Statistics & Data Analysis, 52(1):234–238, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004798.

Khan:2007:BRL

[1535] Jafar A. Khan, Stefan Van Aelst, and Ruben H. Zamar. Building a robust linear model with forward selection and stepwise procedures. Computational Statistics & Data Analysis, 52(1):239– 248, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000096.

McCann:2007:RVS

[1536] Lauren McCann and Roy E. Welsch.

Robust variable selection using least angle regression and elemental set sampling. Computational Statistics & Data Analysis, 52(1):249–257, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000163.

Hennig:2007:CWA

[1537] Christian Hennig. Cluster-wise assessment of cluster stability. Computational Statistics & Data Analysis, 52(1):258–271, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306004622.

Atkinson:2007:ETC

[1538] A. C. Atkinson and M. Riani. Exploratory tools for clustering multivariate data. Computational Statistics & Data Analysis, 52(1):272–285, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004865.

Wang:2007:CNP

[1539] Xiaogang Wang, Weiliang Qiu, and Ruben H. Zamar. CLUES: a nonparametric clustering method based on local shrinking. Computational Statistics & Data Analysis, 52(1):286– 298, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004786.

Neykov:2007:RFM

[1540] N. Neykov, P. Filzmoser, R. Dimova, and P. Neytchev. Robust fitting of mixtures using the trimmed likelihood estimator. Computational Statistics & Data Analysis, 52(1):299–308, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005019.

Dolia:2007:KET

[1541] A. N. Dolia, C. J. Harris, J. S. Shawe-Taylor, and D. M. Titterington. Kernel ellipsoidal trimming. Computational Statistics & Data Analysis, 52(1):309– 324, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001156.

Tsao:2007:SAV

[1542] C. Andy Tsao and Yuan chin Ivan Chang. A stochastic approximation view of boosting. Computational Statistics & Data Analysis, 52(1):325– 334, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002599.

Huang:2007:MSS

[1543] Chien-Ming Huang, Yuh-Jye Lee, Dennis K. J. Lin, and Su-Yun Huang. Model selection for support vector machines via uniform design. Computational Statistics & Data Analysis, 52(1):335–346, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000552.

Christmann:2007:RLB

[1544] Andreas Christmann, Ingo Steinwart, and Mia Hubert. Robust learning from bites for data mining. Computational Statistics & Data Analysis, 52(1):347– 361, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004853.

Croux:2007:TB

[1545] Christophe Croux, Kristel Joossens, and Aurélie Lemmens. Trimmed bagging. Computational Statistics & Data Analysis, 52(1):362–368, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002484.

Bock:2007:SLM

[1546] Hans-Hermann Bock and Maurizio Vichi. Statistical learning methods including dimensionality reduction. Computational Statistics & Data Analysis, 52(1):370–373, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001673.

Meinshausen:2007:RL

[1547] Nicolai Meinshausen. Relaxed lasso. Computational Statistics & Data Analysis, 52(1):374–393, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306004956.

Takane:2007:RLK

[1548] Yoshio Takane and Heungsun Hwang. Regularized linear and kernel redundancy analysis. Computational Statistics & Data Analysis, 52(1):394–405, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000710.

Simila:2007:ISS

[1549] Timo Similä and Jarkko Tikka. Input selection and shrinkage in multiresponse linear regression. Computational Statistics & Data Analysis, 52(1):406–422, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947307000205.

Tebbens:2007:IIL

[1550] Jurjen Duintjer Tebbens and Pavel Schlesinger. Improving implementation of linear discriminant analysis for the high dimension/small sample size problem. Computational Statistics & Data Analysis, 52(1):423–437, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000540.

Scrucca:2007:CPG

[1551] Luca Scrucca. Class prediction and gene selection for DNA microarrays using regularized sliced inverse regression. Computational Statistics & Data Analysis, 52(1):438–451, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000643.

Zhang:2007:HDP

[1552] Chunming Zhang, Haoda Fu, Yuan Jiang, and Tao Yu. High-dimensional pseudo-logistic regression and classification with applications to gene expression data. Computational Statistics & Data Analysis, 52(1):452–470, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005044.

Calo:2007:GMM

[1553] Daniela G. Calò. Gaussian mixture model classification: a projection pursuit approach. Computational Statistics & Data Analysis, 52(1):471–482, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730600510X.

Strobl:2007:USS

[1554] Carolin Strobl, Anne-Laure Boulesteix, and Thomas Augustin. Unbiased split selection for classification trees based on the Gini index. Computational Statistics & Data Analysis, 52(1):483– 501, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306005093.

Bouveyron:2007:HDD

[1555] C. Bouveyron, S. Girard, and C. Schmid. High-dimensional data clustering. Computational Statistics & Data Analysis, 52(1):502–519, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000692.

Galimberti:2007:MBM

[1556] Giuliano Galimberti and Gabriele Soffritti. Model-based methods to identify multiple cluster structures in a data set. Computational Statistics & Data Analysis, 52(1):520–536, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000758.

VanMechelen:2007:UMI

[1557] Iven Van Mechelen and Jan Schepers. A unifying model involving a categorical and/or dimensional reduction for multimode data. Computational Statistics & Data Analysis, 52(1):537– 549, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001028.

Cumming:2007:DRP

[1558] J. A. Cumming and D. A. Wooff.

Dimension reduction via principal variables. Computational Statistics & Data Analysis, 52(1):550–565, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000564.

Lombardo:2007:NSC

[1559] R. Lombardo, E. J. Beh, and L. D’Ambra. Non-symmetric correspondence analysis with ordinal variables using orthogonal polynomials. Computational Statistics & Data Analysis, 52(1):566–577, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306005160.

Quevedo:2007:SEM

[1560] José R. Quevedo, Antonio Bahamonde, and Oscar Luaces. A simple and efficient method for variable ranking according to their usefulness for learning. Computational Statistics & Data Analysis, 52(1):578– 595, September 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000655.

Plasse:2007:CUA

[1561] Marie Plasse, Ndeye Niang, Gilbert Saporta, Alexandre Villeminot, and Laurent Leblond. Combined use of association rules mining and clustering methods to find relevant links between binary rare attributes in a large data set. Computational Statistics & Data Analysis, 52(1):596–613, September 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307000904.

Anonymous:2007:EBj

[1562] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52 (2):iii–v, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003726.

Anonymous:2007:Cj

[1563] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (2):vi–ix, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003738.

Gatu:2007:SIS

[1564] Cristian Gatu, James Gentle, John Hinde, and Moon Huh. Special issue on statistical algorithms and software. Computational Statistics & Data Analysis, 52(2):616–617, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002290.

Kratzig:2007:SFD

[1565] Markus Krätzig. A software framework for data analysis. Computational Statistics & Data Analysis, 52(2):618– 634, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306002672.

Aluja-Banet:2007:GCS

[1566] Tomàs Aluja-Banet, Josep Daunis i Estadella, and David Pellicer. GRAFT, a complete system for data fusion. Computational Statistics & Data Analysis, 52(2):635–649, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004658.

Fujiwara:2007:UME

[1567] Takeshi Fujiwara, Junji Nakano, and Yoshikazu Yamamoto. Using mathematical expressions in a statistical language. Computational Statistics & Data Analysis, 52(2):650–662, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003902.

Massmann:2007:CPC

[1568] Michael Massmann. Cobra: a package for co-breaking analysis. Computational Statistics & Data Analysis, 52(2):663–679, October 15, 2007.

CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947306003987.

Hohle:2007:RRP

[1569] Michael Höhle and Ulrike Feldmann. RLadyBug — an R package for stochastic epidemic models. Computational Statistics & Data Analysis, 52(2):680– 686, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004221.

Chavent:2007:DMD

[1570] Marie Chavent, Yves Lechevallier, and Olivier Briant. DIVCLUS-T: a monothetic divisive hierarchical clustering method. Computational Statistics & Data Analysis, 52(2):687–701, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001181.

Fernando:2007:CSA

[1571] W. T. P. S. Fernando and D. D. S. Kulatunga. On the computation and some applications of multivariate isotonic regression. Computational Statistics & Data Analysis, 52(2):702–712, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000138.

Tvrdik:2007:APB

[1572] Josef Tvrdík, Ivan Křivý, and Ladislav Mišík. Adaptive populationbased search: Application to estimation of nonlinear regression parameters. Computational Statistics & Data Analysis, 52(2):713–724, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003835.

Gramm:2007:ACL

[1573] Jens Gramm, Jiong Guo, Falk Hüffner, Rolf Niedermeier, Hans-Peter Piepho, and Ramona Schmid. Algorithms for compact letter displays: Comparison and evaluation. Computational Statistics & Data Analysis, 52(2):725– 736, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306003598.

Beninel:2007:EAC

[1574] Farid Beninel and Gérard Grélaud.

Exact algorithms for computing pvalues of statistics-linear combination of 3-nomial variables. Computational Statistics & Data Analysis, 52(2):737– 749, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700134X.

Park:2007:SSD

[1575] Byung-Hoon Park, George Ostrouchov, and Nagiza F. Samatova. Sampling streaming data with replacement. Computational Statistics & Data Analysis, 52(2):750–762, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307001089.

Bernholt:2007:CLQ

[1576] Thorsten Bernholt, Robin Nunkesser, and Karen Schettlinger. Computing the least quartile difference estimator in the plane. Computational Statistics & Data Analysis, 52(2):763– 772, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306005159.

Iacus:2007:MDI

[1577] Stefano M. Iacus and Giuseppe Porro. Missing data imputation, matching and other applications of random recursive partitioning. Computational Statistics & Data Analysis, 52(2):773– 789, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306005135.

Consonni:2007:MFV

[1578] Guido Consonni and Jean-Michel Marin. Mean-field variational approximate Bayesian inference for latent variable models. Computational Statistics & Data Analysis, 52(2):790– 798, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003951.

Gatu:2007:GAG

[1579] Cristian Gatu, Petko I. Yanev, and Erricos J. Kontoghiorghes. A graph approach to generate all possible regression submodels. Computational Statistics & Data Analysis, 52(2):799– 815, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000746.

Pollock:2007:SIS

[1580] D. S. G. Pollock and Tommaso Proietti.

2nd special issue on statistical signal extraction and filtering. Computational Statistics & Data Analysis, 52(2):817– 820, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003015.

Bakker:2007:LAI

[1581] Bart Bakker and Tom Heskes. Learning and approximate inference in dynamic hierarchical models. Computational Statistics & Data Analysis, 52(2):821– 839, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000047.

Shin:2007:NFF

[1582] Vladimir Shin, Georgy Shevlyakov, and Kiseon Kim. A new fusion formula and its application to continuoustime linear systems with multisensor environment. Computational Statistics & Data Analysis, 52(2):840–854, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004671.

Klauenberg:2007:HMR

[1583] Katy Klauenberg and Francesco Lagona. Hidden Markov random field models for TCA image analysis. Computational Statistics & Data Analysis, 52(2):855–868, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002575.

Howlett:2007:OME

[1584] P. G. Howlett, A. Torokhti, and C. E. M. Pearce. Optimal multilinear estimation of a random vector under constraints of causality and limited memory. Computational Statistics & Data Analysis, 52(2):869–878, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947306003884.

Maharaj:2007:DLS

[1585] Elizabeth A. Maharaj and Andrés M. Alonso. Discrimination of locally stationary time series using wavelets. Computational Statistics & Data Analysis, 52(2):879–895, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002083.

Cubadda:2007:UFA

[1586] Gianluca Cubadda. A unifying framework for analysing common cyclical features in cointegrated time series. Computational Statistics & Data Analysis, 52(2):896–906, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002654.

Mastromarco:2007:RBC

[1587] Camilla Mastromarco and Ulrich Woitek. Regional business cycles in Italy. Computational Statistics & Data Analysis, 52(2):907–918, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700165X.

Morana:2007:MML

[1588] Claudio Morana. Multivariate modelling of long memory processes with common components. Computational Statistics & Data Analysis, 52(2):919– 934, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004877.

Proietti:2007:SEF

[1589] Tommaso Proietti. Signal extraction and filtering by linear semiparametric methods. Computational Statistics & Data Analysis, 52(2):935–958, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002964.

Weinert:2007:ECW

[1590] Howard L. Weinert. Efficient computation for Whittaker–Henderson smoothing. Computational Statistics & Data Analysis, 52(2):959–974, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947306004713.

Maravall:2007:TAS

[1591] A. Maravall and A. del Río. Temporal aggregation, systematic sampling, and the Hodrick–Prescott filter. Computational Statistics & Data Analysis, 52(2):975–998, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002976.

Bujosa:2007:LDH

[1592] Marcos Bujosa, Antonio García-Ferrer, and Peter C. Young. Linear dynamic harmonic regression. Computational Statistics & Data Analysis, 52(2):999– 1024, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002666.

Triantafyllopoulos:2007:BAM

[1593] K. Triantafyllopoulos and G. P. Nason. A Bayesian analysis of moving average processes with time-varying parameters. Computational Statistics & Data Analysis, 52(2):1025– 1046, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001569.

Ledesma:2007:TAM

[1594] Sergio Ledesma, Derong Liu, and Donato Hernández. Two approximation methods to synthesize the power spectrum of fractional Gaussian noise. Computational Statistics & Data Analysis, 52(2):1047– 1062, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307000734.

Fried:2007:RDE

[1595] Roland Fried. On the robust detection of edges in time series filtering. Computational Statistics & Data Analysis, 52(2):1063–1074, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002472.

VanHuffel:2007:TLS

[1596] Sabine Van Huffel, Chi-Lun Cheng, Nicola Mastronardi, Chris Paige, and Alexander Kukush. Total least squares and errors-in-variables modeling. Computational Statistics & Data Analysis, 52(2):1076–1079, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002617.

Watson:2007:RCE

[1597] G. A. Watson. Robust counterparts of errors-in-variables problems. Computational Statistics & Data Analysis, 52(2):1080–1089, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730700206X.

Lampe:2007:QEO

[1598] J. Lampe and H. Voss. On a quadratic eigenproblem occurring in regularized total least squares. Computational Statistics & Data Analysis, 52(2):1090– 1102, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002149.

Sima:2007:LCT

[1599] Diana M. Sima and Sabine Van Huffel. Level choice in truncated total least squares. Computational Statistics & Data Analysis, 52(2):1103– 1118, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002071.

Mastronardi:2007:FRR

[1600] Nicola Mastronardi and Dianne P. O’Leary. Fast robust regression algorithms for problems with Toeplitz structure. Computational Statistics & Data Analysis, 52(2):1119–1131, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002101.

deCastro:2007:LIA

[1601] Mário de Castro, Manuel Galea-Rojas, and Heleno Bolfarine. Local influence assessment in heteroscedastic measurement error models. Computational Statistics & Data Analysis, 52(2):1132– 1142, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002137.

Schneeweiss:2007:ELR

[1602] H. Schneeweiß and H. Shalabh. On the estimation of the linear relation when the error variances are known. Computational Statistics & Data Analysis, 52(2):1143–1148, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002563.

Shalabh:2007:RRE

[1603] Shalabh, Gaurav Garg, and Neeraj Misra. Restricted regression estimation in measurement error models. Computational Statistics & Data Analysis, 52(2):1149–1166, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002125.

Kukush:2007:ELM

[1604] A. Kukush, I. Markovsky, and S. Van Huffel. Estimation in a linear multivariate measurement error model with a change point in the data. Computational Statistics & Data Analysis, 52(2):1167–1182, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002460.

Vehkalahti:2007:EME

[1605] Kimmo Vehkalahti, Simo Puntanen, and Lauri Tarkkonen. Effects of measurement errors in predictor selection of linear regression model. Computational Statistics & Data Analysis, 52(2):1183– 1195, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002058.

Barros:2007:HTU

[1606] Michelli Barros, Viviana Giampaoli, and Claudia R. O. P. Lima. Hypothesis testing in the unrestricted and restricted parametric spaces of structural models. Computational Statistics & Data Analysis, 52(2):1196– 1207, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002216.

Kanatani:2007:PEI

[1607] Kenichi Kanatani and Yasuyuki Sugaya. Performance evaluation of iterative geometric fitting algorithms. Computational Statistics & Data Analysis, 52(2):1208–1222, October 15, 2007. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002150.

Petras:2007:SSD

[1608] Ivo Petras and Igor Podlubny. State space description of national economies: the v4 countries. Computational Statistics & Data Analysis, 52(2):1223–1233, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947307002095.

Ramos:2007:ATR

[1609] José A. Ramos. Applications of TLS and related methods in the environmental sciences. Computational Statistics & Data Analysis, 52(2):1234– 1267, October 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700237X.

Anonymous:2008:EBa

[1610] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52(3):iii–v, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004082.

Anonymous:2008:Ca

[1611] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (3):vi–viii, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004094.

Cornillon:2008:FTS

[1612] P.-A. Cornillon, W. Imam, and E. Matzner-Løber. Forecasting time series using principal component analysis with respect to instrumental variables. Computational Statistics & Data Analysis, 52(3):1269–1280, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002411.

Gonzalez:2008:NPB

[1613] M. González, J. Martín, R. Martínez, and M. Mota. Non-parametric Bayesian estimation for multitype branching processes through simulationbased methods. Computational Statistics & Data Analysis, 52(3):1281– 1291, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002459.

Moller:2008:TAQ

[1614] Jan Kloppenborg Møller, Henrik Aalborg Nielsen, and Henrik Madsen. Time-adaptive quantile regression. Computational Statistics & Data Analysis, 52(3):1292–1303, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002502.

Shao:2008:ETP

[1615] Quanxi Shao, Yongqin D. Chen, and Lu Zhang. An extension of threeparameter Burr III distribution for lowflow frequency analysis. Computational Statistics & Data Analysis, 52(3):1304– 1314, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002526.

Wicker:2008:MLA

[1616] Nicolas Wicker, Jean Muller, Ravi Kiran Reddy Kalathur, and Olivier Poch. A maximum likelihood approximation method for Dirichlet’s parameter estimation. Computational Statistics & Data Analysis, 52(3):1315– 1322, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002848.

Somerville:2008:SSP

[1617] Paul N. Somerville and Claudia Hemmelmann. Step-up and stepdown procedures controlling the number and proportion of false positives. Computational Statistics & Data Analysis, 52(3):1323–1334, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002897.

Stamey:2008:CIP

[1618] James D. Stamey, Doyle H. Boese, and Dean M. Young. Confidence intervals for parameters of two diagnostic tests in the absence of a gold standard. Computational Statistics & Data Analysis, 52(3):1335– 1346, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003362.

Ho:2008:FMR

[1619] Remus K. W. Ho and Inchi Hu. Flexible modelling of random effects in linear mixed models — a Bayesian approach. Computational Statistics & Data Analysis, 52(3):1347–1361, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700343X.

Gray:2008:CTA

[1620] J. Brian Gray and Guangzhe Fan. Classification tree analysis using TARGET. Computational Statistics & Data Analysis, 52(3):1362– 1372, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700117X.

Marchette:2008:PUL

[1621] David J. Marchette and Carey E. Priebe. Predicting unobserved links in incompletely observed networks. Computational Statistics & Data Analysis, 52(3):1373–1386, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001193.

Tsai:2008:TSC

[1622] Guei-Feng Tsai and Annie Qu. Testing the significance of cell-cycle patterns in time-course microarray data using nonparametric quadratic inference functions. Computational Statistics & Data Analysis, 52(3):1387– 1398, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001338.

Chen:2008:ERA

[1623] Chu-Chih Chen. Extended rank analysis of covariance adjusting for local correlations. Computational Statistics & Data Analysis, 52(3):1399– 1412, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001351.

Hoshino:2008:BPS

[1624] Takahiro Hoshino. A Bayesian propensity score adjustment for latent variable modeling and MCMC algorithm. Computational Statistics & Data Analysis, 52(3):1413– 1429, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001429.

Sultan:2008:EPG

[1625] K. S. Sultan, G. R. AL-Dayian, and H. H. Mohammad. Estimation and prediction from gamma distribution based on record values. Computational Statistics & Data Analysis, 52(3):1430– 1440, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001557.

DelaCruz-Mesia:2008:MBC

[1626] Rolando De la Cruz-Mesía, Fernando A. Quintana, and Guillermo Marshall. Model-based clustering for longitudinal data. Computational Statistics & Data Analysis, 52(3):1441– 1457, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001570.

Wong:2008:VCS

[1627] Heung Wong, Wai cheung Ip, and Riquan Zhang. Varying-coefficient single-index model. Computational Statistics & Data Analysis, 52(3):1458– 1476, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001594.

Pardo:2008:ELR

[1628] L. Pardo and M. C. Pardo. An extension of likelihood-ratio-test for testing linear hypotheses in the baselinecategory logit model. Computational Statistics & Data Analysis, 52(3):1477– 1489, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001600.

Ong:2008:CDD

[1629] S. H. Ong, K. Shimizu, and Choung Min Ng. A class of discrete distributions arising from difference of two random variables. Computational Statistics & Data Analysis, 52(3):1490– 1499, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001636.

Neto:2008:CRM

[1630] Eufrásio de A. Lima Neto and Francisco de A. T. de Carvalho. Centre and range method for fitting a linear regression model to symbolic interval data. Computational Statistics & Data Analysis, 52(3):1500– 1515, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001934.

Pewsey:2008:WSF

[1631] Arthur Pewsey. The wrapped stable family of distributions as a flexible model for circular data. Computational Statistics & Data Analysis, 52(3):1516– 1523, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001958.

Yu:2008:MCM

[1632] Binbing Yu and Yingwei Peng. Mixture cure models for multivariate survival data. Computational Statistics & Data Analysis, 52(3):1524– 1532, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307001983.

Beunckens:2008:SSC

[1633] Caroline Beunckens, Cristina Sotto, and Geert Molenberghs. A simulation study comparing weighted estimating equations with multiple imputation based estimating equations for longitudinal binary data. Computational Statistics & Data Analysis, 52(3):1533– 1548, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002009.

Giannikis:2008:MNH

[1634] D. Giannikis, I. D. Vrontos, and P. Dellaportas. Modelling nonlinearities and heavy tails via threshold normal mixture GARCH models. Computational Statistics & Data Analysis, 52(3):1549– 1571, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002010.

Bate:2008:CCD

[1635] S. T. Bate, E. J. Godolphin, and J. D. Godolphin. Choosing crossover designs when few subjects are available. Computational Statistics & Data Analysis, 52(3):1572–1586, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002034.

Yu:2008:SIP

[1636] Jun-Wu Yu, Guo-Liang Tian, and Man-Lai Tang. Statistical inference and prediction for the Weibull process with incomplete observations. Computational Statistics & Data Analysis, 52(3):1587–1603, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002022.

Debruyne:2008:CDQ

[1637] M. Debruyne, M. Hubert, S. Portnoy, and K. Vanden Branden. Censored depth quantiles. Computational Statistics & Data Analysis, 52(3):1604– 1614, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002046.

Ausin:2008:BPT

[1638] M. Concepción Ausín, Michael P.

Wiper, and Rosa E. Lillo. Bayesian prediction of the transient behaviour and busy period in short- and long-tailed GI/G/1 queueing systems. Computational Statistics & Data Analysis, 52(3):1615–1635, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002113.

Yu:2008:SEB

[1639] Chang Yu and Daniel Zelterman.

Sums of exchangeable Bernoulli random variables for family and litter frequency data. Computational Statistics & Data Analysis, 52(3):1636– 1649, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002162.

Ricci:2008:ATM

[1640] Lila Ricci and Raúl Martínez. Adjusted R^{2}-type measures for Tweedie models. Computational Statistics & Data Analysis, 52(3):1650–1660, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002174.

Delicado:2008:SSC

[1641] P. Delicado and M. N. Goria. A small sample comparison of maximum likelihood, moments and L-moments methods for the asymmetric exponential power distribution. Computational Statistics & Data Analysis, 52(3):1661– 1673, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002186.

Golightly:2008:BIN

[1642] A. Golightly and D. J. Wilkinson. Bayesian inference for nonlinear multivariate diffusion models observed with error. Computational Statistics & Data Analysis, 52(3):1674– 1693, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002198.

Filzmoser:2008:OIH

[1643] Peter Filzmoser, Ricardo Maronna, and Mark Werner. Outlier identification in high dimensions. Computational Statistics & Data Analysis, 52(3):1694– 1711, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002204.

Serneels:2008:PCA

[1644] Sven Serneels and Tim Verdonck. Principal component analysis for data containing outliers and missing elements. Computational Statistics & Data Analysis, 52(3):1712–1727, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002241.

Tiku:2008:EBN

[1645] Moti L. Tiku, M. Qamarul Islam, and Hakan S. Sazak. Estimation in bivariate nonnormal distributions with stochastic variance functions. Computational Statistics & Data Analysis, 52(3):1728–1745, January 1, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002289.

Mauricio:2008:CUR

[1646] José Alberto Mauricio. Computing and using residuals in time series models. Computational Statistics & Data Analysis, 52(3):1746– 1763, January 1, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002344.

Anonymous:2008:EBb

[1647] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52 (4):iii–v, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004203.

Anonymous:2008:Cb

[1648] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (4):vi–viii, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004215.

Jasra:2008:ISM

[1649] Ajay Jasra, Arnaud Doucet, David A. Stephens, and Christopher C. Holmes. Interacting sequential Monte Carlo samplers for trans-dimensional simulation. Computational Statistics & Data Analysis, 52(4):1765–1791, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003398.

Alonso:2008:TSB

[1650] Andrés M. Alonso and Ana E. Sipols. A time series bootstrap procedure for interpolation intervals. Computational Statistics & Data Analysis, 52(4):1792– 1805, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002307.

Tomita:2008:ILB

[1651] Makoto Tomita, Masahiro Hatsumichi, and Koji Kurihara. Identify LD blocks based on hierarchical spatial data. Computational Statistics & Data Analysis, 52(4):1806–1820, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002319.

Sarhan:2008:BER

[1652] Ammar M. Sarhan and Debasis Kundu. Bayes estimators for reliability measures in geometric distribution model using masked system life test data. Computational Statistics & Data Analysis, 52(4):1821– 1836, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002320.

Timmerman:2008:BCI

[1653] Marieke E. Timmerman and Cajo J. F. Ter Braak. Bootstrap confidence intervals for principal response curves. Computational Statistics & Data Analysis, 52(4):1837– 1849, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002332.

Dannemann:2008:LRT

[1654] Jörn Dannemann and Hajo Holzmann. The likelihood ratio test for hidden Markov models in two-sample problems. Computational Statistics & Data Analysis, 52(4):1850–1859, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002356.

Corduas:2008:TSC

[1655] Marcella Corduas and Domenico Piccolo. Time series clustering and classification by the autoregressive metric. Computational Statistics & Data Analysis, 52(4):1860–1872, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002368.

Kundu:2008:GED

[1656] Debasis Kundu and Rameshwar D. Gupta. Generalized exponential distribution: Bayesian estimations. Computational Statistics & Data Analysis, 52(4):1873–1883, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947307002381.

Meintanis:2008:FMT

[1657] Simos G. Meintanis and George Iliopoulos. Fourier methods for testing multivariate independence. Computational Statistics & Data Analysis, 52(4):1884–1895, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794730700240X.

Orelien:2008:FEV

[1658] Jean G. Orelien and Lloyd J. Edwards. Fixed-effect variable selection in linear mixed models using R^{2} statistics. Computational Statistics & Data Analysis, 52(4):1896– 1907, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002423.

Sheather:2008:SMV

[1659] Simon J. Sheather, Joseph W. McKean, and Kimberly Crimin. Sliced mean variance–covariance inverse regression. Computational Statistics & Data Analysis, 52(4):1908–1927, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002435.

Zhang:2008:LBA

[1660] Chun-Xia Zhang and Jiang-She Zhang.

A local boosting algorithm for solving classification problems. Computational Statistics & Data Analysis, 52(4):1928– 1941, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002447.

Musonda:2008:SCC

[1661] Patrick Musonda, Mounia N. Hocine, Heather J. Whitaker, and C. Paddy Farrington. Self-controlled case series analyses: Small-sample performance. Computational Statistics & Data Analysis, 52(4):1942–1957, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002496.

Roca-Pardinas:2008:ACB

[1662] Javier Roca-Pardiñas, Carmen Cadarso-Suárez, Pablo G. Tahoces, and María J. Lado. Assessing continuous bivariate effects among different groups through nonparametric regression models: an application to breast cancer detection. Computational Statistics & Data Analysis, 52(4):1958– 1970, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002514.

Karvanen:2008:CGL

[1663] Juha Karvanen and Arto Nuutinen. Characterizing the generalized lambda distribution by Lmoments. Computational Statistics & Data Analysis, 52(4):1971–1983, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700254X.

Rocci:2008:TMM

[1664] Roberto Rocci and Maurizio Vichi. Two-mode multi-partitioning. Computational Statistics & Data Analysis, 52(4):1984–2003, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947307002551.

Shang:2008:BVA

[1665] Junfeng Shang and Joseph E. Cavanaugh. Bootstrap variants of the Akaike information criterion for mixed model selection. Computational Statistics & Data Analysis, 52(4):2004– 2021, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002587.

Gomes:2008:STJ

[1666] M. Ivette Gomes, Andreia Hall, and M. Cristina Miranda. Subsampling techniques and the jackknife methodology in the estimation of the extremal index. Computational Statistics & Data Analysis, 52(4):2022– 2041, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002605.

Cheng:2008:MTL

[1667] Tsung-Chi Cheng and Atanu Biswas.

Maximum trimmed likelihood estimator for multivariate mixed continuous and categorical data. Computational Statistics & Data Analysis, 52(4):2042– 2065, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002629.

Al-Saleh:2008:MRS

[1668] Mohammad Fraiwan Al-Saleh and Monjed Hisham Samuh. On multistage ranked set sampling for distribution and median estimation. Computational Statistics & Data Analysis, 52(4):2066– 2078, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002630.

Leiva:2008:LAB

[1669] Víctor Leiva, Marco Riquelme, N. Balakrishnan, and Antonio Sanhueza.

Lifetime analysis based on the generalized Birnbaum–Saunders distribution. Computational Statistics & Data Analysis, 52(4):2079–2097, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002642.

Mulekar:2008:HUA

[1670] Madhuri S. Mulekar, John C. Knutson, and Jyoti A. Champanerkar. How useful are approximations to mean and variance of the index of dissimilarity? Computational Statistics & Data Analysis, 52(4):2098– 2109, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002678.

Krnjajic:2008:PNB

[1671] Milovan Krnjajić, Athanasios Kottas, and David Draper. Parametric and nonparametric Bayesian model specification: a case study involving models for count data. Computational Statistics & Data Analysis, 52(4):2110– 2128, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700268X.

Castillo:2008:SPN

[1672] Enrique Castillo, Ali S. Hadi, Beatriz Lacruz, and Rosa E. Pruneda. Semi-parametric nonlinear regression and transformation using functional networks. Computational Statistics & Data Analysis, 52(4):2129–2157, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002691.

Mandel:2008:SCI

[1673] Micha Mandel and Rebecca A. Betensky. Simultaneous confidence intervals based on the percentile bootstrap approach. Computational Statistics & Data Analysis, 52(4):2158– 2165, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002708.

McLeod:2008:FAM

[1674] A. I. McLeod and Y. Zhang. Faster ARMA maximum likelihood estimation. Computational Statistics & Data Analysis, 52(4):2166–2176, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700285X.

Said:2008:RSN

[1675] Yasmin H. Said, Edward J. Wegman, Walid K. Sharabati, and John T. Rigsby. RETRACTED: Social networks of author–coauthor relationships. Computational Statistics & Data Analysis, 52(4):2177–2184, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002861. See

retraction [2348].

Coin:2008:GFT

[1676] Daniele Coin. A goodness-of-fit test for normality based on polynomial regression. Computational Statistics & Data Analysis, 52(4):2185– 2198, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002873.

SanMartino:2008:PBT

[1677] Silvina San Martino, Julio M. Singer, and Edward J. Stanek III. Performance of balanced two-stage empirical predictors of realized cluster latent values from finite populations: a simulation study. Computational Statistics & Data Analysis, 52(4):2199– 2217, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002903.

deLeon:2008:AEM

[1678] A. R. de Leon and Y. Zhu. ANOVA extensions for mixed discrete and continuous data. Computational Statistics & Data Analysis, 52(4):2218– 2227, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002927.

Dray:2008:NPC

[1679] Stéphane Dray. On the number of principal components: a test of dimensionality based on measurements of similarity between matrices. Computational Statistics & Data Analysis, 52(4):2228– 2237, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002939.

Miao:2008:CID

[1680] Weiwen Miao and Paul Chiou. Confidence intervals for the difference between two means. Computational Statistics & Data Analysis, 52(4):2238– 2248, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002940.

Archer:2008:ECR

[1681] Kellie J. Archer and Ryan V. Kimes.

Empirical characterization of random forest variable importance measures. Computational Statistics & Data Analysis, 52(4):2249–2260, January 10, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003076.

Demirtas:2008:ICD

[1682] Hakan Demirtas. On imputing continuous data when the eventual interest pertains to ordinalized outcomes via threshold concept. Computational Statistics & Data Analysis, 52(4):2261– 2271, January 10, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003581.

Anonymous:2008:EBc

[1683] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52 (5):iii–v, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004756.

Anonymous:2008:Cc

[1684] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (5):vi–viii, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004768.

Yu:2008:ESB

[1685] Yaming Yu. Efficient simulation of a bivariate exponential conditionals distribution. Computational Statistics & Data Analysis, 52(5):2273– 2276, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003970.

Zhou:2008:OSA

[1686] Tianni Zhou and Robert Shumway. One-step approximations for detecting regime changes in the state space model with application to the influenza data. Computational Statistics & Data Analysis, 52(5):2277– 2291, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003325.

Tuncer:2008:OSD

[1687] Yalcin Tuncer, Murat M. Tanik, and David B. Allison. An overview of statistical decomposition techniques applied to complex systems. Computational Statistics & Data Analysis, 52(5):2292– 2310, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003350.

Paroli:2008:BIN

[1688] Roberta Paroli and Luigi Spezia.

Bayesian inference in non-homogeneous Markov mixtures of periodic autoregressions with state-dependent exogenous variables. Computational Statistics & Data Analysis, 52(5):2311– 2330, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003520.

Hartman:2008:FKL

[1689] Linda Hartman and Ola Hössjer. Fast kriging of large data sets with Gaussian Markov random fields. Computational Statistics & Data Analysis, 52(5):2331– 2349, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003568.

Celisse:2008:NDE

[1690] Alain Celisse and Stéphane Robin.

Nonparametric density estimation by exact leave-p-out cross-validation. Computational Statistics & Data Analysis, 52(5):2350–2368, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947307003842.

Kessels:2008:ODC

[1691] Roselinde Kessels, Peter Goos, and Martina Vandebroek. Optimal designs for conjoint experiments. Computational Statistics & Data Analysis, 52(5):2369–2387, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947307003994.

Wang:2008:DRN

[1692] Yong Wang. Dimension-reduced nonparametric maximum likelihood computation for interval-censored data. Computational Statistics & Data Analysis, 52(5):2388–2402, January 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004033.

Berget:2008:NMA

[1693] Ingunn Berget, Bjørn-Helge Mevik, and Tormod Næs. New modifications and applications of fuzzy Cmeans methodology. Computational Statistics & Data Analysis, 52(5):2403– 2418, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730700401X.

Spierdijk:2008:NCH

[1694] Laura Spierdijk. Nonparametric conditional hazard rate estimation: a local linear approach. Computational Statistics & Data Analysis, 52(5):2419– 2434, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700299X.

Giordano:2008:NNB

[1695] F. Giordano and M. L. Parrella.

Neural networks for bandwidth selection in local linear regression of time series. Computational Statistics & Data Analysis, 52(5):2435– 2450, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003003.

Aubin:2008:PDE

[1696] Jean-Baptiste Aubin and Samuela Leoni-Aubin. Projection density estimation under a m-sample semiparametric model. Computational Statistics & Data Analysis, 52(5):2451– 2468, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003039.

daRosa:2008:TSS

[1697] Joel Corrêa da Rosa, Alvaro Veiga, and Marcelo C. Medeiros. Treestructured smooth transition regression models. Computational Statistics & Data Analysis, 52(5):2469– 2488, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003064.

Salvador:2008:RCR

[1698] B. Salvador, M. A. Fernández, I. Martín, and C. Rueda. Robustness of classification rules that incorporate additional information. Computational Statistics & Data Analysis, 52(5):2489– 2495, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003210.

Meintanis:2008:NAG

[1699] Simos G. Meintanis. A new approach of goodness-of-fit testing for exponentiated laws applied to the generalized Rayleigh distribution. Computational Statistics & Data Analysis, 52(5):2496– 2503, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003283.

Semenov:2008:TRW

[1700] Andrei Semenov. Testing the random walk hypothesis through robust estimation of correlation. Computational Statistics & Data Analysis, 52(5):2504– 2513, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003295.

Kauermann:2008:SLE

[1701] Göran Kauermann, Ronghui Xu, and Florin Vaida. Stacked Laplace-EM algorithm for duration models with time-varying and random effects. Computational Statistics & Data Analysis, 52(5):2514–2528, January 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003301.

Muggeo:2008:FGL

[1702] Vito M. R. Muggeo and Giancarlo Ferrara. Fitting generalized linear models with unspecified link function: a P -spline approach. Computational Statistics & Data Analysis, 52(5):2529– 2537, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003313.

Liang:2008:IAS

[1703] Hua Liang and Guohua Zou. Improved AIC selection strategy for survival analysis. Computational Statistics & Data Analysis, 52(5):2538– 2548, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003349.

Sexton:2008:LEV

[1704] Joseph Sexton and Petter Laake. LogitBoost with errors-in-variables. Computational Statistics & Data Analysis, 52(5):2549–2559, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947307003374.

He:2008:PEV

[1705] Yi He and James S. Hodges. Point estimates for variance-structure parameters in Bayesian analysis of hierarchical models. Computational Statistics & Data Analysis, 52(5):2560– 2577, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003386.

Park:2008:CGF

[1706] Changyi Park, Ja-Yong Koo, Sujong Kim, Insuk Sohn, and Jae Won Lee. Classification of gene functions using support vector machine for time-course gene expression data. Computational Statistics & Data Analysis, 52(5):2578– 2587, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003404.

Kondylis:2008:SPK

[1707] Athanassios Kondylis and Joe Whittaker. Spectral preconditioning of Krylov spaces: Combining PLS and PC regression. Computational Statistics & Data Analysis, 52(5):2588– 2603, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003416.

Streftaris:2008:EAA

[1708] George Streftaris and Bruce J. Worton. Efficient and accurate approximate Bayesian inference with an application to insurance data. Computational Statistics & Data Analysis, 52(5):2604– 2622, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003428.

Lopez-de-Ullibarri:2008:NEC

[1709] Ignacio López de Ullibarri, Ricardo Cao, Carmen Cadarso-Suárez, and María J. Lado. Nonparametric estimation of conditional ROC curves: Application to discrimination tasks in computerized detection of early breast cancer. Computational Statistics & Data Analysis, 52(5):2623– 2631, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003465.

Webb:2008:BMD

[1710] Emily L. Webb and Jonathan J.

Forster. Bayesian model determination for multivariate ordinal and binary data. Computational Statistics & Data Analysis, 52(5):2632– 2649, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003477.

Liang:2008:HMA

[1711] Shengde Liang, Sudipto Banerjee, Sally Bushhouse, Andrew O. Finley, and Bradley P. Carlin. Hierarchical multiresolution approaches for dense point-level breast cancer treatment data. Computational Statistics & Data Analysis, 52(5):2650– 2668, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003489.

An:2008:IPK

[1712] Lihua An and S. Ejaz Ahmed. Improving the performance of kurtosis estimator. Computational Statistics & Data Analysis, 52(5):2669– 2681, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003507.

Kim:2008:CER

[1713] Myung Suk Kim and Suojin Wang. Consistent estimation in regression models for the drift function in some continuous time models. Computational Statistics & Data Analysis, 52(5):2682–2691, January 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003532.

Kundu:2008:HFB

[1714] Debasis Kundu, Nandini Kannan, and N. Balakrishnan. On the hazard function of Birnbaum–Saunders distribution and associated inference. Computational Statistics & Data Analysis, 52(5):2692–2702, January 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730700357X.

Xie:2008:IPP

[1715] Xian-Jin Xie, Jane Pendergast, and William Clarke. Increasing the power: a practical approach to goodness-of-fit test for logistic regression models with continuous predictors. Computational Statistics & Data Analysis, 52(5):2703– 2713, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003635.

Watanabe:2008:LSA

[1716] Daisuke Watanabe, Susumu Okada, Yasunori Fujikoshi, and Takakazu Sugiyama. Large sample approximations for the LR statistic for equality of the smallest eigenvalues of a covariance matrix under elliptical population. Computational Statistics & Data Analysis, 52(5):2714– 2724, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003659.

Balakrishnan:2008:ELI

[1717] N. Balakrishnan and Abbas Rasouli.

Exact likelihood inference for two exponential populations under joint Type-II censoring. Computational Statistics & Data Analysis, 52(5):2725– 2738, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003908.

Bertail:2008:ARB

[1718] Patrice Bertail and Stéphan Clémençon.

Approximate regenerative-block bootstrap for Markov chains. Computational Statistics & Data Analysis, 52(5):2739–2756, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016794730700391X.

Aneiros-Perez:2008:LPE

[1719] G. Aneiros-Pérez and J. M. Vilar-Fernández. Local polynomial estimation in partial linear regression models under dependence. Computational Statistics & Data Analysis, 52(5):2757– 2777, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003933.

Buchholz:2008:PPD

[1720] Anika Buchholz, Norbert Holländer, and Willi Sauerbrei. On properties of predictors derived with a two-step bootstrap model averaging approach

— a simulation study in the linear regression model. Computational Statistics & Data Analysis, 52(5):2778– 2793, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003957.

Cressie:2008:CML

[1721] Noel Cressie and Nicolas Verzelen. Conditional-mean least-squares fitting of Gaussian Markov random fields to Gaussian fields. Computational Statistics & Data Analysis, 52(5):2794– 2807, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003969.

Boente:2008:RBS

[1722] Graciela Boente and Daniela Rodriguez. Robust bandwidth selection in semiparametric partly linear regression models: Monte Carlo study and influential analysis. Computational Statistics & Data Analysis, 52(5):2808– 2828, January 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004008.

Fang:2008:SPS

[1723] Yue Fang. Semi-parametric specification tests for mixing distributions. Computational Statistics & Data Analysis, 52(5):2829–2839, January 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004021.

Anonymous:2008:EBd

[1724] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52 (6):iii–v, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800039X.

Anonymous:2008:Cd

[1725] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (6):vi–viii, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000406.

Amendola:2008:SIS

[1726] Alessandra Amendola, David Belsley, Erricos John Kontoghiorghes, Herman K. van Dijk, Yasuhiro Omori, and Eric Zivot. Special issue on statistical and computational methods in finance. Computational Statistics & Data Analysis, 52(6):2842– 2845, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004823.

Ruiz:2008:MLM

[1727] Esther Ruiz and Helena Veiga. Modelling long-memory volatilities with leverage effect: A-LMSV versus FIE-GARCH. Computational Statistics & Data Analysis, 52(6):2846–2862, February 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003544.

Creal:2008:AFS

[1728] Drew D. Creal. Analysis of filtering and smoothing algorithms for Lévy-driven stochastic volatility models. Computational Statistics & Data Analysis, 52(6):2863–2876, February 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700271X.

Lindstrom:2008:SCO

[1729] Erik Lindström, Jonas Ströjby, Mats Brodén, Magnus Wiktorsson, and Jan Holst. Sequential calibration of options. Computational Statistics & Data Analysis, 52(6):2877– 2891, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003027.

Omori:2008:BSP

[1730] Yasuhiro Omori and Toshiaki Watanabe. Block sampler and posterior mode estimation for asymmetric stochastic volatility models. Computational Statistics & Data Analysis, 52(6):2892– 2910, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003337.

Strickland:2008:PEM

[1731] Chris M. Strickland, Gael M. Martin, and Catherine S. Forbes. Parameterisation and efficient MCMC estimation of non-Gaussian state space models. Computational Statistics & Data Analysis, 52(6):2911–2930, February 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003866.

Heinen:2008:MRR

[1732] Andréas Heinen and Erick Rengifo.

Multivariate reduced rank regression in non-Gaussian contexts, using copulas. Computational Statistics & Data Analysis, 52(6):2931– 2944, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003052.

Giet:2008:MHD

[1733] Ludovic Giet and Michel Lubrano. A minimum Hellinger distance estimator for stochastic differential equations: an application to statistical inference for continuous time interest rate models. Computational Statistics & Data Analysis, 52(6):2945– 2965, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003921.

Mancino:2008:RFE

[1734] M. E. Mancino and S. Sanfelici. Robustness of Fourier estimator of integrated volatility in the presence of microstructure noise. Computational Statistics & Data Analysis, 52(6):2966– 2989, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002885.

Chen:2008:VFU

[1735] Cathy W. S. Chen, Richard Gerlach, and Edward M. H. Lin. Volatility forecasting using threshold heteroskedastic models of the intraday range. Computational Statistics & Data Analysis, 52(6):2990– 3010, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002952.

Gallo:2008:VSI

[1736] Giampiero M. Gallo and Edoardo Otranto. Volatility spillovers, interdependence and comovements: a Markov switching approach. Computational Statistics & Data Analysis, 52(6):3011– 3026, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003490.

Francq:2008:DAP

[1737] Christian Francq and Jean-Michel Zakoïan. Deriving the autocovariances of powers of Markov-switching GARCH models, with applications to statistical inference. Computational Statistics & Data Analysis, 52(6):3027– 3046, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307002988.

Amendola:2008:GPC

[1738] Alessandra Amendola and Giuseppe Storti. A GMM procedure for combining volatility forecasts. Computational Statistics & Data Analysis, 52(6):3047– 3060, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003854.

Gallegati:2008:WAS

[1739] Marco Gallegati. Wavelet analysis of stock returns and aggregate economic activity. Computational Statistics & Data Analysis, 52(6):3061– 3074, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307002915.

Coakley:2008:RLM

[1740] Jerry Coakley, Jian Dollery, and Neil Kellard. The role of long memory in hedging effectiveness. Computational Statistics & Data Analysis, 52(6):3075– 3082, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730700415X.

Johnson:2008:MEM

[1741] Lorne D. Johnson and Georgios Sakoulis. Maximizing equity market sector predictability in a Bayesian timevarying parameter model. Computational Statistics & Data Analysis, 52(6):3083–3106, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947307003593.

Valle:2008:BAE

[1742] L. Dalla Valle and P. Giudici. A Bayesian approach to estimate the marginal loss distributions in operational risk management. Computational Statistics & Data Analysis, 52(6):3107–3127, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947307003647.

Debon:2008:MRD

[1743] A. Debón, F. Montes, J. Mateu, E. Porcu, and M. Bevilacqua. Modelling residuals dependence in dynamic life tables: a geostatistical approach. Computational Statistics & Data Analysis, 52(6):3128–3147, February 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003040.

Bolduc:2008:IRS

[1744] Denis Bolduc, Lynda Khalaf, and Érick Moyneur. Identification-robust simulation-based inference in joint discrete/continuous models for energy markets. Computational Statistics & Data Analysis, 52(6):3148–3161, February 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003623.

Bertrand:2008:SRT

[1745] Patrice Bertrand and Gilbert Saporta.

Some recent trends in applied stochastic modeling and multidimensional data analysis. Computational Statistics & Data Analysis, 52(6):3164– 3166, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003982.

Pastor:2008:TRP

[1746] Dominique Pastor. A theoretical result for processing signals that have unknown distributions and priors in white Gaussian noise. Computational Statistics & Data Analysis, 52(6):3167– 3186, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730700388X.

Aguilera:2008:FBL

[1747] Ana M. Aguilera, Manuel Escabias, and Mariano J. Valderrama. Forecasting binary longitudinal data by a functional PC-ARIMA model. Computational Statistics & Data Analysis, 52(6):3187–3197, February 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003519.

Regad:2008:HMM

[1748] L. Regad, F. Guyon, J. Maupetit, P. Tufféry, and A. C. Camproux. A hidden Markov model applied to the protein 3D structure analysis. Computational Statistics & Data Analysis, 52(6):3198–3207, February 20, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003453.

Messaoud:2008:DCV

[1749] Amor Messaoud, Claus Weihs, and Franz Hering. Detection of chatter vibration in a drilling process using multivariate control charts. Computational Statistics & Data Analysis, 52(6):3208– 3219, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003611.

Daudin:2008:ECR

[1750] Jean-Jacques Daudin and Tristan Mary-Huard. Estimation of the conditional risk in classification: the swapping method. Computational Statistics & Data Analysis, 52(6):3220– 3232, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003891.

Govaert:2008:BCB

[1751] Gérard Govaert and Mohamed Nadif.

Block clustering with Bernoulli mixture models: Comparison of different approaches. Computational Statistics & Data Analysis, 52(6):3233– 3245, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003441.

Montanari:2008:IFD

[1752] Angela Montanari, Daniela G. Calò, and Cinzia Viroli. Independent factor discriminant analysis. Computational Statistics & Data Analysis, 52(6):3246– 3254, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307003556.

Becue-Bertaut:2008:MFA

[1753] Mónica Bécue-Bertaut and Jérôme Pagès. Multiple factor analysis and clustering of a mixture of quantitative, categorical and frequency data. Computational Statistics & Data Analysis, 52(6):3255– 3268, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730700360X.

DEnza:2008:EDA

[1754] Alfonso Iodice D’Enza, Francesco Palumbo, and Michael Greenacre. Exploratory data analysis leading towards the most interesting simple association rules. Computational Statistics & Data Analysis, 52(6):3269– 3281, February 20, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003878.

Anonymous:2008:EBe

[1755] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52(7):iii–v, March 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001382.

Anonymous:2008:Ce

[1756] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (7):vi–ix, March 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001394.

Scheipl:2008:SPT

[1757] Fabian Scheipl, Sonja Greven, and Helmut Küchenhoff. Size and power of tests for a zero random effect variance or polynomial regression in additive and linear mixed models. Computational Statistics & Data Analysis, 52(7):3283–3299, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004306.

Chang:2008:ADD

[1758] Fengming M. Chang, Long-Hui Chen, Yueh-Li Chen, and Chien-Yu Huang. Approximate distribution of demerit statistic — a bounding approach. Computational Statistics & Data Analysis, 52(7):3300–3309, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004318.

Olsen:2008:MSA

[1759] Lena Ringstad Olsen, Probal Chaudhuri, and Fred Godtliebsen. Multiscale spectral analysis for detecting short and long range change points in time series. Computational Statistics & Data Analysis, 52(7):3310– 3330, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004380.

Lutz:2008:RB

[1760] Roman Werner Lutz, Markus Kalisch, and Peter Bühlmann. Robustified L_{2} boosting. Computational Statistics & Data Analysis, 52(7):3331– 3341, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004410.

Nikoloulopoulos:2008:CME

[1761] Aristidis K. Nikoloulopoulos and Dimitris Karlis. Copula model evaluation based on parametric bootstrap. Computational Statistics & Data Analysis, 52(7):3342–3353, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004422.

Fuertes:2008:SBT

[1762] Ana-Maria Fuertes. Sieve bootstrap t-tests on long-run average parameters. Computational Statistics & Data Analysis, 52(7):3354– 3370, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004495.

Waegeman:2008:SOM

[1763] Willem Waegeman, Bernard De Baets, and Luc Boullart. On the scalability of ordered multi-class ROC analysis. Computational Statistics & Data Analysis, 52(7):3371–3388, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004525.

Kamdem:2008:AMI

[1764] J. Sadefo Kamdem and A. Genz. Approximation of multiple integrals over hyperboloids with application to a quadratic portfolio with options. Computational Statistics & Data Analysis, 52(7):3389–3407, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004665.

Meyer:2008:ARM

[1765] Renate Meyer, Bo Cai, and François Perron. Adaptive rejection Metropolis sampling using Lagrange interpolation polynomials of degree 2. Computational Statistics & Data Analysis, 52(7):3408– 3423, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800008X.

Winker:2008:BRB

[1766] Peter Winker. Book review: Handbook of Parallel Computing and Statistics, E. J. Kontoghiorghes (Ed.). Chapman & Hall/CRC, Boca Raton (2006). Computational Statistics & Data Analysis, 52(7):3424– 3425, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004331.

Fernandez:2008:RIS

[1767] Arturo J. Fernández. Reliability inference and sample-size determination under double censoring for some twoparameter models. Computational Statistics & Data Analysis, 52(7):3426– 3440, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004288.

Komarek:2008:GLM

[1768] Arnošt Komárek and Emmanuel Lesaffre. Generalized linear mixed model with a penalized Gaussian mixture as a random effects distribution. Computational Statistics & Data Analysis, 52(7):3441–3458, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730700429X.

Siniksaran:2008:GIM

[1769] Enis Siniksaran. A geometric interpretation of Mallows’ C_{p} statistic and an alternative plot in variable selection. Computational Statistics & Data Analysis, 52(7):3459– 3467, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730700432X.

Baklizi:2008:LBE

[1770] Ayman Baklizi. Likelihood and Bayesian estimation of Pr(X < Y) using lower record values from the generalized exponential distribution. Computational Statistics & Data Analysis, 52(7):3468–3473, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004343.

Li:2008:LAM

[1771] Yonghai Li and Daniel W. Schafer. Likelihood analysis of the multivariate ordinal probit regression model for repeated ordinal responses. Computational Statistics & Data Analysis, 52(7):3474–3492, March 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947307004355.

Taylor:2008:ABS

[1772] Charles C. Taylor. Automatic bandwidth selection for circular density estimation. Computational Statistics & Data Analysis, 52(7):3493– 3500, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004367.

Ng:2008:CIE

[1773] Hon Keung Tony Ng, Giovanni Filardo, and Gang Zheng. Confidence interval estimating procedures for standardized incidence rates. Computational Statistics & Data Analysis, 52(7):3501– 3516, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004379.

Noh:2008:HLA

[1774] Maengseok Noh and Youngjo Lee.

Hierarchical-likelihood approach for nonlinear mixed-effects models. Computational Statistics & Data Analysis, 52(7):3517–3527, March 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947307004392.

Tian:2008:EME

[1775] Guo-Liang Tian, Man-Lai Tang, Hong-Bin Fang, and Ming Tan. Efficient methods for estimating constrained parameters with applications to regularized (lasso) logistic regression. Computational Statistics & Data Analysis, 52(7):3528– 3542, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004409.

Hoshino:2008:BST

[1776] Takahiro Hoshino. Bayesian significance testing and multiple comparisons from MCMC outputs. Computational Statistics & Data Analysis, 52(7):3543– 3559, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004446.

Song:2008:CAN

[1777] Juwon Song and Thomas R. Belin. Choosing an appropriate number of factors in factor analysis with incomplete data. Computational Statistics & Data Analysis, 52(7):3560– 3569, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004458.

Dimitrova:2008:GSE

[1778] Dimitrina S. Dimitrova, Vladimir K. Kaishev, and Spiridon I. Penev. GeD spline estimation of multivariate Archimedean copulas. Computational Statistics & Data Analysis, 52(7):3570– 3582, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730700446X.

Pendharkar:2008:MEL

[1779] Parag C. Pendharkar. Maximum entropy and least square error minimizing procedures for estimating missing conditional probabilities in Bayesian networks. Computational Statistics & Data Analysis, 52(7):3583–3602, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004471.

Liang:2008:LOT

[1780] Tachen Liang, Wen-Tao Huang, and Kun-Cheng Yang. Locally optimal tests for exponential distributions with type-I censoring. Computational Statistics & Data Analysis, 52(7):3603– 3615, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004501.

Kumasaka:2008:HDD

[1781] Natsuhiko Kumasaka and Ritei Shibata. High-dimensional data visualisation: the textile plot. Computational Statistics & Data Analysis, 52(7):3616– 3644, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004513.

Hsu:2008:SSV

[1782] Nan-Jung Hsu, Hung-Lin Hung, and Ya-Mei Chang. Subset selection for vector autoregressive processes using lasso. Computational Statistics & Data Analysis, 52(7):3645– 3657, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004549.

Nott:2008:PPD

[1783] David J. Nott. Predictive performance of Dirichlet process shrinkage methods in linear regression. Computational Statistics & Data Analysis, 52(7):3658– 3669, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004562.

Nielsen:2008:AFM

[1784] J. D. Nielsen and C. B. Dean. Adaptive functional mixed NHPP models for the analysis of recurrent event panel data. Computational Statistics & Data Analysis, 52(7):3670– 3685, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004574.

Bougeard:2008:CRP

[1785] S. Bougeard, M. Hanafi, and E. M. Qannari. Continuum redundancy– PLS regression: a simple continuum approach. Computational Statistics & Data Analysis, 52(7):3686– 3696, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004677.

Zhang:2008:BAM

[1786] Xiao Zhang, W. John Boscardin, and Thomas R. Belin. Bayesian analysis of multivariate nominal measures using multivariate multinomial probit models. Computational Statistics & Data Analysis, 52(7):3697– 3708, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004690.

Park:2008:SFS

[1787] Changyi Park, Ja-Yong Koo, Peter T.

Kim, and Jae Won Lee. Stepwise feature selection using generalized logistic loss. Computational Statistics & Data Analysis, 52(7):3709– 3718, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004707.

Tang:2008:TEP

[1788] Nian-Sheng Tang, Man-Lai Tang, and Shi-Fang Qiu. Testing the equality of proportions for correlated otolaryngologic data. Computational Statistics & Data Analysis, 52(7):3719– 3729, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004835.

Fernandez:2008:TTS

[1789] V. Alba Fernández, M. D. Jiménez Gamero, and J. Muñoz García. A test for the two-sample problem based on empirical characteristic functions. Computational Statistics & Data Analysis, 52(7):3730– 3748, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004847.

Sun:2008:BSP

[1790] Xiaoqian Sun, Zhuoqiong He, and John Kabrick. Bayesian spatial prediction of the site index in the study of the Missouri ozark forest ecosystem project. Computational Statistics & Data Analysis, 52(7):3749– 3764, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004872.

Huang:2008:LTH

[1791] Yangxin Huang. Long-term HIV dynamic models incorporating drug adherence and resistance to treatment for prediction of virological responses. Computational Statistics & Data Analysis, 52(7):3765– 3778, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000029.

Wu:2008:IEP

[1792] Shu-Fei Wu. Interval estimation for a Pareto distribution based on a doubly type II censored sample. Computational Statistics & Data Analysis, 52(7):3779–3788, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000030.

Ollila:2008:CVI

[1793] Esa Ollila, Hannu Oja, and Visa Koivunen. Complex-valued ICA based on a pair of generalized covariance matrices. Computational Statistics & Data Analysis, 52(7):3789–3805, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000054.

Shen:2008:LBS

[1794] Xiaojing Shen, Yunmin Zhu, and Lixin Song. Linear B-spline copulas with applications to nonparametric estimation of copulas. Computational Statistics & Data Analysis, 52(7):3806– 3819, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000066.

Silva:2008:LBX

[1795] Giovana Oliveira Silva, Edwin M. M.

Ortega, Vicente G. Cancho, and Mauricio Lima Barreto. Log-Burr XII regression models with censored data. Computational Statistics & Data Analysis, 52(7):3820– 3842, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000078.

Yoo:2008:NMB

[1796] Jae Keun Yoo. A novel momentbased sufficient dimension reduction approach in multivariate regression. Computational Statistics & Data Analysis, 52(7):3843–3851, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000091.

Fernandes:2008:UPA

[1797] Leandro A. F. Fernandes, Manuel M. Oliveira, and Roberto da Silva. Uncertainty propagation: Avoiding the expensive sampling process for real-time image-based measurements. Computational Statistics & Data Analysis, 52(7):3852– 3876, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004550.

Nores:2008:SPR

[1798] María Laura Nores and María del Pilar Díaz. Some properties of regression estimators in GEE models for clustered ordinal data. Computational Statistics & Data Analysis, 52(7):3877– 3888, March 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004689.

Anonymous:2008:EBf

[1799] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52(8):iii–v, April 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001837.

Anonymous:2008:Cf

[1800] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (8):vi–vii, April 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001849.

Tahmasbi:2008:TPL

[1801] Rasool Tahmasbi and Sadegh Rezaei.

A two-parameter lifetime distribution with decreasing failure rate. Computational Statistics & Data Analysis, 52 (8):3889–3901, April 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004537.

Wong:2008:ADG

[1802] A. C. M. Wong. Approximating the F distribution via a general version of the modified signed log-likelihood ratio statistic. Computational Statistics & Data Analysis, 52(8):3902– 3912, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000133.

Ding:2008:EBN

[1803] Chris Ding, Tao Li, and Wei Peng. On the equivalence between non-negative matrix factorization and probabilistic latent semantic indexing. Computational Statistics & Data Analysis, 52 (8):3913–3927, April 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000145.

Carter:2008:GRN

[1804] Nathan Carter, Charles Hadlock, and Dominique Haughton. Generating random networks from a given distribution. Computational Statistics & Data Analysis, 52(8):3928– 3938, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000248.

Zhou:2008:EAE

[1805] Xingcai Zhou and Xinsheng Liu. The EM algorithm for the extended finite mixture of the factor analyzers model. Computational Statistics & Data Analysis, 52(8):3939– 3953, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000327.

Park:2008:SAC

[1806] Cheolwoo Park and Kee-Hoon Kang. SiZer analysis for the comparison of regression curves. Computational Statistics & Data Analysis, 52(8):3954– 3970, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000042.

Huang:2008:PPI

[1807] Yufen Huang, Ching-Ren Cheng, and Tai-Ho Wang. Pair-perturbation influence functions of nongaussianity by projection pursuit. Computational Statistics & Data Analysis, 52(8):3971– 3987, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000108.

Ravishanker:2008:NMM

[1808] Nalini Ravishanker, Zhaohui Liu, and Bonnie K. Ray. NHPP models with Markov switching for software reliability. Computational Statistics & Data Analysis, 52(8):3988– 3999, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800011X.

Shmueli:2008:TSC

[1809] Galit Shmueli, Wolfgang Jank, and Valerie Hyde. Transformations for semi-continuous data. Computational Statistics & Data Analysis, 52(8):4000– 4020, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000169.

Carrasco:2008:LMW

[1810] Jalmar M. F. Carrasco, Edwin M. M.

Ortega, and Gilberto A. Paula. Logmodified Weibull regression models with censored data: Sensitivity and residual analysis. Computational Statistics & Data Analysis, 52(8):4021– 4039, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000224.

Chan:2008:PIE

[1811] P. S. Chan, H. K. T. Ng, N. Balakrishnan, and Q. Zhou. Point and interval estimation for extremevalue regression model under Type-II censoring. Computational Statistics & Data Analysis, 52(8):4040– 4058, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000236.

Hrdlickova:2008:APS

[1812] Zuzana Hrdličková. Approximation of powers of some tests in one-way MANOVA type multivariate generalized linear model. Computational Statistics & Data Analysis, 52(8):4059– 4075, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000261.

Klingenberg:2008:RMB

[1813] Bernhard Klingenberg. Regression models for binary time series with gaps. Computational Statistics & Data Analysis, 52(8):4076– 4090, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000297.

Lui:2008:NIE

[1814] Kung-Jong Lui. Notes on interval estimation of risk difference in stratified noncompliance randomized trials: a Monte Carlo evaluation. Computational Statistics & Data Analysis, 52 (8):4091–4103, April 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000339.

Sohn:2008:NNM

[1815] Insuk Sohn, Sujong Kim, Changha Hwang, and Jae Won Lee. New normalization methods using support vector machine quantile regression approach in microarray analysis. Computational Statistics & Data Analysis, 52 (8):4104–4115, April 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000534.

Coull:2008:SAM

[1816] Scott E. Coull and Boleslaw K. Szymanski. Sequence alignment for masquerade detection. Computational Statistics & Data Analysis, 52(8):4116– 4131, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000170.

Roman:2008:FPT

[1817] P. Román, J. J. Serrano, and F. Torres. First-passage-time location function: Application to determine firstpassage-time densities in diffusion processes. Computational Statistics & Data Analysis, 52(8):4132– 4146, April 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000303.

Anonymous:2008:EBg

[1818] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52(9):iii–v, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002338.

Anonymous:2008:Cg

[1819] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (9):vi–vii, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800234X.

Turner:2008:DML

[1820] Rolf Turner. Direct maximization of the likelihood of a hidden Markov model. Computational Statistics & Data Analysis, 52(9):4147– 4160, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000200.

Verdier:2008:ATC

[1821] Ghislain Verdier, Nadine Hilgert, and Jean-Pierre Vila. Adaptive threshold computation for CUSUM-type procedures in change detection and isolation problems. Computational Statistics & Data Analysis, 52(9):4161– 4174, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000212.

Ong:2008:CGN

[1822] S. H. Ong and Wen-Jau Lee. Computer generation of negative binomial variates by envelope rejection. Computational Statistics & Data Analysis, 52 (9):4175–4183, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000522.

Gschlossl:2008:DGS

[1823] Susanne Gschlößl and Claudia Czado.

Does a Gibbs sampler approach to spatial Poisson regression models outperform a single site MH sampler? Computational Statistics & Data Analysis, 52(9):4184– 4202, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000844.

Danilov:2008:ELS

[1824] Dmitry Danilov and Jan R. Magnus. On the estimation of a large sparse Bayesian system: the snaer program. Computational Statistics & Data Analysis, 52(9):4203– 4224, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000868.

Duong:2008:FSM

[1825] Tarn Duong, Arianna Cowling, Inge Koch, and M. P. Wand. Feature significance for multivariate kernel density estimation. Computational Statistics & Data Analysis, 52(9):4225– 4242, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001503.

Knusel:2008:CRC

[1826] Leo Knüsel. Chisquare as a rotation criterion in factor analysis. Computational Statistics & Data Analysis, 52 (9):4243–4252, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000194.

Stearns:2008:EGP

[1827] Matthew Stearns and Sarjinder Singh. On the estimation of the general parameter. Computational Statistics & Data Analysis, 52(9):4253– 4271, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800056X.

Mongoue-Tchokote:2008:NSS

[1828] Solange Mongoué-Tchokoté and Jong-Sung Kim. New statistical software for the proportional hazards model with current status data. Computational Statistics & Data Analysis, 52 (9):4272–4286, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000625.

Roy:2008:ASS

[1829] Roch Roy and Abdessamad Saidi. Aggregation and systematic sampling of periodic ARMA processes. Computational Statistics & Data Analysis, 52 (9):4287–4304, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000649.

Dette:2008:RDS

[1830] Holger Dette and Douglas P. Wiens.

Robust designs for series estimation. Computational Statistics & Data Analysis, 52(9):4305–4324, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308000650.

Lui:2008:NEP

[1831] Kung-Jong Lui and William G.

Cumberland. Notes on estimation of proportion ratio under a noncompliance randomized trial with missing outcomes. Computational Statistics & Data Analysis, 52(9):4325– 4345, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000686.

Congdon:2008:BFM

[1832] Peter Congdon. A bivariate frailty model for events with a permanent survivor fraction and non-monotonic hazards; with an application to age at first maternity. Computational Statistics & Data Analysis, 52(9):4346– 4356, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000704.

Hohle:2008:CDR

[1833] Michael Höhle and Michaela Paul. Count data regression charts for the monitoring of surveillance time series. Computational Statistics & Data Analysis, 52(9):4357–4368, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000716.

Weiss:2008:DPC

[1834] Christian H. Weiß and Rainer Göb. Discovering patterns in categorical time series using IFS. Computational Statistics & Data Analysis, 52 (9):4369–4379, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800087X.

Reynes:2008:NGA

[1835] Christelle Reynès, Robert Sabatier, Nicolas Molinari, and Sylvain Lehmann. A new genetic algorithm in proteomics: Feature selection for SELDI-TOF data. Computational Statistics & Data Analysis, 52(9):4380– 4394, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000881.

Shu:2008:ACP

[1836] Lianjie Shu, Wei Jiang, and Zhang Wu.

Adaptive CUSUM procedures with Markovian mean estimation. Computational Statistics & Data Analysis, 52 (9):4395–4409, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800090X.

Lee:2008:UIW

[1837] Seung-Hwan Lee, Eun-Joo Lee, and Bernard Oguna Omolo. Using integrated weighted survival difference for the two-sample censored data problem. Computational Statistics & Data Analysis, 52(9):4410– 4416, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000911.

Espinheira:2008:IDB

[1838] Patrícia L. Espinheira, Silvia L. P.

Ferrari, and Francisco Cribari-Neto. Influence diagnostics in beta regression. Computational Statistics & Data Analysis, 52(9):4417– 4431, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001473.

Poulin:2008:PTK

[1839] Jennifer Poulin and Pierre Duchesne. On the power transformation of kernel-based tests for serial correlation in vector time series: Some finite sample results and a comparison with the bootstrap. Computational Statistics & Data Analysis, 52 (9):4432–4457, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001497.

Liu:2008:REF

[1840] Lei Liu, Mark R. Conaway, William A. Knaus, and James D. Bergin. A random effects four-part model, with application to correlated medical costs. Computational Statistics & Data Analysis, 52(9):4458– 4473, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001527.

Alonso:2008:FTD

[1841] A. Alonso, S. Litière, and G. Molenberghs. A family of tests to detect misspecifications in the randomeffects structure of generalized linear mixed models. Computational Statistics & Data Analysis, 52(9):4474– 4486, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001540.

Palma:2008:AIG

[1842] Wilfredo Palma, Pascal Bondon, and José Tapia. Assessing influence in Gaussian long-memory models. Computational Statistics & Data Analysis, 52(9):4487–4501, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001576.

Jukic:2008:ENW

[1843] Dragan Jukić, Mirta Benšić, and Rudolf Scitovski. On the existence of the nonlinear weighted least squares estimate for a three-parameter Weibull distribution. Computational Statistics & Data Analysis, 52(9):4502– 4511, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001588.

Wang:2008:NMD

[1844] Qin Wang and Xiangrong Yin. A nonlinear multi-dimensional variable selection method for high dimensional data: Sparse MAVE. Computational Statistics & Data Analysis, 52(9):4512– 4520, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800159X.

Juszczak:2008:PBC

[1845] Piotr Juszczak, Niall M. Adams, David J. Hand, Christopher Whitrow, and David J. Weston. Off-the-peg and bespoke classifiers for fraud detection. Computational Statistics & Data Analysis, 52(9):4521– 4532, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001710.

Colubi:2008:FFB

[1846] Ana Colubi, Gil González-Rodríguez, María José Domínguez-Cuesta, and Montserrat Jiménez-Sánchez. Favorability functions based on kernel density estimation for logistic models: a case study. Computational Statistics & Data Analysis, 52(9):4533– 4543, May 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001746.

Moskovitch:2008:DUC

[1847] Robert Moskovitch, Yuval Elovici, and Lior Rokach. Detection of unknown computer worms based on behavioral classification of the host. Computational Statistics & Data Analysis, 52 (9):4544–4566, May 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000315.

Anonymous:2008:EBh

[1848] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52(10):iii–v, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800265X.

Anonymous:2008:Ch

[1849] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (10):vi–vii, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002661.

McCullough:2008:SSM

[1850] B. D. McCullough. Special section on Microsoft Excel 2007. Computational Statistics & Data Analysis, 52(10): 4568–4569, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001655.

McCullough:2008:ASP

[1851] B. D. McCullough and David A.

Heiser. On the accuracy of statistical procedures in Microsoft Excel 2007. Computational Statistics & Data Analysis, 52(10):4570– 4578, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001606.

Yalta:2008:ASD

[1852] A. Talha Yalta. The accuracy of statistical distributions in Microsoft^{(}R) Excel 2007. Computational Statistics & Data Analysis, 52(10):4579– 4586, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001618.

McCullough:2008:MEW

[1853] B. D. McCullough. Microsoft Excel — not the Wichmann–Hill random number generators. Computational Statistics & Data Analysis, 52(10): 4587–4593, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800162X.

Su:2008:EPC

[1854] Yu-Sung Su. It’s easy to produce chartjunk using Microsoft ^{(}R) Excel 2007 but hard to make good graphs. Computational Statistics & Data Analysis, 52(10):4594–4601, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001631.

Nash:2008:TSE

[1855] John C. Nash. Teaching statistics with Excel 2007 and other spreadsheets. Computational Statistics & Data Analysis, 52(10):4602– 4606, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001643.

Fruhwirth-Schnatter:2008:MLN

[1856] Sylvia Frühwirth-Schnatter and Helga Wagner. Marginal likelihoods for non-Gaussian models using auxiliary mixture sampling. Computational Statistics & Data Analysis, 52(10):4608– 4624, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800176X.

Mukhopadhyay:2008:OMG

[1857] S. Mukhopadhyay and A. I. Khuri. Optimization in a multivariate generalized linear model situation. Computational Statistics & Data Analysis, 52(10): 4625–4634, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002028.

Trosset:2008:SPC

[1858] Michael W. Trosset and Carey E.

Priebe. The out-of-sample problem for classical multidimensional scaling. Computational Statistics & Data Analysis, 52(10):4635–4642, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001515.

Trosset:2008:SLD

[1859] Michael W. Trosset, Carey E. Priebe, Youngser Park, and Michael I. Miller. Semisupervised learning from dissimilarity data. Computational Statistics & Data Analysis, 52(10):4643– 4657, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001539.

Tsai:2008:DFW

[1860] Chieh-Yuan Tsai and Chuang-Cheng Chiu. Developing a feature weight selfadjustment mechanism for a K-means clustering algorithm. Computational Statistics & Data Analysis, 52(10): 4658–4672, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001552.

Saburi:2008:MLM

[1861] S. Saburi and N. Chino. A maximum likelihood method for an asymmetric MDS model. Computational Statistics & Data Analysis, 52(10): 4673–4684, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001667.

Otranto:2008:CHT

[1862] Edoardo Otranto. Clustering heteroskedastic time series by model-based procedures. Computational Statistics & Data Analysis, 52(10):4685– 4698, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001679.

Hitchcock:2008:SDC

[1863] David B. Hitchcock and Zhimin Chen. Smoothing dissimilarities to cluster binary data. Computational Statistics & Data Analysis, 52(10):4699– 4711, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001692.

McNicholas:2008:SLA

[1864] P. D. McNicholas, T. B. Murphy, and M. O’Regan. Standardising the lift of an association rule. Computational Statistics & Data Analysis, 52(10): 4712–4721, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001709.

McVinish:2008:SBC

[1865] R. McVinish and K. Mengersen.

Semiparametric Bayesian circular statistics. Computational Statistics & Data Analysis, 52(10):4722– 4730, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001722.

Marrel:2008:EMM

[1866] Amandine Marrel, Bertrand Iooss, François Van Dorpe, and Elena Volkova. An efficient methodology for modeling complex computer codes with Gaussian processes. Computational Statistics & Data Analysis, 52(10): 4731–4744, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001758.

Qin:2008:REE

[1867] Guo You Qin, Zhong Yi Zhu, and Wing K. Fung. Robust estimating equations and bias correction of correlation parameters for longitudinal data. Computational Statistics & Data Analysis, 52(10):4745– 4753, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001771.

Ahlgren:2008:BFD

[1868] N. Ahlgren and J. Antell. Bootstrap and fast double bootstrap tests of cointegration rank with financial time series. Computational Statistics & Data Analysis, 52(10):4754– 4767, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001783.

Tian:2008:ETA

[1869] Guo-Liang Tian, Kai Wang Ng, and Ming Tan. EM-type algorithms for computing restricted MLEs in multivariate normal distributions and multivariate t-distributions. Computational Statistics & Data Analysis, 52(10): 4768–4778, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800193X.

Sabatier:2008:ESC

[1870] Robert Sabatier and Christelle Reynès. Extensions of simple component analysis and simple linear discriminant analysis using genetic algorithms. Computational Statistics & Data Analysis, 52(10):4779–4789, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001941.

Li:2008:CFD

[1871] Bin Li and Qingzhao Yu. Classification of functional data: a segmentation approach. Computational Statistics & Data Analysis, 52(10): 4790–4800, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001953.

Chen:2008:OCI

[1872] Hubert J. Chen, Hsiu-Ling Li, and Miin-Jye Wen. Optimal confidence interval for the largest mean of correlated normal populations and its application to stock fund evaluation. Computational Statistics & Data Analysis, 52(10):4801–4813, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001989.

Lee:2008:BEM

[1873] Sik-Yum Lee and Xin-Yuan Song. On Bayesian estimation and model comparison of an integrated structural equation model. Computational Statistics & Data Analysis, 52(10):4814– 4827, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001990.

Qiu:2008:NPB

[1874] Peihua Qiu. A nonparametric procedure for blind image deblurring. Computational Statistics & Data Analysis, 52(10):4828–4841, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002004.

Yuan:2008:SEM

[1875] Ke-Hai Yuan and Wai Chan. Structural equation modeling with near singular covariance matrices. Computational Statistics & Data Analysis, 52(10):

4842–4858, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002016.

Zhang:2008:EMI

[1876] Chunming Zhang, Yuefeng Lu, Tom Johnstone, Terry Oakes, and Richard J. Davidson. Efficient modeling and inference for event-related fMRI data. Computational Statistics & Data Analysis, 52(10):4859–4871, June 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800203X.

Tian:2008:CIE

[1877] Lili Tian and Gregory E. Wilding. Confidence interval estimation of a common correlation coefficient. Computational Statistics & Data Analysis, 52(10): 4872–4877, June 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002053.

Anonymous:2008:EBi

[1878] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52(11):iii–v, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003083.

Anonymous:2008:Ci

[1879] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (11):vi–vii, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003095.

Martin:2008:WTD

[1880] Donald E. K. Martin and John A. D.

Aston. Waiting time distribution of generalized later patterns. Computational Statistics & Data Analysis, 52 (11):4879–4890, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002077.

Zhang:2008:FHI

[1881] Ji-Hong Zhang, Ling-Yun Wu, Jian Chen, and Xiang-Sun Zhang. A fast haplotype inference method for large population genotype data. Computational Statistics & Data Analysis, 52 (11):4891–4902, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002090.

Nisen:2008:SMC

[1882] Jeffrey A. Nisen and Neil C. Schwertman. A simple method of computing the sample size for chi-square test for the equality of multinomial distributions. Computational Statistics & Data Analysis, 52(11):4903– 4908, July 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002119.

DiNardo:2008:SCM

[1883] E. Di Nardo, G. Guarino, and D. Senato. Symbolic computation of moments of sampling distributions. Computational Statistics & Data Analysis, 52 (11):4909–4922, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002193.

Depril:2008:AAC

[1884] Dirk Depril, Iven Van Mechelen, and Boris Mirkin. Algorithms for additive clustering of rectangular data tables. Computational Statistics & Data Analysis, 52(11):4923– 4938, July 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002247.

Cysneiros:2008:BSI

[1885] Audrey H. M. A. Cysneiros, Francisco Cribari-Neto, and Carlos A. G. Araújo, Jr. On Birnbaum–Saunders inference. Computational Statistics & Data Analysis, 52(11):4939– 4950, July 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002065.

Bathke:2008:HCS

[1886] Arne C. Bathke, Solomon W. Harrar, and Laurence V. Madden. How to compare small multivariate samples using nonparametric tests. Computational Statistics & Data Analysis, 52 (11):4951–4965, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002107.

Cutillo:2008:LED

[1887] L. Cutillo and U. Amato. Localized empirical discriminant analysis. Computational Statistics & Data Analysis, 52 (11):4966–4978, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002120.

Cuesta-Albertos:2008:RTD

[1888] J. A. Cuesta-Albertos and A. Nieto-Reyes. The random Tukey depth. Computational Statistics & Data Analysis, 52(11):4979–4988, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308002144.

Jahn-Eimermacher:2008:CAG

[1889] Antje Jahn-Eimermacher. Comparison of the Andersen–Gill model with Poisson and negative binomial regression on recurrent event data. Computational Statistics & Data Analysis, 52 (11):4989–4997, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002156.

Caporale:2008:MUU

[1890] Guglielmo Maria Caporale and Luis A.

Gil-Alana. Modelling the US, UK and Japanese unemployment rates: Fractional integration and structural breaks. Computational Statistics & Data Analysis, 52(11):4998– 5013, July 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002260.

Randal:2008:RRS

[1891] John A. Randal. A reinvestigation of robust scale estimation in finite samples. Computational Statistics & Data Analysis, 52(11):5014– 5021, July 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002272.

Wang:2008:CTE

[1892] D. Wang and S.-J. Soong. Comparisons of titer estimation methods for multiplexed pneumococcal opsonophagocytic killing assay. Computational Statistics & Data Analysis, 52(11): 5022–5032, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002284.

Jara:2008:LMM

[1893] Alejandro Jara, Fernando Quintana, and Ernesto San Martín. Linear mixed models with skew-elliptical distributions: a Bayesian approach. Computational Statistics & Data Analysis, 52 (11):5033–5045, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002454.

Reiczigel:2008:ECS

[1894] Jenő Reiczigel, Zsolt Abonyi-Tóth, and Júlia Singer. An exact confidence set for two binomial proportions and exact unconditional confidence intervals for the difference and ratio of proportions. Computational Statistics & Data Analysis, 52(11): 5046–5053, July 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002491.

Anonymous:2008:EBj

[1895] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 52 (12):iii–v, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003630.

Anonymous:2008:Cj

[1896] Anonymous. Contents. Computational Statistics & Data Analysis, 52 (12):vi–vii, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003642.

Vera:2008:GSA

[1897] J. Fernando Vera and Jose A. Díaz-García. A global simulated annealing heuristic for the three-parameter lognormal maximum likelihood estimation. Computational Statistics & Data Analysis, 52(12):5055–5065, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002430.

Joe:2008:ALA

[1898] Harry Joe. Accuracy of Laplace approximation for discrete response mixed models. Computational Statistics & Data Analysis, 52(12):5066– 5074, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002533.

Cabral:2008:BDE

[1899] Celso Rômulo Barbosa Cabral, Heleno Bolfarine, and José Raimundo Gomes Pereira. Bayesian density estimation using skew Student-t-normal mixtures. Computational Statistics & Data Analysis, 52(12):5075–5090, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002545.

Kwong:2008:EJD

[1900] Koon-Shing Kwong and Yiu-Man Chan. On the evaluation of the joint distribution of order statistics. Computational Statistics & Data Analysis, 52(12):5091–5099, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002557.

Botts:2008:FAB

[1901] Carsten H. Botts and Michael J. Daniels. A flexible approach to Bayesian multiple curve fitting. Computational Statistics & Data Analysis, 52(12):5100–5120, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308002570.

Salibian-Barrera:2008:RMS

[1902] Matias Salibian-Barrera and Stefan Van Aelst. Robust model selection using fast and robust bootstrap. Computational Statistics & Data Analysis, 52(12):5121–5135, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002600.

Hu:2008:BQE

[1903] Jiaqiao Hu and Zheng Su. Bootstrap quantile estimation via importance resampling. Computational Statistics & Data Analysis, 52(12):5136– 5142, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800282X.

Nashimoto:2008:BMC

[1904] Kane Nashimoto and F. T. Wright.

Bayesian multiple comparisons of simply ordered means using priors with a point mass. Computational Statistics & Data Analysis, 52(12):5143– 5153, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002843.

Cizek:2008:SER

[1905] P. Čižek, J. Tamine, and W. Härdle. Smoothed L-estimation of regression function. Computational Statistics & Data Analysis, 52(12):5154–5162, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002909.

Hofert:2008:SAC

[1906] Marius Hofert. Sampling Archimedean copulas. Computational Statistics & Data Analysis, 52(12):5163–5174, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002910.

Singh:2008:CBE

[1907] Bhupendra Singh, K. K. Sharma, and Anuj Kumar. A classical and Bayesian estimation of a k-components loadsharing parallel system. Computational Statistics & Data Analysis, 52(12): 5175–5185, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002958.

Hubert:2008:ABS

[1908] M. Hubert and E. Vandervieren. An adjusted boxplot for skewed distributions. Computational Statistics & Data Analysis, 52(12):5186–5201, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004434.

Yang:2008:NBE

[1909] Mingan Yang, Timothy Hanson, and Ronald Christensen. Nonparametric Bayesian estimation of a bivariate density with interval censored data. Computational Statistics & Data Analysis, 52(12):5202– 5214, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002089.

Last:2008:PA

[1910] Michael Last, Gheorghe Luta, Alex Orso, Adam Porter, and Stan Young. Pooled ANOVA. Computational Statistics & Data Analysis, 52(12):5215– 5228, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002168.

Takai:2008:TIC

[1911] Keiji Takai and Yutaka Kano. Test of independence in a 2 × 2 contingency table with nonignorable nonresponse via constrained EM algorithm. Computational Statistics & Data Analysis, 52(12):5229–5241, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002417.

Gonzalez-Manteiga:2008:ABA

[1912] W. González-Manteiga, M. J. Lombardía, I. Molina, D. Morales, and L. Santamaría. Analytic and bootstrap approximations of prediction errors under a multivariate Fay– Herriot model. Computational Statistics & Data Analysis, 52(12):5242– 5252, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002429.

Tekle:2008:MOD

[1913] Fetene B. Tekle, Frans E. S. Tan, and Martijn P. F. Berger. Maximin Doptimal designs for binary longitudinal responses. Computational Statistics & Data Analysis, 52(12):5253– 5262, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002442.

Forcina:2008:IEL

[1914] Antonio Forcina. Identifiability of extended latent class models with individual covariates. Computational Statistics & Data Analysis, 52(12): 5263–5268, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002478.

Nguyen:2008:AAC

[1915] Nam-Ky Nguyen and Min-Qian Liu. An algorithmic approach to constructing mixed-level orthogonal and nearorthogonal arrays. Computational Statistics & Data Analysis, 52(12): 5269–5276, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002569.

Wang:2008:NAG

[1916] Hansheng Wang and Chenlei Leng. A note on adaptive group lasso. Computational Statistics & Data Analysis, 52(12):5277–5286, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308002582.

Tscheschel:2008:QPS

[1917] André Tscheschel and Sung Nok Chiu.

Quasi-plus sampling edge correction for spatial point patterns. Computational Statistics & Data Analysis, 52(12): 5287–5295, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002727.

Zou:2008:RSM

[1918] Hui Zou and Ming Yuan. Regularized simultaneous model selection in multiple quantiles regression. Computational Statistics & Data Analysis, 52(12): 5296–5304, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002739.

Binder:2008:IUA

[1919] Harald Binder and Willi Sauerbrei. Increasing the usefulness of additive spline models by knot removal. Computational Statistics & Data Analysis, 52(12):5305–5318, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002740.

Prendergast:2008:TIO

[1920] Luke A. Prendergast. Trimming influential observations for improved singleindex model estimated sufficient summary plots. Computational Statistics & Data Analysis, 52(12):5319– 5327, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002806.

Chernov:2008:FCD

[1921] N. Chernov and P. N. Sapirstein. Fitting circles to data with correlated noise. Computational Statistics & Data Analysis, 52(12):5328– 5337, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002831.

Best:2008:CST

[1922] D. J. Best, J. C. W. Rayner, and O. Thas. Comparison of some tests of fit for the Laplace distribution. Computational Statistics & Data Analysis, 52(12):5338–5343, August 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002855.

Gueorguieva:2008:DCR

[1923] Ralitza Gueorguieva, Robert Rosenheck, and Daniel Zelterman. Dirichlet component regression and its applications to psychiatric data. Computational Statistics & Data Analysis, 52(12):5344–5355, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308002971.

Liu:2008:AAC

[1924] Xueli Liu, Sheng-Chien Lee, George Casella, and Gary F. Peter. Assessing agreement of clustering methods with gene expression microarray data. Computational Statistics & Data Analysis, 52(12):5356– 5366, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003022.

Fridley:2008:RGM

[1925] Brooke L. Fridley, Stephen T. Turner, Arlene B. Chapman, Andrei S. Rodin, Eric Boerwinkle, and Kent R. Bailey. Reproducibility of genotypes as measured by the Affymetrix GeneChip^{(}R) 100K Human Mapping Array set. Computational Statistics & Data Analysis, 52(12):5367– 5374, August 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002818. See

comments [2345].

Anonymous:2008:EBk

[1926] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53 (1):v–vii, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004118.

Anonymous:2008:Ck

[1927] Anonymous. Contents. Computational Statistics & Data Analysis, 53(1):viii– ix, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800412X.

Bentarzi:2008:MPA

[1928] M. Bentarzi and F. Hamdi. Mixture periodic autoregressive conditional heteroskedastic models. Computational Statistics & Data Analysis, 53(1):1– 16, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800323X.

Wang:2008:CPP

[1929] Hsiuying Wang. Coverage probability of prediction intervals for discrete random variables. Computational Statistics & Data Analysis, 53(1):17– 26, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003502.

Yu:2008:CMS

[1930] Q. Yu, W. Tang, Y. Ma, S. A. Gamble, and X. M. Tu. Comparing multiple sensitivities and specificities with different diagnostic criteria: Applications to sexual abuse and sexual health research. Computational Statistics & Data Analysis, 53(1):27– 37, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800296X.

Lau:2008:BMA

[1931] John W. Lau and Mike K. P.

So. Bayesian mixture of autoregressive models. Computational Statistics & Data Analysis, 53(1):38–60, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002983.

Belitz:2008:SSV

[1932] Christiane Belitz and Stefan Lang. Simultaneous selection of variables and smoothing parameters in structured additive regression models. Computational Statistics & Data Analysis, 53(1):61–81, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308003009.

Josse:2008:TSR

[1933] J. Josse, J. Pagès, and F. Husson. Testing the significance of the RV coefficient. Computational Statistics & Data Analysis, 53(1):82–91, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003186.

Vidal:2008:CBM

[1934] Ignacio Vidal and Pilar Iglesias. Comparison between a measurement error model and a linear model without measurement error. Computational Statistics & Data Analysis, 53(1):92– 102, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003277.

Magnussen:2008:MBR

[1935] Steen Magnussen and Rob Reeves. A method for bias-reduction of sample-based MLE of the autologistic model. Computational Statistics & Data Analysis, 53(1):103–111, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003289.

Liu:2008:SNM

[1936] Wei Liu and Lang Wu. A semiparametric nonlinear mixed-effects model with non-ignorable missing data and measurement errors for HIV viral data. Computational Statistics & Data Analysis, 53(1):112–122, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003290.

Gramacy:2008:GPL

[1937] Robert B. Gramacy and Herbert K. H. Lee. Gaussian processes and limiting linear models. Computational Statistics & Data Analysis, 53(1):123– 136, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003307.

Herwartz:2008:HPU

[1938] H. Herwartz and F. Siedenburg. Homogeneous panel unit root tests under cross sectional dependence: Finite sample modifications and the wild bootstrap. Computational Statistics & Data Analysis, 53(1):137–150, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003319.

Aguilera:2008:DDL

[1939] Ana M. Aguilera, Manuel Escabias, and Mariano J. Valderrama. Discussion of different logistic models with functional data. Application to systemic Lupus Erythematosus. Computational Statistics & Data Analysis, 53(1):151– 163, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003332.

Zheng:2008:SCD

[1940] Shurong Zheng. Selection of components and degrees of smoothing via lasso in high dimensional nonparametric additive models. Computational Statistics & Data Analysis, 53(1):164– 175, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003356.

Sun:2008:PSE

[1941] Jie Sun, Karen A. Kopciuk, and Xuewen Lu. Polynomial spline estimation of partially linear singleindex proportional hazards regression models. Computational Statistics & Data Analysis, 53(1):176–188, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003381.

Fernandez-Macho:2008:SES

[1942] Javier Fernández-Macho. Spectral estimation of a structural thin-plate smoothing model. Computational Statistics & Data Analysis, 53(1):189– 195, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003447.

Dorta-Guerra:2008:TLE

[1943] Roberto Dorta-Guerra, Enrique González-Dávila, and Josep Ginebra. Twolevel experiments for binary response data. Computational Statistics & Data Analysis, 53(1):196–208, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003472.

Lui:2008:THR

[1944] Kung-Jong Lui and Kuang-Chao Chang. Testing homogeneity of risk difference in stratified randomized trials with noncompliance. Computational Statistics & Data Analysis, 53(1):209– 221, September 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/

article/pii/S0167947308003514. See

corrigendum [2067].

Oh:2008:MDP

[1945] Sohee Oh, K. C. Carriere, and Taesung Park. Model diagnostic plots for repeated measures data using the generalized estimating equations approach. Computational Statistics & Data Analysis, 53(1):222–232, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003526.

Wan:2008:CCR

[1946] Shuwen Wan and Biao Zhang. Comparing correlated ROC curves for continuous diagnostic tests under density ratio models. Computational Statistics & Data Analysis, 53(1):233–245, September 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800354X.

Anonymous:2008:EBl

[1947] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53 (2):iii–v, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005136.

Anonymous:2008:Cl

[1948] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (2):vi–vii, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005148.

Harman:2008:COE

[1949] Radoslav Harman and Tomáš Jurík. Computing c-optimal experimental designs using the simplex method of linear programming. Computational Statistics & Data Analysis, 53(2):247– 254, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003320.

Rivero:2008:ARL

[1950] Carlos Rivero and Teofilo Valdes. An algorithm for robust linear estimation with grouped data. Computational Statistics & Data Analysis, 53(2):255– 271, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003344.

Liesenfeld:2008:IMU

[1951] Roman Liesenfeld and Jean-François Richard. Improving MCMC, using efficient importance sampling. Computational Statistics & Data Analysis, 53(2):272–288, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003770.

Stout:2008:URP

[1952] Quentin F. Stout. Unimodal regression via prefix isotonic regression. Computational Statistics & Data Analysis, 53(2):289–297, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003964.

Schmid:2008:BAM

[1953] Matthias Schmid and Torsten Hothorn.

Boosting additive models using componentwise P -splines. Computational Statistics & Data Analysis, 53(2):298–311, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004416.

Dette:2008:IUR

[1954] Holger Dette, Andrey Pepelyshev, and Anatoly Zhigljavsky. Improving updating rules in multiplicative algorithms for computing D-optimal designs. Computational Statistics & Data Analysis, 53(2):312–320, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004659.

Chang:2008:TCM

[1955] Ching-Hui Chang and Nabendu Pal.

Testing on the common mean of several normal distributions. Computational Statistics & Data Analysis, 53(2):321– 333, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003538.

Yoo:2008:RDR

[1956] Jae Keun Yoo and R. Dennis Cook. Response dimension reduction for the conditional mean in multivariate regression. Computational Statistics & Data Analysis, 53(2):334–343, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003721.

Tsiaplias:2008:FEU

[1957] Sarantis Tsiaplias. Factor estimation using MCMC-based Kalman filter methods. Computational Statistics & Data Analysis, 53(2):344–353, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003769.

Conti:2008:EMN

[1958] Pier Luigi Conti, Daniela Marella, and Mauro Scanu. Evaluation of matching noise for imputation techniques based on nonparametric local linear regression estimators. Computational Statistics & Data Analysis, 53(2):354– 365, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003794.

Balakrishnan:2008:FIB

[1959] N. Balakrishnan, Marco Burkschat, Erhard Cramer, and Glenn Hofmann. Fisher information based progressive censoring plans. Computational Statistics & Data Analysis, 53(2):366–380, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003836.

Rigby:2008:FMO

[1960] R. A. Rigby, D. M. Stasinopoulos, and C. Akantziliotou. A framework for modelling overdispersed count data, including the Poisson-shifted generalized inverse Gaussian distribution. Computational Statistics & Data Analysis, 53(2):381–393, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003861.

Park:2008:MMR

[1961] Mingue Park and HyungJun Cho.

Minimum MSE regression estimator with estimated population quantities of auxiliary variables. Computational Statistics & Data Analysis, 53(2):394– 404, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003903.

Siddique:2008:UAB

[1962] Juned Siddique and Thomas R. Belin. Using an approximate Bayesian bootstrap to multiply impute nonignorable missing data. Computational Statistics & Data Analysis, 53(2):405– 415, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003915.

Ahmad:2008:AHD

[1963] M. Rauf Ahmad, C. Werner, and E. Brunner. Analysis of highdimensional repeated measures designs: the one sample case. Computational Statistics & Data Analysis, 53(2):416–427, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947308004003.

Georgiou:2008:MSD

[1964] Stelios D. Georgiou. Modelling by supersaturated designs. Computational Statistics & Data Analysis, 53(2):428– 435, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004015.

DelaCruz:2008:BNL

[1965] Rolando De la Cruz. Bayesian nonlinear regression models with skewelliptical errors: Applications to the classification of longitudinal profiles. Computational Statistics & Data Analysis, 53(2):436–449, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004027.

Carrasco:2008:GMW

[1966] Jalmar M. F. Carrasco, Edwin M. M. Ortega, and Gauss M. Cordeiro. A generalized modified Weibull distribution for lifetime modeling. Computational Statistics & Data Analysis, 53(2):450– 462, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004192.

Hayter:2008:CPC

[1967] A. J. Hayter, Jongphil Kim, and W. Liu. Critical point computations for one-sided and two-sided pairwise comparisons of three treatment means. Computational Statistics & Data Analysis, 53(2):463–470, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004209.

Subramanian:2008:MCI

[1968] Sundarraman Subramanian and Derek Bean. The missing censoring indicator model and the smoothed bootstrap. Computational Statistics & Data Analysis, 53(2):471–476, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004246.

Kim:2008:SBW

[1969] Jin Seon Kim and Bong-Jin Yum. Selection between Weibull and lognormal distributions: a comparative simulation study. Computational Statistics & Data Analysis, 53(2):477–485, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004258.

Yamano:2008:SAI

[1970] Takuya Yamano, Kodai Sato, Taisei Kaizoji, Jan-Michael Rost, and Lukáš Pichl. Symbolic analysis of indicator time series by quantitative sequence alignment. Computational Statistics & Data Analysis, 53(2):486– 495, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004325.

Wu:2008:ODC

[1971] Zhang Wu, Mei Yang, Wei Jiang, and Michael B. C. Khoo. Optimization designs of the combined Shewhart-CUSUM control charts. Computational Statistics & Data Analysis, 53(2):496– 506, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004337.

Mealli:2008:CPS

[1972] Fabrizia Mealli and Barbara Pacini. Comparing principal stratification and selection models in parametric causal inference with nonignorable missingness. Computational Statistics & Data Analysis, 53(2):507–516, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004374.

vanderLinde:2008:VBF

[1973] Angelika van der Linde. Variational Bayesian functional PCA. Computational Statistics & Data Analysis, 53(2):517–533, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004404.

Kapetanios:2008:BBT

[1974] George Kapetanios. Bootstrapbased tests for deterministic timevarying coefficients in regression models. Computational Statistics & Data Analysis, 53(2):534–545, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004428.

Liang:2008:CIC

[1975] Hua Liang, Haiyan Su, and Guohua Zou. Confidence intervals for a common mean with missing data with applications in an AIDS study. Computational Statistics & Data Analysis, 53(2):546– 553, December 15, 2008. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004556.

Chen:2008:CIC

[1976] Chia-Cheng Chen and Huiman X.

Barnhart. Comparison of ICC and CCC for assessing agreement for data without and with replications. Computational Statistics & Data Analysis, 53(2):554–564, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800457X.

Sakhanenko:2008:TES

[1977] Lyudmila Sakhanenko. Testing for ellipsoidal symmetry: a comparison study. Computational Statistics & Data Analysis, 53(2):565–581, December 15, 2008. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004295.

Anonymous:2009:EBa

[1978] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53 (3):iii–v, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800529X.

Anonymous:2009:Ca

[1979] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (3):vi–vii, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005306.

Edler:2009:CSW

[1980] Lutz Edler, Jae Won Lee, Martina Mittlböck, Joyce Niland, and Norbert Victor. Computational statistics within clinical research. Computational Statistics & Data Analysis, 53(3):583– 585, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004684.

Sill:2009:DLD

[1981] Michael W. Sill and Allan R. Sampson. Drop-the-losers design: Binomial case. Computational Statistics & Data Analysis, 53(3):586–595, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003812.

Ahn:2009:EII

[1982] Chul Ahn, Fan Hu, and Celette Sugg Skinner. Effect of imbalance and intracluster correlation coefficient in cluster randomized trials with binary outcomes. Computational Statistics & Data Analysis, 53(3):596–602, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004477.

Leon:2009:SSR

[1983] Andrew C. Leon and Moonseong Heo.

Sample sizes required to detect interactions between two binary fixedeffects in a mixed-effects linear regression model. Computational Statistics & Data Analysis, 53(3):603–608, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003241.

Gillen:2009:RWA

[1984] Daniel L. Gillen. A random walk approach for quantifying uncertainty in group sequential survival trials. Computational Statistics & Data Analysis, 53(3):609–620, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004222.

Tang:2009:ACI

[1985] Man-Lai Tang and Maozai Tian. Asymptotic confidence interval construction for risk difference under inverse sampling. Computational Statistics & Data Analysis, 53(3):621–631, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003575.

Parsons:2009:RMP

[1986] Nick R. Parsons, Matthew L. Costa, Juul Achten, and Nigel Stallard. Repeated measures proportional odds logistic regression analysis of ordinal score data in the statistical software package R. Computational Statistics & Data Analysis, 53(3):632–641, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003939.

Antic:2009:CNM

[1987] J. Antic, C. M. Laffont, D. Chafaï, and D. Concordet. Comparison of nonparametric methods in nonlinear mixed effects models. Computational Statistics & Data Analysis, 53(3):642– 656, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004210.

Costa:2009:PPS

[1988] M. J. Costa and J. E. H. Shaw.

Parametrization and penalties in spline models with an application to survival analysis. Computational Statistics & Data Analysis, 53(3):657–670, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003745.

Frobish:2009:PEC

[1989] Daniel Frobish and Nader Ebrahimi.

Parametric estimation of change-points for actual event data in recurrent events models. Computational Statistics & Data Analysis, 53(3):671–682, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004180.

Lo:2009:SAS

[1990] Serigne N. Lô, Stephane Heritier, and Malcolm Hudson. Saddlepoint approximation for semi-Markov processes with application to a cardiovascular randomised study. Computational Statistics & Data Analysis, 53(3):683– 698, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004398.

Ghosh:2009:BAL

[1991] Pulak Ghosh and Paul S. Albert.

A Bayesian analysis for longitudinal semicontinuous data with an application to an acupuncture clinical trial. Computational Statistics & Data Analysis, 53(3):699–706, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004465.

Page:2009:EDS

[1992] John H. Page and Andrea Rotnitzky. Estimation of the disease-specific diagnostic marker distribution under verification bias. Computational Statistics & Data Analysis, 53(3):707–717, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003265.

Adler:2009:BET

[1993] Werner Adler and Berthold Lausen. Bootstrap estimated true and false positive rates and ROC curve. Computational Statistics & Data Analysis, 53(3):718–729, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016794730800460X.

Konietschke:2009:NAC

[1994] Frank Konietschke and Edgar Brunner. Nonparametric analysis of clustered data in diagnostic trials: Estimation problems in small sample sizes. Computational Statistics & Data Analysis, 53(3):730–741, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800399X.

Nofuentes:2009:DSS

[1995] J. A. Roldán Nofuentes, J. D. Luna del Castillo, and M. A. Montero Alonso. Determining sample size to evaluate and compare the accuracy of binary diagnostic tests in the presence of partial disease verification. Computational Statistics & Data Analysis, 53(3):742– 755, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003782.

Yang:2009:EMC

[1996] Tae Young Yang. Efficient multiclass cancer diagnosis algorithm, using a global similarity pattern. Computational Statistics & Data Analysis, 53(3):756–765, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308004271.

Kitchen:2009:TWB

[1997] Christina M. R. Kitchen, Jing Kroll, Daniel R. Kuritzkes, Erik Bloomquist, Steven G. Deeks, and Marc A. Suchard. Two-way Bayesian hierarchical phylogenetic models: an application to the co-evolution of gp120 and gp41 during and after enfuvirtide treatment. Computational Statistics & Data Analysis, 53(3):766–775, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003204.

Kropf:2009:CEP

[1998] Siegfried Kropf, Jürgen Läuter, Daniela Kose, and Dietrich von Rosen. Comparison of exact parametric tests for high-dimensional data. Computational Statistics & Data Analysis, 53(3):776–787, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003459.

Cheng:2009:IVI

[1999] Cheng Cheng. Internal validation inferences of significant genomic features in genome-wide screening. Computational Statistics & Data Analysis, 53(3):788– 800, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800340X.

Sexton:2009:SEB

[2000] Joseph Sexton and Petter Laake. Standard errors for bagged and random forest estimators. Computational Statistics & Data Analysis, 53(3):801–811, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003988.

Rosenkranz:2009:MLD

[2001] Gerd K. Rosenkranz. Modeling laboratory data from clinical trials. Computational Statistics & Data Analysis, 53(3):812–819, January 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003411.

Aban:2009:IPA

[2002] Inmaculada B. Aban, Gary R. Cutter, and Nsoki Mavinga. Inferences and power analysis concerning two negative binomial distributions with an application to MRI lesion counts data. Computational Statistics & Data Analysis, 53(3):820– 833, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800371X.

OConnell:2009:RMA

[2003] Rachel L. O’Connell and H. Malcolm Hudson. Risk of mortality after acute myocardial infarction: Performance of model updating methods for application in different geographical regions. Computational Statistics & Data Analysis, 53(3):834– 846, January 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004234.

Anonymous:2009:EBb

[2004] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53 (4):iii–v, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000139.

Anonymous:2009:Cb

[2005] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (4):vi–ix, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000140.

Jamalizadeh:2009:RRD

[2006] A. Jamalizadeh, M. Khosravi, and N. Balakrishnan. Recurrence relations for distributions of a skew-t and a linear combination of order statistics from a bivariate-t. Computational Statistics & Data Analysis, 53(4):847– 852, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005367.

Liu:2009:NCS

[2007] Huan Liu, Yongqiang Tang, and Hao Helen Zhang. A new chisquare approximation to the distribution of non-negative definite quadratic forms in non-central normal variables. Computational Statistics & Data Analysis, 53(4):853–856, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005653.

Ahmad:2009:TAC

[2008] Ibrahim A. Ahmad and Mohammad B. Sepehrifar. On testing alternative classes of life distribution with guaranteed survival times. Computational Statistics & Data Analysis, 53(4):857– 864, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004660.

Liang:2009:LBN

[2009] Faming Liang and Jian Zhang.

Learning Bayesian networks for discrete data. Computational Statistics & Data Analysis, 53(4):865–876, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004805.

Klugkist:2009:OSN

[2010] Irene Klugkist and Herbert Hoijtink.

Obtaining similar null distributions in the normal linear model using computational methods. Computational Statistics & Data Analysis, 53(4):877– 888, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004842.

Lim:2009:MLE

[2011] Johan Lim, Xinlei Wang, and Wanseok Choi. Maximum likelihood estimation of ordered multinomial probabilities by geometric programming. Computational Statistics & Data Analysis, 53(4):889–893, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004891.

Karlis:2009:NEP

[2012] Dimitris Karlis. A note on the exponential Poisson distribution: a nested EM algorithm. Computational Statistics & Data Analysis, 53(4):894–899, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004908.

Papastamoulis:2009:RJM

[2013] Panagiotis Papastamoulis and George Iliopoulos. Reversible jump MCMC in mixtures of normal distributions with the same component means. Computational Statistics & Data Analysis, 53(4):900–911, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004933.

Shim:2009:SVC

[2014] Jooyong Shim and Changha Hwang. Support vector censored quantile regression under random censoring. Computational Statistics & Data Analysis, 53(4):912–919, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004957.

Welham:2009:NBL

[2015] S. J. Welham and R. Thompson. A note on bimodality in the loglikelihood function for penalized spline mixed models. Computational Statistics & Data Analysis, 53(4):920–931, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005021.

Weinert:2009:FCA

[2016] Howard L. Weinert. A fast compact algorithm for cubic spline smoothing. Computational Statistics & Data Analysis, 53(4):932–940, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005045.

Murakami:2009:BPM

[2017] Junko Murakami. Bayesian posterior mean estimates for Poisson hidden Markov models. Computational Statistics & Data Analysis, 53(4):941– 955, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005379.

Kundu:2009:EPM

[2018] Debasis Kundu and Arabin Kumar Dey. Estimating the parameters of the Marshall–Olkin bivariate Weibull distribution by EM algorithm. Computational Statistics & Data Analysis, 53(4):956–965, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005392.

Endo:2009:IPS

[2019] Yushi Endo and Akimichi Takemura. Iterative proportional scaling via decomposable submodels for contingency tables. Computational Statistics & Data Analysis, 53(4):966–978, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005483.

Trendafilov:2009:SID

[2020] Nickolay T. Trendafilov and Karen Vines. Simple and interpretable discrimination. Computational Statistics & Data Analysis, 53(4):979–989, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005495.

Shi:2009:SCR

[2021] Xiaoping Shi, Yuehua Wu, and Baiqi Miao. Strong convergence rate of estimators of change point and its application. Computational Statistics & Data Analysis, 53(4):990–998, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005513.

Oman:2009:ESM

[2022] Samuel D. Oman. Easily simulated multivariate binary distributions with given positive and negative correlations. Computational Statistics & Data Analysis, 53(4):999–1005, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005598.

Hara:2009:MBC

[2023] Hisayuki Hara, Akimichi Takemura, and Ruriko Yoshida. A Markov basis for conditional test of common diagonal effect in quasi-independence model for square contingency tables. Computational Statistics & Data Analysis, 53(4):1006–1014, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005689.

Rokach:2009:CAB

[2024] Lior Rokach. Collective-agreement-based pruning of ensembles. Computational Statistics & Data Analysis, 53(4):1015–1026, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308005707.

Feddag:2009:PLL

[2025] M.-L. Feddag and S. Bacci. Pairwise likelihood for the longitudinal mixed Rasch model. Computational Statistics & Data Analysis, 53(4):1027– 1037, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004301.

Tenreiro:2009:CSP

[2026] Carlos Tenreiro. On the choice of the smoothing parameter for the BHEP goodness-of-fit test. Computational Statistics & Data Analysis, 53(4):1038– 1053, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004362.

Chakraborti:2009:PIN

[2027] S. Chakraborti, S. Eryilmaz, and S. W. Human. A phase II nonparametric control chart based on precedence statistics with runs-type signaling rules. Computational Statistics & Data Analysis, 53(4):1054– 1065, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004532.

Ghosh:2009:RBA

[2028] Pulak Ghosh, C. L. Bayes, and V. H. Lachos. A robust Bayesian approach to null intercept measurement error model with application to dental data. Computational Statistics & Data Analysis, 53(4):1066–1079, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004581.

Zou:2009:NCI

[2029] Guang Yong Zou, Wenyi Huang, and Xiaohe Zhang. A note on confidence interval estimation for a linear function of binomial proportions. Computational Statistics & Data Analysis, 53(4):1080– 1085, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004623.

Wilderjans:2009:SAC

[2030] Tom Wilderjans, Eva Ceulemans, and Iven Van Mechelen. Simultaneous analysis of coupled data blocks differing in size: a comparison of two weighting schemes. Computational Statistics & Data Analysis, 53(4):1086– 1098, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004647.

Jonkman:2009:EPU

[2031] Jeffrey N. Jonkman, Patrick D. Gerard, and William H. Swallow. Estimating probabilities under the threeparameter gamma distribution using composite sampling. Computational Statistics & Data Analysis, 53(4):1099– 1109, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004738.

Fan:2009:MES

[2032] Juanjuan Fan, Martha E. Nunn, and Xiaogang Su. Multivariate exponential survival trees and their application to tooth prognosis. Computational Statistics & Data Analysis, 53(4):1110– 1121, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004787.

Kim:2009:EBG

[2033] Sung-Ho Kim, Hyemi Choi, and Sangjin Lee. Estimate-based goodnessof-fit test for large sparse multinomial distributions. Computational Statistics & Data Analysis, 53(4):1122– 1131, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004817.

Xu:2009:NGF

[2034] Xingzhong Xu, Xiaobo Ding, and Shuran Zhao. New goodness-of-fit tests based on fiducial empirical distribution function. Computational Statistics & Data Analysis, 53(4):1132– 1141, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004829.

Proust-Lima:2009:JMM

[2035] Cécile Proust-Lima, Pierre Joly, Jean-François Dartigues, and Hélène Jacqmin-Gadda. Joint modelling of multivariate longitudinal outcomes and a time-to-event: a nonlinear latent class approach. Computational Statistics & Data Analysis, 53(4):1142– 1154, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004830.

Cordeiro:2009:PSG

[2036] Gauss M. Cordeiro, Marinho G. Andrade, and Mário de Castro. Power series generalized nonlinear models. Computational Statistics & Data Analysis, 53(4):1155–1166, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004866.

Karvanen:2009:ACE

[2037] Juha Karvanen. Approximate costefficient sequential designs for binary response models with application to switching measurements. Computational Statistics & Data Analysis, 53(4):1167–1176, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004878.

Liang:2009:GTM

[2038] Jiajuan Liang and Man-Lai Tang. Generalized F -tests for the multivariate normal mean. Computational Statistics & Data Analysis, 53(4):1177– 1190, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800488X.

Prescott:2009:OCL

[2039] Philip Prescott. Orthogonal-column Latin hypercube designs with small samples. Computational Statistics & Data Analysis, 53(4):1191–1200, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004945.

Song:2009:SSS

[2040] James X. Song. Sample size for simultaneous testing of rate differences in non-inferiority trials with multiple endpoints. Computational Statistics & Data Analysis, 53(4):1201– 1207, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004969.

Briand:2009:SMA

[2041] Bénédicte Briand, Gilles R. Ducharme, Vanessa Parache, and Catherine Mercat-Rommens. A similarity measure to assess the stability of classification trees. Computational Statistics & Data Analysis, 53(4):1208– 1217, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004970.

Zhang:2009:UBP

[2042] Chun-Xia Zhang, Jiang-She Zhang, and Gai-Ying Zhang. Using boosting to prune double-bagging ensembles. Computational Statistics & Data Analysis, 53(4):1218–1231, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004982.

Thomas:2009:EDN

[2043] Alun Thomas and Peter J. Green. Enumerating the decomposable neighbors of a decomposable graph under a simple perturbation scheme.

Computational Statistics & Data Analysis, 53(4):1232–1238, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005008.

Moghimbeigi:2009:STZ

[2044] Abbas Moghimbeigi, Mohammad Reza Eshraghian, Kazem Mohammad, and Brian McArdle. A score test for zero-inflation in multilevel count data. Computational Statistics & Data Analysis, 53(4):1239–1248, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800501X.

Osorio:2009:EID

[2045] Felipe Osorio, Gilberto A. Paula, and Manuel Galea. On estimation and influence diagnostics for the Grubbs’ model under heavy-tailed distributions. Computational Statistics & Data Analysis, 53(4):1249–1263, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005057.

Broda:2009:EDR

[2046] S. Broda and M. S. Paolella. Evaluating the density of ratios of noncentral quadratic forms in normal variables. Computational Statistics & Data Analysis, 53(4):1264–1270, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005070.

Sepulveda:2009:BAA

[2047] Nuno Sepúlveda, Carlos Daniel Paulino, and Carlos Penha-Gonçalves. Bayesian analysis of allelic penetrance models for complex binary traits. Computational Statistics & Data Analysis, 53(4):1271– 1283, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005082.

Daye:2009:SMS

[2048] Z. John Daye and X. Jessie Jeng. Shrinkage and model selection with correlated variables via weighted fusion. Computational Statistics & Data Analysis, 53(4):1284–1298, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005355.

Nielsen:2009:SCU

[2049] Jens D. Nielsen, Rafael Rumí, and Antonio Salmerón. Supervised classification using probabilistic decision graphs. Computational Statistics & Data Analysis, 53(4):1299–1311, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005380.

Shin:2009:RST

[2050] Dong Wan Shin, Soo Jung Park, and Man-Suk Oh. A robust sign test for panel unit roots under cross sectional dependence. Computational Statistics & Data Analysis, 53(4):1312– 1327, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005409.

Abdel-Hamid:2009:ESS

[2051] Alaa H. Abdel-Hamid and Essam K.

AL-Hussaini. Estimation in stepstress accelerated life tests for the exponentiated exponential distribution with type-I censoring. Computational Statistics & Data Analysis, 53(4):1328– 1338, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800546X.

Dette:2009:BTC

[2052] Holger Dette and Rafael Weißbach. A bootstrap test for the comparison of nonlinear time series. Computational Statistics & Data Analysis, 53(4):1339– 1349, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005562.

Meza:2009:EPN

[2053] Cristian Meza, Florence Jaffrézic, and Jean-Louis Foulley. Estimation in the probit normal model for binary outcomes using the SAEM algorithm. Computational Statistics & Data Analysis, 53(4):1350–1360, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005574.

Liu:2009:SCR

[2054] Xueli Liu and Mark C. K. Yang. Simultaneous curve registration and clustering for functional data. Computational Statistics & Data Analysis, 53(4):1361– 1376, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005586.

Martin:2009:BAG

[2055] J. Martín and C. J. Pérez. Bayesian analysis of a generalized lognormal distribution. Computational Statistics & Data Analysis, 53(4):1377– 1387, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005604.

Lambert:2009:BDE

[2056] Philippe Lambert and Paul H. C. Eilers. Bayesian density estimation from grouped continuous data. Computational Statistics & Data Analysis, 53(4):1388–1399, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005616.

Ferraty:2009:APB

[2057] Frédéric Ferraty and Philippe Vieu. Additive prediction and boosting for functional data. Computational Statistics & Data Analysis, 53(4):1400– 1413, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005628.

Douzal-Chouakria:2009:ACT

[2058] Ahlame Douzal-Chouakria, Alpha Diallo, and Françoise Giroud. Adaptive clustering for time series: Application for identifying cell cycle expressed genes. Computational Statistics & Data Analysis, 53(4):1414– 1426, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005677.

Pinto:2009:SAP

[2059] D. Leão Pinto, Jr., R. Aoki, and G. F. Silva. Statistical analysis of proficiency testing results under elliptical distributions. Computational Statistics & Data Analysis, 53(4):1427– 1439, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005720.

Zhang:2009:CCB

[2060] Jiujun Zhang, Zhonghua Li, and Zhaojun Wang. Control chart based on likelihood ratio for monitoring linear profiles. Computational Statistics & Data Analysis, 53(4):1440– 1448, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005744.

Bijwaard:2009:ERP

[2061] Govert E. Bijwaard and Philip Hans Franses. The effect of rounding on payment efficiency. Computational Statistics & Data Analysis, 53(4):1449– 1461, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004672.

Chakraborty:2009:SCC

[2062] Sounak Chakraborty. Simultaneous cancer classification and gene selection with Bayesian nearest neighbor method: an integrated approach. Computational Statistics & Data Analysis, 53(4):1462–1474, February 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004726.

Reiter:2009:VSE

[2063] Jerome P. Reiter, Anna Oganian, and Alan F. Karr. Verification servers: Enabling analysts to assess the quality of inferences from public use data. Computational Statistics & Data Analysis, 53(4):1475– 1482, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004751.

Menahem:2009:IMD

[2064] Eitan Menahem, Asaf Shabtai, Lior Rokach, and Yuval Elovici. Improving malware detection by applying multiinducer ensemble. Computational Statistics & Data Analysis, 53(4):1483– 1494, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004763.

Gupta:2009:ERP

[2065] Ramesh C. Gupta and Cheng Peng.

Estimating reliability in proportional odds ratio models. Computational Statistics & Data Analysis, 53(4):1495– 1510, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004799.

Barros:2009:RIG

[2066] Michelli Barros, Gilberto A. Paula, and Víctor Leiva. An R implementation for generalized Birnbaum– Saunders distributions. Computational Statistics & Data Analysis, 53(4):1511– 1528, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005434.

Lui:2009:CTH

[2067] Kung-Jong Lui and Kuang-Chao Chang. Corrigendum to: “Testing homogeneity of risk difference in stratified randomized trials with noncompliance” [comput. statist. data anal. 53 (2008) 209–221]. Computational Statistics & Data Analysis, 53(4): 1529, February 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/

                                        article/pii/S0167947308004520. See

[1944].

Anonymous:2009:EBc

[2068] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53(5):iii–v, March 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000528.

Anonymous:2009:Cc

[2069] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (5):vi–viii, March 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900053X.

Allison:2009:SGS

[2070] David B. Allison, Peter M. Visscher, Guilherme J. M. Rosa, and Christopher I. Amos. Statistical genetics & statistical genomics: Where biology, epistemology, statistics, and computation collide. Computational Statistics & Data Analysis, 53(5):1531– 1534, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000255.

Salas-Gonzalez:2009:HTE

[2071] Diego Salas-Gonzalez, Ercan E. Kuruoglu, and Diego P. Ruiz. A heavytailed empirical Bayes method for replicated microarray data. Computational Statistics & Data Analysis, 53(5):1535– 1546, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003952.

Angelini:2009:BMT

[2072] Claudia Angelini, Daniela De Canditiis, and Marianna Pensky. Bayesian models for two-sample time-course microarray experiments. Computational Statistics & Data Analysis, 53(5):1547– 1565, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003496.

Muir:2009:MMA

[2073] W. M. Muir, G. J. M. Rosa, B. R.

Pittendrigh, Z. Xu, S. D. Rider, M. Fountain, and J. Ogas. A mixture model approach for the analysis of small exploratory microarray experiments. Computational Statistics & Data Analysis, 53(5):1566– 1576, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003162.

Chan:2009:RRE

[2074] Shih-Huang Chan and Wan-Chi Chang. A robust ratio estimator of gene expression via inversevariance weighting for cDNA microarray images. Computational Statistics & Data Analysis, 53(5):1577– 1589, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002995.

vanWieringen:2009:SPU

[2075] Wessel N. van Wieringen, David Kun, Regina Hampel, and Anne-Laure Boulesteix. Survival prediction using gene expression data: a review and comparison. Computational Statistics & Data Analysis, 53(5):1590– 1603, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002946.

Pounds:2009:AAA

[2076] Stan Pounds and Shesh N. Rai. Assumption adequacy averaging as a concept for developing more robust methods for differential gene expression analysis. Computational Statistics & Data Analysis, 53(5):1604– 1612, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002752.

Leiva:2009:GND

[2077] Víctor Leiva, Antonio Sanhueza, Diana M. Kelmansky, and Elena J. Martínez. On the glog-normal distribution and its application to the gene expression problem. Computational Statistics & Data Analysis, 53(5):1613– 1621, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002223.

Gordon:2009:BTO

[2078] Alexander Gordon, Linlin Chen, Galina Glazko, and Andrei Yakovlev.

Balancing type one and two errors in multiple testing for differential expression of genes. Computational Statistics & Data Analysis, 53(5):1622– 1629, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002181.

Marot:2009:SMM

[2079] Guillemette Marot, Jean-Louis Foulley, and Florence Jaffrézic. A structural mixed model to shrink covariance matrices for time-course differential gene expression studies. Computational Statistics & Data Analysis, 53(5):1630– 1638, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800217X.

Steibel:2009:ODT

[2080] Juan P. Steibel, Guilherme J. M.

Rosa, and Robert J. Tempelman. Optimizing design of two-stage experiments for transcriptional profiling. Computational Statistics & Data Analysis, 53(5):1639–1649, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001977.

Parrish:2009:DMS

[2081] Rudolph S. Parrish, Horace J. Spencer III, and Ping Xu. Distribution modeling and simulation of gene expression data. Computational Statistics & Data Analysis, 53(5):1650– 1660, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001965.

Wang:2009:CQT

[2082] Dong Wang and Song Xi Chen. Combining quantitative trait loci analyses and microarray data: an empirical likelihood approach. Computational Statistics & Data Analysis, 53(5):1661– 1673, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000923.

Xu:2009:MLD

[2083] Ping Xu, Guy N. Brock, and Rudolph S. Parrish. Modified linear discriminant analysis approaches for classification of high-dimensional microarray data. Computational Statistics & Data Analysis, 53(5):1674– 1687, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000558.

Hunt:2009:BBD

[2084] Daniel L. Hunt, Cheng Cheng, and Stanley Pounds. The beta-binomial distribution for estimating the number of false rejections in microarray gene expression studies. Computational Statistics & Data Analysis, 53(5):1688– 1700, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800025X.

Khalili:2009:RUA

[2085] Abbas Khalili, Tim Huang, and Shili Lin. A robust unified approach to analyzing methylation and gene expression data. Computational Statistics & Data Analysis, 53(5):1701– 1710, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003393.

Bingham:2009:CCM

[2086] Ella Bingham and Heikki Mannila.

Complexity control in a mixture model by the Hardy–Weinberg equilibrium. Computational Statistics & Data Analysis, 53(5):1711–1719, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003484.

Sharp:2009:SAP

[2087] Julia L. Sharp, John J. Borkowski, Denise Schmoyer, Don S. Daly, Samuel Purvine, William R. Cannon, and Gregory B. Hurst. Statistically appraising process quality of affinity isolation experiments. Computational Statistics & Data Analysis, 53(5):1720– 1726, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002594.

Sohn:2009:ITF

[2088] Insuk Sohn, Jooyong Shim, Changha Hwang, Sujong Kim, and Jae Won Lee. Informative transcription factor selection using support vector machinebased generalized approximate cross validation criteria. Computational Statistics & Data Analysis, 53(5):1727– 1735, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002521.

Shim:2009:SMG

[2089] Jooyong Shim, Insuk Sohn, Sujong Kim, Jae Won Lee, Paul E. Green, and Changha Hwang. Selecting marker genes for cancer classification using supervised weighted kernel clustering and the support vector machine. Computational Statistics & Data Analysis, 53(5):1736– 1742, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002466.

Yang:2009:SBB

[2090] Tae Young Yang. Simple Bayesian binary framework for discovering significant genes and classifying cancer diagnosis. Computational Statistics & Data Analysis, 53(5):1743– 1754, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002132.

Vaughan:2009:UPS

[2091] Laura K. Vaughan, Jasmin Divers, Miguel A. Padilla, David T. Redden, Hemant K. Tiwari, Daniel Pomp, and David B. Allison. The use of plasmodes as a supplement to simulations: a simple example evaluating individual admixture estimation methodologies. Computational Statistics & Data Analysis, 53(5):1755– 1766, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000856.

Li:2009:EFP

[2092] Zhaohai Li, Hong Zhang, Gang Zheng, Joseph L. Gastwirth, and Mitchell H. Gail. Excess false positive rate caused by population stratification and disease rate heterogeneity in case–control association studies. Computational Statistics & Data Analysis, 53(5):1767– 1781, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000893.

Karvanen:2009:ODS

[2093] Juha Karvanen, Sangita Kulathinal, and Dario Gasbarra. Optimal designs to select individuals for genotyping conditional on observed binary or survival outcomes and nongenetic covariates. Computational Statistics & Data Analysis, 53(5):1782– 1793, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000674.

Sun:2009:FIS

[2094] Yan V. Sun, Douglas M. Jacobsen, Stephen T. Turner, Eric Boerwinkle, and Sharon L. R. Kardia. Fast implementation of a scan statistic for identifying chromosomal patterns of genome wide association studies. Computational Statistics & Data Analysis, 53(5):1794– 1801, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002235.

Hartman:2009:UID

[2095] Linda Hartman, Ola Hössjer, and Keith Humphreys. Utilizing identity-by-descent probabilities for genetic finemapping in population based samples, via spatial smoothing of haplotype effects. Computational Statistics & Data Analysis, 53(5):1802– 1817, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002922.

Thomas:2009:EGM

[2096] Alun Thomas. Estimation of graphical models whose conditional independence graphs are interval graphs and its application to modelling linkage disequilibrium. Computational Statistics & Data Analysis, 53(5):1818– 1828, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000546.

Sung:2009:PLA

[2097] Heejong Sung, Fei Ji, Deborah L. Levy, Steven Matthysse, and Nancy Role Mendell. The power of linkage analysis of a disease-related endophenotype using asymmetrically ascertained sib pairs. Computational Statistics & Data Analysis, 53(5):1829– 1842, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004040.

Cheng:2009:PSM

[2098] Jung-Yu Cheng and Shinn-Jia Tzeng.

Parametric and semiparametric methods for mapping quantitative trait loci. Computational Statistics & Data Analysis, 53(5):1843–1849, March 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308004283.

Shriner:2009:DIC

[2099] Daniel Shriner and Nengjun Yi. Deviance information criterion (DIC) in Bayesian multiple QTL mapping. Computational Statistics & Data Analysis, 53(5):1850–1860, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000285.

Usotskaya:2009:AIT

[2100] N. Usotskaya and B. Ryabko. Application of information-theoretic tests for the analysis of DNA sequences based on Markov chain models. Computational Statistics & Data Analysis, 53(5):1861– 1872, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800337X.

Nur:2009:BHM

[2101] Darfiana Nur, David Allingham, Judith Rousseau, Kerrie L. Mengersen, and Ross McVinish. Bayesian hidden Markov model for DNA sequence segmentation: a prior sensitivity analysis. Computational Statistics & Data Analysis, 53(5):1873– 1882, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003368.

Narain:2009:TDT

[2102] P. Narain. Transmission disequilibrium test (TDT) for a pair of linked marker loci. Computational Statistics & Data Analysis, 53(5):1883– 1893, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001734.

Verzelen:2009:TGG

[2103] N. Verzelen and F. Villers. Tests for Gaussian graphical models. Computational Statistics & Data Analysis, 53(5):1894–1905, March 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947308004507.

Lama:2009:EDM

[2104] Nicola Lama, Patrizia Boracchi, and Elia Biganzoli. Exploration of distributional models for a novel intensity-dependent normalization procedure in censored gene expression data. Computational Statistics & Data Analysis, 53(5):1906– 1922, March 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800563X.

Anonymous:2009:EBd

[2105] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53(6):iii–v, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000796.

Anonymous:2009:Cd

[2106] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (6):vi–viii, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000802.

Belsley:2009:FSI

[2107] David A. Belsley, Russell Davidson, Erricos John Kontoghiorghes, James G. MacKinnon, and Herman K. van Dijk. The fourth special issue on computational econometrics. Computational Statistics & Data Analysis, 53 (6):1923–1924, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000322.

Ando:2009:BIH

[2108] Tomohiro Ando. Bayesian inference for the hazard term structure with functional predictors using Bayesian predictive information criteria. Computational Statistics & Data Analysis, 53 (6):1925–1939, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947307004859.

Arteche:2009:UBF

[2109] Josu Arteche and Jesus Orbe. Using the bootstrap for finite sample confidence intervals of the log periodogram regression. Computational Statistics & Data Analysis, 53(6):1940– 1953, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002508.

Bauer:2009:USA

[2110] Dietmar Bauer and Martin Wagner. Using subspace algorithm cointegration analysis: Simulation performance and application to the term structure. Computational Statistics & Data Analysis, 53(6):1954– 1973, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005033.

Bauwens:2009:EIS

[2111] L. Bauwens and F. Galli. Efficient importance sampling for ML estimation of SCD models. Computational Statistics & Data Analysis, 53(6):1974– 1992, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000637.

Chen:2009:BCE

[2112] Cathy W. S. Chen, Richard Gerlach, and D. C. M. Wei. Bayesian causal effects in quantiles: Accounting for heteroscedasticity. Computational Statistics & Data Analysis, 53 (6):1993–2007, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005811.

Beaulieu:2009:FSM

[2113] Marie-Claude Beaulieu, Jean-Marie Dufour, and Lynda Khalaf. Finite sample multivariate tests of asset pricing models with coskewness. Computational Statistics & Data Analysis, 53 (6):2008–2021, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002041.

Borovkova:2009:IVO

[2114] Svetlana Borovkova and Ferry J.

Permana. Implied volatility in oil markets. Computational Statistics & Data Analysis, 53(6):2022– 2039, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000662.

Bouezmarni:2009:SMD

[2115] T. Bouezmarni and J. V. K. Rombouts. Semiparametric multivariate density estimation for positive data using copulas. Computational Statistics & Data Analysis, 53(6):2040– 2054, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003010.

tenCate:2009:SMI

[2116] Arie ten Cate. Solving models with inequalities using standard econometric software. Computational Statistics & Data Analysis, 53(6):2055– 2060, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000698.

Chen:2009:MSP

[2117] Baoline Chen and Peter A. Zadrozny. Multi-step perturbation solution of nonlinear differentiable equations applied to an econometric analysis of productivity. Computational Statistics & Data Analysis, 53(6):2061– 2074, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003976.

Chiarella:2009:VSF

[2118] Carl Chiarella, Hing Hung, and Thuy-Duong Tô. The volatility structure of the fixed income market under the HJM framework: a nonlinear filtering approach. Computational Statistics & Data Analysis, 53(6):2075– 2088, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003733.

Davidson:2009:TTI

[2119] James Davidson and Nigar Hashimzade.

Type I and type II fractional Brownian motions: a reconsideration. Computational Statistics & Data Analysis, 53 (6):2089–2106, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005458.

Elezovic:2009:FMV

[2120] Suad Elezović. Functional modelling of volatility in the Swedish limit order book. Computational Statistics & Data Analysis, 53(6):2107– 2118, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000121.

Godolphin:2009:NFR

[2121] J. D. Godolphin. New formulations for recursive residuals as a diagnostic tool in the fixed-effects linear model with design matrices of arbitrary rank. Computational Statistics & Data Analysis, 53(6):2119– 2128, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004544.

Haas:2009:AMN

[2122] Markus Haas, Stefan Mittnik, and Marc S. Paolella. Asymmetric multivariate normal mixture GARCH. Computational Statistics & Data Analysis, 53(6):2129–2154, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947307004860.

Herwartz:2009:NAB

[2123] H. Herwartz and F. Xu. A new approach to bootstrap inference in functional coefficient models. Computational Statistics & Data Analysis, 53 (6):2155–2167, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004519.

Fantazzini:2009:EMM

[2124] Dean Fantazzini. The effects of misspecified marginals and copulas on computing the value at risk: a Monte Carlo study. Computational Statistics & Data Analysis, 53(6):2168– 2188, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000340.

Febrer:2009:FEW

[2125] Antonia Febrer and Juan Mora. Flexible estimation of wage distributions in the presence of covariates. Computational Statistics & Data Analysis, 53 (6):2189–2200, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003435.

Figa-Talamanca:2009:TVA

[2126] Gianna Figà-Talamanca. Testing volatility autocorrelation in the constant elasticity of variance stochastic volatility model. Computational Statistics & Data Analysis, 53(6):2201– 2218, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004039.

Foschi:2009:CPD

[2127] Paolo Foschi and Andrea Pascucci. Calibration of a path-dependent volatility model: Empirical tests. Computational Statistics & Data Analysis, 53 (6):2219–2235, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005069.

Gimeno:2009:GAE

[2128] Ricardo Gimeno and Juan M. Nave.

A genetic algorithm estimation of the term structure of interest rates. Computational Statistics & Data Analysis, 53(6):2236–2250, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947308004994.

Grossi:2009:RFW

[2129] Luigi Grossi and Fabrizio Laurini.

A robust forward weighted Lagrange multiplier test for conditional heteroscedasticity. Computational Statistics & Data Analysis, 53(6):2251– 2263, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004593.

Hubert:2009:RPS

[2130] Mia Hubert, Peter Rousseeuw, and Tim Verdonck. Robust PCA for skewed data and its outlier map. Computational Statistics & Data Analysis, 53 (6):2264–2274, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002879.

Lee:2009:OTA

[2131] Yonghee Lee and Dong Wan Shin. Optimal tests against the alternative hypothesis of panel unit roots. Computational Statistics & Data Analysis, 53 (6):2275–2283, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003198.

Liu:2009:EEC

[2132] Yan Liu and Richard Luger. Efficient estimation of copula-GARCH models. Computational Statistics & Data Analysis, 53(6):2284–2297, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308000182.

Lombardi:2009:IES

[2133] Marco J. Lombardi and Giorgio Calzolari. Indirect estimation of α-stable stochastic volatility models. Computational Statistics & Data Analysis, 53 (6):2298–2308, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005537.

Lombardi:2009:IEE

[2134] Marco J. Lombardi and David Veredas.

Indirect estimation of elliptical stable distributions. Computational Statistics & Data Analysis, 53(6):2309– 2324, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800248X.

Murphy:2009:BLM

[2135] A. Murphy and M. Izzeldin. Bootstrapping long memory tests: Some Monte Carlo results. Computational Statistics & Data Analysis, 53(6):2325– 2334, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800385X.

Nakajima:2009:LHT

[2136] Jouchi Nakajima and Yasuhiro Omori.

Leverage, heavy-tails and correlated jumps in stochastic volatility models. Computational Statistics & Data Analysis, 53(6):2335– 2353, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001680.

Nunkesser:2009:OAT

[2137] Robin Nunkesser, Roland Fried, Karen Schettlinger, and Ursula Gather. Online analysis of time series by the Q_{n} estimator. Computational Statistics & Data Analysis, 53(6):2354– 2362, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001485.

Retzer:2009:IIP

[2138] J. J. Retzer, E. S. Soofi, and R. Soyer. Information importance of predictors: Concept, measures, Bayesian inference, and applications. Computational Statistics & Data Analysis, 53 (6):2363–2377, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308001564.

Swamy:2009:EME

[2139] P. A. V. B. Swamy, Jatinder S. Mehta, I-Lok Chang, and T. S. Zimmerman. An efficient method of estimating the true value of a population characteristic from its discrepant estimates. Computational Statistics & Data Analysis, 53(6):2378– 2389, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307004483.

Takada:2009:SMH

[2140] Teruko Takada. Simulated minimum Hellinger distance estimation of stochastic volatility models. Computational Statistics & Data Analysis, 53 (6):2390–2403, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003253.

Takahashi:2009:ESV

[2141] Makoto Takahashi, Yasuhiro Omori, and Toshiaki Watanabe. Estimating stochastic volatility models using daily returns and realized volatility simultaneously. Computational Statistics & Data Analysis, 53(6):2404– 2426, April 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003824.

Taufer:2009:SLD

[2142] Emanuele Taufer and Nikolai Leonenko. Simulation of Lévy-driven Ornstein–Uhlenbeck processes with given marginal distribution. Computational Statistics & Data Analysis, 53 (6):2427–2437, April 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308001461.

Anonymous:2009:EBe

[2143] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53(7):iii–v, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000887.

Anonymous:2009:Ce

[2144] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (7):vi–vii, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000899.

Kokolakis:2009:IPM

[2145] G. Kokolakis and D. Fouskakis. Importance partitioning in microaggregation. Computational Statistics & Data Analysis, 53(7):2439– 2445, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004611.

Bangalore:2009:HAE

[2146] Sai Santosh Bangalore, Jelai Wang, and David B. Allison. How accurate are the extremely small P -values used in genomic research: an evaluation of numerical libraries. Computational Statistics & Data Analysis, 53 (7):2446–2452, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005501.

Shafik:2009:BNA

[2147] Nivien Shafik and Gerhard Tutz. Boosting nonlinear additive autoregressive time series. Computational Statistics & Data Analysis, 53(7):2453– 2464, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005719.

Liu:2009:ACM

[2148] Min-Qian Liu and Li Zhang. An algorithm for constructing mixedlevel k-circulant supersaturated designs. Computational Statistics & Data Analysis, 53(7):2465–2470, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005793.

deLeeuw:2009:SQM

[2149] Jan de Leeuw and Kenneth Lange.

Sharp quadratic majorization in one dimension. Computational Statistics & Data Analysis, 53(7):2471– 2484, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000036.

Hong:2009:MFU

[2150] Tzung-Pei Hong, Chun-Wei Lin, and Yu-Lung Wu. Maintenance of fast updated frequent pattern trees for record deletion. Computational Statistics & Data Analysis, 53(7):2485– 2499, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000279.

Mosler:2009:CZT

[2151] Karl Mosler, Tatjana Lange, and Pavel Bazovkin. Computing zonoid trimmed regions of dimension d > 2. Computational Statistics & Data Analysis, 53 (7):2500–2510, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000334.

Abdel-Hamid:2009:CPA

[2152] Alaa H. Abdel-Hamid. Constantpartially accelerated life tests for Burr type-XII distribution with progressive type-II censoring. Computational Statistics & Data Analysis, 53(7):2511– 2523, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000401.

Wu:2009:SBT

[2153] Edmond H. C. Wu, Philip L. H. Yu, and W. K. Li. A smoothed bootstrap test for independence based on mutual information. Computational Statistics & Data Analysis, 53(7):2524– 2536, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005641.

An:2009:SEG

[2154] Lihua An, Sévérien Nkurunziza, Karen Y. Fung, Daniel Krewski, and Isaac Luginaah. Shrinkage estimation in general linear models. Computational Statistics & Data Analysis, 53 (7):2537–2549, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005665.

Cao:2009:GPE

[2155] Jiguo Cao and James O. Ramsay. Generalized profiling estimation for global and adaptive penalized spline smoothing. Computational Statistics & Data Analysis, 53(7):2550– 2562, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005690.

Wu:2009:MMA

[2156] Qiang Wu and Allan R. Sampson.

Mixture modeling with applications in schizophrenia research. Computational Statistics & Data Analysis, 53 (7):2563–2572, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005732.

Melo:2009:ITI

[2157] Tatiane F. N. Melo, Silvia L. P. Ferrari, and Francisco Cribari-Neto. Improved testing inference in mixed linear models. Computational Statistics & Data Analysis, 53(7):2573– 2582, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005756.

Kenward:2009:IAP

[2158] Michael G. Kenward and James H.

Roger. An improved approximation to the precision of fixed effects from restricted maximum likelihood. Computational Statistics & Data Analysis, 53 (7):2583–2595, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005768.

Berger:2009:VES

[2159] Yves G. Berger, Juan F. Muñoz, and Eric Rancourt. Variance estimation of survey estimates calibrated on estimated control totals — an application to the extended regression estimator and the regression composite estimator. Computational Statistics & Data Analysis, 53(7):2596– 2604, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800577X.

Malloy:2009:CMM

[2160] Elizabeth J. Malloy, Donna Spiegelman, and Ellen A. Eisen. Comparing measures of model selection for penalized splines in Cox models. Computational Statistics & Data Analysis, 53 (7):2605–2616, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005781.

Wang:2009:BEV

[2161] Hai-Bin Wang. Bayesian estimation and variable selection for single index models. Computational Statistics & Data Analysis, 53(7):2617– 2627, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800580X.

Baillo:2009:TZI

[2162] A. Baíllo, J. R. Berrendero, and J. Cárcamo. Tests for zero-inflation and overdispersion: a new approach based on the stochastic convex order. Computational Statistics & Data Analysis, 53 (7):2628–2639, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005823.

Sulieman:2009:PSC

[2163] H. Sulieman, I. Kucuk, and P. J.

McLellan. Parametric sensitivity: a case study comparison. Computational Statistics & Data Analysis, 53 (7):2640–2652, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000048.

Maravelakis:2009:ECM

[2164] Petros E. Maravelakis and Philippe Castagliola. An EWMA chart for monitoring the process standard deviation when parameters are estimated. Computational Statistics & Data Analysis, 53(7):2653– 2664, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900005X.

Li:2009:EPD

[2165] Lexin Li. Exploiting predictor domain information in sufficient dimension reduction. Computational Statistics & Data Analysis, 53(7):2665– 2672, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000061.

Hsieh:2009:SAR

[2166] S. H. Hsieh, S. M. Lee, and P. S.

Shen. Semiparametric analysis of randomized response data with missing covariates in logistic regression. Computational Statistics & Data Analysis, 53 (7):2673–2692, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000073.

Luo:2009:ACC

[2167] Yunzhao Luo, Zhonghua Li, and Zhaojun Wang. Adaptive CUSUM control chart with variable sampling intervals. Computational Statistics & Data Analysis, 53(7):2693– 2701, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000085.

Wang:2009:TVA

[2168] Bing Xing Wang. Testing for the validity of the assumptions in the exponential step-stress accelerated lifetesting model. Computational Statistics & Data Analysis, 53(7):2702– 2709, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000267.

Lu:2009:CSM

[2169] Zeng-Hua Lu. Covariate selection in mixture models with the censored response variable. Computational Statistics & Data Analysis, 53(7):2710– 2723, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000280.

Eleftheraki:2009:BAT

[2170] Anastasia G. Eleftheraki, Maria Kateri, and Ioannis Ntzoufras. Bayesian analysis of two dependent 2 × 2 contingency tables. Computational Statistics & Data Analysis, 53(7):2724– 2732, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000292.

Wang:2009:CSC

[2171] Hongkun Wang and Yichuan Zhao. A comparison of some confidence intervals for the mean quality-adjusted lifetime with censored data. Computational Statistics & Data Analysis, 53 (7):2733–2739, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000309.

Thandrayen:2009:LVR

[2172] Joanne Thandrayen and Yan Wang.

A latent variable regression model for capture–recapture data. Computational Statistics & Data Analysis, 53 (7):2740–2746, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000310.

Krzanowski:2009:SMS

[2173] Wojtek J. Krzanowski and David J.

Hand. A simple method for screening variables before clustering microarray data. Computational Statistics & Data Analysis, 53(7):2747– 2753, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900036X.

Lui:2009:IEO

[2174] Kung-Jong Lui and Kuang-Chao Chang. Interval estimation of odds ratio in a stratified randomized clinical trial with noncompliance. Computational Statistics & Data Analysis, 53 (7):2754–2766, May 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000425.

Ambrogi:2009:ECC

[2175] Federico Ambrogi, Elia Biganzoli, and Patrizia Boracchi. Estimating crude cumulative incidences through multinomial logit regression on discrete cause-specific hazards. Computational Statistics & Data Analysis, 53(7):2767– 2779, May 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000024.

Anonymous:2009:EBf

[2176] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53(8):iii–v, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001121.

Anonymous:2009:Cf

[2177] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (8):vi–viii, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001133.

LeSage:2009:SSM

[2178] James LeSage, Sudipto Banerjee, Manfred M. Fischer, and Peter Congdon. Spatial statistics: Methods, models & computation. Computational Statistics & Data Analysis, 53 (8):2781–2785, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005446.

Ceyhan:2009:OPS

[2179] Elvan Ceyhan. Overall and pairwise segregation tests based on nearest neighbor contingency tables. Computational Statistics & Data Analysis, 53 (8):2786–2808, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003897.

Wu:2009:SSM

[2180] Liu-Cang Wu and Hui-Qiong Li. Summary statistics for measuring the relationship among three types of points in multivariate point patterns. Computational Statistics & Data Analysis, 53 (8):2809–2816, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004702.

Assuncao:2009:SDE

[2181] Renato Assunção and Thais Correa.

Surveillance to detect emerging space– time clusters. Computational Statistics & Data Analysis, 53(8):2817– 2830, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004921.

Hossain:2009:AMB

[2182] Md. Monir Hossain and Andrew B. Lawson. Approximate methods in Bayesian point process spatial models. Computational Statistics & Data Analysis, 53(8):2831–2842, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002880.

Cucala:2009:FSS

[2183] Lionel Cucala. A flexible spatial scan test for case event data. Computational Statistics & Data Analysis, 53 (8):2843–2850, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004714.

Zhang:2009:SSS

[2184] Tonglin Zhang and Ge Lin. Spatial scan statistics in loglinear models. Computational Statistics & Data Analysis, 53 (8):2851–2858, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004490.

Bivand:2009:PCG

[2185] Roger Bivand, Werner G. Müller, and Markus Reder. Power calculations for global and local Moran’s I. Computational Statistics & Data Analysis, 53 (8):2859–2872, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003551.

Finley:2009:IPP

[2186] Andrew O. Finley, Huiyan Sang, Sudipto Banerjee, and Alan E. Gelfand. Improving the performance of predictive process modeling for large datasets. Computational Statistics & Data Analysis, 53(8):2873– 2884, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004386.

Bolin:2009:FES

[2187] David Bolin, Johan Lindström, Lars Eklundh, and Finn Lindgren. Fast estimation of spatially dependent temporal vegetation trends using Gaussian Markov random fields. Computational Statistics & Data Analysis, 53 (8):2885–2896, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004453.

Baltagi:2009:THS

[2188] Badi H. Baltagi, Seuck Heun Song, and Jae Hyeok Kwon. Testing for heteroskedasticity and spatial correlation in a random effects panel data model. Computational Statistics & Data Analysis, 53(8):2897– 2922, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003228.

Haining:2009:MSA

[2189] Robert Haining, Jane Law, and Daniel Griffith. Modelling small area counts in the presence of overdispersion and spatial autocorrelation. Computational Statistics & Data Analysis, 53 (8):2923–2937, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003940.

Ugarte:2009:EBF

[2190] M. D. Ugarte, T. Goicoa, and A. F. Militino. Empirical Bayes and fully Bayes procedures to detect high-risk areas in disease mapping. Computational Statistics & Data Analysis, 53 (8):2938–2949, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003034.

MacNab:2009:EBP

[2191] Ying C. MacNab and Yi Lin. On empirical Bayes penalized quasilikelihood inference in GLMMs and in Bayesian disease mapping and ecological modeling. Computational Statistics & Data Analysis, 53(8):2950– 2967, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003800.

Lee:2009:SCM

[2192] Dae-Jin Lee and María Durbán. Smooth-CAR mixed models for spatial count data. Computational Statistics & Data Analysis, 53(8):2968– 2979, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003563.

Smirnov:2009:PMC

[2193] Oleg A. Smirnov and Luc E. Anselin.

An O(N) parallel method of computing the log-Jacobian of the variable transformation for models with spatial interaction on a lattice. Computational Statistics & Data Analysis, 53 (8):2980–2988, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004854.

Choi:2009:STA

[2194] Jungsoon Choi, Montserrat Fuentes, and Brian J. Reich. Spatial–temporal association between fine particulate matter and daily mortality. Computational Statistics & Data Analysis, 53 (8):2989–3000, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308002892.

Hatfield:2009:SST

[2195] Laura A. Hatfield, Richard W. Hoffbeck, Bruce H. Alexander, and Bradley P. Carlin. Spatiotemporal and spatial threshold models for relating UV exposures and skin cancer in the central united states. Computational Statistics & Data Analysis, 53 (8):3001–3015, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730800474X.

Kang:2009:SAS

[2196] Emily L. Kang, Desheng Liu, and Noel Cressie. Statistical analysis of small-area data based on independence, spatial, non-hierarchical, and hierarchical models. Computational Statistics & Data Analysis, 53(8):3016– 3032, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003587.

White:2009:SNC

[2197] Gentry White and Sujit K. Ghosh. A stochastic neighborhood conditional autoregressive model for spatial data. Computational Statistics & Data Analysis, 53(8):3033– 3046, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003885.

Congdon:2009:MIS

[2198] Peter Congdon. Modelling the impact of socioeconomic structure on spatial health outcomes. Computational Statistics & Data Analysis, 53 (8):3047–3056, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947307003945.

Wall:2009:SLC

[2199] Melanie M. Wall and Xuan Liu. Spatial latent class analysis model for spatially distributed multivariate binary data. Computational Statistics & Data Analysis, 53(8):3057– 3069, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003757.

Bonneu:2009:SPP

[2200] Florent Bonneu and Christine Thomas-Agnan. Spatial point process models for location–allocation problems. Computational Statistics & Data Analysis, 53(8):3070–3081, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004696.

Bel:2009:CAS

[2201] L. Bel, D. Allard, J. M. Laurent, R. Cheddadi, and A. Bar-Hen. CART algorithm for spatial data: Application to environmental and ecological data. Computational Statistics & Data Analysis, 53(8):3082– 3093, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800443X.

Wu:2009:IDE

[2202] Jincao Wu, Tasneem H. Patwa, David M. Lubman, and Debashis Ghosh. Identification of differentially expressed spatial clusters using humoral response microarray data. Computational Statistics & Data Analysis, 53(8):3094– 3102, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002259.

Blasius:2009:SIC

[2203] J. Blasius, M. Greenacre, P. J. F. Groenen, and M. van de Velden. Special issue on correspondence analysis and related methods. Computational Statistics & Data Analysis, 53 (8):3103–3106, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308005525.

Greenacre:2009:PTC

[2204] Michael Greenacre. Power transformations in correspondence analysis. Computational Statistics & Data Analysis, 53(8):3107–3116, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947308004350.

Polak:2009:TTS

[2205] Marike Polak, Willem J. Heiser, and Mark de Rooij. Two types of singlepeaked data: Correspondence analysis as an alternative to principal component analysis. Computational Statistics & Data Analysis, 53(8):3117– 3128, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004441.

vandeVelden:2009:SCC

[2206] Michel van de Velden, Patrick J. F. Groenen, and Jeroen Poblome. Seriation by constrained correspondence analysis: a simulation study. Computational Statistics & Data Analysis, 53 (8):3129–3138, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003927.

Warrens:2009:DRD

[2207] Matthijs J. Warrens and Willem J.

Heiser. Diagnostics for regression dependence in tables re-ordered by the dominant correspondence analysis solution. Computational Statistics & Data Analysis, 53(8):3139– 3144, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003848.

Blasius:2009:BB

[2208] Jörg Blasius, Paul H. C. Eilers, and John Gower. Better biplots. Computational Statistics & Data Analysis, 53 (8):3145–3158, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003174.

Takane:2009:RNC

[2209] Yoshio Takane and Sunho Jung.

Regularized nonsymmetric correspondence analysis. Computational Statistics & Data Analysis, 53(8):3159– 3170, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004313.

Zarraga:2009:SAM

[2210] A. Zárraga and B. Goitisolo. Simultaneous analysis and multiple factor analysis for contingency tables: Two methods for the joint study of contingency tables. Computational Statistics & Data Analysis, 53(8):3171– 3182, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002211.

terBraak:2009:ASM

[2211] Cajo J. F. ter Braak, Yiannis Kourmpetis, Henk A. L. Kiers, and Marco C. A. M. Bink. Approximating a similarity matrix by a latent class model: a reappraisal of additive fuzzy clustering. Computational Statistics & Data Analysis, 53(8):3183– 3193, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308004775.

Vichi:2009:CDP

[2212] Maurizio Vichi and Gilbert Saporta. Clustering and disjoint principal component analysis. Computational Statistics & Data Analysis, 53(8):3194– 3208, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308002934.

deRooij:2009:TVM

[2213] Mark de Rooij. Trend vector models for the analysis of change in continuous time for multiple groups. Computational Statistics & Data Analysis, 53 (8):3209–3216, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308004635.

DeSarbo:2009:TWC

[2214] Wayne S. DeSarbo, A. Selin Atalay, and Simon J. Blanchard. A three-way clusterwise multidimensional unfolding procedure for the spatial representation of context dependent preferences. Computational Statistics & Data Analysis, 53(8):3217– 3230, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730800220X.

Vera:2009:DLC

[2215] J. Fernando Vera, Rodrigo Macías, and Willem J. Heiser. A dual latent class unfolding model for two-way twomode preference rating data. Computational Statistics & Data Analysis, 53 (8):3231–3244, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003460.

Caussinus:2009:BR

[2216] Henri Caussinus. Book review. Computational Statistics & Data Analysis, 53 (8):3245–3246, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308003150.

Reed:2009:BR

[2217] Ken Reed. Book review. Computational Statistics & Data Analysis, 53 (8):3247–3248, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003216.

leRoux:2009:BR

[2218] Niël J. le Roux. Book review. Computational Statistics & Data Analysis, 53 (8):3249–3251, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000157.

deFalguerolles:2009:BR

[2219] Antoine de Falguerolles. Book review. Computational Statistics & Data Analysis, 53(8):3252–3255, June 15, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308000273.

Lebart:2009:BR

[2220] Ludovic Lebart. Book review. Computational Statistics & Data Analysis, 53 (8):3256–3257, June 15, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947308003423.

Anonymous:2009:EBg

[2221] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53(9):iii–v, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001480.

Anonymous:2009:Cg

[2222] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (9):vi–vii, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001492.

Zeileis:2009:ERS

[2223] Achim Zeileis, Kurt Hornik, and Paul Murrell. Escaping RGBland: Selecting colors for statistical graphics. Computational Statistics & Data Analysis, 53 (9):3259–3270, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947308005549.

Rufo:2009:IEF

[2224] M. J. Rufo, J. Martín, and C. J. Pérez.

Inference on exponential families with mixture of prior distributions. Computational Statistics & Data Analysis, 53 (9):3271–3280, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000413.

Lin:2009:ECN

[2225] Chien-Tai Lin and N. Balakrishnan.

Exact computation of the null distribution of a test for multiple outliers in an exponential sample. Computational Statistics & Data Analysis, 53 (9):3281–3290, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000590.

Carbonell:2009:GTV

[2226] F. Carbonell, K. J. Worsley, N. J. Trujillo-Barreto, and M. Vega-Hernandez. The geometry of time-varying crosscorrelation random fields. Computational Statistics & Data Analysis, 53 (9):3291–3304, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000619.

Saha:2009:THM

[2227] Krishna K. Saha and Roger Bilisoly. Testing the homogeneity of the means of several groups of count data in the presence of unequal dispersions. Computational Statistics & Data Analysis, 53(9):3305–3313, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000358.

Tian:2009:NIS

[2228] Guo-Liang Tian, Kai Wang Ng, Kai-Can Li, and Ming Tan. Non-iterative sampling-based Bayesian methods for identifying changepoints in the sequence of cases of haemolytic uraemic syndrome. Computational Statistics & Data Analysis, 53(9):3314– 3323, July 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000395.

Su:2009:CIQ

[2229] Steve Su. Confidence intervals for quantiles using generalized lambda distributions. Computational Statistics & Data Analysis, 53(9):3324– 3333, July 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000437. See

erratum [2346].

Yu:2009:EIB

[2230] Lei Yu, Suzanne L. Tyas, David A.

Snowdon, and Richard J. Kryscio. Effects of ignoring baseline on modeling transitions from intact cognition to dementia. Computational Statistics & Data Analysis, 53(9):3334– 3343, July 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000449.

Martinez-Camblor:2009:NPS

[2231] Pablo Martínez-Camblor and Jacobo de Uña-Álvarez. Non-parametric ksample tests: Density functions vs distribution functions. Computational Statistics & Data Analysis, 53(9): 3344–3357, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000450.

Gosoniu:2009:MMR

[2232] L. Gosoniu, P. Vounatsou, N. Sogoba, N. Maire, and T. Smith. Mapping malaria risk in West Africa using a Bayesian nonparametric nonstationary model. Computational Statistics & Data Analysis, 53(9): 3358–3371, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000462.

Bodnar:2009:SCM

[2233] Olha Bodnar, Taras Bodnar, and Yarema Okhrin. Surveillance of the covariance matrix based on the properties of the singular Wishart distribution. Computational Statistics & Data Analysis, 53(9):3372– 3385, July 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000589.

Quenneville:2009:NPI

[2234] B. Quenneville, S. Fortier, and C. Gagné. A non-parametric iterative smoothing method for benchmarking and temporal distribution. Computational Statistics & Data Analysis, 53 (9):3386–3396, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000607.

Cordeiro:2009:DPR

[2235] Gauss M. Cordeiro and Alexandre B. Simas. The distribution of Pearson residuals in generalized linear models. Computational Statistics & Data Analysis, 53(9):3397– 3411, July 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000620.

Liu:2009:SCA

[2236] Yafen Liu, Zhen He, Lianjie Shu, and Zhang Wu. Statistical computation and analyses for attribute events. Computational Statistics & Data Analysis, 53 (9):3412–3425, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000644.

Henze:2009:IDF

[2237] Norbert Henze, Yakov Nikitin, and Bruno Ebner. Integral distributionfree statistics of L_{p}-type and their asymptotic comparison. Computational Statistics & Data Analysis, 53 (9):3426–3438, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000656.

Tiku:2009:EHT

[2238] Moti L. Tiku and Birdal Şenoğlu.

Estimation and hypothesis testing in BIB design and robustness. Computational Statistics & Data Analysis, 53 (9):3439–3451, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000668.

Park:2009:SCF

[2239] Juhyun Park, Theo Gasser, and Valentin Rousson. Structural components in functional data. Computational Statistics & Data Analysis, 53 (9):3452–3465, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000681.

Wong:2009:ELB

[2240] Heung Wong, Feng Liu, Min Chen, and Wai Cheung Ip. Empirical likelihood based diagnostics for heteroscedasticity in partial linear models. Computational Statistics & Data Analysis, 53 (9):3466–3477, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000693.

Xie:2009:STZ

[2241] Feng-Chang Xie, Bo-Cheng Wei, and Jin-Guan Lin. Score tests for zeroinflated generalized Poisson mixed regression models. Computational Statistics & Data Analysis, 53(9):3478– 3489, July 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900070X.

Alvarez:2009:AMC

[2242] Jacobo de Uña Álvarez and Javier Roca Pardiñas. Additive models in censored regression. Computational Statistics & Data Analysis, 53(9):3490– 3501, July 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000474.

Soofi:2009:EPE

[2243] E. S. Soofi, P. C. Nystrom, and M. Yasai-Ardekani. Executives’ perceived environmental uncertainty shortly after 9/11. Computational Statistics & Data Analysis, 53(9): 3502–3515, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000577.

Mostafa:2009:NCI

[2244] Mohamed M. Mostafa and Rajan Nataraajan. A neuro-computational intelligence analysis of the ecological footprint of nations. Computational Statistics & Data Analysis, 53 (9):3516–3531, July 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000991.

Anonymous:2009:EBh

[2245] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53(10):iii–v, August 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900190X.

Anonymous:2009:Ch

[2246] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (10):vi–vii, August 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001819.

Scheipl:2009:LAB

[2247] Fabian Scheipl and Thomas Kneib.

Locally adaptive Bayesian P -splines with a Normal–Exponential–Gamma prior. Computational Statistics & Data Analysis, 53(10):3533–3552, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001030.

Lai:2009:GIG

[2248] Yongzeng Lai. Generating inverse Gaussian random variates by approximation. Computational Statistics & Data Analysis, 53(10):3553– 3559, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001078.

Parsi:2009:EVN

[2249] Safar Parsi and Ismihan Bairamov. Expected values of the number of failures for two populations under joint Type-II progressive censoring. Computational Statistics & Data Analysis, 53(10): 3560–3570, August 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001236.

Loisel:2009:FIR

[2250] Sébastien Loisel and Marina Takane. Fast indirect robust generalized method of moments. Computational Statistics & Data Analysis, 53(10):3571– 3579, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001285.

Basak:2009:ETP

[2251] Prasanta Basak, Indrani Basak, and N. Balakrishnan. Estimation for the three-parameter lognormal distribution based on progressively censored data. Computational Statistics & Data Analysis, 53(10):3580– 3592, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001303.

Perez:2009:NPP

[2252] Ana Pérez, Esther Ruiz, and Helena Veiga. A note on the properties of power-transformed returns in long-memory stochastic volatility models with leverage effect. Computational Statistics & Data Analysis, 53(10): 3593–3600, August 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000711.

Cao:2009:MCG

[2253] Kim-Anh Lê Cao, Agnès Bonnet, and Sébastien Gadat. Multiclass classification and gene selection with a stochastic algorithm. Computational Statistics & Data Analysis, 53(10): 3601–3615, August 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309000723.

Ugarte:2009:SSS

[2254] M. D. Ugarte, T. Goicoa, A. F. Militino, and M. Durbán. Spline smoothing in small area trend estimation and forecasting. Computational Statistics & Data Analysis, 53(10):3616– 3629, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309000747.

Cui:2009:GST

[2255] Yin Cui, Yuejiao Fu, and Abdulkadir Hussein. Group sequential testing of homogeneity in genetic linkage analysis. Computational Statistics & Data Analysis, 53(10):3630– 3639, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900098X.

White:2009:CNB

[2256] Laura Forsberg White, Marco Bonetti, and Marcello Pagano. The choice of the number of bins for the M statistic. Computational Statistics & Data Analysis, 53(10):3640–3649, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001017.

Wong:2009:NIE

[2257] Augustine C. M. Wong and Yan Yan Wu. A note on interval estimation of P (X < Y) using lower record data from the generalized exponential distribution. Computational Statistics & Data Analysis, 53(10):3650– 3658, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001029.

Wu:2009:EBP

[2258] Shuo-Jye Wu and Coşkun Kuş. On estimation based on progressive firstfailure-censored sampling. Computational Statistics & Data Analysis, 53(10):3659–3670, August 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309001054.

Fu:2009:LIA

[2259] Ying-Zi Fu, Nian-Sheng Tang, and Xing Chen. Local influence analysis of nonlinear structural equation models with nonignorable missing outcomes from reproductive dispersion models. Computational Statistics & Data Analysis, 53(10):3671– 3684, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001066.

Hossain:2009:FAL

[2260] Ahmed Hossain, Joseph Beyene, Andrew R. Willan, and Pingzhao Hu. A flexible approximate likelihood ratio test for detecting differential expression in microarray data. Computational Statistics & Data Analysis, 53(10): 3685–3695, August 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001212.

Wang:2009:QRC

[2261] You-Gan Wang, Quanxi Shao, and Min Zhu. Quantile regression without the curse of unsmoothness. Computational Statistics & Data Analysis, 53(10): 3696–3705, August 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001224.

Chen:2009:RPP

[2262] Tao Chen, Elaine Martin, and Gary Montague. Robust probabilistic PCA with missing data and contribution analysis for outlier detection. Computational Statistics & Data Analysis, 53(10):3706–3716, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001248.

Rodriguez-Avi:2009:GWR

[2263] J. Rodríguez-Avi, A. Conde-Sánchez, A. J. Sáez-Castillo, M. J. Olmo-Jiménez, and A. M. Martínez-Rodríguez. A generalized Waring regression model for count data. Computational Statistics & Data Analysis, 53(10):3717–3725, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900125X.

Jones:2009:MLK

[2264] M. C. Jones and D. A. Henderson. Maximum likelihood kernel density estimation: On the potential of convolution sieves. Computational Statistics & Data Analysis, 53(10):3726– 3733, August 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001261.

Anonymous:2009:EBi

[2265] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53 (11):iii–v, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002102.

Anonymous:2009:Ci

[2266] Anonymous. Contents. Computational Statistics & Data Analysis, 53(11): vi–vii, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002114.

Kim:2009:ECE

[2267] Ji-Hyun Kim. Estimating classification error rate: Repeated crossvalidation, repeated hold-out and bootstrap. Computational Statistics & Data Analysis, 53(11):3735–3745, September 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001601.

Sorbye:2009:BMF

[2268] Sigrunn H. Sørbye, Kristian Hindberg, Lena R. Olsen, and Håvard Rue. Bayesian multiscale feature detection of log-spectral densities. Computational Statistics & Data Analysis, 53(11):3746–3754, September 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001273.

Zou:2009:CIE

[2269] Guang Yong Zou, Julia Taleban, and Cindy Y. Huo. Confidence interval estimation for lognormal data with application to health economics. Computational Statistics & Data Analysis, 53(11):3755–3764, September 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001315.

Crespi:2009:BMC

[2270] Catherine M. Crespi and W. John Boscardin. Bayesian model checking for multivariate outcome data. Computational Statistics & Data Analysis, 53(11):3765–3772, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309001339.

Rosen:2009:BRM

[2271] Ori Rosen and Wesley K. Thompson. A Bayesian regression model for multivariate functional data. Computational Statistics & Data Analysis, 53(11):3773–3786, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309001352.

Vasconcellos:2009:IAH

[2272] Klaus L. P. Vasconcellos and L. M. Zea Fernandez. Influence analysis with homogeneous linear restrictions. Computational Statistics & Data Analysis, 53(11):3787–3794, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309001364.

Dine:2009:MTM

[2273] Abdessamad Dine, Denis Larocque, and François Bellavance. Multivariate trees for mixed outcomes. Computational Statistics & Data Analysis, 53(11):3795–3804, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309001376.

Rosychuk:2009:PEM

[2274] Rhonda J. Rosychuk and Shofiqul Islam. Parameter estimation in a model for misclassified Markov data

— a Bayesian approach. Computational Statistics & Data Analysis, 53 (11):3805–3816, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309001388.

Chiu:2009:GCM

[2275] Sung Nok Chiu and Kwong Ip Liu. Generalized Cramér–von Mises goodness-of-fit tests for multivariate distributions. Computational Statistics & Data Analysis, 53(11):3817– 3834, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900139X. See

corrigendum [2344].

Villasenor-Alva:2009:BGF

[2276] José A. Villaseñor-Alva and Elizabeth González-Estrada. A bootstrap goodness of fit test for the generalized Pareto distribution. Computational Statistics & Data Analysis, 53(11):3835–3841, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947309001406.

Laurini:2009:SSE

[2277] Fabrizio Laurini and Francesco Pauli. Smoothing sample extremes: the mixed model approach. Computational Statistics & Data Analysis, 53(11):3842– 3854, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001418.

Serneels:2009:PCR

[2278] Sven Serneels and Tim Verdonck. Principal component regression for data containing outliers and missing elements. Computational Statistics & Data Analysis, 53(11):3855– 3863, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001595.

Shih:2009:SEN

[2279] Joanna H. Shih and Shou-En Lu. Semiparametric estimation of a nested random effects model for the analysis of multi-level clustered failure time data. Computational Statistics & Data Analysis, 53(11):3864– 3871, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001662.

Maugis:2009:VSM

[2280] C. Maugis, G. Celeux, and M.- L. Martin-Magniette. Variable selection in model-based clustering: a general variable role modeling. Computational Statistics & Data Analysis, 53(11):3872–3882, September 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001686.

Liang:2009:GSW

[2281] Jiajuan Liang, Man-Lai Tang, and Ping Shing Chan. A generalized Shapiro–Wilk W statistic for testing high-dimensional normality. Computational Statistics & Data Analysis, 53(11):3883–3891, September 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001704.

Temiyasathit:2009:SPO

[2282] Chivalai Temiyasathit, Seoung Bum Kim, and Sun-Kyoung Park. Spatial prediction of ozone concentration profiles. Computational Statistics & Data Analysis, 53(11):3892– 3906, September 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001340.

Anonymous:2009:EBj

[2283] Anonymous. Editorial Board. Computational Statistics & Data Analysis, 53 (12):iii–v, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002850.

Anonymous:2009:Cj

[2284] Anonymous. Contents. Computational Statistics & Data Analysis, 53 (12):vi–ix, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002862.

Wang:2009:PPP

[2285] Xiaoning Wang, Alan Schumitzky, and David Z. D’Argenio. Population pharmacokinetic/pharmacodynamic mixture models via maximum a posteriori estimation. Computational Statistics & Data Analysis, 53(12):3907– 3915, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001649.

You:2009:HLM

[2286] Na You and Chang Xuan Mao. On hierarchical loglinear models in capture– recapture studies. Computational Statistics & Data Analysis, 53(12): 3916–3920, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001650.

Brombin:2009:MAP

[2287] Chiara Brombin and Luigi Salmaso.

Multi-aspect permutation tests in shape analysis with small sample size. Computational Statistics & Data Analysis, 53(12):3921–3931, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001741.

Komarek:2009:NRP

[2288] Arnošt Komárek. A new R package for Bayesian estimation of multivariate normal mixtures allowing for selection of the number of components and interval-censored data. Computational Statistics & Data Analysis, 53(12): 3932–3947, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001935.

Hatjispyros:2009:BNS

[2289] Spyridon J. Hatjispyros, Theodoros Nicoleris, and Stephen G. Walker. A Bayesian nonparametric study of a dynamic nonlinear model. Computational Statistics & Data Analysis, 53(12): 3948–3956, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001972.

Jimenez-Gamero:2009:GFT

[2290] M. D. Jiménez-Gamero, V. Alba-Fernández, J. Muñoz-García, and Y. Chalco-Cano. Goodness-of-fit tests based on empirical characteristic functions. Computational Statistics & Data Analysis, 53(12):3957– 3971, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002308.

Melo:2009:SRT

[2291] Tatiane F. N. Melo, Klaus L. P. Vasconcellos, and Artur J. Lemonte. Some restriction tests in a new class of regression models for proportions. Computational Statistics & Data Analysis, 53(12):3972–3979, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002333.

Caudle:2009:NDE

[2292] Kyle A. Caudle and Edward Wegman. Nonparametric density estimation of streaming data using orthogonal series. Computational Statistics & Data Analysis, 53(12):3980– 3986, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002394.

vonBorries:2009:PCH

[2293] George von Borries and Haiyan Wang. Partition clustering of high dimensional low sample size data based on p-values. Computational Statistics & Data Analysis, 53(12):3987– 3998, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002412.

Heath:2009:NGO

[2294] Jeffrey W. Heath, Michael C. Fu, and Wolfgang Jank. New global optimization algorithms for modelbased clustering. Computational Statistics & Data Analysis, 53(12):3999– 4017, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002473.

Jamalizadeh:2009:RRB

[2295] A. Jamalizadeh, Y. Mehrali, and N. Balakrishnan. Recurrence relations for bivariate t and extended skew-t distributions and an application to order statistics from bivariate t. Computational Statistics & Data Analysis, 53(12):4018– 4027, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002643.

Calderhead:2009:EBF

[2296] Ben Calderhead and Mark Girolami.

Estimating Bayes factors via thermodynamic integration and population MCMC. Computational Statistics & Data Analysis, 53(12):4028– 4045, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002722.

Rokach:2009:TCE

[2297] Lior Rokach. Taxonomy for characterizing ensemble methods in classification tasks: a review and annotated bibliography. Computational Statistics & Data Analysis, 53(12): 4046–4072, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002631.

Ghosh:2009:SBA

[2298] Pulak Ghosh, Lan Huang, Binbing Yu, and Ram C. Tiwari. Semiparametric Bayesian approaches to joinpoint regression for population-based cancer survival data. Computational Statistics & Data Analysis, 53(12): 4073–4082, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001431.

Pareek:2009:PCC

[2299] Bhuvanesh Pareek, Debasis Kundu, and Sumit Kumar. On progressively censored competing risks data for Weibull distributions. Computational Statistics & Data Analysis, 53(12): 4083–4094, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001613.

Bianco:2009:RTL

[2300] Ana M. Bianco and Elena Martínez. Robust testing in the logistic regression model. Computational Statistics & Data Analysis, 53(12):4095– 4105, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001637.

Li:2009:LDA

[2301] Lexin Li and Xiangrong Yin. Longitudinal data analysis using sufficient dimension reduction method. Computational Statistics & Data Analysis, 53(12):4106–4115, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167947309001674.

Strickland:2009:EBE

[2302] Chris M. Strickland, Ian. W. Turner, Robert Denham, and Kerrie L. Mengersen. Efficient Bayesian estimation of multivariate state space models. Computational Statistics & Data Analysis, 53(12):4116–4125, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001698.

Frey:2009:CBC

[2303] Jesse Frey. Confidence bands for the CDF when sampling from a finite population. Computational Statistics & Data Analysis, 53(12):4126– 4132, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001716.

Pai:2009:MLE

[2304] Jeffrey Pai and Nalini Ravishanker.

Maximum likelihood estimation in vector long memory processes via EM algorithm. Computational Statistics & Data Analysis, 53(12):4133– 4142, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900173X.

Russo:2009:IDN

[2305] Cibele M. Russo, Gilberto A. Paula, and Reiko Aoki. Influence diagnostics in nonlinear mixed-effects elliptical models. Computational Statistics & Data Analysis, 53(12):4143– 4156, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001753.

Liu:2009:TSH

[2306] Qin Liu, Nancy R. Cook, Anna Bergström, and Chung-Cheng Hsieh. A two-stage hierarchical regression model for meta-analysis of epidemiologic nonlinear dose–response data. Computational Statistics & Data Analysis, 53(12):4157–4167, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001765.

Schnabel:2009:OES

[2307] Sabine K. Schnabel and Paul H. C. Eilers. Optimal expectile smoothing. Computational Statistics & Data Analysis, 53(12):4168–4177, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001911.

Minoda:2009:ECS

[2308] Yuta Minoda and Takemi Yanagimoto. Estimation of a common slope in a gamma regression model with multiple strata: an empirical Bayes method. Computational Statistics & Data Analysis, 53(12):4178– 4185, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001947.

Li:2009:IUC

[2309] Jialiang Li, Yingcun Xia, Mari Palta, and Anoop Shankar. Impact of unknown covariance structures in semiparametric models for longitudinal data: an application to Wisconsin diabetes data. Computational Statistics & Data Analysis, 53(12):4186– 4197, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001959.

Chakraborty:2009:BBK

[2310] Sounak Chakraborty. Bayesian binary kernel probit model for microarray based cancer classification and gene selection. Computational Statistics & Data Analysis, 53(12):4198– 4209, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001960.

Wang:2009:CCC

[2311] Hsiuying Wang. Comparison of p control charts for low defective rate. Computational Statistics & Data Analysis, 53(12):4210– 4220, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309001996.

Jhun:2009:BLD

[2312] Myoungshic Jhun and Inkyung Choi. Bootstrapping least distance estimator in the multivariate regression model. Computational Statistics & Data Analysis, 53(12):4221– 4227, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016794730900200X.

Woodcock:2009:DPS

[2313] Simon D. Woodcock and Gary Benedetto. Distribution-preserving statistical disclosure limitation. Computational Statistics & Data Analysis, 53(12):4228–4242, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167947309002011.

Umashanger:2009:EMC

[2314] T. Umashanger and T. N. Sriram. L_{2} estimation of mixture complexity for count data. Computational Statistics & Data Analysis, 53(12): 4243–4254, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002023.

Gutierrez-Pena:2009:OPM

[2315] Eduardo Gutiérrez-Peña, Raúl Rueda, and Alberto Contreras-Cristán. Objective parametric model selection procedures from a Bayesian nonparametric perspective. Computational Statistics & Data Analysis, 53(12):4255– 4265, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002035.

Chiu:2009:HTS

[2316] Sung Nok Chiu and Ling Wang. Homogeneity tests for several Poisson populations. Computational Statistics & Data Analysis, 53(12):4266– 4278, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002047.

Pincon:2009:NLT

[2317] Claire Pinçon and Odile Pons. Nonparametric linear tests with multiple events. Computational Statistics & Data Analysis, 53(12):4279– 4289, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002199.

Park:2009:PTD

[2318] P. J. Park, J. Manjourides, M. Bonetti, and M. Pagano. A permutation test for determining significance of clusters with applications to spatial and gene expression data. Computational Statistics & Data Analysis, 53(12): 4290–4300, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002217.

Galimberti:2009:PFM

[2319] Giuliano Galimberti, Angela Montanari, and Cinzia Viroli. Penalized factor mixture analysis for variable selection in clustered data. Computational Statistics & Data Analysis, 53(12): 4301–4310, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002229.

McGrory:2009:FBA

[2320] C. A. McGrory, A. N. Pettitt, and M. J. Faddy. A fully Bayesian approach to inference for Coxian phasetype distributions with covariate dependent mean. Computational Statistics & Data Analysis, 53(12):4311– 4321, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002242.

Fernandez-Alcala:2009:SID

[2321] R. M. Fernández-Alcalá, J. Navarro-Moreno, and J. C. Ruiz-Molina. Statistical inference for doubly stochastic multichannel Poisson processes: a PCA approach. Computational Statistics & Data Analysis, 53(12):4322– 4331, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002254.

Jeske:2009:CTT

[2322] Daniel R. Jeske, Veronica Montes De Oca, Wolfgang Bischoff, and Mazda Marvasti. Cusum techniques for timeslot sequences with applications to network surveillance. Computational Statistics & Data Analysis, 53(12): 4332–4344, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900228X.

DeOliveira:2009:SPI

[2323] Victor De Oliveira and Changxiang Rui. On shortest prediction intervals in log-Gaussian random fields. Computational Statistics & Data Analysis, 53(12):4345– 4357, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002291.

Chung:2009:CCI

[2324] Hie-Choon Chung and Chien-Pai Han.

Conditional confidence intervals for classification error rate. Computational Statistics & Data Analysis, 53(12): 4358–4369, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002357.

Yousef:2009:NET

[2325] Waleed A. Yousef, Subrata Kundu, and Robert F. Wagner. Nonparametric estimation of the threshold at an operating point on the ROC curve. Computational Statistics & Data Analysis, 53(12):4370– 4383, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002369.

Scarrott:2009:SMT

[2326] C. J. Scarrott and G. Tunnicliffe Wilson. Spatial multi-taper spectrum estimation for nuclear reactor modelling. Computational Statistics & Data Analysis, 53(12):4384– 4402, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002370.

Xie:2009:DSN

[2327] Feng-Chang Xie, Jin-Guan Lin, and Bo-Cheng Wei. Diagnostics for skewnormal nonlinear regression models with AR(1) errors. Computational Statistics & Data Analysis, 53(12): 4403–4416, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002400.

Shen:2009:SAS

[2328] Pao sheng Shen. Semiparametric analysis of survival data with left truncation and right censoring. Computational Statistics & Data Analysis, 53(12):4417–4432, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002424.

Chahkandi:2009:SLD

[2329] M. Chahkandi and M. Ganjali. On some lifetime distributions with decreasing failure rate. Computational Statistics & Data Analysis, 53(12): 4433–4440, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002436.

Lemonte:2009:BSN

[2330] Artur J. Lemonte and Gauss M.

Cordeiro. Birnbaum–Saunders nonlinear regression models. Computational Statistics & Data Analysis, 53(12): 4441–4452, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016794730900245X.

Peltonen:2009:VAC

[2331] Jaakko Peltonen, Jarkko Venna, and Samuel Kaski. Visualizations for assessing convergence and mixing of Markov chain Monte Carlo simulations. Computational Statistics & Data Analysis, 53(12):4453–4470, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002461.

Arbia:2009:NCS

[2332] Giuseppe Arbia and Gianfranco Piras. A new class of spatial concentration measures. Computational Statistics & Data Analysis, 53(12):

4471–4481, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002485.

Silva:2009:LEW

[2333] Giovana O. Silva, Edwin M. M. Ortega, and Gauss M. Cordeiro. A log-extended Weibull regression model. Computational Statistics & Data Analysis, 53(12):4482–4489, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002503.

Richard:2009:MFD

[2334] Patrick Richard. Modified fast double sieve bootstraps for ADF tests. Computational Statistics & Data Analysis, 53(12):4490–4499, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002515.

Hawkins:2009:PRF

[2335] Douglas M. Hawkins and Dost Muhammad Khan. A procedure for robust fitting in nonlinear regression. Computational Statistics & Data Analysis, 53(12):4500–4507, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309002527.

Hirotsu:2009:CRC

[2336] C. Hirotsu. Clustering rows and/ or columns of a two-way contingency table and a related distribution theory. Computational Statistics & Data Analysis, 53(12):4508– 4515, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002539.

Rosadi:2009:TIH

[2337] Dedi Rosadi. Testing for independence in heavy-tailed time series using the codifference function. Computational Statistics & Data Analysis, 53(12): 4516–4529, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002540.

Chan:2009:NDM

[2338] Jennifer S. K. Chan, Doris Y. P. Leung, S. T. Boris Choy, and Wai Y. Wan. Nonignorable dropout models for longitudinal binary data with random effects: an application of Monte Carlo approximation through the Gibbs output. Computational Statistics & Data Analysis, 53(12): 4530–4545, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002552.

Wieringa:2009:GPA

[2339] Jaap Wieringa, Garmt Dijksterhuis, John Gower, and Frederieke van Perlo. Generalised Procrustes analysis with optimal scaling: Exploring data from a power supplier. Computational Statistics & Data Analysis, 53(12):4546– 4554, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001327.

Arnold:2009:MOS

[2340] Barry C. Arnold, Enrique Castillo, and José María Sarabia. On multivariate order statistics. application to ranked set sampling. Computational Statistics & Data Analysis, 53(12): 4555–4569, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309001984.

Johnson:2009:BAD

[2341] Timothy D. Johnson and Morand Piert. A Bayesian analysis of dual autoradiographic images. Computational Statistics & Data Analysis, 53(12): 4570–4583, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002205.

Ahmed:2009:EGN

[2342] Rashid Ahmed, Munir Akhtar, and M. H. Tahir. Economical generalized neighbor designs for use in serology. Computational Statistics & Data Analysis, 53(12):4584–4589, October 1, 2009. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002230.

Jones:2009:NBM

[2343] Mark Jones, Annette Dobson, Mark Onslow, and Brenda Carey. Negative binomial mixed models for analysis of stuttering rates. Computational Statistics & Data Analysis, 53(12): 4590–4600, October 1, 2009. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309002680.

Chiu:2010:CGC

[2344] Sung Nok Chiu and Kwong Ip Liu. Corrigendum to “Generalized Cramér–von Mises goodness-of-fit tests for multivariate distributions” [comput. statist. data anal. 53 (2009) 3817–3834]. Computational Statistics & Data Analysis, 54(2):607, February 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309003296. See

[2275].

Vonthein:2010:UTR

[2345] Reinhard Vonthein and Andreas Ziegler. On the use of the terms repeatability and reproducibility regarding “Reproducibility of genotypes as measured by the Affymetrix GeneChip(R) 100K Human Mapping Array Set” by Fridley and colleagues (2008) Comput. Stat. Data Anal. 52:5367–74. Computational Statistics & Data Analysis, 54(4): 803, April 1, 2010. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167947309004447. See

[1925].

Su:2010:ECI

[2346] Steve Su. Erratum to: “Confidence intervals for quantiles using generalized lambda distributions” [comput. statist. data anal. 53 (2009) 3324–3333]. Computational Statistics & Data Analysis, 54(5):1404, May 1, 2010. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947309004241. See

[2229].

Ospina:2011:EIP

[2347] Raydonal Ospina, Francisco Cribari-Neto, and Klaus L. P. Vasconcellos. Erratum to “Improved point and interval estimation for a beta regression model” [comput. statist. data anal. 51 (2006) 960–981]. Computational Statistics & Data Analysis, 55(7):2445, July 1, 2011. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311000740. See

[1079].

Said:2011:RNS

[2348] Y. H. Said, E. J. Wegman, W. K. Sharabati, and J. T. Rigsby. Retraction notice to “Social networks of author–coauthor relationships” [comput. statist. data anal. 52 (4) 2177–2184]. Computational Statistics & Data Analysis, 55(12):3386, December 1, 2011. CODEN CS-DADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167947311002799. See

Ferraty:2006:NFD

[2349] Frédéric Ferraty and Philippe Vieu.

Nonparametric functional data analysis: theory and practice. Springer series in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-30369-3. xx + 258 pp. LCCN QA278.8 .F47 2006. URL http:

//www.loc.gov/catdir/toc/fy0709/
2005936519.html;   http://www.loc.
gov/catdir/enhancements/fy0663/
2005936519-d.html.